### v1

# Union of Finitely Generated Congruences on Ground Term Algebra

Abstract
We show that for any ground term equation systems $E$ and $F$, (1) the union of the generated congruences by $E$ and $F$ is a congruence on the ground term algebra if and only if there exists a ground term equation system $H$ such that the congruence generated by $H$ is equal to the union of the congruences generated by $E$ and $F$ if and only if the congruence generated by the union of $E$ and $F$ is equal to the union of the congruences generated by $E$ and $F$, and (2) it is decidable in square time whether the congruence generated by the union of $E$ and $F$ is equal to the union of the congruences generated by $E$ and $F$, where the size of the input is the number of occurrences of symbols in $E$ plus the number of occurrences of symbols in $F$.
keywords:
finitely generated congruence on the ground term algebra, reduced ground term rewrite system , bottom-up tree automaton , square time decision algorithm
journal: some journal
[…]
Department of Foundations of Computer Science University of Szeged, addressline=Arpad ter 2, city=Szeged, postcode=H-6720, country=Hungary
{highlights}
We can decide in square time whether the union of two finitely generated congruences on the ground term algebra is also a congruence.
If the answer is ’yes’, then the union of the two generator sets of ground equations generates the union of the congruences.
1 Introduction
Vágvölgyi [1] showed that it is decidable for any given ground term equation systems (GTESs for short) $E$ and $F$ whether there is a GTES $H$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cap\mathop{\leftrightarrow}^{*}_{F}$. If the answer is yes, then we can construct such a GTES $H$. We show the dual to the above result, first we show that for any GTESs $E$ and $F$, $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra if and only if there exists a GTES $H$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Second, we show that it is decidable in $O(n^{2})$ time for any given GTESs $E$ and $F$ whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n$ is the number of occurrences of symbols in $E$ plus the number of occurrences of symbols in $F$.
For each signature $\Sigma$ without constants, the set of ground terms over $\Sigma$ is the empty set, each GTES over $\Sigma$ is the empty set, and our results above are obvious. Thus, throughout the paper we consider signatures with constants. We now sketch the proof of our main result. Let $E$ and $F$ be GTESs over $\Sigma$, and let $W$ denote the set of subterms appearing in the GTESs $E$ and $F$. Fülöp and Vágvölgyi [2] presented a fast ground completion algorithm which, given GTESs $E$ and $F$ over a signature $\Sigma$, produces in $O(n\,\mathrm{log}\;n)$ time a finite set $C\langle E,F\rangle$ of new constants and a reduced ground term rewrite system (GTRS for short) $R\langle E,F\rangle$ over $\Sigma$ extended with the new constants such that $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$. Here $n$ stands for the number of occurrences of symbols in $E$ plus number of occurrences of symbols in $F$. Furthermore, the set $C\langle E,F\rangle$ of new constants is the set of the equivalence classes of the restriction $\Theta\langle E,F\rangle$ of $\mathop{\leftrightarrow}^{*}_{E}$ to the set $W$. For this purpose we proceed as follows:
[…]
We compute the following sets and relations and dpwpa in $O(n\,\mathrm{log}\,n)$ time:
  • the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$;
  • the set of nullary symbols $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$;
  • the GTRS $R\langle E\cup F,\emptyset\rangle$ stored in a red-black tree;
  • the GTRS $R\langle E,F\rangle$ stored in a red-black tree;
  • the GTRS $R\langle F,E\rangle$ stored in a red-black tree;
  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
Then we decide in $O(n)$ time whether $R\langle E,F\rangle$ and $R\langle F,E\rangle$ are total. Accordingly, we distinguish the following four main cases.
Main Case 1: $\Sigma$ is a unary signature. In this main case $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if for each $a\in C\langle E\cup F,\emptyset\rangle$, Condition (i) or Condition (ii) hold:
  • (i)
    • (a)
      $a\in C\langle E,F\rangle$ and
    • (b)
      for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]\rangle=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
  • (ii)
    • (a)
      $a\in C\langle F,E\rangle$ and
    • (b)
      for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
We decide the above formulas by carrying out a nested partial depth first search on the auxiliary dpwpa $AUX[E;F]$. In the outer loop we carry out a partial depth first search on $AUX[E;F]$ and decide Conditions (i).(a) and (ii).(a). In the inner loop we visit those vertices which are positive-step reachable from the current vertex of the outer loop, and examine at each visited vertex $a$ the value of the Boolean attribute $keeps_{E}$ (respectively $keeps_{F}$), which tells us whether $R\langle E,F\rangle$ (resp. $R\langle F,E\rangle$) keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $a$. In this way we decide Condition (ii).(a) (resp. Condition (ii).(b)). In such manner, we decide in additional $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Main Case 2: Both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total. Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class. We decide in additional $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Main Case 3: $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total. Without loss of generality we may assume that $R\langle E,F\rangle$ is total. Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if for each $a\in C\langle E\cup F,\emptyset\rangle$,
  • (1)
    $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$ and
  • (2)
    if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
In $O(n)$ time, we decide set memberships $a\in C\langle E,F\rangle$ and $a\in C\langle F,E\rangle$ for all $a\in C\langle E\cup F,\emptyset\rangle$. Therefore we decide in $O(n)$ time whether Condition (1) holds for all $a\in C\langle E\cup F,\emptyset\rangle$. We decide in $O(n)$ time whether Condition (2) holds for all $a\in C\langle E\cup F,\emptyset\rangle$ by carrying out a partial depth first search on the auxiliary dpwpa $AUX[E;F]$ by narrowing down the set of vertices from which we start a session of the search. We start the search sessions from those $C\langle E\cup F,\emptyset\rangle$ vertices $a$ which are not in $C\langle E,F\rangle$ and have not been visited previously. If we start a search session from a vertex $a$ and there exists a cycle leading from the vertex $a$ to itself, then along this cycle Algorithm PDFS visits vertex $a$ for a second time. During the current session of the search we find a vertex $b$ with false value for the attribute $keeps_{F}$ if and only if Condition (2) does not hold for the vertex $a\in C\langle E\cup F,\emptyset\rangle$ from which we started the current session of the search. We decide whether Condition (1) holds for all $a\in C\langle E\cup F,\emptyset\rangle$ and whether Conditions (2) holds for all $a\in C\langle E\cup F,\emptyset\rangle$ in $O(n)$ time, in this way, we decide in additional $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Main Case 4: $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total. Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$ or $F\subseteq\mathop{\leftrightarrow}^{*}_{E}$ if and only if $E\subseteq\Theta\langle F,E\rangle$ or $F\subseteq\Theta\langle E,F\rangle$. In this case, by direct inspection of the representations $\tau\langle E,F\rangle$ of $E$, $\tau\langle F,E\rangle$ of $F$, $\rho\langle E,F\rangle$ of ${\Theta\langle E,F\rangle}$ and $\rho\langle F,E\rangle$ of ${\Theta\langle F,E\rangle}$ in the subterm dag for $E$ and $F$, we decide in additional $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Summing up the previous thoughts, our decision algorithm goes as follows: We compute the following sets and relations and dpwpa in $O(n\,\mathrm{log}\,n)$ time:
  • the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$;
  • the set of nullary symbols $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$;
  • the GTRS $R\langle E\cup F,\emptyset\rangle$ stored in a red-black tree;
  • the GTRS $R\langle E,F\rangle$ stored in a red-black tree;
  • the GTRS $R\langle F,E\rangle$ stored in a red-black tree;
  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
We decide whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total.
  • If both of them are total, then we proceed as in main Case 2.
  • Otherwise,
    • if $\Sigma$ is a unary alphabet, then we procced as in Main Case 1.
    • If $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total, the we proceed as in Main Case 3.
    • If $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total, then we proceed as in Main Case 4.
In Main Case 1, the overall complexity of our computation is $O(n^{2})$. In each of Main Cases 2, 3, and 4, the overall complexity of our computation is $O(n\,\mathrm{log}\,n)$. Main Case 2 overlaps with Main Case 1 and Main Case 3, and is a subcase of the union of Main Case 1 and Main Case 3.
We now describe an alternative approach for our decision problem. We can think of the GTRSs $R\langle E,F\rangle$, $R\langle F,E\rangle$, and $R\langle E\cup F,\emptyset\rangle$ as the sets of rewrite rules of bottom-up tree automata. Thence we can also apply the decision algorithm of Champavère et al [5]. Their algorithm decides inclusion between two tree languages recognized by given bottom-up tree automata. Applying their algorithm to our decision problem in each of Main Cases 1, 2, and 3 yields an alternative decision algorithm, which is more sophisticated than ours, and in general takes more time than ours, it runs in $O(n^{3})$ time.
For any given GTESs $E$ and $F$, we can construct the GTES $E\cup F$ in $O(n\,\mathrm{log}\,n)$ time, where $n$ is as above.
The rest of this paper is organized as follows: In Section 2, we recall some basic definitions and fix the notations. In Section 3, we start our running examples. We consider in each main case an example of GTESs $E$ and $F$ such that $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ and an example for GTESs $E$ and $F$ such that $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$. In Section 4, we recall the fast ground completion algorithm of Fülöp and Vágvölgyi [2]. In Section 5, we continue our running examples. In Section 6, we define the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Then we compute $AUX[E;F]$ for our running examples. In Section 7, we further continue our running examples. In Section 8, we construct the auxiliary dpwpa for the GTESs $E$ and $F$. In Section 9, we show some preparatory results. In Section 10, we define the notion that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. In Section 11, we define the four main cases of our decision algorithm. We show that we can decide which Main Cases hold in $O(n\,\mathrm{log}\,n)$ time. In Sections 12, 13, 14, and 15, we study Main Cases 1, 2, 3, and 4, respectively one by one. For each main case, we present our decidability and complexity results, and we continue our relevant running examples. In Section 16, we sum up our decidability and complexity results. In Section 17 we present our concluding remarks.
2 Preliminaries
In this section, we present a brief review of the notions, notation and preliminary results used in the paper.
Algorithm concepts
[…]
Proposition 2.1.
[8, Section 13] The search operation takes $O(\mathrm{log}\;n)$ time on a red-black tree, where $n$ is the size of the tree.
Proposition 2.2.
[8, Section 13] The time-complexity of inserting $n\in{\mathbb{N}}$ elements into an empty red-black tree, not allowing duplicate values in the resulting red-black tree, is $O(n\,\mathrm{log}\;n)$.
Proof.
The time-complexity of inserting a single element into a red-black tree is $O(\mathrm{log}\;n)$, where $n$ is the size of the tree [8]. Consequently it takes $O(n\,\mathrm{log}\;n)$ time to insert $n\in{\mathbb{N}}$ elements into an empty red-black tree. We modify the insertion routine of the red-black tree to not allow duplicate values. The insertion function takes the same amount of time i.e. $O(\mathrm{log}\;n)$, as before, and does not insert a value if it is already in the tree. ∎
Basic notions
The set of nonnegative integers is denoted by ${\mathbb{N}}$, and ${\mathbb{N}}^{*}$ stands for the free monoid generated by ${\mathbb{N}}$ with empty word $\varepsilon$ as identity element. The length of a word $\alpha\in{\mathbb{N}}^{*}$ is defined as the number of integers it contains and is denoted by $\mathrm{length}(\alpha)$. A word $\alpha\in{\mathbb{N}}^{*}$ is a prefix of a word $\beta\in{\mathbb{N}}^{*}$ if there is a word $\gamma\in{\mathbb{N}}^{*}$ such that $\alpha\gamma=\beta$. Moreover, if $\gamma\neq\varepsilon$, then $\alpha$ is a proper prefix of $\beta$.
A binary sequence is a string of ones and zeros, $BIN$ stands for the set of binary sequences. For an integer $\ell\in{\mathbb{N}}$, we write $BIN_{\ell}$ to denote the set of binary sequences of length $\ell$. We define the binary sequence $0^{k}$, $k\in{\mathbb{N}}$ by recursion on $k$. Let $0^{0}=\varepsilon$. For each $k\in{\mathbb{N}}$, let $0^{k+1}=00^{k}$. Let $0\leq k\leq\ell$ and $\alpha\in BIN_{k}$. We pad the binary sequence $\alpha$ with leading zeros to length $\ell$ by producing the binary sequence $0^{\ell-k}\alpha$. We now recall the definition of the lexicographic order on $BIN$. For $\alpha,\beta\in BIN$, we say that $\alpha$ is lexicographically less than $\beta$, denoted by $\alpha<_{lex}\beta$ if either $\alpha$ is a proper prefix of $\beta$, or there exist $\gamma_{1},\gamma_{2},\gamma_{3}\in BIN$ such that $\alpha=\gamma_{1}0\gamma_{2}$ and $\beta=\gamma_{1}1\gamma_{3}$.
[…]
  • Noetherian if there exists no infinite sequence of elements $a_{1},a_{2},a_{3},\ldots$ in $A$ such that $a_{1}{\rightarrow}a_{2}{\rightarrow}a_{3}{\rightarrow}\ldots$,
  • confluent if for any elements $a_{1},a_{2},a_{3}$ in $A$, whenever $a_{1}{\rightarrow}^{*}a_{2}$ and $a_{1}{\rightarrow}^{*}a_{3}$, there exists an element $a_{4}$ in $A$ such that $a_{2}{\rightarrow}^{*}a_{4}$ and $a_{3}{\rightarrow}^{*}a_{4}$,
  • convergent if it is Noetherian and confluent.
[…]
Proposition 2.3.
[…]
If $a\in A$ has a unique normalform, then we denote the latter by $a{\downarrow}$.
Let $\rho$ be an equivalence relation on $A$. Then for every $a\in A$, we denote by $a/_{\rho}$ the $\rho$ equivalence class containing $a$, i.e. $a/_{\rho}=\{\,b\mid a\rho b\,\}$. The classes associated with an equivalence relation $\rho$ are called equivalence classes. For each $B\subseteq A$, let $B/_{\rho}=\{\,b/_{\rho}\mid b\in B\,\}$.
The empty product $\Pi_{i=1}^{0}n_{i}=1$ over ${\mathbb{N}}$ is the result of multiplying no factors. It is by convention equal to $1$. The ceiling function maps a real number $x$ to the least integer greater than or equal to $x$, denoted $\lceil x\rceil$.
Ground terms
A signature $\Sigma$ is a finite set of symbols in which every element $\sigma$ has an arity, $\mathrm{arity}(\sigma)\in{\mathbb{N}}$. For each integer $m\in{\mathbb{N}}$, $\Sigma_{m}$ denotes the elements of $\Sigma$ which have arity $m$. We lose no generality assuming that $\Sigma_{0}\not=\emptyset$, see the Introduction. If $\Sigma=\Sigma_{0}\cup\Sigma_{1}$, then $\Sigma$ is a unary signature. $mar_{\Sigma}$ denotes the maximal arity of the symbols in $\Sigma$.
We denote by $T_{\Sigma}$ the set of ground terms over $\Sigma$. It is the smallest set $U$ for which
  • (i)
    $\Sigma_{0}\subseteq U$ and
  • (ii)
    $\sigma(t_{1},\ldots,t_{m})\in U$ whenever $\sigma\in\Sigma_{m}$ with $m\in{\mathbb{N}}$ and $t_{1},\ldots,t_{m}\in U$.
Ground terms are also called trees. A set of ground terms is also called a tree language. For a ground term $t\in T_{\Sigma}$, the set $sub(t)$ of subterms of $t$ is defined by recursion as follows:
  • (i)
    if $t\in\Sigma_{0}$, then $sub(t)=\{\,t\,\}$,
  • (ii)
    if $t=\sigma(t_{1},\ldots,t_{m})$ for some $m\geq 1$, $\sigma\in\Sigma_{m}$, and $t_{1},\ldots,t_{m}\in T_{\Sigma}$, then we have $sub(t)=\bigcup\{\,sub(t_{i})\mid 1\leq i\leq m\,\}\cup\{\,t\,\}$.
For a finite set $Z\subseteq T_{\Sigma}$ of ground terms,
$sub(Z)=\bigcup\{\,sub(t)\mid t\in Z\,\}$.
For a ground term $t\in T_{\Sigma}$, the height of $t$ is denoted by $height(t)$ and is defined by recursion:
  • (i)
    if $t\in\Sigma_{0}$, then $height(t)=0$, and
  • (ii)
    if $t=\sigma(t_{1},\ldots,t_{m})$ with $m\geq 1$ and $\sigma\in\Sigma_{m}$, then $height(t)=max\{\,height(t_{i})\mid 1\leq i\leq m\,\}+1$.
For a ground term $t\in T_{\Sigma}$, the set of positions $\mathrm{POS}(t)\subseteq{\mathbb{N}}^{*}$ of $t$ is defined by tree induction.
  • If $t\in\Sigma_{0},$ then $\mathrm{POS}(t)=\{\,\varepsilon\,\}$.
  • If $t=\sigma(t_{1},\ldots,t_{m})$ with $\sigma\in\Sigma_{m},\ m>0,$ then $\mathrm{POS}(t)=\{\,\varepsilon\,\}\cup\{\,i\alpha\mid 1\leq i\leq m,\alpha\in\mathrm{POS}(t_{i})\,\}$.
The size of a ground term $t\in T_{\Sigma}$ is $size(t)=|POS(t)|$. For a finite set $Z\subseteq T_{\Sigma}$ of ground terms, $size(Z)$ is the number of occurrences of symbols of $\Sigma$ in $Z$, i.e., $size(Z)=\sum_{t\in Z}size(t)$. For any ground term $t\in T_{\Sigma}$ and position $\alpha\in\mathrm{POS}(t)$, we introduce the subterm $t|_{\alpha}\in T_{\Sigma}$ of $t$ at $\alpha$ as follows:
  • for $t\in\Sigma_{0}$, $t|_{\varepsilon}=t$;
  • for $t=\sigma(t_{1},\ldots,t_{m})$ with $m\geq 1$ and $\sigma\in\Sigma_{m}$ if $\alpha=\varepsilon$ then $t|_{\alpha}=t$, otherwise if $\alpha=i\beta$ with $1\leq i\leq m$, then $t|_{\alpha}=t_{i}|_{\beta}$.
A context $\delta$ over $\Sigma$ can be viewed as an incomplete term, which may contain several empty places or holes. Formally, a context over $\Sigma$ is a ground term containing zero, one or more occurrences of a special constant symbol $\diamondsuit\not\in\Sigma$ denoting holes, i.e., a ground term over the extended signature $\Sigma\cup\{\,\diamondsuit\,\}$. If $\delta$ contains exactly $n\in{\mathbb{N}}$ holes, then we say that $\delta$ is an $n$-context. The set of all $n$-contexts over $\Sigma$ is denoted by $CON_{\Sigma,n}$. For example, let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\$\,\}$, and $\Sigma_{2}=\{\,f,g\,\}$. Then $g(f(\diamondsuit,\#),g(\diamondsuit,\diamondsuit))\in CON_{\Sigma,3}$. Let $\delta$ be an $n$-context over $\Sigma$, and let $t_{1},\ldots,t_{n}\in T_{\Sigma}$ be ground terms. Then $\delta[t_{1},\ldots,t_{n}]$ denotes the result of replacing the holes of $\delta$ from left to right by $t_{1},\ldots,t_{n}$. For example,
$g(f(\diamondsuit,\#),f(\diamondsuit,\diamondsuit))[g(\$,\$),\#,\$]=g(f(g(\$,\$),\#),f(\#,\$))$.
Let $\delta$ be a $1$-context over $\Sigma$, and let $\alpha\in POS(\delta)$ be such that $t|_{\alpha}=\diamondsuit$. Then we call $\alpha$ the address of $\diamondsuit$ in $\delta$, and $addr(\delta)$ stands for $\alpha$. Let $\delta$ be a $1$-context over $\Sigma$ such that $\delta\neq\diamondsuit$, and let $c\in\Sigma_{0}$ be a constant. Then we say that the ground term $\delta[c]$ is a proper extension of the constant $c\in\Sigma_{0}$.
Ground term rewrite systems and equation systems [3]
Let $\Sigma$ be a signature, and $R$ be a finite subset of $T_{\Sigma}\times T_{\Sigma}$. The size $size(R)$ of $R$ is the number of occurrences of symbols of $\Sigma$ in $R$, ie., $size(R)=\sum_{(l,r)\in R}size(l)+size(r)$. Let $\Sigma\langle R\rangle$ denote the signature of those symbols which occur in $R$. For any ground terms $s,t\in T_{\Sigma}$, we say that $s$ rewrites to $t$, and denote this by $s\mathop{\rightarrow}_{R}t$ if there exist an ordered pair $(l,r)$ in $R$, a context $\delta\in CON_{\Sigma,1}$ such that $s=\delta[l]$ and $t=\delta[r]$. We fully specify such a rewrite step by $s\rightarrow_{[\alpha,(l,r)]}t$, where $\alpha=addr(\delta)$. Furthermore, $s\leftrightarrow_{[\alpha,(l,r)]}t$ stands for ($s\rightarrow_{[\alpha,(l,r)]}t$ or $t\rightarrow_{[\alpha,(l,r)]}s$). When we focus on the relations $\mathop{\rightarrow}_{R}$ and $\mathop{\rightarrow}_{R}^{*}$, we call $R$ a ground term rewrite system (GTRS for short), and we call an ordered pair $(l,r)$ a rewrite rule, and denote it as $l\dot{\rightarrow}r$. Moreover, we say that $l$ is the left-hand side and $r$ is the right-hand side of the rewrite rule $l\dot{\rightarrow}r$. When we focus on the symmetric relations $\mathop{\leftrightarrow}_{R}$ and $\mathop{\leftrightarrow}^{*}_{R}$, we call $R$ a ground term equation system (GTES for short), and we denote $(l,r)$ as $l\doteq r$, and call it an equation. Moreover, we denote a GTES as $E$ rather than $R$ [3].
Statement 2.4.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. We can construct the GTES $E\cup F$ in $O(n\,\mathrm{log}\,n)$ time.
Proof.
In $O(n\,\mathrm{log}\,n)$ time we mergesort the GTES $E$ and the GTES $F$ [10, Section 2.2]. Then in $O(n)$ time, scan through each of the sorted relations to produce the result. When the concurrent scan of both GTESs reveals the same equation in both relations, only one of the equations is retained. ∎
Definition 2.5.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then let $\Sigma\langle E,F\rangle=\Sigma\langle E\rangle\cup\Sigma\langle F\rangle$. That is, $\Sigma\langle E,F\rangle$ denotes the signature of those symbols which occur in $E$ or $F$. The Boolean value $FLAG\langle E,F\rangle$ indicates whether $\Sigma=\Sigma\langle E,F\rangle$.
Statement 2.6.
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=size(E)+size(F)$. We can count the elements of $\Sigma\langle E,F\rangle$ and compute the Boolean value $FLAG\langle E,F\rangle$ in $O(n)$ time.
Let $E$ be a GTES over $\Sigma$. Then
$ST\langle E\rangle=\{\,t\in T_{\Sigma}\mid t\in sub(u)\mbox{ or }t\in sub(v)\mbox{ for some }u\doteq v\mbox{ in }E\,\}$
is the set of subterms occurring $E$.
Statement 2.7.
1. Let $Z\subseteq T_{\Sigma}$ be a finite set of ground terms over a signature $\Sigma$. Then $|sub(Z)|\leq size(Z)$.
2. Let $E$ be a GTES over a signature $\Sigma$. Then $|ST\langle E\rangle|\leq size(E)$.
Proof.
By definition. ∎
Let $\Sigma$ be a signature. A $\Sigma$-algebra ${\cal A}$ consists of a carrier set $A$, and a mapping that associates with each symbol $\sigma\in\Sigma_{m}$ a mapping $\sigma^{\cal A}:A^{m}\rightarrow A$ for every $m\in{\mathbb{N}}$ [9]. The $\Sigma$-ground-term algebra ${\cal T}(\Sigma)$ has $T_{\Sigma}$ as carrier set, and interprets the symbols $\sigma\in\Sigma_{m}$ for every $m\in{\mathbb{N}}$ as follows:
$\sigma^{{\cal T}(\Sigma)}:T_{\Sigma}^{m}\rightarrow T_{\Sigma}:(t_{1},\ldots,t_{m})\mapsto\sigma(t_{1},\ldots,t_{m})$.
Let $E$ be a GTES over $\Sigma$. The congruence relation generated by $E$ on the $\Sigma$-ground term algebra ${\cal T}(\Sigma)$ is $\mathop{\leftrightarrow}^{*}_{E}$. It is well known that $\mathop{\leftrightarrow}^{*}_{E}$ is the least congruence on ${\cal T}(\Sigma)$ containing $E$. Let $R$ be a GTRS $R$ over $\Sigma$. We say that GTRS $R$ is equivalent to the GTES $E$ if $\mathop{\leftrightarrow}^{*}_{R}=\mathop{\leftrightarrow}^{*}_{E}$. We say that GTRS $R$ reaches a ground term $s\in T_{\Sigma}$ from a ground term $t\in T_{\Sigma}$ if $t\mathop{\rightarrow}_{R}^{*}s$ holds. We say that GTRS $R$ is Noetherian, (confluent, etc.) if $\mathop{\rightarrow}_{R}$ is Noetherian (confluent, etc.). A term $t\in T_{\Sigma}$ is irreducible with respect to $R$ if it is irreducible with respect to $\mathop{\rightarrow}_{R}$. If a term $t\in T_{\Sigma}$ has a unique $\mathop{\rightarrow}_{R}$ normal form, then the latter is denoted by $t{\downarrow}_{R}$. A GTRS $R$ is reduced if for every rewrite rule $l\dot{\rightarrow}r$ in $R$, $l$ is irreducible with respect to $R-\{\,l\dot{\rightarrow}r\,\}$ and $r$ is irreducible for $R$. We recall the following important result.
Proposition 2.8.
[3] Every reduced GTRS $R$ is convergent.
Definition 2.9.
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $W=ST\langle E\cup F\rangle$. Then let $\Theta\langle E,F\rangle=\mathop{\leftrightarrow}^{*}_{E}\cap W\times W$.
It should be clear that $\Theta\langle E,F\rangle$ is an equivalence relation on $W$. For each $t\in W$, we have $t/_{\Theta\langle E,F\rangle}=t/_{\mathop{\leftrightarrow}^{*}_{E}}\cap W\times W$.
Bottom-up tree automata [11], [12]
Let $\Sigma$ be a signature, a bottom-up tree automaton (bta for short) over $\Sigma$ is a quadruple ${\cal A}=(\Sigma,A,R,A_{f})$, where $A$ is a finite set of states of arity 00, $\Sigma\cap A=\emptyset$, $A_{f}(\subseteq A)$ is the set of final states, $R$ is a finite set of rewrite rules of the following form:
$\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a$ with $m\in{\mathbb{N}}$, $\sigma\in\Sigma_{m}$, $a_{1},\ldots,a_{m},a\in A$.
We consider $R$ as a GTRS over $\Sigma\cup A$. The size $size({\cal A})$ of $\cal A$ is $|A|+size(R)$. The tree language recognized by $\cal A$ is $L({\cal A})=\{\,t\in T_{\Sigma}\mid(\exists a\in A_{f})\;t\mathop{\rightarrow}_{R}^{*}a\}$. We say that ${\cal A}=(\Sigma,A,R,A_{f})$ is a deterministic bta if for any $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, $a_{1},\ldots,a_{m}\in A$, there is at most one rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R$.
Proposition 2.10.
[13] Let $E$ and $F$ be GTESs over a signature $\Sigma$, and let $W=ST(E\cup F)$. For every $p\in W$, we can construct in $O(n\,\mathrm{log}\,n)$ time the deterministic bottom-up tree automaton ${\cal A}=(\Sigma,C\langle E,F\rangle,R\langle E,F\rangle,\{\,p/_{\Theta\langle E,F\rangle}\,\})$, and ${\cal A}$ recognizes the tree language $p/_{\mathop{\leftrightarrow}^{*}_{E}}$.
Proposition 2.11.
[5] Given two btas ${\cal A}=(\Sigma,A,R_{\cal A},A_{f})$ and ${\cal B}=(\Sigma,B,R_{\cal B},B_{f})$, we can decide whether $L({\cal A})\subseteq L({\cal B})$ in $O(size({\cal A})\cdot size({\cal B}))$ time.
Directed pseudographs without parallel arcs [14]
A directed pseudograph without parallel arcs (dpwpa for short) $G$ is an ordered pair $(V,A)$, where $V$ is a finite set, and $A\subseteq V\times V$ is a relation on $V$. The set $V$ is called the vertex set of $G$, and its elements are called vertices. The set $A$ is called the directed arc set of $G$, and its elements are called directed arcs. A walk in $G$ is an alternating sequence $w=x_{1}\,(x_{1},x_{2})\,x_{2}\,(x_{2},x_{3})\,x_{3}\,\ldots\,x_{k-1}\,(x_{k-1},x_{k})\,x_{k}$, $k\geq 1$, of vertices $x_{i}$ and arcs $(x_{j},x_{j+1})$. We say that $w$ is from vertex $x_{1}$ to vertex $x_{k}$. The length of $w$ is the number $k-1$ of its arcs. A trail is a walk $w$ in which all arcs are distinct. If the vertices of a trail $w$ are distinct, then $w$ is a path. If the vertices $x_{1},x_{2},\ldots,x_{k-1}$ of a trail $w$ are distinct, $k\geq 2$, and $x_{1}=x_{k}$, then $w$ is a cycle. We say that a vertex $x$ is reachable from a vertex $y$ if there exists a walk $w$ from vertex $y$ to vertex $x$. In particular, a vertex is reachable from itself. We say that a vertex $x$ is positive-step reachable from a vertex $y$ if there exists a walk $w$ of positive length from vertex $y$ to vertex $x$. Observe that for any vertices $x$ and $y$, the vertex $x$ is positive-step reachable from the vertex $y$ if and only if there exists a path of positive length or a cycle from the vertex $x$ to the vertex $y$.
The adjacency-list representation of a dpwpa $G=(V,A)$ is an array $Adj$ of $|V|$ lists, one list for each vertex in $V$ [8]. For each $x\in V$, the adjacency list $Adj(x)$ contains all the vertices $y$ such that there is a directed arc $(x,y)\in A$. That is, $Adj(x)$ consists of all the vertices adjacent to $x$ in $G$. In pseudocode we treat the list $Adj$ as an attribute of the graph. Let us assume that we are given the set $V$ and the set of arcs. It is well known that there exists an algorithm that builds the adjacency-list representation of the dpwpa $G$ with a time complexity of $O(V)+O(A)$.
Congruence closure on subterm dags [3]
Let $\Sigma$ be a signature. A subterm dag over $\Sigma$ is a directed, labeled and acyclic multigraph $G=(V,A,\lambda)$, where $V$ is the set of vertices, $A\subseteq V\times V$ is the multiset of directed arcs and $\lambda:V\rightarrow\Sigma$ is the labeling function such that if $outdegree(v)=m$ for some vertex $v\in V$, then $\lambda(v)\in\Sigma_{m}$. The multiset of directed arcs leaving a vertex is ordered. We denote by $v/i$ the $i$th successor of a vertex $v$. The vertices of $G$ can be interpreted as ground terms over $\Sigma$ by the function ^$\widehat{\;}:V\rightarrow W$ which is defined by recursion: for a vertex $v\in V$, we have $\widehat{v}=\lambda(v)$ if $outdegree(v)=0$ and we have $\widehat{v}=\sigma(\widehat{v/1},\ldots,\widehat{v/m})$ if $outdegree(v)=m$ for some $m>0$ and $\lambda(v)=\sigma$. Let $E$ be a GTES over $\Sigma$. We say that $G$ is a subterm dag for a GTES $E$ if   $\widehat{\;}$  is a bijection between $V$ and $ST(E)$. Let $E$ and $F$ be GTESs over $\Sigma$. We say that $G$ is a subterm dag for the GTES $E$ and GTES $F$ if   $\widehat{\;}$  is a bijection between $V$ and $ST(E)\cup ST(F)$.
Definition 2.12.
Let $E$ and $F$ be GTESs over a signature $\Sigma$, $G=(V,A,\lambda)$ be a subterm dag for $E$ and $F$. Then let $\tau\langle E,F\rangle=\{\,(u,v)\in V\times V\mid(\hat{u},\hat{v})\in E\,\}$ and let $\rho\langle E,F\rangle$ be the congruence closure of $\tau\langle E,F\rangle$.
We now cite Proposition 3.4 in [3].
Proposition 2.13.
[3] Let $E$ be a GTES over a signature $\Sigma$, and $n=size(E)$. We can construct in $O(n)$ time a subterm dag for the GTES $E$ and the relation $\tau\langle E\rangle$.
By Proposition 2.13, we have the following result.
Statement 2.14.
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=size(E)+size(F)$. We can construct in $O(n)$ time a subterm dag for the GTESs $E$ and $F$.
Definition 2.15.
Let $G=(V,A,\lambda)$ be a subterm dag over $\Sigma$. An equivalence relation $\rho$ on $V$ is a congruence on $V$ if for each ordered pair of vertices $u,v\in V$, the conditions $\lambda(u)=\lambda(v)$ and $u/i\rho v/i$ for each $1\leq i\leq outdegree(v)$ imply that $u\rho v$. Given a relation $\tau$ on $V$, the smallest congruence on $V$ containing $\tau$ is called the congruence closure of $\tau$.
Let $\Sigma$ be a signature, and $G=(V,A,\lambda)$ be a subterm dag. Then for any relation $\tau$ on $V$, there exists a unique congruence closure of $\tau$ on $V$, and algorithms for calculating congruence closures are given in [15] and [4]. The latter can be implemented with a worst-case time of $O(|A|\,\mathrm{log}\,|A|)$; the former runs in $O(|A|^{2})$ time but seems to be faster in practice. We now apply the congruence closure algorithm of [4]. Assume that it performs the weighted union operation [16] on equivalence classes. Then the congruence closure algorithm keeps track the cardinality of each equivalence class during its run, and can output the cardinality of all $\rho$-classes. We rename the congruence classes produced by the congruence closure algorithm from $1$ to the number of the congruence classes in $O(|V|)$ time.
Proposition 2.16.
[4] Let $E$ be a GTES over $\Sigma$. There exists an algorithm which, given a subterm dag $G=(V,A,\lambda)$ for $E$, and a relation $\tau$ on $V$, produces the congruence closure $\rho$ of $\tau$ on $V$ in $O(n\,\mathrm{log}\,n)$ time, where $n=size(E)$. The algorithm gives each $\rho$-class a unique name in the set { $\{\,1,\ldots,|V/\rho|\,\}$, and also outputs an array of the cardinalities of all $\rho$-classes.
Let $E$ be a GTES over $\Sigma$, and $G=(V,A,\lambda)$ be a subterm dag for $E$, and $\rho$ be a congruence on $V$. We shall use a function $\mathrm{FIND}_{\rho}(x)$ that returns a unique name of the congruence class of $\rho$ containing $x$ for each $x\in V$. A possible implementation of such a function was described in [16].
3 Running Examples
We now start our eight running examples. Throughout this section, $E$ and $F$ are GTESs over a signature $\Sigma$, and $W$ stands for $ST\langle E\cup F\rangle$.
Example 3.17.
Let $\Sigma=\Sigma_{0}\cup\Sigma_{1}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{1}=\{\,f,\;g\,\}$, $E=\{\,f(\#)\doteq g(\#)\,\}$, and $F=\{\,f(\$)\doteq g(\$)\,\}$. Observe the following equations:
$\mathop{\leftrightarrow}^{*}_{E}=\{\,(\delta(s),\delta(t))\mid\delta\in CON_{\Sigma,1},s,t\in\{\,f(\#),\;g(\#)\,\}\,\}$,
$\mathop{\leftrightarrow}^{*}_{F}=\{\,(\delta(s),\delta(t))\mid\delta\in CON_{\Sigma,1},s,t\in\{\,f(\$),\;g(\$)\,\}\,\}$,
$\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$,
$ST\langle E\rangle=\{\,\#,\;f(\#),\;g(\#)\,\}$,   $ST\langle F\rangle=\{\,\$,\;f(\$),\;g(\$)\,\}$, and $W=\{\,\#,\;\$,\;f(\#),\;f(\$),\;g(\#),\;g(\$)\,\}$.
[…]
We note that for each $t\in W$,
$t/_{\Theta\langle{E\cup F},\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}$ or $t/_{\Theta\langle{E\cup F},\emptyset\rangle}=t/_{\Theta\langle{F},E\rangle}$.
That is, each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class.
Example 3.18.
Let $\Sigma=\Sigma_{0}\cup\Sigma_{1}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{1}=\{\,f\,\}$, $E=\{\,f(f(\#))\doteq\#\,\}$, and $F=\{\,f(f(f(\#)))\doteq\#\,\}$.
For every $n\in{\mathbb{N}}$, we define $f^{n}(\#)$ as follows:
  • (i)
    $f^{0}(\#)=\#$ and
  • (ii)
    $f^{n}(\#)=f(f^{n-1}(\#))$ for $n\geq 1$.
Observe the following equations:
$\mathop{\leftrightarrow}^{*}_{E}=\{\,f^{m}(\#),f^{n}(\#))\mid m,n\in{\mathbb{N}},\;m-n\mbox{ is divisible by }2\,\}$,
$\mathop{\leftrightarrow}^{*}_{F}=\{\,f^{m}(\#),f^{n}(\#))\mid m,n\in{\mathbb{N}},\;m-n\mbox{ is divisible by }3\,\}$,
$\mathop{\leftrightarrow}^{*}_{E\cup F}=T_{\Sigma}\times T_{\Sigma}$,
$\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$,
$ST\langle E\rangle=\{\,\#,\;f(\#),\;f^{2}(\#)\,\}$,  $ST\langle F\rangle=\{\,\#,\;f(\#),\;f^{2}(\#),\;f^{3}(\#)\,\}$,  and $W=ST\langle F\rangle$.
The equivalence classes of $\Theta\langle E,F\rangle$ are {#$\{\,\#,\;f^{2}(\#)\}$, $\{\,f(\#),\;f^{3}(\#)\,\}$.
The equivalence classes of $\Theta\langle F,E\rangle$ are {#$\{\,\#,\;f^{3}(\#)\}$, $\{\,f(\#)\,\}$,   $\{\,f^{2}(\#)\,\}$.
The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is {#$\{\,\#,\;f(\#),\;f^{2}(\#),\;f^{3}(\#)\,\}$.
We note that
#$\#/_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and #$\#/_{\Theta\langle F,E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
Consequently,
for each $t\in W$, $\left(t/_{\Theta\langle{E\cup F},\emptyset\rangle}\neq t/_{\Theta\langle E,F\rangle}\mbox{ and }t/_{\Theta\langle{E\cup F},\emptyset\rangle}\neq t/_{\Theta\langle{F},E\rangle}\right)$.
Example 3.19.
Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\,\$,\;\mathsterling\,\}$, $\Sigma_{2}=\{\,f\,\}$.
Intuitively, we now define GTESs $E$ and $F$ such that Statements 1 – 4 hold.
1. $\mathop{\leftrightarrow}^{*}_{E}$ has four congruence classes:
[…]
  • the set of all ground terms which are different from the constant ##\##, and contain only the symbols $f$ and ##\##,
  • the set of all ground terms containing only the symbols $f$ and \$currency-dollar\\$\$, and
[…]
2. We obtain GTES $F$ from GTES $E$ by swapping the constants ##\## and \$currency-dollar\\$\$.
3. $\mathop{\leftrightarrow}^{*}_{F}$ has four congruence classes:
[…]
  • the set of all ground terms which are different from the constant \$currency-dollar\\$\$, and contain only the symbols $f$ and \$currency-dollar\\$\$,
  • the set of all ground terms containing only the symbols $f$ and ##\##, and
[…]
4. $\mathop{\leftrightarrow}^{*}_{E\cup F}$ has three congruence classes:
  • the set of all ground terms which contain only the symbols $f$ and \$currency-dollar\\$\$,
  • the set of all ground terms containing only the symbols $f$ and ##\##, and
[…]
Let
$E=\{\,f(\#,f(\#,\#))\doteq f(\#,\#),\;f(f(\#,\#),\#)\doteq f(\#,\#),\;f(f(\#,\#),f(\#,\#))\doteq f(\#,\#),\;\newline f(\$,\$)\doteq\$,\;\newline f(\#,\$)\doteq\mathsterling,\;f(\#,\mathsterling)\doteq\mathsterling,\;\newline f(\$,\#)\doteq\mathsterling,\;f(\$,\mathsterling)\doteq\mathsterling,\;\newline f(\mathsterling,\#)\doteq\mathsterling,\;f(\mathsterling,\$)\doteq\mathsterling,\;f(\mathsterling,\mathsterling)\doteq\mathsterling\newline f(\$,f(\#,\#))\doteq\mathsterling,\;f(\mathsterling,f(\#,\#))\doteq\mathsterling,\;\newline f(f(\#,\#),\$)\doteq\mathsterling,\;f(f(\#,\#),\mathsterling)\doteq\mathsterling\,\}$
$F=\{\,f(\$,f(\$,\$))\doteq f(\$,\$),\;f(f(\$,\$),\$)\doteq f(\$,\$),\;f(f(\$,\$),f(\$,\$))\doteq f(\$,\$),\;\newline f(\#,\#)\doteq\#,\;\newline f(\$,\#)\doteq\mathsterling,\;f(\$,\mathsterling)\doteq\mathsterling,\;\newline f(\#,\$)\doteq\mathsterling,\;f(\#,\mathsterling)\doteq\mathsterling,\;\newline f(\mathsterling,\#)\doteq\mathsterling,\;f(\mathsterling,\$)\doteq\mathsterling,\;f(\mathsterling,\mathsterling)\doteq\mathsterling\newline f(\#,f(\$,\$))\doteq\mathsterling,\;f(\mathsterling,f(\$,\$))\doteq\mathsterling,\;\newline f(f(\$,\$),\#)\doteq\mathsterling,\;f(f(\$,\$),\mathsterling)\doteq\mathsterling\,\}$.
GTES $E$ is symmetric to GTES $F$. Note the following equations:
  • $ST\langle E\rangle=\{\,\#,\;\$,\;\mathsterling,\;\newline f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;\newline f(f(\#,\#),f(\#,\#)),\;f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;\newline f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling)\,\}$,
  • $ST\langle F\rangle=\{\,\$,\;\#,\;\mathsterling,\;\newline f(\$,\$),\;f(\$,\#),\;f(\$,\mathsterling),\;\newline f(\#,\$),\;f(\#,\#),\;f(\#,\mathsterling),\;\newline f(\mathsterling,\$),\;f(\mathsterling,\#),\;f(\mathsterling,\mathsterling),\;\newline f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$)),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;\newline f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}$,
  • $W=\{\,\#,\;\$,\;\mathsterling,\;\newline f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;\newline f(f(\#,\#),f(\#,\#)),\;f(\$,f(\$,\$)),\;\newline f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$)),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}$.
The equivalence classes of $\Theta\langle E,F\rangle$ are
{#},{\$$\{\,\#\,\},\;\newline \{\,\$,\;f(\$,\$)\,\},\;\newline \{\,f(\#,\#),\;f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#))\,\},\newline \;\{\,\mathsterling,\;f(\#,\$),\;f(\#,\mathsterling),\;f(\$,\#),\;f(\$,\mathsterling),\;f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}$.
The equivalence classes of $\Theta\langle F,E\rangle$ are
{#$\{\,\#,\;f(\#,\#)\,\},\;\newline \{\,\$\,\},\newline \{\,f(\$,\$),\;f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$))\},\;\newline \{\,\mathsterling,\;f(\#,\$),\;f(\#,\mathsterling),\;f(\$,\#),\;f(\$,\mathsterling),\;f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}$.
The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is $W$.
$\mathop{\leftrightarrow}^{*}_{E}=\{\,(\#,\#)\,\}\cup(T_{\{\,\#,f\,\}}\setminus\{\,\#\,\})\times(T_{\{\,\#,f\,\}}\setminus\{\,\#\,\})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\$,f\,\}})\right)\times\left(T_{\Sigma}\setminus(T_{\{\,\#,f\,\}})\cup T_{\{\,\$,f\,\}})\right)\cup\newline (T_{\{\,\$,f\,\}}\times T_{\{\,\$,f\,\}})$.
[…]
$\mathop{\leftrightarrow}^{*}_{F}=(T_{\{\,\#,f\,\}})\times T_{\{\,\#,f\,\}})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\$,f\,\}}\cup T_{\{\,\#,f\,\}}\right)\times\left(T_{\Sigma}\setminus(T_{\{\,\$,f\,\}}\cup T_{\{\,\#,f\,\}}))\right)\cup\newline \{\,(\$,\$)\,\}\cup\left(T_{\{\,\$,f\,\}}\setminus\{\,\$\,\}\right)\times\left(T_{\{\,\$,f\,\}}\setminus\{\,\$\,\}\right)$.
[…]
$\mathop{\leftrightarrow}^{*}_{E\cup F}=(T_{\{\,\#,f\,\}})\times T_{\{\,\#,f\,\}})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\$,f\,\}}\cup T_{\{\,\#,f\,\}})\right)\times\left((T_{\Sigma}\setminus(T_{\{\,\$,f\,\}}\cup T_{\{\,\#,f\,\}})\right)\cup\newline (T_{\{\,\$,f\,\}}\times T_{\{\,\$,f\,\}})$.
$\mathop{\leftrightarrow}^{*}_{E\cup F}$ has the congruence classes $T_{\{\,\#,f\,\}}$,   $T_{\{\,\$,f\,\}}$,   and $T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\$,f\,\}})$.
Observe that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$ and $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}=\mathop{\leftrightarrow}^{*}_{E\cup F}$.
Example 3.20.
[…]
Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\$,\;\mathsterling,\;\flat\,\}$, $\Sigma_{2}=\{\,f\,\}$,
$E=\{\,\#\doteq\$,\;\#\doteq\mathsterling,\;\#\doteq f(\#,\#),\;\#\doteq f(\#,\flat),\;\#\doteq f(\flat,\#)\,\}$,
and
$F=\{\,\mathsterling\doteq\flat,\newline \mathsterling\doteq f(\#,\#),\;\mathsterling\doteq f(\#,\$),\;\mathsterling\doteq f(\#,\mathsterling),\;\newline \mathsterling\doteq f(\$,\#),\;\mathsterling\doteq f(\$,\$),\;\mathsterling\doteq f(\$,\mathsterling),\;\newline \mathsterling\doteq f(\mathsterling,\#),\;\mathsterling\doteq f(\mathsterling,\$),\;\mathsterling\doteq f(\mathsterling,\mathsterling)\,\}$.
We note that for each ground term $t\in T_{\Sigma}$ with $height(t)\geq 1$,
$t\mathop{\leftrightarrow}^{*}_{E}\#$ and $t\mathop{\leftrightarrow}^{*}_{F}\mathsterling$.
Therefore,
$\mathop{\leftrightarrow}^{*}_{E}=\{\,(\flat,\flat)\,\}\cup((T_{\Sigma}\setminus\{\,\flat\,\})\times(T_{\Sigma}\setminus\{\,\flat\,\}))$,
$\mathop{\leftrightarrow}^{*}_{F}=\{\,(\#,\#),\;(\$,\$)\,\}\cup((T_{\Sigma}\setminus\{\,\#,\;\$\,\})\times(T_{\Sigma}\setminus\{\,\#,\;\$\,\}))$, and
$\mathop{\leftrightarrow}^{*}_{E\cup F}=T_{\Sigma}\times T_{\Sigma}$.
Note that (#$(\#,\flat)\not\in\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Consequently,
$\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$.
We notice the following equations:
$W=\{\,\#,\;\$,\;\mathsterling,\;\flat,\;\newline f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;f(\$,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;f(\mathsterling,\flat),\;\newline f(\flat,\#),\;f(\flat,\$),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\}$.
The equivalence classes of $\Theta\langle E,F\rangle$ are
{#,\$,£$\{\,\#,\;\$,\;\mathsterling,\;\newline \,f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;f(\$,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\flat),f(\mathsterling,\mathsterling),\;\newline f(\flat,\#),\;f(\flat,\$),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\},\;\newline \{\,\flat\,\}$.
The equivalence classes of $\Theta\langle F,E\rangle$ are
{£$\{\,\mathsterling,\;\flat,\;\newline \,f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;f(\$,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\flat),f(\mathsterling,\mathsterling),\;\newline f(\flat,\#),\;f(\flat,\$),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\},\newline \{\,\#\,\},\{\,\$\,\}$.
The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is $W$. We note that
#$\#/_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and #$\#/_{\Theta\langle{F},E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
Example 3.21.
Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{2}=\{\,f\,\}$,
$E=\{\,f(\#,\#)\doteq\$,\;f(\#,\$)\doteq\$,\;f(\$,\#)\doteq\$,\;f(\$,\$)\doteq\$\,\}$ and $F=\{\,f(\#,\#)\doteq\#\,\}$.
Note the following equations:
$ST\langle E\rangle=\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$,  $ST\langle F\rangle=\{\,\#,\;f(\#,\#)\,\}$,  $W=ST\langle E\rangle$.
[…]
The equivalence classes of $\Theta\langle F,E\rangle$ are {#$\{\,\#,\;f(\#,\#)\,\},\;\{\,\$\,\},\;\{\,f(\#,\$)\,\},\;\{\,f(\$,\#)\,\},\;\{\,f(\$,\$)\,\}$.
The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is {#,\$$\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$.
Observe that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$. Since (#,\$$(\#,\$)\not\in\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, and (#,\$$(\#,\$)\in\mathop{\leftrightarrow}^{*}_{E\cup F}$, we have $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$.
Example 3.22.
Let $\Sigma=\Sigma_{0}\cup\Sigma_{1}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{1}=\{\,g\,\}$, $\Sigma_{2}=\{\,f\,\}$,
$E=\{\,f(\#,\#)\doteq g(\#),\;f(\#,g(\#))\doteq g(\#),\;f(g(\#),\#)\doteq g(\#),\;f(g(\#),g(\#))\doteq g(\#),\;g(g(\#))\doteq g(\#),\;\newline f(\#,\$)\doteq\$,\;f(\$,\#)\doteq\$,\;f(\$,\$)\doteq\$,\;g(\$)\doteq\$\,\}$.
$F=\{\,f(\#,\#)\doteq\#,\;g(\#)\doteq\#\,\}$.
Intuitively, $F$ can simulate the first five equations of $E$, that is, the congruence generated by the first five equations of $E$ is a proper subset of $\mathop{\leftrightarrow}^{*}_{F}$. We need the constant \$currency-dollar\\$\$ and the last four equations of $E$ to ensure that $\mathop{\leftrightarrow}^{*}_{E}\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$. Note the following equations:
  • $ST\langle E\rangle=\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\#),\;g(\$),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(g(\#))\,\}$,
  • $ST\langle F\rangle=\{\,\#,\;f(\#,\#),\;g(\#)\,\}$,
  • $W=ST\langle E\rangle$.
The equivalence classes of $\Theta\langle E,F\rangle$ are
{#},$\{\,\#\,\},\{\,f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\},\;\newline \{\,\$,\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\$),\;g(g(\$))\,\}$,
The equivalence classes of $\Theta\langle F,E\rangle$ are
{#$\{\,\#,\;f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\}$,
{\$},$\{\,\$\,\},\;\{\,f(\#,\$)\,\},\;\{\,f(\$,\#)\,\},\;\{\,f(\$,\$)\,\},\;\{\,g(\$)\,\},\;\{\,g(g(\$))\,\}$.
The equivalence classes of $\Theta\langle E\cup F,\emptyset\rangle$ are
{#$\{\,\#,\;f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\},\;\newline \{\,\$,\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\$),\;g(g(\$))\,\}$.
Observe that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$.
\$$\$/_{\mathop{\leftrightarrow}^{*}_{E}}=\$/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=\{\,t\in T_{\Sigma}\mid\mbox{ there exists }\alpha\in POS(t)\mbox{ such that }t|_{\alpha}=\$\,\}$.
#$\#/_{\mathop{\leftrightarrow}^{*}_{F}}=\#/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=\{\,t\in T_{\Sigma}\mid\mbox{ there does not exist }\alpha\in POS(t)\mbox{ such that }t|_{\alpha}=\$\,\}$.
Accordingly $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Example 3.23.
Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{2}=\{\,f\,\}$,
$E=\{\,f(\#,\#)\doteq f(\#,\$),\;f(\#,\#)\doteq f(\$,\#),\;f(\#,\#)\doteq f(\$,\$)\,\}$ and $F=\{\,\#\doteq\$\,\}$.
Observe the following equations and inclusion:
$ST\langle E\rangle=\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$,  $ST\langle F\rangle=\{\,\#,\$\,\}$,   $W=ST\langle E\rangle$,  $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$,   $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{F}$.
Since $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$, each equivalence class of $\Theta\langle E,F\rangle$ is a subset of some equivalence class of $\Theta\langle F,E\rangle$.
The equivalence classes of $\Theta\langle E,F\rangle$ are {#},{\$},$\{\,\#\,\},\;\{\,\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$.
The equivalence classes of $\Theta\langle F,E\rangle$ are {#,\$},$\{\,\#,\;\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$.
The equivalence classes of $\Theta\langle E\cup F,\emptyset\rangle$ are {#,\$},$\{\,\#,\;\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$.
Example 3.24.
Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\flat,\;\$,\;\mathsterling\,\}$, $\Sigma_{2}=\{\,f\,\}$,
$E=\{\,\#\doteq\$\,\}$, and $F=\{\,\mathsterling\doteq\flat\,\}$.
We remark that $f(\#,\mathsterling)\mathop{\leftrightarrow}^{*}_{E\cup F}f(\$,\flat)$,  $(f(\#,\mathsterling),f(\$,\flat))\not\in\mathop{\leftrightarrow}^{*}_{E}$,   and $(f(\#,\mathsterling),f(\$,\flat))\not\in\mathop{\leftrightarrow}^{*}_{F}$. Thence
$\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$.
Observe the following equations:
$ST\langle E\rangle=\{\,\#,\;\$\,\}$,   $ST\langle F\rangle=\{\,\mathsterling,\;\flat\,\}$,   $W=\{\,\#,\;\$,\;\mathsterling,\;\flat\,\}$.
[…]
Notice that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$.
4 Fast Ground Completion Introducing Additional Constants
From now on throughout this section, let $\Sigma$ be a signature, $E$ and $F$ be GTESs over $\Sigma$, $W=ST(E\cup F)$, and $n=size(E)+size(F)$. We recall the fast ground completion algorithm of Fülöp and Vágvölgyi [2]. Its input is $\Sigma$, $E$, and $F$. Their fast ground completion algorithm [2] produces in $O(n\,\mathrm{log}\;n)$ time a finite set of new constants and a reduced GTRS $R\langle E,F\rangle$ over $\Sigma$ extended with the new constants such that $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$. Their algorithm calls a congruence closure algorithm on the subterm dag for $E\cup F$ [4], and runs in $O(n\,\mathrm{log}\,n)$ time by Proposition 2.16.
We now cite the Claim on [17, p. 167].
Proposition 4.25.
[17] $\Theta\langle E,F\rangle=\{\,(\widehat{x},\widehat{y})\mid(x,y)\in\rho\langle E,F\rangle\,\}$.
Definition 4.26.
We put $C\langle E,F\rangle=\{\,t/_{\Theta\langle E,F\rangle}\mid t\in W\,\}$.
By Proposition 4.25, for each equivalence class $\vartheta\in W/_{\Theta\langle E,F\rangle}$, there exists exactly one equivalence class $\eta\in V/_{\rho\langle E,F\rangle}$ such that $\vartheta=\{\,\widehat{x}\mid x\in\eta\,\}$. To simplify our notations, we give $\vartheta$ the unique name of $\eta$. The elements of $C\langle E,F\rangle$ are viewed as constants, and we define the signature $\Sigma\cup C\langle E,F\rangle$.
Statement 4.27.
$|C\langle E,F\rangle|\leq|W|\leq size(W)\leq n$.
Proof.
By Definition 4.26. ∎
Definition 4.28.
[2] Let
$R\langle E,F\rangle=\newline \{\,\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\sigma(t_{1},\ldots,t_{m})/_{\Theta\langle E,F\rangle}\mid\sigma\in\Sigma_{m},m\in{\mathbb{N}},t_{1},\ldots,t_{m}\in T_{\Sigma},\sigma(t_{1},\ldots,t_{m})\in W\,\}$
be a GTRS over the ranked alphabet $\Sigma\cup C\langle E,F\rangle$.
Note that for any $\sigma(t_{1},\ldots,t_{m})\in W$ with $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, and $t_{1},\ldots,t_{m}\in T_{\Sigma}$ there exists at most one rewrite rule with left-hand side $\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})$ in $R\langle E,F\rangle$.
Definition 4.29.
We say that GTRS $R\langle E,F\rangle$ is total if for any $\sigma\in\Sigma_{m}$ with $m\in{\mathbb{N}}$, and $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$, there exists a rewrite rule in GTRS $R\langle E,F\rangle$ with left-hand side $\sigma(a_{1},\ldots,a_{m})$.
Fülöp and Vágvölgyi [2] showed the following properties of $R\langle E,F\rangle$.
Proposition 4.30.
[2] $R\langle E,F\rangle$ is reduced.
By Proposition 4.30 and Proposition 2.8, we have the following result.
Proposition 4.31.
[2] $R\langle E,F\rangle$ is convergent.
Proposition 4.32.
[2] For every $t\in W$, we have $t\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
Proposition 4.33.
[2] Every $t\in W$ has a unique $R\langle E,F\rangle$ normal form. For every $t\in W$, we have $t{\downarrow}_{R\langle E,F\rangle}=t/_{\Theta\langle E,F\rangle}$.
Proof.
Let $t\in W$. By Proposition 4.32, we have $t\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$. By direct inspection of the rewrite rules of $R\langle E,F\rangle$, we get that for each $t\in W$, $t/_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$. Accordingly $t/_{\Theta\langle E,F\rangle}$ is a $\mathop{\rightarrow}_{R\langle E,F\rangle}$ normal form of $t$. By Proposition 2.3, Proposition 4.31 $t\in W$ has a unique normal form, $t{\downarrow}_{R\langle E,F\rangle}$. Then we have $t{\downarrow}_{R\langle E,F\rangle}=t/_{\Theta\langle E,F\rangle}$. ∎
Proposition 4.34.
[2] For all $s\in T_{\Sigma}$ and $t\in W$,
if $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$, then $s\mathop{\leftrightarrow}^{*}_{E}t$.
Proposition 4.35.
[2] For all $s,t\in T_{\Sigma}$,
$s\mathop{\leftrightarrow}^{*}_{E}t$ if and only if $s{\downarrow}_{R\langle E,F\rangle}=t{\downarrow}_{R\langle E,F\rangle}$.
Proposition 4.36.
[2] For all $s\in T_{\Sigma}$ and $t\in W$,
if $s\mathop{\leftrightarrow}^{*}_{E}t$, then $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
Proposition 4.37.
[2] $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$.
By Proposition 4.34 and Proposition 4.36, we have the following result.
Proposition 4.38.
For all $s\in T_{\Sigma}$ and $t\in W$,
$s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$ if and only if $s\mathop{\leftrightarrow}^{*}_{E}t$.
Statement 4.39.
If $R\langle E,F\rangle=\emptyset$, then $R\langle E,F\rangle$ is not total.
Proof.
Let $\sigma\in\Sigma_{0}$. There does not exists a rule in $R\langle E,F\rangle$ with left-hand side $\sigma$. Consequently, $R\langle E,F\rangle$ is not total. ∎
Lemma 4.40.
If  $\bigcup W/_{\mathop{\leftrightarrow}^{*}_{E}}=T_{\Sigma}$, then $R\langle E,F\rangle$ is total.
Proof.
Let $\sigma\in\Sigma_{m}$ with $m\in{\mathbb{N}}$, and let $t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle}$ be equivalence classes of $\Theta\langle E,\;F\rangle$ for some $t_{1},\ldots t_{m}\in W$. By our assumption, there exists $s\in W$ such that $\sigma(t_{1},\ldots,t_{m})\mathop{\leftrightarrow}^{*}_{E}s$. By Proposition 4.35, $\sigma(t_{1},\ldots,t_{m}){\downarrow}_{R\langle E,F\rangle}=s{\downarrow}_{R\langle E,F\rangle}$. By Proposition 4.33 $s{\downarrow}_{R\langle E,F\rangle}=s/_{\Theta\langle E,F\rangle}$. Thus $\sigma(t_{1},\ldots,t_{m}){\downarrow}_{R\langle E,F\rangle}=s/_{\Theta\langle E,F\rangle}$. Therefore,
$\sigma(t_{1},\ldots,t_{m})\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\sigma(t_{1}{\downarrow}_{R\langle E,F\rangle},\ldots,t_{m}{\downarrow}_{R\langle E,F\rangle})=$ (by Proposistion 4.32)
$\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})\mathop{\rightarrow}_{R\langle E,F\rangle}s/_{\Theta\langle E,F\rangle}$     (by direct inspection of the form of the rewrite rules of $R\langle E,F\rangle$).
By direct inspection of the rewrite rules of $R$, we get that in the last step of the sequence of reductions we applied the rewrite rule
$\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})\dot{\rightarrow}s/_{\Theta\langle E,F\rangle}$ in $R\langle E,F\rangle$. ∎
Lemma 4.41.
If $R\langle E,F\rangle$ is a total GTRS, then for each $s\in T_{\Sigma}$, there exists $t\in W$ such that $s{\downarrow}_{R\langle E,F\rangle}=t/_{\Theta\langle E,F\rangle}$.
Proof.
We proceed by induction on $height(s)$.
Base case: Let $height(s)=0$. Then $s=\sigma$ for some $\sigma\in\Sigma_{0}$. By our assumption that $R\langle E,F\rangle$ is a total GTES and by Definitions 4.28 and 4.29, there is a rewrite rule $\sigma\dot{\rightarrow}\sigma/_{\Theta\langle E,F\rangle}$ in $R\langle E,F\rangle$. Then $\sigma\in W$ and $s{\downarrow}_{R\langle E,F\rangle}=\sigma/_{\Theta\langle E,F\rangle}$.
[…]
$\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\sigma(p_{1},\ldots,p_{m})/_{\Theta\langle E,F\rangle}$
[…]
$s=\sigma(s_{1},\ldots,s_{m})\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}s(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})\mathop{\rightarrow}_{R\langle E,F\rangle}\sigma(p_{1},\ldots,p_{m})/_{\Theta\langle E,F\rangle}$.
By direct inspection of Definition 4.28, we get that $p_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$. Thus $s{\downarrow}_{R\langle E,F\rangle}=\sigma(p_{1},\ldots,p_{m})/_{\Theta\langle E,F\rangle}$. ∎
Proposition 4.42.
$\bigcup W/_{\mathop{\leftrightarrow}^{*}_{E}}=T_{\Sigma}$ if and only if $R\langle E,F\rangle$ is total.
Proof.
It is a consequence of Lemma 4.40 and Lemma 4.41 and Proposition 4.35. ∎
Statement 4.43.
GTRS $R\langle E,F\rangle$ reach a constant $a\in C\langle E,F\rangle$ from a ground term $t$ over the ranked alphabet $\Sigma\cup C\langle E,F\rangle$ if and only if $t{\downarrow}_{R\langle E,F\rangle}=a$.
Proof.
()($\Rightarrow$)( ⇒ ) The constant $a\in C\langle E,F\rangle$ is irreducibe for $R\langle E,F\rangle$ by direct inspection of Definition 4.28. By Proposition 4.31, ${R\langle E,F\rangle}$ is convergent. Accordingly $a\in C\langle E,F\rangle$ is the unique $R\langle E,F\rangle$-normal form of $t$.
()($\Leftarrow$)( ⇐ ) This direction is trivial. ∎
Definition 4.44.
Let
$BSTEP\langle E,F\rangle=\{\,(a,b)\in C\langle E,F\rangle\times C\langle E,F\rangle\mid\mbox{ there exists a rewrite rule }\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E,F\rangle\mbox{ and }1\leq i\leq m\mbox{ such that }b=a_{i}\,\}$.
We call an element of $BSTEP\langle E,F\rangle$ a transition backstep.
Statement 4.45.
Let $n=size(E)+size(F)$. Then
1. $|R\langle E,F\rangle|\leq|W|\leq size(W)\leq size(E\cup F)\leq size(E)+size(F)=n$.
2. $size(R\langle E,F\rangle)\leq size(W)+|W|\leq 2\cdot n$  and
3. $|BSTEP\langle E,F\rangle|\leq size(R\langle E,F\rangle)\leq 2\cdot n$.
Proof.
By direct inspection of Definition 4.28 and by Statement 2.7 we have Statement 1.
[…]
By direct inspection of Definition 4.44 and by Statement 2 we have Statement 3.
We now adopt the fast ground completion algorithm of Fülöp and Vágvölgyi [2] that produces the GTRS $R\langle E,F\rangle$.
Definition 4.46.
Algorithm Fast Ground Completion Introducing Additional Constants (FGC for short) [2]
Input: A signature $\Sigma$, GTES $E$ and GTES $F$ over $\Sigma$.
Output:
  • the relation $\rho\langle E,F\rangle$,
  • the set of nullary symbols $C\langle E,F\rangle$,
  • the GTRS $R\langle E,F\rangle$ stored without duplicates in a red-black tree $RBT1$, and
  • the relation $BSTEP\langle E,F\rangle$ stored without duplicates in a red-black tree $RBT2$.
Data structures: The congruence closure algorithm appearing in Proposition 2.16, gives each $\rho$-class a unique name, which is an element of the set { $\{\,1,\ldots,|W|\,\}$. We assign to each symbol of $\Sigma$ a unique number in the set
$\{\,|W|+1,\ldots,2\cdot|W|\,\}$.
Let
$\ell=\lceil\mathrm{log}_{2}|W|\rceil+1$.
For each symbol $\sigma\in\Sigma\langle E,F\rangle$, we pad the binary form of its number with leading zeros to length $\ell$. The binary sequence obtained in this way is called the binary representation of the symbol $\sigma\in\Sigma\langle E,F\rangle$, and is denoted as $\phi(\sigma)$. Similarly, for each constant $a\in C\langle E,F\rangle$, we pad the binary form of its number with leading zeros to length $\ell$. The binary sequence obtained in this way is called the binary representation of the constant $a\in C\langle E,F\rangle$, and is denoted as $\phi(a)$.
The binary representation of a rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E,F\rangle$ is the string $\phi(\sigma)\phi(a_{1}),\ldots,\phi(a_{m})\phi(a)$. We store the binary representations of the rewrite rules in $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT1$ in lexicographic order. The binary representation of an ordered pair $(a,b)\in BSTEP\langle E,F\rangle$ is the ordered pair $(\phi(a),\phi(b))$. We store the binary representations of the elements of $BSTEP\langle E,F\rangle$ without duplicates in the red-black tree $RBT2$ in lexicographic order. Here for every ordered pair $(a,b)\in BSTEP\langle E,F\rangle$, the search key is the first component $\phi(a)$, the second component, $\phi(b)$ is satellite data.
var $i$, $m$: ${\mathbb{N}}$; $\sigma$: $\Sigma$; $a,a_{1},\ldots,a_{mar_{\Sigma}}:C\langle E,F\rangle$; $RBT1$, $RBT2$: red-black tree;
1. Create a subterm dag $G=(V,A,\lambda)$ for $E$ and $F$.
By Proposition 2.16, compute the congruence closure $\rho\langle E,F\rangle$ of the relation $\tau\langle E,F\rangle=\{\,(u,v)\mid(\hat{u},\hat{v})\in E\,\}.$
Compute $C\langle E,F\rangle=\{\,x/_{\rho\langle E,F\rangle}\mid x\in V\,\}$.
2. { Computing the GTRS $R\langle E,F\rangle$
$R\langle E,F\rangle:=\emptyset$;
$BSTEP\langle E,F\rangle:=\emptyset$;
For each $x\in V$ do
begin
$\sigma:=\lambda(x)$;
$m:=outdegree(x);$
$a:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x)$;
for each $1\leq i\leq m$ do
$a_{i}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x/i)$;
$R\langle E,F\rangle:=R\langle E,F\rangle\cup\{\,\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\,\}$;
$BSTEP\langle E,F\rangle:=BSTEP\langle E,F\rangle\cup\{\,(a_{1},a),\ldots,(a_{m},a)\,\}$;
end
By Statement 4.45 and Proposition 2.2, the time complexity of storing $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT1$ is $O(n\,\mathrm{log}\,n)$, where $n=size(E)+size(F)$. By Statement 4.45 and Proposition 2.2, the time complexity of storing $BSTEP\langle E,F\rangle$ without duplicates in the red-black tree $RBT2$ is $O(n\,\mathrm{log}\,n)$, where $n=size(E)+size(F)$.
Proposition 4.47.
[2] Algorithm FGC takes as input a signature $\Sigma$, and GTESs $E$ and $F$ over $\Sigma$. Let $n=size(E)+size(F)$.
1. Algorithm FGC produces in $O(n\,\mathrm{log}\,n)$ time the relation $\rho\langle E,F\rangle$ and the set of nullary symbols $C\langle E,F\rangle$.
2. In $O(n\,\mathrm{log}\,n)$ time, Algorithm FGC computes the GTRS $R\langle E,F\rangle$ over $\Sigma\cup C\langle E,F\rangle$, and stores $R\langle E,F\rangle$ without duplicates in the lexical order $<_{lex}$ on binary sequences in the red-black tree $RBT1$.
3. In $O(n\,\mathrm{log}\,n)$ time, Algorithm FGC computes the relation $BSTEP\langle E,F\rangle$, and stores $BSTEP\langle E,F\rangle$ without duplicates in the red-black tree $RBT2$. The search key is the first component, the second component is satellite data.
Proof.
Fülöp and Vágvölgyi [2] showed that we can construct the GTRS $R\langle E,F\rangle$ in $O(n\,\mathrm{log}\;n)$ time. Hence we can construct the relation $BSTEP\langle E,F\rangle$ in $O(n\,\mathrm{log}\;n)$ time as well. In addition to the result of Fülöp and Vágvölgyi [2], we now compute the cost of storing the GTRS $R\langle E,F\rangle$. and the cost of storing the relation $BSTEP\langle E,F\rangle$. First we show Statement 1. Algorithm FGC produces the set of nullary symbols $C\langle E,F\rangle$. By Definition 4.26, for every equivalence class $\vartheta\in\Theta\langle E,F\rangle$, Algorithm FGC gives a unique name for $\vartheta$ which is the unique name of the $\rho\langle E,F\rangle$ equivalence class $\eta$ such that $\vartheta=\{\,\widehat{x}\mid x\in\eta\,\}$.
We now show Statement 2. Algorithm FGC produces the GTRS $R\langle E,F\rangle$ apart from that, for every equivalence class $\vartheta\in\Theta\langle E,F\rangle$, Algorithm FGC writes the unique name for $\vartheta$ [2]. By Statement 4.45, $size(R\langle E,F\rangle)\leq 2\cdot|W|\leq 2\cdot n$. Then by Proposition 2.2, the time-complexity of inserting the elements of $R\langle E\cup F,\emptyset\rangle$ into an empty red-black tree, not allowing duplicate values in the resulting red-black tree $RBT1$, is $O(n\,\mathrm{log}\;n)$,
We now show Statement 3. In the relation $BSTEP\langle E,F\rangle\subseteq C\langle E,F\rangle\times C\langle E,F\rangle$, we collect all pairs $(a_{i},a)$ such that there is a rewrite rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a$ in $R\langle E,F\rangle$. By Statement 6.63, $|BSTEP\langle E,F\rangle|\leq n$. Then by Proposition 2.2, the time-complexity of inserting the elements of $BSTEP\langle E,F\rangle$ into an empty red-black tree, not allowing duplicate values in the resulting red-black tree, is $O(n\,\mathrm{log}\;n)$.
We give the GTRS $R\langle E,F\rangle$ as input to an algorithm by giving the red-black tree $RBT1$ output by Algorithm FGC storing the GTRS $R\langle E,F\rangle$. We now define an algorithm deciding whether $R\langle E,F\rangle$ is total.
Definition 4.48.
Let $E$ and $F$ be GTESs over a ranked alphabet $\Sigma$.
Algorithm Deciding whether GTRS $R\langle E,F\rangle$ is Total (DTOT for short).
Input: The signature $\Sigma$, the Boolean value $FLAG\langle E,F\rangle$, the set of constants $C\langle E,F\rangle$, and the red-black tree $RBT$ storing the GTRS $R\langle E,F\rangle$.
Output: Boolean. { $\mathrm{true}$ if $R\langle E,F\rangle$ total. Otherwise $\mathrm{false}$.  }
var $i$, $m,count:{\mathbb{N}}$;
If $FLAG\langle E,F\rangle=\mathrm{false}$, then return $\mathrm{false}$. If $R\langle E,F\rangle=\emptyset$, then return $\mathrm{false}$. Otherwise, traverse the red-black tree $RBT$ in in-order traversal, and read all binary sequences stored in the red-black tree $RBT$ in the lexical order $<_{lex}$ on binary sequences. For every $m\in{\mathbb{N}}$, $\sigma\in\Sigma\langle R\langle E,F\rangle_{m}$, read the rules with $\sigma$ appering on the left-hand side one after another. During this process,
for every $\sigma\in\Sigma\langle E,F\rangle$ do
begin
count all rules with $\sigma$ appearing on the left-hand side, and store the result in the variable $count$;
$m:=\mathrm{arity}(\sigma)$;
${\bf if}\;count<|C\langle E,F\rangle|^{m}$
then return $\mathrm{false}$
end
${\bf output}(\mathrm{true})$
Lemma 4.49.
Algorithm DTOT decides in $O(n)$ time whether $R\langle E,F\rangle$ is total, where $n=size(E)+size(F)$.
Proof.
First, assume that $R\langle E,F\rangle=\emptyset$. Recall that there exists a constant $\sigma\in\Sigma_{0}$. Then there does not exist a rewrite rule with left-hand side $\sigma$ in $R\langle E,F\rangle$. Consequently, $R\langle E,F\rangle$ is not total.
Second, assume that $R\langle E,F\rangle\neq\emptyset$. If $FLAG\langle E,F\rangle=\mathrm{false}$, that is, $|\Sigma\langle E,F\rangle|<|\Sigma|$, then $R\langle E,F\rangle$ is not total. Assume that $FLAG\langle E,F\rangle=\mathrm{true}$, that is, $|\Sigma\langle E,F\rangle|=|\Sigma|$. Then for each symbol $\sigma\in\Sigma$, we count the rewrite rules in $R\langle E,F\rangle$ with $\sigma$ appearing in the left-hand side and store the result in the variable $count$. If there exists a symbol $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, such that for $\sigma$, $count<|\langle C\langle E,F\rangle\rangle|^{m}$, then there exist constants $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$ such that there does not exist a rewrite rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R\langle E,F\rangle$. Therefore, $R\langle E,F\rangle$ is not total. Conversely if $R\langle E,F\rangle$ is not total, then there exists a symbol $\sigma\in\Sigma_{m}$ and constants $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$ such that there does not exist a rewrite rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R\langle E,F\rangle$. Then for $\sigma$, $count<|\langle C\langle E,F\rangle\rangle|^{m}$. Consequently if $R\langle E,F\rangle$ is total, then Algorithm DTOT outputs true; otherwise it returns false.
We now consider the time complexity of the Algorithm DTOT. If $|\Sigma\langle E,F\rangle|<|\Sigma|$ or $R\langle E,F\rangle=\emptyset$, then Algorithm DTOT takes less than $O(n)$ time. Assume that $|\Sigma\langle E,F\rangle|=|\Sigma|$ and $R\langle E,F\rangle\neq\emptyset$. Then Algorithm DTOT traverse the red-black tree $RBT$ in in-order traversal, and it reads all binary sequences stored in the red-black tree $RBT$ in the lexical order $<_{lex}$ on binary sequences. For every $\sigma\in\Sigma\langle R\langle E,F\rangle$, it reads the rules with $\sigma$ appering on the left-hand side one after another, and counts the rewrite rules in $R\langle E,F\rangle$ with $\sigma$ appearing in the left-hand side for all symbols $\sigma\in\Sigma$. This takes $O(n)$ time.
If there exists a symbol $\sigma\in\Sigma$ such that $\sigma$ does not appear in $R\langle E,F\rangle$, then Algorithm DTOT outpouts false, and does not evaluate the expression $count<|C\langle E,F\rangle|^{m}$. Otherwise, for every symbol $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, we carry out $m$ multiplications when computing $|C\langle E,F\rangle|^{m}$. Because each symbol $\sigma\in\Sigma\langle R\langle E,F\rangle_{m}$, $m\in{\mathbb{N}}$, has $m$ constants in its arguments in the left-hand side of some rewrite rule in $R\langle E,F\rangle$.
The number of multiplications $\leq$ $\sum_{m\in{\mathbb{N}}}\sum_{\sigma\in\Sigma\langle R\langle E,F\rangle_{m}}m\leq$ $size(R\langle E,F\rangle)\leq n$.
by Statement 4.45. Consequently, the number of multiplications is less than or equal to $n$. The number of comparisons of nonnegative integers is $|\Sigma|\leq n$. Therefore the time complexity of algorithm DTOT is $O(n)$. ∎
5 Running Examples Continued
We now illustrate the concepts of Section 4 by continuing our running examples. As in Section 3, throughout this section $E$ and $F$ are GTESs over a signature $\Sigma$, and $W$ stands for $ST\langle E\cup F\rangle$.
Example 5.50.
We continue Example 3.17.
[…]
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;g(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow},\;g(\$)/_{\Theta\langle E,F\rangle}\;\}$,
[…]
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;g(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E,F\rangle}\;\}$,
[…]
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline g(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},g(\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
Example 5.51.
We continue Example 3.18.
$C\langle E,F\rangle=\{\,\{\,\#,\;f^{2}(\#)\},\;\{\,f(\#),\;f^{3}(\#)\,\}\,\}$,
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;f(f(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle}\,\}$,
$C\langle F,E\rangle=\{\,\{\,\#,\;f^{3}(\#)\,\}$, $\{\,f(\#)\,\}$, $\{\,f^{2}(\#)\,\}\}$,
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle F,E\rangle},\;f(f(\#)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f^{2}(\#)/_{\Theta\langle F,E\rangle},\;\;\newline f(f^{2}(\#)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle}\,\}$,
$C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;f(\#),\;f^{2}(\#),\;f^{3}(\#)\,\}\,\}=\{\,W\,\}$,
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
We note that $R\langle E,F\rangle$, $R\langle F,E\rangle$, and $R\langle E\cup F,E\rangle$ are total.
Example 5.52.
We continue Example 3.19.
$C\langle E,F\rangle=\{\,\{\,\#\,\},\;\newline \{\,\$,\;f(\$,\$),\;f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$))\,\},\;\newline \{\,f(\#,\#),\;f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#))\,\},\newline \;\{\,\mathsterling,\;f(\#,\$),\;f(\#,\mathsterling),\;f(\$,\#),\;f(\$,\mathsterling),\;f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}\,\}$,
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;f(\#/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\#/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\$/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\$/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\;f(\$/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\mathsterling/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\#/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;\newline f(f(\#,\#)/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;\,f(f(\#,\#)/_{\Theta\langle E,F\rangle},f(\#,\#))/_{\Theta\langle E,F\rangle}\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle}\;\;\newline f(\$/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;f(\mathsterling/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(f(\#,\#)/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;f(f(\#,\#)/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle}\,\}$.
$C\langle F,E\rangle=\{\,\{\,\$\,\},\;\newline \{\,\#,\;f(\#,\#),\,f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#))\,\},\;\newline \{\,f(\$,\$),\;f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$))\,\},\newline \;\{\,\mathsterling,\;f(\$,\#),\;f(\$,\mathsterling),\;f(\#,\$),\;f(\#,\mathsterling),\;f(\mathsterling,\$),\;f(\mathsterling,\#),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling)\,\}\,\}$,
$R\langle F,E\rangle=\{\,\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\;\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\$/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;f(\$/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\$/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\;f(\#/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\mathsterling/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\$/_{\Theta\langle F,E\rangle},f(\$,\$)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;\newline f(f(\$,\$)/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;\,f(f(\$,\$)/_{\Theta\langle F,E\rangle},f(\$,\$))/_{\Theta\langle F,E\rangle}\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle}\;\;\newline f(\#/_{\Theta\langle F,E\rangle},f(\$,\$)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;f(\mathsterling/_{\Theta\langle F,E\rangle},f(\$,\$)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(f(\$,\$)/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;f(f(\$,\$)/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$.
$C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;\$,\;\mathsterling,\;\newline f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#)),\;\;\newline f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$)),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}\,\}$,
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
We observe that $R\langle E,F\rangle$, $R\langle F,E\rangle$, and $R\langle E\cup F,\emptyset\rangle$ are total.
Example 5.53.
We now continue Example 3.20.
$C\langle E,F\rangle=\{\,\#/_{\Theta\langle E,F\rangle},\;\flat/_{\Theta\langle E,F\rangle}\,\}$,
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\flat\dot{\rightarrow}\flat/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\flat/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;f(\flat/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#\,\}$.
$C\langle F,E\rangle=\{\,\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle}\,\}$,
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\flat\dot{\rightarrow}\mathsterling_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\newline f(\$/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\$/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\$/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathsterling/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathsterling/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$.
$C\langle E\cup F,\emptyset\rangle=\{\,\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\flat\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
We note that $R\langle E,F\rangle$, $R\langle F,E\rangle$, and $R\langle E\cup F,\emptyset\rangle$ are total.
Example 5.54.
We now continue Example 3.21.
$C\langle E,F\rangle=\{\,\#/_{\Theta\langle E,F\rangle},\;\$/_{\Theta\langle E,F\rangle}\,\}$.
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\$/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle}\;\}$.
Observe that $R\langle E,F\rangle$ is total.
$C\langle F,E\rangle=\{\,\{\,\#,\;f(\#,\#)\,\},\;\{\,\$\,\},\;\{\,f(\#,\$)\,\},\;\{\,f(\$,\#)\,\},\;\{\,f(\$,\$)\,\}\,\}$.
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},f(\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\$)/_{\Theta\langle F,E\rangle},\;\newline f(\$/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\#)/_{\Theta\langle F,E\rangle},\;f(\$/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle}\,\}$.
Note that $R\langle F,E\rangle$ is not total.
$C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}\,\}$.
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,E\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
We see that $R\langle E\cup F,\emptyset\rangle$ is total.
Example 5.55.
We now continue Example 3.22. Recall that the set of equivalence classes of $\Theta\langle E,F\rangle$ is:
$C\langle E,F\rangle=\{\,\{\,\#\,\},\;\{\,f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\},\;\newline \{\,\$,\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\$),\;g(g(\$))\,\}\,\}$.
That is, $C\langle E,F\rangle=\{\,\#/_{\Theta\langle E,F\rangle},\;\$/_{\Theta\langle E,F\rangle},\,f(\#,\#)/_{\Theta\langle E,F\rangle}\,\}$.
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline f(g(\#)/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;f(g(\#)/_{\Theta\langle E,F\rangle},g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;g(g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\$/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;g(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\$)/_{\Theta\langle E,F\rangle},\;g(g(\$)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle}$.
Note that $R\langle E,F\rangle$ is total.
Recall that the set of equivalence classes of $\Theta\langle F,E\rangle$ is:
$C\langle F,E\rangle=\{\,\#,\;f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\}$,
{\$},$\{\,\$\,\},\;\{\,g(\$)\,\},\;\{\,f(\#,\$)\,\},\;\{\,f(\$,\#)\,\},\;\{\,f(\$,\$)\,\},\;\{\,g(g(\$))\,\}$.
That is,
$C\langle F,E\rangle=\{\,\#/_{\Theta\langle F,E\rangle},\;\$/_{\Theta\langle F,E\rangle},\;g(\$)/_{\Theta\langle E,E\rangle},\;f(\#,\$)/_{\Theta\langle E,F\rangle},\;f(\$,\#)/_{\Theta\langle E,F\rangle},\;\newline f(\$,\$)/_{\Theta\langle E,F\rangle},\;g(g(\$))/_{\Theta\langle E,F\rangle}\;\,\}$.
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},g(\#)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\newline f(g(\#)/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(g(\#)/_{\Theta\langle F,E\rangle},g(\#)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\newline g(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;g(\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}g(\$)/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\$)/_{\Theta\langle F,E\rangle},\,f(\$/_{\Theta\langle F,E\rangle},\#_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\#)/_{\Theta\langle F,E\rangle}\newline f(\$/_{\Theta\langle F,E\rangle},\$_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;g(g(\$)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}g(g(\$))/_{\Theta\langle F,E\rangle}\,\}$.
Note that $R\langle F,E\rangle$ is not total.
Recall that the set of equivalence classes of $\Theta\langle E\cup F,\emptyset\rangle$ is:
$C\langle E\cup F,\emptyset\rangle=\ \{\,\#,\;f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\},\;\newline \{\,\$,\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\$),\;g(g(\$))\,\}$.
That is, $C\langle E\cup F,\emptyset\rangle=\{\,\#/_{\Theta\langle E\cup F,\emptyset\rangle},\$/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\$_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\$/_{\Theta\langle E\cup F,\emptyset\rangle},\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline g(\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
Observe that $R\langle E\cup F,\emptyset\rangle$ is total.
Example 5.56.
We continue Example 3.23.
$C\langle E,F\rangle=\{\,\{\,\#\,\},\;\{\,\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}\,\}$.
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\newline f(\$/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle}\,\},$
$C\langle F,E\rangle=\{\,\{\,\#,\;\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}\,\}$,
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle F,E\rangle}\,\}$,
$C\langle E\cup F,\emptyset\rangle=C\langle F,E\rangle$,   $R\langle E\cup F,\emptyset\rangle=R\langle F,E\rangle$.
We notice that $R\langle E,F\rangle$, $R\langle F,E\rangle$, and $R\langle E\cup F,\emptyset\rangle$ are not total.
Example 5.57.
We now continue Example 3.24.
$C\langle E,F\rangle=\{\,\{\,\#,\;\$\,\},\;\{\,\mathsterling\,\},\;\{\,\flat\,\}\,\}$,
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\flat\dot{\rightarrow}\flat/_{\Theta\langle E,F\rangle}\,\}$,
$C\langle F,E\rangle=\{\,\{\,\#\,\},\;\{\,\$\,\},\;\{\,\mathsterling,\;\flat\,\}\,\}$,
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\flat\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$,
$C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;\$\,\},\;\,\{\,\mathsterling,\;\flat\,\}\,\}$, and
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E\cup F,\emptyset\rangle},\;\flat\dot{\rightarrow}\mathsterling/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
Observe that GTRS $R\langle E,F\rangle$, GTRS $R\langle F,E\rangle$, and GTRS $R\langle E\cup F,\emptyset\rangle$ are not total.
6 The Auxiliary dpwpa for the GTESs $E$ and $F$
From now on throughout the paper, $E$ and $F$ are GTESs over a signature $\Sigma$, $n=size(E)+size(F)$, and $W$ stands for $ST\langle E\cup F\rangle$. We define the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F]\rangle]$ for the GTESs $E$ and $F$. To simplify our notations, each vertex of the dpwpa $AUX[E;F]\rangle$ is a constant $a\in C\langle E\cup F,\emptyset\rangle$, i.e, an equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$. Furthermore, we study $R\langle{E\cup F},\emptyset\rangle$ reachability of a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$.
Definition 6.58.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then let
$INC(E\cup F,E)=\{\,(a,b)\mid a\in C\langle E\cup F,\emptyset\rangle,b\in C\langle E,F\rangle,\mbox{ and }b\subseteq a\,\}$,
$NUM(E\cup F,E)=\{\,(a,k)\mid a\in C\langle E\cup F,\emptyset\rangle,k=|\{b\in C\langle E,F\rangle\mid b\subseteq a\,\}|\,\}$,
$INC(E\cup F,F)=\{\,(a,b)\mid a\in C\langle E\cup F,\emptyset\rangle,b\in C\langle F,E\rangle\mbox{ and }b\subseteq a\,\}$, and
$NUM(E\cup F,F)=\{\,(a,k)\mid a\in C\langle E\cup F,\emptyset\rangle,k=|\{b\in C\langle F,E\rangle\mid b\subseteq a\,\}|\,\}$.
Intuitively, $INC(E\cup F,E)$ shows the inclusions between the elements of $C\langle E\cup F,\emptyset\rangle$ and $C\langle E,F\rangle$. $NUM(E\cup F,E)$ consists of all pairs $(a,k)$, where the second component $k$ is the number of all pairs in $INC(E\cup F,E)$, with first component $a$.
Statement 6.59.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then $|INC(E\cup F,E)|\leq n$ and $|INC(E\cup F,F)|\leq n$.
Proof.
By Statement 4.27, $|C\langle E,F\rangle|\leq n$. For each $b\in C\langle E,F\rangle$, there exists at most one $a\in C\langle E\cup F,\emptyset\rangle$ such that $(a,b)\in INC(E\cup F,E)$. Therefore, $|INC(E\cup F,E)|\leq|C\langle E,F\rangle|\leq n$.
We can show similarly that $|INC(E\cup F,F)|\leq|C\langle F,E\rangle|\leq n$. ∎
Statement 6.60.
1. For every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,b)\in INC(E\cup F,E)$ if and only if $a\in C\langle E,F\rangle$. Similarly, for every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,b)\in INC(E\cup F,F)$ if and only if $a\in C\langle F,E\rangle$.
2. For every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,k)\in NUM(E\cup F,E)$. Here $k=1$ if and only if $a\in C\langle E,F\rangle$. Similarly, for every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,\ell)\in NUM(E\cup F,F)$. Here $\ell=1$ if and only if $a\in C\langle F,E\rangle$.
Proof.
By direct inspection of Definition 6.58
Definition 6.61.
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $a\in C\langle E\cup F,\emptyset\rangle$. We say that GTRS $R\langle E,F\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant $a\in C\langle E\cup F,\emptyset\rangle$ if
for every rewrite rule of the form $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle{E\cup F},\emptyset\rangle$ with $m\in{\mathbb{N}}$,
[…]
there exists a rewrite rule in $R\langle E,F\rangle$ of the form $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b$ for some $b\in C\langle E,F\rangle$.
Note that for the constant $b\in C\langle E,F\rangle$ appearing in the definition above, we have $b\subseteq a$. Definition 6.61 also defines the notion that GTRS $R\langle F,E\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant $a\in C\langle E\cup F,\emptyset\rangle$.
Definition 6.62.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. The auxiliary dpw for the GTESs $E$ and $F$ is $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, where $A[E;F]=BSTEP\langle E\cup F,\emptyset\rangle$. We assign the attributes
$equal_{E},\;equal_{F},\;keeps_{E},\;keeps_{F},\;visited\_outer\_loop,\;visited\_inner\_loop_{E},\;visited\_inner\_loop_{F},\;and\;visited:\mathrm{Boolean}$
to the vertices of $C\langle E\cup F,\emptyset\rangle$. Furthermore, we assign the satellite data $counter_{E},counter_{F}:{\mathbb{N}}$ to the rewrite rules of $R\langle E\cup F,\emptyset\rangle$.
For each vertex $a\in$, our decision algorithm computes the values of the above attributes such that the following conditions hold:
  • 1.
    $a.equal_{E}=\mathrm{true}$ if and only if $a\in C\langle E,F\rangle$.
  • 2.
    $a.equal_{F}=\mathrm{true}$ if and only if $a\in C\langle F,E\rangle$.
  • 3.
    $a.keeps_{E}=\mathrm{true}$ if and only if $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $a$.
  • 4.
    $a.keeps_{F}=\mathrm{true}$ if and only if $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $a$.
[…]
Statement 6.63.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then $|A[E;F]|\leq n$.
Proof.
By Definition 6.62, Statement 4.45, we have $|A[E;F]|=|BSTEP\langle E\cup F,\emptyset\rangle|\leq n$. ∎
Intuitively if $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$, then there exists a $1$-context $\zeta$ over $\Sigma$ such that such that
  • $\zeta\neq\diamondsuit$,
  • $R\langle E\cup F,\emptyset\rangle$ reach $a$ from $\zeta[b]$, and
  • $R\langle E\cup F,\emptyset\rangle$ reduces the subterms of $\zeta[b]$ at different prefixes of $\mathrm{addr}(\zeta)$ – except possibly for the pair of $\varepsilon$ and $\mathrm{addr}(\zeta)$ – to different constants in $C\langle E\cup F,\emptyset\rangle$.
Thus the length of $\mathrm{addr}(\zeta)$ is less than equal to $|C\langle E\cup F,\emptyset\rangle|$.
Lemma 6.64.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Let $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$. Then there exists a $1$-context $\zeta$ over $\Sigma$, where
  • (a)
    $\zeta\neq\diamondsuit$,
  • (b)
    $k\geq 1$ denotes $\mathrm{length}(\mathrm{addr}(\zeta))$,
  • (c)
[…]
such that
  • (d)
    $\zeta[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=a$,
  • (e)
    $\zeta[b]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}\neq\zeta[b]|_{\alpha_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$ for any $0\leq i<j\leq k$ such that $1\leq i$ or $j\leq k-1$, and
  • (f)
    $k\leq|C\langle E\cup F,\emptyset\rangle|$.
Proof.
Intuitively, the basic idea of the proof is similar to that of the pumping lemma for regular languages [18], context-free languages [18], and regular tree languages [12], [19]. The main difference from the proof of the pumping lemma is that we iterate contracting the ground term rather than pumping it. We stop when we get such a small ground term that we cannot contract it any more.
It is sufficient to prove that there exists a $1$-context $\zeta$ over $\Sigma$ such that Conditions (a)–(e) hold. Because for any $1$-context $\zeta$ over $\Sigma$ satisfying Conditions (a)–(e), Condition (f) holds as well. By Statement 4.43, there exists a $1$-context $\delta$ over $\Sigma$ such that $\delta\neq\diamondsuit$ and $\delta[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=a$, i.e. Conditions (a)–(d) hold for $\zeta=\delta$. We now distinguish two cases:
Case 1: Conditions (e) holds for $\zeta=\delta$. Then we are done.
[…]
$\left(0\leq i<j\leq k-1\mbox{ or }1\leq i<j\leq k\right)$ and $\zeta[b]|_{\beta_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=\zeta[b]|_{\beta_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$.
We define $\delta^{\prime}\in CON_{\Sigma,1}$ from $\delta$ by replacing the subterm $\delta|_{\beta_{i}}$ at $\beta_{i}$ by the subterm $\delta|_{\beta_{j}}$ at $\beta_{j}$. Then
$\delta^{\prime}\neq\diamondsuit$, $\mathrm{length}(\mathrm{addr}(\delta^{\prime}))<\mathrm{length}(\mathrm{addr}(\delta))$, and $\delta^{\prime}[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=a$ i.e. Conditions (a)–(d) hold for $\zeta=\delta^{\prime}$.
We now distinguish two cases:
Case 2.1: Condition (e) holds for $\zeta=\delta^{\prime}$. Then Conditions (a)–(e) hold for $\zeta=\delta^{\prime}$. In this case we are done.
Case 2.2: Condition (e) does not hold for $\zeta=\delta^{\prime}$. In this case we let $\delta=\delta^{\prime}$ and repeat the procedure of contracting $\delta$ described above finitely many times until Condition (e) holds for $\zeta=\delta^{\prime}$. Then Conditions (a)–(e) hold for $\zeta=\delta^{\prime}$. ∎
Intuitively, we now consider two constants, $a$ and $b$ in $C\langle E\cup F,\emptyset\rangle$ and a sequence of rewrite rules in $R\langle{E\cup F},\emptyset\rangle$ of which consecutive applications leads us from $b$ to $a$. Let $a,b\in C\langle E\cup F,\emptyset\rangle$, $k\geq 1$, and for each $j=1,\ldots,k$, let
  • $\sigma_{j}(a_{j1},\ldots,a_{jm_{j}})\dot{\rightarrow}a_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$, be a rewrite rule and
[…]
such that
  • $b=a_{1i_{1}}$,
  • for each $j=1,\ldots,k-1$, $a_{j}=a_{(j+1)i_{j+1}}$, and
  • $a=a_{k}$.
The above sequence of rewrite rules is connected in the sense that we can apply them one after another, the right-hand side of each rule appears on the left-hand side of the next rule. Furthermore, the consecutive applications of these rewrite rules leads us from the constant $b$ to the constant $a$ since $b$ appears on the left-hand side of the first rule, and $a$ is the right-hand side of the last rule. We can apply the above rewrite rules one after another in the following way:
$b$ appears in the left-hand side of the first rewrite rule,
we apply the first rewrite rule to its left-hand side, $\sigma_{1}(a_{11},\ldots,a_{1m_{1}})$, we get the right-hand side $a_{1}$, which appears on the left-hand side of the second rewrite rule,
we apply the second rewrite rule to its left-hand side, $\sigma_{2}(a_{21},\ldots,a_{2m_{2}})$, we get the right-hand side $a_{2}$, which appears on the left-hand side of the third rewrite rule,
we apply the third rewrite rule to its left-hand side, $\sigma_{3}(a_{31},\ldots,a_{3m_{3}})$, we get the right-hand side $a_{3}$, which appears on the left-hand side of the fourth rewrite rule, and son on.
Finally, the right-hand side $a_{k-1}$ of the last but one rewrite rule appears in the left-hand side of the last rewrite rule,
we apply the last rewrite rule to its left-hand side, $\sigma_{k}(a_{k1},\ldots,a_{km_{k}})$, we get its right-hand side $a_{k}=a$.
We now state formally and show the following intuitive statement: if $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$, then there exists a connected sequence of rewrite rules in $R\langle E\cup F,\emptyset\rangle$ that leads us from the constant $b$ to the constant $a$. Moreover, $b$ and the intermediate constants between $b$ and $a$, appearing as the right-hand sides of the applied rewrite rules except for the last one, are pairwise different.
Lemma 6.65.
[…]
  • $b=a_{1i_{1}}$,
  • for each $j=1,\ldots,k-1$, $a_{j}=a_{(j+1)i_{j+1}}$,
  • $b\not\in\{\,a_{1},\ldots,a_{k-1}\,\}$ and for any $11\leq i<j\leq k-$ 1, $a_{i}\neq a_{j}$, and
  • $a=a_{k}$.
Proof.
By Lemma 6.64, there exists a $1$-context $\zeta$ over $\Sigma$, where
  • (a)
    $\zeta\neq\diamondsuit$,
  • (b)
    $k\geq 1$ denotes $\mathrm{length}(\mathrm{addr}(\zeta))$,
  • (c)
[…]
such that
  • (d)
    $\zeta[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=a$,
  • (e)
    $\zeta[b]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}\neq\zeta[b]|_{\alpha_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$ for any $0\leq i<j\leq k$ such that $1\leq i$ or $j\leq k-1$, and
  • (f)
    $k\leq|C\langle E\cup F,\emptyset\rangle|$.
As $\mathrm{addr}(\xi)=\alpha_{k}$, we have $\xi[b]|_{\alpha_{k}}=b$.
For each $i\in\{1,\ldots,k-1\,\}$, let $a_{i}=\zeta[b]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$.
Since $b=\zeta[b]|_{\alpha_{k}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$, by Condition (e) we have
$b\not\in\{\,a_{1},\ldots,a_{k-1}\,\}$ and for any $11\leq i<j\leq k-$ 1, $a_{i}\neq a_{j}$.
Then consider a sequence of $R\langle{E\cup F},\emptyset\rangle$ reductions, where we indicate the rewrite rules applied at the positions $\alpha_{k-1}$, $\alpha_{k-2}$, $\ldots$, $\alpha_{0}$:
$\zeta[b]=t_{0}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{1}\mathop{\rightarrow}_{[\alpha_{k-1},\sigma_{1}(a_{11},\ldots,a_{1m_{1}})\dot{\rightarrow}a_{1}]}t_{1}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{2}\mathop{\rightarrow}_{[\alpha_{k-2},\sigma_{2}(a_{21},\ldots,a_{2m_{2}})\dot{\rightarrow}a_{2}]}t_{2}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}\newline \cdots\mathop{\rightarrow}_{[\alpha_{1},\sigma_{k-1}(a_{(k-1)1},\ldots,a_{(k-1)m_{1}})\dot{\rightarrow}a_{k-1}]}t_{k-1}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{k}\mathop{\rightarrow}_{[\alpha_{0},\sigma_{k}(a_{k1},\ldots,a_{km_{k}})\dot{\rightarrow}a_{k}]}a_{k}=a$.
Here
  • $s_{i}\in T_{\Sigma\cup C\langle E\cup F,\emptyset\rangle}$ for every $i\in\{1,\ldots,k\,\}$, and $t_{i}\in T_{\Sigma\cup C\langle E\cup F,\emptyset\rangle}$ for every $i\in\{0,\ldots,k-1\,\}$,
  • for each $i\in\{1,\ldots,k\,\}$, $a_{i}=\zeta[b]|_{\alpha_{k-i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$,
[…]
  • for each $j\in\{1,\ldots,k\,\}$, $\sigma_{j}(a_{j1},\ldots,a_{jm_{j}})\dot{\rightarrow}a_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$,
[…]
  • $a=a_{k}$. ∎
Lemma 6.66.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. If $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$, then there exists a path of positive length or a cycle from the vertex $a$ to the vertex $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
Proof.
We have the lemma by Lemma 6.65, Statement 4.43, and by the definition of $A[E;F]$ in Definition 6.62. ∎
We now show the converse of Lemma 6.66. Intuitively, if there exists a connected sequence of rewrite rules in $R\langle E\cup F,\emptyset\rangle$ that leads us from a constant $b\in C\langle E\cup F,\emptyset\rangle$ to a constant $a\in C\langle E\cup F,\emptyset\rangle$, then $R\langle E\cup F,\emptyset\rangle$ reach the constant $a$ from a proper extension of the constant $b$.
Lemma 6.67.
[…]
$\sigma(s_{1},\ldots,s_{i-1},a_{i},s_{i+1},\ldots,s_{m})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\sigma(a_{1},\ldots,a_{m})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a$.
Proof.
By Definition 4.26, Definition 4.28, and Proposition 4.32 we have the lemma. ∎
Lemma 6.68.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Let $k\geq 1$, and $(a_{k},a_{k-1}),\,(a_{k-1},a_{k-2}),\,\ldots,\,(a_{1},a_{0})\in A[E;F]$. Then there exists a $1$-context $\delta$ over $\Sigma$ such that $\delta\neq\diamondsuit$ and $\delta[a_{0}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}a_{k}$.
Proof.
By Definition 6.62 and Definition 4.44, for each $j=1,\ldots,k$, there exists a rewrite rule $\sigma_{j}(a_{j1},\ldots,a_{jm_{j}})\dot{\rightarrow}a_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$, such that
[…]
We proceed by induction on $k$.
Base case: $k=1$. By Lemma 6.67, we are done.
Induction Step: Let $k\geq 1$, and assume that the statement of the lemma is true for $k$. We now show that it is true for $k+1$. By the induction hypothesis, there exists a $1$-context $\delta$ over $\Sigma$ such that
$\delta\neq\diamondsuit$ and $\delta[a_{0}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}a_{k}$.
By Lemma 6.67, there exist ground terms $s_{1},\ldots,s_{i-1},s_{i+1},\ldots,s_{m_{k+1}}\in T_{\Sigma}$ such that
$\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},a_{k},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})=\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},a_{(k+1)i_{k+1}},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\newline \sigma_{k+1}(a_{(k+1)1},\ldots,a_{(k+1)i_{k+1}-1},a_{(k+1)i_{k+1}},a_{(k+1)i_{k+1}+1},\ldots,a_{(k+1)m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a_{k+1}$.
Thence we have
$\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},\delta[a_{0}],s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},a_{k},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})=\newline \sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},a_{(k+1)i_{k+1}},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\newline \sigma_{k+1}(a_{(k+1)1},\ldots,a_{(k+1)i_{k+1}-1},a_{(k+1)i_{k+1}},a_{(k+1)i_{k+1}+1},\ldots,a_{(k+1)m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a_{k+1}$. ∎
Lemma 6.69.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. If there exists a path of positive length or a cycle from a vertex $a$ to a vertex $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E\cup F,\emptyset\rangle$ reach the constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of the constant $b\in C\langle E\cup F,\emptyset\rangle$.
Proof.
We have the lemma by Lemma 6.68. ∎
Lemma 6.70.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$ if and only if there is a path of positive length or a cycle from the vertex $a$ to the vertex $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
Proof.
By Lemma 6.66 and Lemma 6.69, we have the lemma. ∎
7 Running Examples Further Continued
We continue our running examples, for each one of them we present the dpwpa $AUX[E;F]$.
Example 7.71.
We now continue Examples 3.17 and 5.50. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F]$ for the GTESs $E$ and $F$. Recall that
[…]
Observe that $A[E;F]=\{\,(f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$/_{\Theta\langle E\cup F,\emptyset\rangle})\,\}$.
  • 1.
    $equal_{E},\;equal_{F},\;keeps_{E},\;keeps_{F}$ are attributes of the vertices in $C\langle E\cup F,\emptyset\rangle$.
  • 2.
    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$,  \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, and
    $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$.
    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,  \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and
    $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$.
  • 3.
    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$,  \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}).keeps_{E}=\mathrm{true}$, and
    $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}).keeps_{E}=\mathrm{true}$.
    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$,  \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$, and
    $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
Example 7.72.
We now continue Examples 3.18 and 5.51. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Recall that
$C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;f(\#),\;f^{2}(\#),\;f^{3}(\#)\,\}\,\}=\{\,W\,\}$.
Observe that $A[E;F]=\{\,(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\,\}$.
  • 1.
    $equal_{E},\;equal_{F},\;keeps_{E},\;keeps_{F}$ are attributes of the vertices in $(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
  • 2.
    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$   and   #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$.
  • 3.
    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{false}$   and   #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
Example 7.73.
We now continue Example 3.21 and its sequel 5.54. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. By definition,
$C\langle E\cup F,\emptyset\rangle,A[E;F]\rangle=C\langle E\cup F,\emptyset\rangle$.
Observe that
$A[E;F])=\{\,(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\,\}$.
By direct inspection of $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$, we get that
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}\not\in C\langle E,F\rangle$ hence #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$,
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}\not\in C\langle F,E\rangle$ hence #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$.
By direct inspection of $C\langle E,F\rangle$, $R\langle E,F\rangle$, $C\langle E\cup F,\emptyset\rangle$, and $R\langle E\cup F,\emptyset\rangle$, we get that $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$, hence #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$.
The rewrite rule $f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ is in $R\langle E\cup F,\emptyset\rangle$ and $f(\#,\$)/_{\Theta\langle F,\emptyset\rangle}\subseteq\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. However, in $R\langle F,E\rangle$ there does not exist a rewrite rule with left-hand side $f(f(\#,\$)/_{\Theta\langle F,\emptyset\rangle},f(\#,\$)/_{\Theta\langle F,\emptyset\rangle})$. Hence $R\langle F,E\rangle$ does not keep up with $R\langle E\cup F,\emptyset\rangle$ writing #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
Example 7.74.
We now continue Examples 3.22 and 5.55. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Observe that
$A[E;F]=\{\,(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(\$/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(\$/_{\Theta\langle E\cup F,\emptyset\rangle},\$/_{\Theta\langle E\cup F,\emptyset\rangle})\,\}$.
By direct inspection of $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$, we get that
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,
\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$,
We now show that GTRS $R\langle F,\emptyset\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
  • #$\#/_{\Theta\langle F,E\rangle}=\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Furthermore, #$\#/_{\Theta\langle F,E\rangle}$ is the only element of $C\langle F,E\rangle$ which is a subset of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
  • For the rewrite rule #˙#$\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\in R\langle E\cup F,\emptyset\rangle$, there is a rewrite rule with left-hand side ##\## in $R\langle F,E\rangle$.
  • For the rewrite rule $f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\in R\langle E\cup F,\emptyset\rangle$, there is a rewrite rule with left-hand side $f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})$ in $R\langle F,E\rangle$.
Consequently, GTRS $R\langle F,E\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing ##\##, and
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
We now show that GTRS $R\langle F,E\rangle$ does not keep up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
  • $g(\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle}\in R\langle E\cup F,\emptyset\rangle$,
  • $g(g(\$))/_{\Theta\langle F,E\rangle})\subseteq\$/_{\Theta\langle E\cup F,\emptyset\rangle})$, and
  • in $R\langle E\cup F,\emptyset\rangle$, there does not exist a rewrite rule with left-hand side $g(g(g(\$))/_{\Theta\langle F,E\rangle})$.
Thus GTRS $R\langle F,E\rangle$ does not keep up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$, and
\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
8 Constructing the Auxiliary dpwpa for the GTESs $E$ and $F$
We construct the sets and relations and dpwpa, which we introduced in Section 6 and our decision procedure needs.
[…]
Definition 8.75.
Algorithm Constructing Auxiliary dpwpa $AUX[E;F]$ (CAD for short)
Input: A signature $\Sigma$ and GTESs $E$ and $F$ over $\Sigma$.
Output:
  • the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$;
  • the set of nullary symbols $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$;
  • $INC(E\cup F,E)$ stored without duplicates in a red-black tree saved in a variable $RBT1$, for every ordered pair$(a,b)\in INC\langle E\cup F,E\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
  • $INC(E\cup F,F)$ stored in a red-black tree saved without duplicates in a variable $RBT2$, for every ordered pair$(a,b)\in INC(E\cup F,F)$, the search key is the first component $a$, the second component, $b$ is satellite data;
  • $NUM(E\cup F,E)$ stored in a red-black tree saved without duplicates in a variable $RBT3$, for every ordered pair$(a,k)\in NUM(E\cup F,E)$, the search key is the first component $a$, the second component, $k$ is satellite data;
  • $NUM(E\cup F,F)$ stored in a red-black tree saved without duplicates in a variable $RBT4$ for every ordered pair$(a,k)\in NUM(E\cup F,F)$, the search key is the first component $a$, the second component, $k$ is satellite data;
  • the GTRS $R\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved without duplicates in a variable $RBT5$; each node in $RBT5$ contains a serach key which is a rule in $R\langle E\cup F,\emptyset\rangle$, and satellite data $counter_{E}$ and $counter_{F}$;
  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved without duplicates in a variable $RBT6$; for every ordered pair$(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
  • the GTRS $R\langle E,F\rangle$ stored in a red-black tree saved without duplicates in a variable $RBT7$;
  • the GTRS $R\langle F,E\rangle$ stored in a red-black tree saved without duplicates in a variable $RBT8$;
  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$;
var $i$, $m$: integer; $\sigma$: symbol in $\Sigma$;
$a,a_{1},\ldots,a_{mar_{\Sigma}}:C\langle E\cup F,\emptyset\rangle$;  $b,b_{1},\ldots,b_{mar_{\Sigma}}:C\langle E,F\rangle$;   $c,c_{1},\ldots,c_{mar_{\Sigma}}:C\langle F,E\rangle$;
$RBT1$, $RBT2$, $RBT3$, $RBT4$, $RBT6$, $RBT5$, $RBT7$, $RBT8$: red-black trees.
1. Create a subterm dag $G=(V,A,\lambda)$ for $E$ and $F$.
By Proposition 2.16, compute the congruence closure $\rho\langle E,F\rangle$ of the relation $\tau\langle E,F\rangle=\{\,(u,v)\mid(\hat{u},\hat{v})\in E\,\}.$
$C\langle E,F\rangle:=\{\,x/_{\rho\langle E,F\rangle}\mid x\in V\,\}$.
By Proposition 2.16, compute the congruence closure $\rho\langle F,E\rangle$ of the relation $\tau\langle F,E\rangle=\{\,(u,v)\mid(\hat{u},\hat{v})\in W\,\}.$
$C\langle F,E\rangle:=\{\,x/_{\rho\langle F,E\rangle}\mid x\in V\,\}$.
Compute $\tau\langle E\cup F,\emptyset\rangle$ using the equality $\tau\langle E\cup F,\emptyset\rangle=\tau\langle E,F\rangle\cup\tau\langle F,E\rangle$.
By Proposition 2.16, compute the congruence closure $\rho\langle E\cup F,\emptyset\rangle$ of the relation $\tau\langle E\cup F,\emptyset\rangle=\{\,(u,v)\mid(\hat{u},\hat{v})\in E\cup F\,\}.$
$C\langle E\cup F,\emptyset\rangle:=\{\,x/_{\rho\langle E\cup F,\emptyset\rangle}\mid x\in V\,\}$.
2.{ Computing the relations $INC(E\cup F,E)$, $INC(E\cup F,F)$, $NUM(E\cup F,E)$, $NUM(E\cup F,F)$, $BSTEP\langle E\cup F,\emptyset\rangle$, the GTRSs $R\langle E\cup F,\emptyset\rangle$, $R\langle E,F\rangle$, $R\langle F,E\rangle$, the values of the attributes $equal_{E}$, $equal_{F}$ for every vertex in $C\langle E,F\rangle$, and the satellite data $counter_{E}$, and $counter_{F}$ for every rule in $R\langle E\cup F,\emptyset\rangle$. }
begin
$R\langle E,F\rangle:=\emptyset$;
$R\langle F,E\rangle:=\emptyset$;
$R\langle E\cup F,\emptyset\rangle:=\emptyset$;
$INC(E\cup F,E):=\emptyset$;
$INC(E\cup F,F):=\emptyset$;
$NUM(E\cup F,E):=\emptyset$;
$NUM(E\cup F,F):=\emptyset$;
$BSTEP\langle E\cup F,\emptyset\rangle:=\emptyset$
end
for each $c\in C\langle E\cup F,\emptyset\rangle$ do
begin
$c.equal_{E}=\mathrm{\mathrm{false}};$;
$c.equal_{F}=\mathrm{\mathrm{false}};$
end
for each $x\in V$ do
begin
$\sigma:=\lambda(x)$;
$m:=outdegree(x);$
$a:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x)$;
$b:=\mathrm{FIND}_{\rho\langle F,E\rangle}(x)$;
$c:=\mathrm{FIND}_{\rho\langle E\cup F,\emptyset\rangle}(x)$;
$INC(E\cup F,E):=INC(E\cup F,E)\cup\{\,(c,a)\,\}$;
{ store $INC(E\cup F,E)$ without duplicates in the red-black tree $RBT1$.}
$INC(E\cup F,F):=INC(E\cup F,F(\cup\{\,(c,b)\,\}$;
{ store $INC(E\cup F,F)$ without duplicates in the red-black tree $RBT2$.}
if $|a|=|c|$
${\bf then}$ $c.equal_{E}:=\mathrm{\mathrm{true}};$
if $|b|=|c|$
${\bf then}$ $c.equal_{F}:=\mathrm{\mathrm{true}};$
{ By Proposition 2.16 we know the cardinalities of all $W/_{\Theta\langle E,F\rangle}$-classes, $W/_{\Theta\langle F,E\rangle}$-classes, and $W/_{\Theta\langle E\cup F,\emptyset\rangle}$-classes. }
for each $1\leq i\leq m$ do
begin
$a_{i}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x/i)$;
$b_{i}:=\mathrm{FIND}_{\rho\langle F,E\rangle}(x/i)$;
$c_{i}:=\mathrm{FIND}_{\rho\langle E\cup F,\emptyset\rangle}(x/i)$;
end
if $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c\not\in R\langle E\cup F,\emptyset\rangle$
then begin
$R\langle E\cup F,\emptyset\rangle:=R\langle E\cup F,\emptyset\rangle\cup\{\,\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c\,\}$;
{ Store $R\langle E\cup F,\emptyset\rangle$ without duplicates in the red-black tree $RBT5$. }
$\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}:=0$;
$\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{F}:=0$;
$BSTEP\langle E\cup F,\emptyset\rangle:=BSTEP\langle E\cup F,\emptyset\rangle\cup\{\,(c_{1},c),\ldots,(c_{m},c)\,\}$;
{ Store $BSTEP\langle E\cup F,\emptyset\rangle$ without duplicates in the red-black tree $RBT6$}
end
if $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\not\in R\langle E,F\rangle$
then begin
$R\langle E,F\rangle:=R\langle E,F\rangle\cup\{\,\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\,\}$;
{ Store $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT7$. }
$\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}:=\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}+1$;
end
if $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b\not\in R\langle F,E\rangle$
then begin
$R\langle F,E\rangle:=R\langle F,E\rangle\cup\{\,\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b\,\}$;
{ Store $R\langle F,E\rangle$ without duplicates in the red-black tree $RBT8$. }
$\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{F}:=\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{F}+1$;
end
end
3. { Computing the relations $NUM(E\cup F,E)$, $NUM(E\cup F,F)$. }
Read the nodes of the red-black tree $RBT1$ in search key order, during this process create the red-black tree $RBT3$ containing without duplicates the relation $NUM(E\cup F,E)$;
Read the nodes of the red-black tree $RBT2$ in search key order, during this process create the red-black tree $RBT4$ containing without duplicates the relation $NUM(E\cup F,F)$;
4. { Creating the adjacency-list representation of the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$
We build the adjacency-list representation of the dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ by reading in search key order the nodes of the black-red tree $RBT6$ storing the relation $BSTEP\langle E\cup F,\emptyset\rangle$. Recall that for every ordered pair $(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data.
5. { Computing the values of the attribute $keeps_{E}$. }
for each $a\in C\langle E\cup F,\emptyset\rangle$ do
$a.keeps_{E}:=\mathrm{true}$;
for each rewrite rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle$ with $m\in{\mathbb{N}}$ do
if $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a.counter_{E}\neq\Pi_{i=1}^{m}|a_{i}/_{\Theta\langle E,F\rangle}|$
then $a.keeps_{E}:=\mathrm{false}$
6. { Computing the values of the attribute $keeps_{F}$. }
for each $a\in C\langle E\cup F,\emptyset\rangle$ do
$a.keeps_{F}:=\mathrm{true}$;
for each rewrite rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle$ with $m\in{\mathbb{N}}$ do
if $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a.counter_{F}\neq\Pi_{i=1}^{m}|a_{i}/_{\Theta\langle F,E\rangle}|$
then $a.keeps_{F}:=\mathrm{false}$
Lemma 8.76.
Let $E$ and $F$ be GTESs over a signature $\Sigma$.
1. Algorithm CAD on $\Sigma$, $E$, and $F$ produces
  • the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$;
  • the set of nullary symbols $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$;
  • $INC(E\cup F,E)$ stored without duplicate values in a red-black tree saved in a variable $RBT1$;
  • $INC(E\cup F,F)$ stored without duplicate values in a red-black tree saved in a variable $RBT2$;
  • $NUM(E\cup F,E)$ stored without duplicate values in a red-black tree saved in a variable $RBT3$;
  • $NUM(E\cup F,F)$ stored without duplicate values in a red-black tree saved in a variable $RBT4$;
  • the GTRS $R\langle E\cup F,\emptyset\rangle$ stored without duplicate values in a red-black tree saved in a variable $RBT5$, each node in $RBT5$ contains a search key which is a rule in $R\langle E\cup F,\emptyset\rangle$, and satellite data $counter_{E}$ and $counter_{F}$;
  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored without duplicate values in a red-black tree saved in a variable $RBT6$; for every ordered pair$(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
  • the GTRS $R\langle E,F\rangle$ stored without duplicate values in a red-black tree saved in a variable $RBT7$;
  • the GTRS $R\langle F,E\rangle$ stored without duplicate values in a red-black tree saved in a variable $RBT8$;
  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$.
2. Algorithm CAD on $\Sigma$, $E$, and $F$ decides for every constant $a\in C\langle E\cup F,\emptyset\rangle$, whether $a\in C\langle E,F\rangle$ and whether $a\in C\langle F,E\rangle$, and sets the value of the attribute $a.equal_{E}$ to true if and only if $a\in C\langle E,F\rangle$, and sets the value of the attribute $a.equal_{F}$ to true if and only if $a\in C\langle F,E\rangle$.
[…]
Proof.
By Proposition 4.47 and by direct inspection of Definition 8.75, we obtain Statement 1.
By Proposition 2.16, we know the cardinality of all $W/_{\Theta\langle E,F\rangle}$-classes, $W/_{\Theta\langle F,E\rangle}$-classes, and $W/_{\Theta\langle E\cup F,\emptyset\rangle}$-classes. For all $a\in R\langle E,F\rangle$ and $c\in R\langle E\cup F,\emptyset\rangle$ such that $a\cap c\neq\emptyset$, we have $a=c$ if and only if $|a|=|c|$. Hence we have Statement 2.
We now show Statement 3. For every rewrite rule $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c\in R\langle E\cup F,\emptyset\rangle$, we decide whether for all $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$ with $a_{1}\subseteq c_{1},\ldots,a_{m}\subseteq c_{m}$, there exists a rewrite rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R\langle E,F\rangle$. When constructing the GTRS $R\langle E,F\rangle$, whenever we add to $R\langle E,F\rangle$ a new rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$, where $a_{1}\subseteq a_{1},\ldots,a_{m}\subseteq c_{m}$, we add $1$ to the attribute $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}$. In this way, we count the rules in $R\langle E,F\rangle$ with left-hand side $\sigma(a_{1},\ldots,a_{m})$, where $a_{1}\subseteq c_{1},\ldots,a_{m}\subseteq c_{m}$, and store the result in the attribute $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}$. Then $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}=\Pi_{i=1}^{m}|c_{i}/_{\Theta\langle E,F\rangle}|$ if and only if for all $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$ with $a_{1}\subseteq c_{1},\ldots,a_{m}\subseteq c_{m}$, there exists a rewrite rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R\langle E,F\rangle$. When Algorithm CAD terminates, for every vertex $c\in C\langle E\cup F,\emptyset\rangle$,
$c.keeps_{E}=\mathrm{true}$ if and only if
for every rewrite rule of the form $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c\in R\langle{E\cup F},\emptyset\rangle$,
[…]
there exists a rewrite rule in $R\langle E,F\rangle$ of the form $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a$ for some $a\in C\langle E,F\rangle$.
The proof of Satement 4 is similar to that of Statement 3. ∎
Lemma 8.77.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Algorithm CAD runs in $O(n\,\mathrm{log}\,n)$ time on $\Sigma$, $E$, and $F$, where $n=size(E)+size(F)$.
Proof.
The proof consists of showing the following nine statements:
1. The time complexity of constructing the red-black trees $RBT1$ and $RBT2$ is $O(n)$.
2. The time complexity of constructing the relation $\tau\langle E\cup F,\emptyset\rangle$ and the GTES $E\cup F$ stored without duplicate values in a red-black tree saved in the variable $RBT5$, is $O(n^{2})$ time.
3. The time complexity of computing the value attributes $a.equal_{E}$ and $a.equal_{F}$ for all constants $a\in C\langle E\cup F,\emptyset\rangle$, is $O(n)$.
4. The time complexity of constructing the following data is $O(n\,\mathrm{log}\,n)$.
  • the GTRS $R\langle E\cup F,\emptyset\rangle$ stored without duplicate values in a red-black tree saved in the variable $RBT5$, each node in $RBT5$ contains a search key which is a rule in $R\langle E\cup F,\emptyset\rangle$;
  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in the variable $RBT6$;
  • the GTRS $R\langle E,F\rangle$ stored without duplicate values in a red-black tree saved in the variable $RBT7$;
  • the GTRS $R\langle F,E\rangle$ stored without duplicate values in a red-black tree saved in the variable $RBT8$;
5. The time complexity of constructing
  • $NUM(E\cup F,E)$ stored without duplicate values in a red-black tree saved in a variable $RBT3$, and
  • $NUM(E\cup F,F)$ stored without duplicate values in a red-black tree saved in a variable $RBT4$
is $O(n)$.
6. The time complexitity of computing the values of the satellite data $counter_{E}$, and $counter_{F}$ for all rules in $R\langle E\cup F,\emptyset\rangle$ is $O(n)$.
7. The time complexity of constructing the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ is $O(n)$.
8. The time complexitity of computing the values of the attributes $equal_{E}$ and $equal_{F}$ of all constants $a\in C\langle E\cup F,\emptyset\rangle$ is $O(n)$.
9. The time complexitity of computing the values of the attributes $keeps_{E}$ and $keeps_{F}$ of all constants $a\in C\langle E\cup F,\emptyset\rangle$ is $O(n\,\mathrm{log}\,n)$ time.
First we now show Statement 1. By Statement 6.59, $INC(E\cup F,E)leqn$. Consequently, when we construct the red-black tree $RBT1$, we carry out at most $n$ insertion. By Proposition 2.2, we construct the red-black tree $RBT1$ in $O(n\,\mathrm{log}\,n)$ time. Similarly, we construct the red-black tree $RBT2$ in $O(n\,\mathrm{log}\,n)$ time.
Second, by Statement 2.4, we have Statement 2.
Third, we carry out $2n$ comparisons of nonnegative numbers in $O(n)$ time.
Fourth, By Statement 2 and Statement 3 of 4.47, we have Statement 4.
Fifth, we now show Statement 5. When we construct the red-black tree $RBT3$, we read the whole red-black tree $RBT1$ in search key order in $O(n)$ time. Each modification takes constant time. By Statement 6.59, $|INC(E\cup F,E)|\leq n$. Consequently, we construct the red-black tree $RBT3$ in $O(n)$ time. Similarly, we construct the red-black tree $RBT4$ in $O(n)$ time.
Sixth, we now show Statement 6. When Algorithm CAD computes the values of the satellite data $counter_{E}$ and $counter_{F}$ for all rules of the GTRS $R\langle E\cup F,\emptyset\rangle$, it carries out at most $2n$ additions, since $|C\langle E,F\rangle|\leq|W|\leq n$ and $|C\langle F,E\rangle|\leq|W|\leq n$.
Seventh we now show Statement 7. We build the adjacency-list representation of the dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ by reading in search key order the nodes of the black-red tree $RBT6$ storing the relation $BSTEP\langle E\cup F,\emptyset\rangle$. Recall that for every ordered pair $(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data. Then the time complexity of building the adjacency-list representation of the dpwpa $AUX[E;F]$ is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$. By Statement 4.27 and Statement 4.45 $|C\langle E\cup F,\emptyset\rangle|\leq|W|\leq n$. By Statement 6.63, $|A[E;F]|=|BSTEP\langle E\cup F,\emptyset\rangle|\leq size(E)+size(F)=n$. Therefore the time complexity of building the adjacency-list representation of the dpwpa $AUX[E;F]$ is $O(n)$.
Eighth, Statement 8.78 implies Statement 8.
Nineth, we now show Statement 9. We now estimate the number of operations when computing the values of the attributes $a.keeps_{E}$ for all $a\in C\langle E\cup F,\emptyset\rangle$. For each rewrite rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle$, Algorithm CAD carries out $m$ searches in the red-black tree $RBT3$ and less than or equal to $m$ multiplications. By Statement 4.45, $size(R\langle E\cup F,\emptyset\rangle)\leq 2\cdot n$. In this way, when we compute the values of the attribute $keeps_{E}$ for all rewrite rules $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle$, the number of searches in the red-black tree $RBT3$ is less than equal to $n$. Consequently, by Proposition 2.1, the searches in the red-black tree $RBT3$ take $O(n\,\mathrm{log}\,n)$ time.
Furthermore,
the number of multiplications $\leq\sum_{\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle}m\leq size(R\langle E\cup F,\emptyset\rangle)\leq 2n$
by Statement 4.45. Thus, the number of multiplications is less than or equal to $n$.
For each $a\in C\langle E\cup F,\emptyset\rangle$, the number of comparisons of nonnegative integers during computing the value of the attribute $a.keeps_{E}$ is equal to the number of rules in $R\langle E\cup F,\emptyset\rangle$ with right-hand side $a$. Consequently, the number of comparisons of nonnegative integers during computing the values of the attributes $a.keeps_{E}$ for all $a\in C\langle E\cup F,\emptyset\rangle$ is equal to $|R\langle E\cup F,\emptyset\rangle|$. By Statement 4.45
$|R\langle E\cup F,\emptyset\rangle|\leq n$.
Consequently the time complexity of computing for all $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attribute $a.keeps_{E}$ is $O(n\,\mathrm{log}\,n)$. Similarly, the time complexity of computing for all $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attribute $a.keeps_{F}$ is $O(n\,\mathrm{log}\,n)$.
The Lemma follows from Statements 1, 2, 3, 4, 5, 6, 7, 8, and 9. ∎
Statement 8.78.
1. We can decide in $O(n)$ time whether for all $a\in C\langle E\cup F,\emptyset\rangle$, $a\in C\langle E,F\rangle$.
2. We can decide in $O(n)$ time whether for all $a\in C\langle E\cup F,\emptyset\rangle$, $a\in C\langle F,E\rangle$.
Proof.
First we show Statement 1. We read each nodes of the red-black tree $RBT3$ in lexicographic order. For every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,k)\in NUM(E\cup F,E)$. If $k=1$, then by Statement 6.60, $a\in C\langle E,F\rangle$. Otherwise, $a\not\in C\langle E,F\rangle$.
We can show Statement 2 similarly to Statement 1. ∎
9 Preparatory Results
The decidability and complexity results of the following sections require some preparatory results which we group together here.
Proposition 9.79.
Let $A$ be a set, $a\in A$, and let $\rho,\tau,\omega$ be equivalence relations on $A$ such that $a/_{\rho}\subset a/_{\omega}$ and $a/_{\tau}\subset a/_{\omega}$. Then $a/_{\rho}\cup a/_{\tau}\subset a/_{\omega}$.
Proof.
There exist $b,c\in A$, such that $b\in a/_{\omega}\setminus a/_{\rho}$ and $c\in a/_{\omega}\setminus a/_{\tau}$. We now distinguish three cases.
Case 1: $(b,c)\in\rho$. Then by the definition of $b$, $(a,c)\not\in\rho$. By the definition of $c$, $(a,c)\not\in\tau$. Thence $(a,c)\not\in\rho\cup\tau$.
Case 2: $(b,c)\in\tau$. Then by the definition of $c$, $(a,b)\not\in\tau$. By the definition of $b$, $(a,b)\not\in\rho$. Consequently $(a,b)\not\in\rho\cup\tau$.
Case 3: $(b,c)\not\in\rho\cup\tau$. Then by the definition of $b$ and $c$, $(b,c)\in\omega$, and hence $(b,c)\in\omega\setminus(\rho\cup\tau)$.
In all three cases we have $a/_{\rho}\cup a/_{\tau}\subset a/_{\omega}$. ∎
Proposition 9.80.
Let $A$ be a set, $a\in A$, and let $\rho,\tau,\omega$ be equivalence relations on $A$ such that $a/_{\rho}\cup a/_{\tau}=a/_{\omega}$. Then $a/_{\rho}=a/_{\omega}$ or $a/_{\tau}=a/_{\omega}$.
Proof.
By contradiction, assume that $a/_{\rho}\neq a/_{\omega}$ and $a/_{\tau}\neq a/_{\omega}$. Then $a/_{\rho}\subset a/_{\omega}$ and $a/_{\tau}\subset a/_{\omega}$. Consequently, by Proposition 9.79, $a/_{\rho}\cup a/_{\tau}\subset a/_{\omega}$. ∎
Lemma 9.81.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. If GTRS $R\langle E,F\rangle$ is total, then $R\langle E\cup F,\emptyset\rangle$ is total as well.
Proof.
[…]
Proposition 9.82.
For any GTESs $E$ and $F$ over a signature $\Sigma$, the following four statements are equivalent:
  • 1.
    $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
  • 2.
    $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
  • 3.
    There exists a GTES $H$ over $\Sigma$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
  • 4.
    $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$.
Proof.
$(1\Rightarrow 2)$ For the GTES $E\cup F$, we have $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subseteq\mathop{\leftrightarrow}^{*}_{E\cup F}$. Thence $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
$(2\Rightarrow 3)$ For the GTES $E\cup F$, we have $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
$(3\Rightarrow 4)$ $\mathop{\leftrightarrow}^{*}_{H}$ is a congrunce on the ground term algebra ${\cal T}(\Sigma)$, hence $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$.
$(4\Rightarrow 1)$ $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$. Since $\mathop{\leftrightarrow}^{*}_{E\cup F}$ is the least congruence on $T_{\Sigma}$ containing $E\cup F$, we have $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. ∎
In light of Proposition 9.82, from now on we decide for GTESs $E$ and $F$ over a signature $\Sigma$, whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Lemma 9.83.
Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. Let $W=ST\langle E\cup F\rangle$.
$\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if for each $p\in W$, $\left(p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{E}}\mbox{ or }p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{F}}\right)$.
Proof.
($\Rightarrow$) If $W=\emptyset$, then we are done. Assume that $W\neq\emptyset$. Let $p\in W$, $s\in T_{\Sigma}$, and $p\mathop{\leftrightarrow}^{*}_{E\cup F}s$. Then $p\mathop{\leftrightarrow}^{*}_{E}s$ or $p\mathop{\leftrightarrow}^{*}_{F}s$. Consequently $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\subseteq p/_{\mathop{\leftrightarrow}^{*}_{E}}\cup p/_{\mathop{\leftrightarrow}^{*}_{F}}$. Since $p/_{\mathop{\leftrightarrow}^{*}_{E}}\cup p/_{\mathop{\leftrightarrow}^{*}_{F}}\subseteq p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$, we have $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{E}}\cup p/_{\mathop{\leftrightarrow}^{*}_{F}}$. By Proposition 9.80, $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{E}}$ or $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{F}}$.
($\Leftarrow$) Let $s,t\in T_{\Sigma}$ be arbitrary such that $s\mathop{\leftrightarrow}^{*}_{E\cup F}t$. Then
$s=q_{0}\mathop{\leftrightarrow}_{[\alpha_{1},(l_{1},r_{1})]}q_{1}\mathop{\leftrightarrow}_{[\alpha_{2},(l_{2},r_{2})]}q_{2}\mathop{\leftrightarrow}_{[\alpha_{3},(l_{3},r_{3})]}\cdots\mathop{\leftrightarrow}_{[\alpha_{n},(l_{n},r_{n})]}q_{n}=t$ for some $n\in{\mathbb{N}}$,
[…]
$s\mathop{\leftrightarrow}^{*}_{E}t$ or $s\mathop{\leftrightarrow}^{*}_{F}t$.
[…]
  • $addr(\delta)=\alpha_{i}$,
  • $s=\delta[v]$, $t=\delta[z]$, and
  • $q_{i-1}=\delta[l_{i}]$, $q_{i}=\delta[r_{i}]$.
So,
$v\mathop{\leftrightarrow}^{*}_{E\cup F}l_{i}\mathop{\leftrightarrow}_{E\cup F}r_{i}\mathop{\leftrightarrow}^{*}_{E\cup F}z$  or   $v\mathop{\leftrightarrow}^{*}_{E\cup F}r_{i}\mathop{\leftrightarrow}_{E\cup F}l_{i}\mathop{\leftrightarrow}^{*}_{E\cup F}z$.
Consequently, $v,z\in l_{i}/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. As $l_{i}/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=l_{i}/_{\mathop{\leftrightarrow}^{*}_{E}}$ or $l_{i}/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=l_{i}/_{\mathop{\leftrightarrow}^{*}_{F}}$, we have $v,z\in l_{i}/_{\mathop{\leftrightarrow}^{*}_{E}}$ or $v,z\in l_{i}/_{\mathop{\leftrightarrow}^{*}_{F}}$. Therefore
$v\mathop{\leftrightarrow}^{*}_{E}z$ or $v\mathop{\leftrightarrow}^{*}_{F}z$.
Then
$s=\delta[v]\mathop{\leftrightarrow}^{*}_{E}\delta[z]=t$ or $s=\delta[v]\mathop{\leftrightarrow}^{*}_{F}\delta[z]=t$. ∎
Proposition 9.84.
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=size(E)+size(F)$. Let $W=ST\langle E\cup F\rangle$. We can decide in $O(n^{3})$ time whether for each $p\in W$,
$p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{E}}$ or $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{F}}$.
Proof.
By Proposition 2.10 and Proposition 2.11 we have the result. ∎
Lemma 9.85.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E\cup F,\emptyset\rangle$ is total. Let $W=ST\langle E\cup F\rangle$. Then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if for each $p\in W$,
$p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{E}}$ or $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{F}}$.
Proof.
($\Rightarrow$) Let $p\in W$, $s\in T_{\Sigma}$, and $p\mathop{\leftrightarrow}^{*}_{E\cup F}s$. Then $p\mathop{\leftrightarrow}^{*}_{E}s$ or $p\mathop{\leftrightarrow}^{*}_{F}s$. Consequently $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\subseteq p/_{\mathop{\leftrightarrow}^{*}_{E}}\cup p/_{\mathop{\leftrightarrow}^{*}_{F}}$. Observe that $p/_{\mathop{\leftrightarrow}^{*}_{E}}\cup p/_{\mathop{\leftrightarrow}^{*}_{F}}\subseteq p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. Thus $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{E}}\cup p/_{\mathop{\leftrightarrow}^{*}_{F}}$. By Proposition 9.80, $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{E}}$ or $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{F}}$.
($\Leftarrow$) Let $s,t\in T_{\Sigma}$ be arbitrary such that $s\mathop{\leftrightarrow}^{*}_{E\cup F}t$. Then by Lemma 4.41, $s\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}p/_{\Theta\langle E\cup F,\emptyset\rangle}$ for some $p\in W$. Consequently, $s\mathop{\leftrightarrow}^{*}_{E\cup F}p$. Since $s\mathop{\leftrightarrow}^{*}_{E\cup F}t$, we have $t\mathop{\leftrightarrow}^{*}_{E\cup F}p$. Then by our assumption,
$s\mathop{\leftrightarrow}^{*}_{E}p$ and $t\mathop{\leftrightarrow}^{*}_{E}p$ or $s\mathop{\leftrightarrow}^{*}_{F}p$ and $t\mathop{\leftrightarrow}^{*}_{F}p$.
For this reason $s\mathop{\leftrightarrow}^{*}_{E}t$ or $s\mathop{\leftrightarrow}^{*}_{F}t$. Thence $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. ∎
Lemma 9.86.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total. Let $W=ST\langle E\cup F\rangle$. Then for each $p\in W$,
$p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{E}}\mbox{ if and only if }p/_{\Theta\langle E\cup F,\emptyset\rangle}=p/_{\Theta\langle E,F\rangle}$.
Proof.
($\Rightarrow$) Let $p\in W$. By the definition of ${\Theta\langle E,F\rangle}$ and ${\Theta\langle E\cup F,\emptyset\rangle}$,
$p/_{\Theta\langle E\cup F,\emptyset\rangle}=p_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\cap T_{\Sigma}\times T_{\Sigma}$ and $p/_{\Theta\langle E,F\rangle}=p_{\mathop{\leftrightarrow}^{*}_{E}}\cap T_{\Sigma}\times T_{\Sigma}$.
Thus,
$p/_{\Theta\langle E\cup F,\emptyset\rangle}=p/_{\Theta\langle E,F\rangle}$.
($\Leftarrow$) Let $p\in W$. Since $\mathop{\leftrightarrow}^{*}_{E}\subseteq\mathop{\leftrightarrow}^{*}_{E\cup F}$, we have $p/_{\mathop{\leftrightarrow}^{*}_{E}}\subseteq p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. Our aim is to show that $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\subseteq p/_{\mathop{\leftrightarrow}^{*}_{E}}$. To this end, let $s\in T_{\Sigma}$ be such that $s\mathop{\leftrightarrow}^{*}_{E\cup F}p$. We now show that $s\mathop{\leftrightarrow}^{*}_{E}p$ as well. As $R\langle E,F\rangle$ is total, by Lemma 4.41,
$s{\downarrow}_{R\langle E,F\rangle}=t/_{\Theta\langle E,F\rangle}$ for some $t\in W$.
By Propositon 4.34, $s\mathop{\leftrightarrow}^{*}_{E}t$. Then
$t\mathop{\leftrightarrow}^{*}_{E}s\mathop{\leftrightarrow}^{*}_{E\cup F}p$.
Thus
$t\mathop{\leftrightarrow}^{*}_{E\cup F}p$.
Since $t,p\in W$, by the definition of $\Theta\langle E\cup F,\emptyset\rangle$, we have
$(t,p)\in{\Theta\langle E\cup F,\emptyset\rangle}$.
As $\mathop{\leftrightarrow}^{*}_{E}\subseteq\mathop{\leftrightarrow}^{*}_{E\cup F}$, we have
$t_{\Theta\langle E,F\rangle}\subseteq t_{\Theta\langle E\cup F,\emptyset\rangle}=p_{\Theta\langle E\cup F,\emptyset\rangle}=p/_{\Theta\langle E,F\rangle}$.
Consequently
$t/_{\Theta\langle E,F\rangle}\subseteq p_{\Theta\langle E,F\rangle}$.
Thence
$t/_{\Theta\langle E,F\rangle}=p_{\Theta\langle E,F\rangle}$.
Consequently, $s\mathop{\leftrightarrow}^{*}_{E}t\mathop{\leftrightarrow}^{*}_{E}p$. For this reason, by the definition of $s$,
$p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\subseteq p/_{\mathop{\leftrightarrow}^{*}_{E}}$.
Thus, $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{E}}$. ∎
Lemma 9.87.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total. Then
$\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if for each $p\in W$, $\left(p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{E}}\mbox{ or }p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{F}}\right)$.
Proof.
By Lemma 9.81, $R\langle E\cup F,\emptyset\rangle$ is total as well. Consequently, by Lemma 9.85 and Lemma 9.86, we have the lemma. ∎
10 GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$
We now introduce the notion of GTRS $R\langle E,F\rangle$ simulating the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant $a\in C\langle E\cup F,\emptyset\rangle$, then the concept of GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Then we describe the connection between the latter concept and the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
Definition 10.88.
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and let $a\in C\langle E\cup F,\emptyset\rangle$. We say that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$ if Conditions 1 and 2 hold:
  • 1.
    $a\in C\langle E,F\rangle$.
  • 2.
    For every $b\in C\langle E\cup F,\emptyset\rangle$, if $R\langle{E\cup F},\emptyset\rangle$ reach $a$ from a proper extension of the constant $b$, then
    for every rewrite rule of the form $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b\in R\langle{E\cup F},\emptyset\rangle$ with $m\in{\mathbb{N}}$,
[…]
    there exists $c\in C\langle E,F\rangle$ such that the rewrite rule $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c$ is in $R\langle E,F\rangle$.
Statement 10.89.
$c\subseteq b$ for $c,b$ appearing in Condition 2 in Definition 10.88.
Intuitively, in Condition 2 in Definition 10.88, the rewrite rule $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c$ in $R\langle E,F\rangle$ fits to the rewrite rule $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b\in R\langle{E\cup F},\emptyset\rangle$.
Statement 10.90.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total, and let $a\in C\langle E\cup F,\emptyset\rangle$. GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$ if and only if $a\in C\langle E,F\rangle$.
Proof.
By direct inspection of Definition 4.29 and Definition 10.88. ∎
Definition 10.91.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. We say that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if for each constant $a\in C\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on $a$.
Statement 10.92.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total. GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant $a\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
1. $a\in C\langle E,F\rangle$.
2. $a\in C\langle F,E\rangle$ and GTRS $R\langle F,E\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on $a$.
Proof.
We have the satement by direct inspection of Definition 10.91 and Statement 10.90. ∎
Lemma 10.93.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant $a\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
1. $a\in C\langle E,F\rangle$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
2. $a\in C\langle F,E\rangle$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
Proof.
By Definitions 10.88 and 10.91, and Lemma 6.70, and Definition 6.61, we have the lemma. ∎
Lemma 10.94.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that $R\langle E,F\rangle$ is total. GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant $a\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
1. $a\in C\langle E,F\rangle$.
2. $a\in C\langle F,E\rangle$ and for each constant $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
Proof.
By Statement 10.92, Definition 10.88, Lemma 6.70, and Definition 6.61 we have the lemma. ∎
Example 10.95.
We now continue Example 3.17, Example 5.50, and Example 7.71. There exist two paths of positive length in $AUX[E;F]$. There do not exist cycles in $AUX[E;F]$.
The length of the first path is $1$, and it starts from the vertex $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
$f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$   and   #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$.
The length of the second path is $1$, and it starts from the vertex $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
$f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$   and   \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
[…]
Example 10.96.
We now continue Example 3.18, Example 5.51, and Example 7.72. There does not exists a path of positive length in $AUX[E;F]$. There exists one cycle in $AUX[E;F]$. It is of length $1$, and it starts from the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
Observe that neither Condition 1 in nor Condition 2 in Lemma 10.93 holds for the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Thus by Lemma 10.93, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
Intuitively, assume that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a vertex $c\in C\langle E\cup F,\emptyset\rangle$ of $AUX[E;F]$. Consider a term $\delta[a_{1},\ldots,a_{k}]$ for some $\delta\in CON_{\Sigma,k}$, $a_{1},\ldots,a_{k}\in C\langle{E\cup F},\emptyset\rangle$ and a sequence of reductions
(i)   $\delta[a_{1},\ldots,a_{k}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}c$.
[…]
(ii)   $\delta[b_{1},\ldots,b_{k}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}d$
for some $d\in C\langle E,F\rangle$. We obtain (ii) following step by step the reductions in (i). For each reduction of (i), we apply a fitting rule at the same position in (ii). Then for each reduction step, for the two terms $s$ and $t$ obtained in (i) and (ii), respectively, each constant $b\in C\langle E,F\rangle$ appearing in $t$ corresponds to the constant $a\in C\langle E\cup F,\emptyset\rangle$ appearing at the same position in $s$. Moreover $b$, as congruence class, is subset of the corresponding constant $a$. Finally, constant $d$ corresponds to constant $c$, and $d\subseteq c$. Formally, we state and prove the following result.
Lemma 10.97.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant $c\in C\langle E\cup F,\emptyset\rangle$. Let Conditions (a) and (b) hold.
  • (a)
    $\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-2}[a_{(n-2)1},\ldots,a_{(n-2)k_{n-2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[a_{11},\ldots,a_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}c$,
[…]
  • (b)
[…]
$\delta_{n}[b_{n_{1}},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-2}[b_{(n-2)1},\ldots,b_{(n-2)k_{n-2}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\newline \delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}d$.
Proof.
We proceed by induction on $n$.
Base Case: $n=1$. Assume that Conditions (a) and (b) hold. Then $\delta_{1}=\sigma(\diamondsuit,\ldots,\diamondsuit)$ for some $\sigma\in\Sigma_{k_{1}}$, $k_{1}\in{\mathbb{N}}$, where
$\delta_{1}[a_{11},\ldots,a_{1k_{1}}]=\sigma(a_{11},\ldots,a_{1k_{1}})$,   $\sigma(a_{11},\ldots,a_{1k_{1}})\dot{\rightarrow}c\in R\langle{E\cup F},\emptyset\rangle$,   and   $\sigma(a_{11},\ldots,a_{1k_{1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}c$.
[…]
$\sigma(b_{11},\ldots,b_{1k_{1}})\dot{\rightarrow}b\in R\langle E,F\rangle$, where $b\subseteq c$.
Since $\delta_{1}[b_{11},\ldots,b_{1k_{1}}]=\sigma(b_{11},\ldots,b_{1k_{1}})$, we have
$\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}b$, where $b\in C\langle E,F\rangle$ and $b\subseteq c$.
Induction Step: Let $n\geq 1$, and assume that the statement of the claim is true for $n$. We now show that it is also true for $n+1$ as well. Assume that Conditions ($\mathrm{a}^{\prime}$) and ($\mathrm{b}^{\prime}$) hold:
  • ($\mathrm{a}^{\prime}$)
    $\delta_{n+1}[a_{(n+1)1},\ldots,a_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[a_{11},\ldots,a_{1k_{0}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}c$,
[…]
  • ($\mathrm{b}^{\prime}$)
[…]
Then Conditions (c)–(f) hold:
  • (c)
[…]
  • (d)
    $\delta_{n+1}[a_{(n+1)1},\ldots,a_{(n+1)k_{n+1}}]=\delta_{n}[a_{(n+1)1},\ldots,a_{(n+1)(j-1)},\sigma(a_{(n+1)j},\ldots,a_{(n+1)(j+m-1)}),a_{(n+1)(j+m)},\ldots,a_{(n+1)k_{n+1}}]$.
  • (e)
    For every $\ell\in\{1,\ldots,j-1\,\}$, $a_{n\ell}=a_{(n+1)\ell}$ and for every $\ell\in\{j+1,\ldots,k_{n}\,\}$, $a_{n\ell}=a_{(n+1)(\ell+m-1)}$.
  • (f)
    In the last reduction step of (a’) we apply the rule $\sigma(a_{(n+1)j},\ldots,a_{(n+1)(j+m-1)})\dot{\rightarrow}a_{nj}\in R\langle{E\cup F},\emptyset\rangle$.
By Condition ($\mathrm{a}^{\prime}$), $R\langle E\cup F,\emptyset\rangle$ reach $c$ from a proper extension of $a_{nj}$. By Condition ($\mathrm{b}^{\prime}$), for every $\ell\in\{\,j,\ldots,j+m-1\,\}$, $b_{(n+1)\ell}\in C\langle E,F\rangle$ and $b_{(n+1)\ell}\subseteq a_{n\ell}$. We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $c\in C\langle E\cup F,\emptyset\rangle$. Therefore by the rewrite rule appearing in Condition (f) there exists a rewrite rule
$\sigma(b_{(n+1)j},\ldots,b_{(n+1)(j+m-1)})\dot{\rightarrow}b$ in $R\langle E,F\rangle$ with $b\subseteq a_{nj}$.
For every $\ell\in\{1,\ldots,j-1\,\}$, let $b_{n\ell}=b_{(n+1)\ell}$,
let $b_{nj}=b$,
for every $\ell\in\{j+1,\ldots,k_{n}\,\}$, let $b_{n\ell}=b_{(n+1)(\ell+j-1)}$.
Thence
for every { $\ell\in\{\,1,\ldots,k_{n}\,\}$, $b_{n\ell}\subseteq a_{n\ell}$ and $\delta_{n+1}[b_{(n+1)1},\ldots,b_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]$.
By the sequence of reductions $(a^{\prime})$ we have
$\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{1}[a_{11},\ldots,a_{1k_{0}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}c$.
Consequently by the induction hypothesis,
[…]
there exists $d\in C\langle E,F\rangle$ with $d\subseteq c$
such that the following sequence of reductions holds:
$\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}d$.
Then we have
$\delta_{n+1}[b_{(n+1)1},\ldots,b_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\newline \cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}d$.∎
Lemma 10.98.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant $c\in C\langle E\cup F,\emptyset\rangle$. Let $s\in T_{\Sigma}$ and $s\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}c$. Then there exists $d\in C\langle E,F\rangle$ with $d\subseteq c$ such that $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}d$.
Proof.
Lemma 10.97 implies the lemma. ∎
Lemma 10.99.
Let $E$ and $F$ be GTESs over a signature $\Sigma$.
For each $t\in W$, $\left(t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{{\mathop{\leftrightarrow}^{*}_{E}}}\mbox{ or }t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{{\mathop{\leftrightarrow}^{*}_{F}}}\right)$
if and only if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
Proof.
()($\Rightarrow$)( ⇒ ) Let $t\in W$. Then by our assumption,
$t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}\mbox{ or }t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle F,E\rangle}$.
[…]
$t_{1},\ldots,t_{m}\in W$,
$t_{1}/_{\Theta\langle E,F\rangle}=b_{1},\ldots,t_{m}/_{\Theta\langle E,F\rangle}=b_{m}$,
$t_{1}/_{\Theta\langle{E\cup F},\emptyset\rangle}=a_{1},\ldots,t_{m}/_{\Theta\langle{E\cup F},\emptyset\rangle}=a_{m}$.
Hence $\sigma(t_{1}/_{\Theta\langle{E\cup F},\emptyset\rangle},\ldots,t_{m}/_{\Theta\langle{E\cup F},\emptyset\rangle})\dot{\rightarrow}c\in R\langle{E\cup F},\emptyset\rangle$.
Then by Proposition 4.32,
$\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\delta[\sigma(a_{1},\ldots,a_{m})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[c]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
By Proposition 4.34, $\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\leftrightarrow}^{*}_{E\cup F}t$. By the equation $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$, we have $\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\leftrightarrow}^{*}_{E}t$. By Proposition 4.36,
$\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t_{\Theta\langle E,F\rangle}$.
By Proposition 4.31, $R\langle E,F\rangle$ is convergent. Observe that $t_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$. Thence we have
$\delta[\sigma(t_{1},\ldots,t_{m})]{\downarrow}_{R\langle E,F\rangle}=t/_{\Theta\langle E,F\rangle}$.
[…]
$\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\delta[\sigma(t_{1}/_{\Theta\langle{E},F\rangle},\ldots,t_{m}/_{\Theta\langle{E},F\rangle})]$.
By Proposition 4.31, $R\langle E,F\rangle$ is convergent. Thence we have
$\delta[\sigma(t_{1}/_{\Theta\langle{E},F\rangle},\ldots,t_{m}/_{\Theta\langle{E},F\rangle})]{\downarrow}_{R\langle E,F\rangle}=t/_{\Theta\langle E,F\rangle}$.
Therefore, there exists a rewrite rule with left-hand side $\sigma(t_{1}/_{\Theta\langle{E},F\rangle},\ldots,t_{m}/_{\Theta\langle{E},F\rangle})$ in $R\langle{E},F\rangle$.
()($\Leftarrow$)( ⇐ ) Let $t\in W$. We assumed that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. By Definition 10.91 GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\in C\langle E\cup F,\emptyset\rangle$. Assume that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. Then $t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}$. Let $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. Consequently by Proposition 4.36, $s\mathop{\rightarrow}_{R\langle E\cup F,E\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. By Lemma 10.98, there exists $d\in C\langle E,F\rangle$ such that $d\subseteq t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}d$. Then $d=p/_{\Theta\langle E,F\rangle}$ for some $p\in W$. By Proposition 4.34, $s\mathop{\leftrightarrow}^{*}_{E\cup F}t$ and $s\mathop{\leftrightarrow}^{*}_{E}p$. Then $s\mathop{\leftrightarrow}^{*}_{E\cup F}p$. Thus $d=p/_{\Theta\langle E,F\rangle}\subseteq t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}$. Consequently $d=t/_{\Theta\langle E,F\rangle}$. Hence $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. Consequently, for any $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$, we have $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. Thus $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\subseteq t/_{\mathop{\leftrightarrow}^{*}_{E}}$, and hence $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$. We showed that if GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$, then $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$.
We can show similarly that if GTRS $R\langle F,E\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$, then $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}$. Consequently,
for each $t\in W$, $\left(t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}\mbox{ or }t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}\right)$. ∎
Example 10.100.
We continue Example 3.21 and its sequels, Examples 5.54, 7.73. There does not exist a path of positive length in $AUX[E;F]$. There exists one cycle in $AUX[E;F]$. The cycle is of length $1$, it starts from the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$ and #$\#/_{\Theta\langle F,E\rangle}.keeps_{F}=\mathrm{false}$. Then the formula
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$ or #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$
does not hold. Thence the following statement does not hold: for each $a\in C\langle E\cup F,\emptyset\rangle$,
  • $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$ and
  • if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
By Lemma 10.94, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
Example 10.101.
We continue Example 3.22, and its sequels, Examples 5.55, and 7.74. There exist one path of positive length and two cycles in $AUX[E;F]$.
The first cycle: Its length is $1$, and it starts from the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
The path of positive length: Its length is $1$, and it starts from the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
The second cycle: Its length is $1$, and it starts from the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
Observe that the following statement holds: for each $a\in C\langle E\cup F,\emptyset\rangle$,
  • $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$ and
  • if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
By Lemma 10.94, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
11 Main Cases of the Decision Algorithm
We now define the four main cases of our decision algorithm. We show that we can decide which Main Cases hold in $O(n\,\mathrm{log}\,n)$ time.
  • Main Case 1:
    $\Sigma$ is a unary signature. Example 3.17 and its sequels, Examples 5.50, 7.71, 10.95. and 12.109, and 12.114, also Example 3.18 and its sequels, 5.51, 7.72, 10.96, 12.110, and 12.115 illustrate this case.
  • Main Case 2:
    Both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total. Example 3.19 and its sequels Examples 5.52, 13.117, further Example 3.20 and its sequels, Examples 5.53, 13.118, exemplify this case.
  • Main Case 3:
    $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total. Example 3.21 and its sequels, Examples 5.54, 7.73, 10.100, 14.125, and 14.132, moreover Example 3.22, and its sequels, Examples 5.55, 7.74, 10.101, 14.126, and 14.133 present this case.
  • Main Case 4:
    $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total. Example 3.23 and its sequel Example 5.56, furthermore, Example 3.24 and its sequels Example 5.57 and Example 15.134 demonstrate this case.
Main Case 2 overlaps with Main Case 1 and Main Case 3, and is a subcase of the union of Main Case 1 and Main Case 3.
Lemma 11.102.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. We can decide which Main Cases hold in $O(n\,\mathrm{log}\,n)$ time.
Proof.
We run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.47, we get the following output in $O(n\,\mathrm{log}\,n)$ time:
  • the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$;
  • the set of nullary symbols $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$;
  • $INC(E\cup F,E)$ stored in a red-black tree saved in a variable $RBT1$;
  • $INC(E\cup F,F)$ stored in a red-black tree saved in a variable $RBT2$;
  • $NUM(E\cup F,E)$ stored in a red-black tree saved in a variable $RBT3$;
  • $NUM(E\cup F,F)$ stored in a red-black tree saved in a variable $RBT4$;
  • the GTRS $R\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT5$; each node in $RBT5$ contains a search key which is a rule in $R\langle E\cup F,\emptyset\rangle$, and satellite data $counter_{E}$ and $counter_{F}$;
  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT6$; for every ordered pair $(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
  • the GTRS $R\langle E,F\rangle$ stored in a red-black tree saved in a variable $RBT7$;
  • the GTRS $R\langle F,E\rangle$ stored in a red-black tree saved in a variable $RBT8$;
  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
By Lemma 4.49, we decide in $O(n)$ time whether $R\langle E,F\rangle$ is total, and whether $R\langle F,E\rangle$ is total. Then we decide in a total of $O(n\,\mathrm{log}\,n)$ time which Main Cases hold. ∎
We study Main Cases 1, 2, 3, and 4 one by one in Sections 12, 13, 14, and 15, respectively. First we consider Main Case 1.
12 Main Case 1
In this main case, $\Sigma$ is a unary signature, and $E$ and $F$ are GTESs over $\Sigma$.
Theorem 12.103.
Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. We can decide in $O(n^{3})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Proof.
Lemma 9.83, Proposition 9.84, and Proposition 9.82 imply the Theorem. ∎
We present a faster and simpler algorithm for solving our decidability problem. We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.47. Example 3.17 and its sequel, Example 5.50, furthermore Example 3.18 and its sequel, 5.51 illustrate this main case. We assume that we have run Algorithm CAD on $\Sigma$, $E$, and $F$, and study its output, see Proposition 4.47.
Lemma 12.104.
Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. Let $a\in C\langle E\cup F,\emptyset\rangle$ be such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle F$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$. Assume that Conditions (a) and (b) hold:
  • (a)
    $\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[a_{11},\ldots,a_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a$,
[…]
  • (b)
[…]
$\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}b$,
Proof.
We proceed by induction on $n$.
Base Case: $n=1$. Then $\delta_{1}=\sigma(\diamondsuit,\ldots,\diamondsuit)$ for some $\sigma\in\Sigma_{k_{1}}$, $\delta_{1}[a_{11},\ldots,{k_{1}}]=\sigma(a_{11},\ldots,{k_{1}})$, and $\sigma(a_{11},\ldots,a_{1k_{1}})\dot{\rightarrow}a\in R\langle{E\cup F},\emptyset\rangle$. We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$. Hence by Condition (b), and the rewrite rule above there exists a rewrite rule
$\sigma(b_{11},\ldots,b_{1k_{1}})\dot{\rightarrow}b\in R\langle E,F\rangle$
for some $b\in C\langle E,F\rangle$ with $b\subseteq a$. Then we have $\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}b$.
Induction Step: Let $n\geq 1$, and assume that the statement of the lemma is true for $n$. We now show that it is also true for $n+1$ as well. Assume that Conditions $(\mathrm{a}^{\prime})$ and $(\mathrm{b}^{\prime})$ hold:
  • ($\mathrm{a}^{\prime}$)
    $\delta_{n+1}[a_{(n+1)1},\ldots,a_{(n+1)k_{(n+1)}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[a_{11},\ldots,a_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a$,
[…]
  • ($\mathrm{b}^{\prime}$)
[…]
Then we obtain $\delta_{n+1}$ from $\delta_{n}$ replacing the $\ell$th occurrence from left-to right of $\diamondsuit$ with the term $\sigma(\diamondsuit,\ldots,\diamondsuit)$ for some $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$. Consequently,
$k_{n+1}-m+1=k_{n}$,
$\delta_{n+1}[a_{(n+1)1},\ldots,a_{(n+1)k_{(n+1)}}]=\delta_{n}[a_{(n+1)1},\ldots,a_{(n+1)(\ell-1)},\sigma(a_{(n+1)\ell},\ldots,a_{(n+1)(\ell+m)}),a_{(n+1)(\ell+1)},\ldots a_{(n+1)k_{(n+1)}}]$,
$\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]=\delta_{n}[a_{(n+1)1},\ldots,a_{(n+1)(\ell-1)},a_{n,\ell},a_{(n+1)(\ell+1)},\ldots a_{(n+1)k_{(n+1)}}]$,
$a_{n1}=a_{(n+1)1},\ldots,a_{n(\ell-1)}=a_{(n+1)(\ell-1)},a_{n\ell}=a_{n\ell}a_{n(\ell+1)}=a_{(n+1)(\ell-m+1)},\ldots,a_{nk_{n}}=a_{(n+1)k_{(n+1)}}$
and
$\sigma(a_{(n+1)\ell},\ldots,a_{(n+1)(\ell+m)})\dot{\rightarrow}a_{n\ell}\in R\langle{E\cup F},\emptyset\rangle.$ (1)
We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$. By Condition ($\mathrm{a}^{\prime}$), $R\langle E\cup F,\emptyset\rangle$ reach $a$ from a proper extension of $a_{n\ell}$. Therefore by the rewrite rule (1) and Condition ($\mathrm{b}^{\prime}$), there exists a rewrite rule
$\sigma(b_{(n+1)\ell},\ldots,b_{(n+1)(\ell+m)})\dot{\rightarrow}b_{n\ell}\in R\langle{E\cup F},\emptyset\rangle.$ (2)
for some $b_{n\ell}\in C\langle E\cup F,E\rangle$ with $b_{n\ell}\subseteq a_{n\ell}$. Let
$b_{n1}=b_{(n+1)1},\ldots,b_{n(l-1)}=b_{(n+1)(l-1)},\ldots,b_{n(l+1)}=b_{(n+1)(l-m+1)},\ldots,b_{nk_{n}}=b_{(n+1)k_{(n+1)}}$
By the induction hypothesis, there exists $b\in C\langle E,F\rangle$ with $b\subseteq a$ such that the following sequence of reductions holds:
$\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}b$,
In this way
$\delta_{n+1}[b_{(n+1)1},\ldots,b_{(n+1)k_{(}n+1)}]=\newline \delta_{n}[b_{(n+1)1},\ldots,b_{(n+1)(\ell-1)},\sigma(b_{(n+1)\ell},\ldots,b_{(n+1)(\ell+m)}),b_{(n+1)(\ell+1)},\ldots b_{(n+1)k_{(n+1)}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{n}[b_{(n+1)1},\ldots,b_{(n+1)(\ell-1)},b_{n,\ell},b_{(n+1)(\ell+1)},\ldots b_{(n+1)k_{(n+1)}}]=\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}b$, ∎
Lemma 12.105.
Let $E$ and $F$ be GTESs over a unary signature $\Sigma$.
For each $t\in W$, $\left(t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}\mbox{ or }t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}\right)$
if and only if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
Proof.
()($\Rightarrow$)( ⇒ ) Let $t\in W$. Then by our assumption, $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$ or $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}$. Assume that $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$. The case $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}$ is similar. Let $b\in C\langle E\cup F,\emptyset\rangle$ be such that $R\langle{E\cup F},\emptyset\rangle$ reach $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ from a proper extension of the constant $b$. Then there exists $\delta\in C_{\Sigma,1}$ such that $\delta[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. Consider any rewrite rule $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b$ in $R\langle E\cup F,\emptyset\rangle$. with right-hand side $b$. We now distinguish two cases depending on the value of $m$.
Case 1: $m=1$. Then we consider the rewrite rule $\sigma(b_{1})\dot{\rightarrow}b\in R\langle{E\cup F},\emptyset\rangle$. Then let $c_{1}\in C\langle E,F\rangle$ with $c_{1}\subseteq b_{1}$, and let $t_{1}\in T_{\Sigma}$ such that $t_{1}\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}c_{1}$. Then we have
$\delta[\sigma(t_{1})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\delta[\sigma(b_{1})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[b]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle E\cup F,\emptyset\rangle}$.
By Proposition 4.34, $\delta[\sigma(t_{1})]\mathop{\leftrightarrow}^{*}_{E\cup F}t$. As a consequence, by our assumption that $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$, we have $\delta[\sigma(t_{1})]\mathop{\leftrightarrow}^{*}_{E}t$. Consequently, by Proposition 4.36,
$\delta[\sigma(t_{1})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
Hence
$t_{1}\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}c_{1}$,
$\sigma(t_{1})\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\sigma(c_{1})\mathop{\rightarrow}_{R\langle E,F\rangle}c$ for some $c\in C\langle E,F\rangle$, where the rewrite rule $\sigma(c_{1})\dot{\rightarrow}c$ is in $R\langle E,F\rangle$, and
$\delta[\sigma(t_{1})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\delta[\sigma(c_{1})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\delta[c]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
Case 2: $m=0$. Then we consider the rewrite rule $\sigma\dot{\rightarrow}b\in R\langle{E\cup F},\emptyset\rangle$. Then we have
$\delta[\sigma]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[b]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle E\cup F,\emptyset\rangle}$.
By Proposition 4.34, $\delta[\sigma]\mathop{\leftrightarrow}^{*}_{E\cup F}t$. As a consequence, by our assumption that $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$, we have $\delta[\sigma]\mathop{\leftrightarrow}^{*}_{E}t$. Consequently, by Proposition 4.36,
$\delta[\sigma]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
Therefore
$\sigma\mathop{\rightarrow}_{R\langle E,F\rangle}c$ for some $c\in C\langle E,F\rangle$, where the rewrite rule $\sigma\dot{\rightarrow}c$ is in $R\langle E,F\rangle$, and
$\delta[\sigma]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta[c]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
()($\Leftarrow$)( ⇐ ) Let $t\in W$ and $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. Then $s\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. We assumed that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Consequently, $R\langle E,F\rangle$ or $R\langle F,E\rangle$ simulate $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. Assume that $R\langle E,F\rangle$ simulates $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. The case $R\langle F,E\rangle$ simulates $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ is similar. By Lemma 12.104, there exists $d\in C\langle E,F\rangle$ with $d\subseteq t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ such that $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}d$. By Proposition 4.38, $d=t/_{\Theta\langle E,F\rangle}$. For this reason $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. Consequently, for any $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$, we have $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. As a consequence $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\subseteq t/_{\mathop{\leftrightarrow}^{*}_{E}}$, and hence $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$. ∎
12.1 Nested Partial Depth First Search on the Auxiliary dpwpa $AUX[E;F]$ for the GTESs $E$ and $F$
Algorithm Nested Partial Depth First Search (NPDFS for short) on the Auxiliary dpwpa $AUX[E;F]$ for the GTESs $E$ and $F$ implements a nested loop. Algorithm NPDFS calls the function NPDFS-VISIT, and in this way it implements the outer loop, which is a prefix of a depth first search on $AUX[E;F]$, and can terminate before visiting all vertices of $AUX[E;F]$ in case that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
  • Algorithm NPDFS calls the functions INNER-LOOP-E, INNER-LOOP-F, and NPDFS-VISIT.
  • Function NPDFS-VISIT calls the function INNER-LOOP-E, INNER-LOOP-F, and NPDFS-VISIT.
In this way, algorithm NPDFS decides the following formula: For each $a\in C\langle E\cup F,\emptyset\rangle$,
  • $a.equal_{E}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{E}=\mathrm{true}$, or
  • $a.equal_{F}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
Function INNER-LOOP-E is an inner loop inside the body of the outer loop, it also implements a prefix of a depth first search on those vertices $b$ which are positive-step reachable from the current node $a$ of the outer loop, it can terminate before visiting all vertices of $AUX[E;F]$ which are positive-step reachable from $a$ in case that it visits a vertex $b$ such that $b.keeps_{E}=\mathrm{false}$. It checks whether for each vertex $b$ which are positive-step reachable from the current node $a$ of the outer loop, $b.keeps_{E}=\mathrm{true}$. INNER-LOOP-E decides the following formula:
For each vertex $b\in C\langle E\cup F,\emptyset\rangle$ which is positive-step reachable from the current node $a$ of the outer loop, $b.keeps_{E}=\mathrm{true}$.
Function INNER-LOOP-F is an inner loop inside the body of the outer loop, it also implements a prefix of a depth first search on those vertices $b$ which are positive-step reachable from the current node $a$ of the outer loop, it can terminate before visiting all vertices of $AUX[E;F]$ which are positive-step reachable from $a$ in case that in case that it visits a vertex $b$ such that $b.keeps_{F}=\mathrm{false}$. It checks whether for each vertex $b$ which are positive-step reachable from the current node $a$ of the outer loop, $b.keeps_{F}=\mathrm{true}$. INNER-LOOP-F decides the following formula:
For each vertex $b\in C\langle E\cup F,\emptyset\rangle$ which is positive-step reachable from the current node $a$ of the outer loop, $b.keeps_{F}=\mathrm{true}$.
Definition 12.106.
Let $\Sigma$ be a signature, $E$ and $F$ be GTESs over $\Sigma$.
Algorithm Nested Partial Depth First Search on Auxiliary dpwpa $AUX[E;F]$ (NPDFS for short)
Input: the auxiliary dpwpa $AUX[E;F]$.
Output: Boolean.
$\mathrm{NPDFS}(AUX[E;F])=\mathrm{true}$ if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$, otherwise $\mathrm{NPDFS}(AUX[E;F])=\mathrm{false}$. }
var $a:C\langle E\cup F,\emptyset\rangle$;
for each vertex $a\in C\langle E\cup F,\emptyset\rangle$ do
begin
$a.visited\_outer\_loop:=\mathrm{false}$
$a.visited\_inner\_loop_{E}:=\mathrm{false}$
$a.visited\_inner\_loop_{F}:=\mathrm{false}$
end
for each vertex $a\in C\langle E\cup F,\emptyset\rangle$ do
if not $a.visited\_outer\_loop$
then begin
$a.visited\_outer\_loop:=\mathrm{true}$
${\bf if}$ not ($a.equal_{E}$ or $a.equal_{F}$)
then return $\mathrm{false}$
${\bf for}$ ${\bf each}$ vertex $b\in Adj(a)$ do
${\bf if}$ not (($a.equal_{E}$ and $\mathrm{INNER-LOOP-E}(AUX[E;F],b)$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],b)$
or ($a.equal_{F}$ and $\mathrm{INNER-LOOP-F}(AUX[E;F],b)$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],b))$
then return $\mathrm{false}$
end
output $(\mathrm{true})$
function $\mathrm{NPDFS-VISIT}(AUX[E;F],a):\;\mathrm{Boolean}$
if not $a.visited\_outer\_loop$
then begin
$a.visited\_outer\_loop:=\mathrm{true}$
${\bf if}$ not ($a.equal_{E}$ or $a.equal_{F}$)
then return $\mathrm{false}$
${\bf for}$ ${\bf each}$ vertex $b\in Adj(a)$ do
${\bf if}$ not (($a.equal_{E}$ and $\mathrm{INNER-LOOP-E}(AUX[E;F],b)$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],b)$
or ($a.equal_{F}$ and $\mathrm{INNER-LOOP-F}(AUX[E;F],b)$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],b))$
then return $\mathrm{false}$
end
output $(\mathrm{true})$
function $\mathrm{INNER-LOOP-E}(AUX[E;F],a):\;\mathrm{Boolean}$
$a.visited\_inner\_loop_{E}:=\mathrm{true}$;
if $a.keeps_{E}$
${\bf then}$ for each vertex $b\in Adj(a)$ do
[…]
${\bf then}\;{\bf begin}$
$a.visited\_inner\_loop_{E}:=\mathrm{false}$;
${\bf return}\;\mathrm{false}$
end
${\bf else}\;{\bf begin}$
$a.visited\_inner\_loop_{E}:=\mathrm{false}$;
${\bf return}\;\mathrm{false}$
end
$a.visited\_inner\_loop_{E}:=\mathrm{false}$;
${\bf output}(\mathrm{true})$
function $\mathrm{INNER-LOOP-F}(AUX[E;F],a):\;\mathrm{Boolean}$
$a.visited\_inner\_loop_{F}:=\mathrm{true}$;
if $a.keeps_{F}$
${\bf then}$ for each vertex $b\in Adj(a)$ do
[…]
${\bf then}\;{\bf begin}$
$a.visited\_inner\_loop_{F}:=\mathrm{false}$;
${\bf return}\;\mathrm{false}$
end
${\bf else}\;{\bf begin}$
$a.visited\_inner\_loop_{F}:=\mathrm{false}$;
${\bf return}\;\mathrm{false}$
end
$a.visited\_inner\_loop_{F}:=\mathrm{false}$;
${\bf output}(\mathrm{true})$
Lemma 12.107.
Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. NPDFS outputs true on $AUX[E;F]$ if and only if for each $a\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
  • 1.
    $a.equal_{E}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{E}=\mathrm{true}$.
  • 2.
    $a.equal_{F}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
Proof.
By direct inspection of Definition 12.106. ∎
Lemma 12.108.
Let $E$ and $F$ be GTESs over a unary signature $\Sigma$.
If GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$, then Conditions 1–3 hold:
  • 1.
    Algorithm NPDFS on $AUX[E;F]$ in the outer loop visits once each vertex $a\in C\langle E\cup F,\emptyset\rangle$.
  • 2.
    For each vertex $a\in C\langle E\cup F,\emptyset\rangle$ visited in the outer loop,
    • (a)
      $a.equal_{E}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $b$ to $a$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{E}=\mathrm{true}$, or
    • (b)
      $a.equal_{F}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $b$ to $a$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
  • 3.
    Algorithm NPDFS on $AUX[E;F]$ outputs true.
If GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$, then Conditions 4–6 hold:
  • 4.
    Algorithm NPDFS on $AUX[E;F]$ visits at most once each vertex $a\in C\langle E\cup F,\emptyset\rangle$ in the outer loop.
  • 5.
    There exists a vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that Algorithm NPDFS on $AUX[E;F]$ visits $a\in C\langle E\cup F,\emptyset\rangle$ in the outer loop, and Conditions 2.(a) and 2.(b) are false for $a$.
  • 6.
    Algorithm NPDFS on $AUX[E;F]$ outputs false.
Proof.
By direct inspection of Definition 12.106. ∎
Example 12.109.
We now continue Example 3.17, Example 5.50, Example 7.71, and Example 10.95. We run Algorithm NPDFS on $AUX[E;F]$. Then Algorithm NPDFS outputs true. By the last line of Example 10.95, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Consequently Algorithm NPDFS outputs the answer to the question whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
Example 12.110.
We now continue Examples 3.18, 5.51, 7.72, and 10.96. We run Algorithm NPDFS on $AUX[E;F]$. Then Algorithm NPDFS outputs false. By the last lines of Example 10.96, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Thus Algorithm NPDFS outputs the answer to the question whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
Lemma 12.111.
[8] Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. Algorithm NPDFS runs on $AUX[E;F]$ in $O(n^{2})$ time.
Proof.
Recall that $n=size(E)+size(F)$. By Statement 4.27, $|C\langle E\cup F,\emptyset\rangle|\leq n$. By Statement 6.63, $|A[E;F]|\leq n$. The running time of depth first search on $AUX[E;F]$ is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$ [8, Section 22.3]. Consequently, the running time of depth first search on $AUX[E;F]$ is $O(n)$.
The outer loop of Algorithm NPDFS on $AUX[E;F]$ implements a prefix of the depth first search on $AUX[E;F]$ [8], it can terminate before visiting all vertices of $AUX[E;F]$. Consequently, the outer loop on $AUX[E;F]$ takes $O(n)$ time.
Function INNER-LOOP-E on $AUX[E;F]$ and any vertex $a$ implements a prefix of a depth first search on $AUX[E;F]$ starting from $a$. Accordingly the running time of function INNER-LOOP-E is $O(n)$. Similarly, the running time of function INNER-LOOP-F is $O(n)$. By Statement 4.27, $|C\langle E\cup F,\emptyset\rangle|\leq n$, therefore INNER-LOOP-E and INNER-LOOP-F run at most $n$ times. Accordingly, Algorithm NPDFS runs on $AUX[E;F]$ in $O(n^{2})$ time. ∎
12.2 Decision Algorithm and its verification
We present our decision algorithm and show its correctnes.
Lemma 12.112.
Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. Then the following conditions are pairwise equivalent:
  • (i)
    $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$,
  • (ii)
    For each $t\in W$, $(t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}\mbox{ or }t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}})$.
  • (iii)
    GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle)$.
  • (iv)
    For each $a\in C\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a$.
  • (v)
    For each $a\in C\langle E\cup F,\emptyset\rangle$,
    • $a.equal_{E}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{E}=\mathrm{true}$, or
    • $a.equal_{F}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
  • (vi)
    $NPDFS(AUX[E;F])=\mathrm{true}$.
Proof.
$\mathrm{(i)}\Leftrightarrow\mathrm{(ii)}$: By Lemma 9.83.
$\mathrm{(ii)}\Leftrightarrow\mathrm{(iii)}$: By Lemma 12.105.
$\mathrm{(iii)}\Leftrightarrow\mathrm{(iv)}$: By Definition 10.91.
$\mathrm{(iv)}\Leftrightarrow\mathrm{(v)}$: By Definition 10.88 and Definition 6.58 and Lemma 8.76 and Lemma 6.70.
$\mathrm{(v)}\Leftrightarrow\mathrm{(vi)}$: By Lemma 12.107. ∎
Lemma 12.113.
Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. Then we can decide in $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Proof.
Recall that we have run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.47, we have obtained in $O(n\,\mathrm{log}\,n)$ time the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
3. We run Algorithm NPDFS on $AUX[E;F]$ in $O(n^{2})$ time by Lemma 12.111. Then by Lemma 12.112, NPDFS outputs true if and only if $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Summing up the time complexities of the three steps above, we get that the time complexity of our entire computation is $O(n^{2})$. ∎
Example 12.114.
We now continue Examples 3.17, 5.50, 7.71, 10.95. and 12.109. Recall that we run Algorithm NPDFS on the auxiliary dpwpa $AUX[E;F]$, and Algorithm NPDFS outputs true to the question whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Hence Condition (iii) in Lemma 12.112 holds. Consequently, by Lemma 12.112, $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Example 12.115.
We now continue Examples 3.18, 5.51, 7.72, 10.96, and 12.110. Recall that we run Algorithm NPDFS on the auxiliary dpwpa $AUX[E;F]$, and Algorithm NPDFS outputs false to the question whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. For this reason Condition (iii) in Lemma 12.112 does not hold. Consequently, by Lemma 12.112, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
13 Main Case 2
In this main case, $E$ and $F$ are GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total.
Theorem 13.116.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total. We can decide in $O(n^{3})$ time whether
$\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Proof.
Lemma 9.87 and Proposition 9.84, and Proposition 9.82 imply the Theorem. ∎
We present a faster and simpler algorithm for solving our decidability problem. We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.47. We showed that $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class. By direct inspection of $INC(E\cup F,E)$ stored in a red-black tree, and $INC(E\cup F,F)$ stored in a red-black tree we decide in $O(n)$ time whether each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class. Thence we decide in $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Example 3.19 and its sequels, Examples 5.52, 13.117, in addition Example 3.20 and its sequels, Examples 5.53, 13.118. exemplify this case.
Example 13.117.
We now continue Examples 3.19 and 5.52. Recall that
  • both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total,
  • #$\#/_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • \$$\$/_{\Theta\langle E,F\rangle}=\$/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • £$\mathsterling/_{\Theta\langle E,F\rangle}=\mathsterling/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • #$\#/_{\Theta\langle F,E\rangle}=\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • \$$\$/_{\Theta\langle F,E\rangle}\subset\$/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
  • £$\mathsterling/_{\Theta\langle F,E\rangle}=\mathsterling/_{\Theta\langle{E\cup F},\emptyset\rangle}$, and
  • $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Example 13.118.
We now continue Examples 3.20 and 5.53. Recall that
  • both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total,
  • the only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is $W$,
  • #$\#/_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and #$\#/_{\Theta\langle{F},E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$, and
  • $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$.
Lemma 13.119.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total. Then
for each $t\in W$, $\left(t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}\mbox{ or }t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}\right)$ if and only if
for each $t\in W$, $\left(t/_{\Theta\langle{E\cup F},\emptyset\rangle}=t/_{\Theta\langle{E,F\rangle}}\mbox{ or }t/_{\Theta\langle{E\cup F},\emptyset\rangle}=t/_{\Theta\langle{F},E\rangle}\right)$.
Proof.
($\Rightarrow$) If $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$, then by Definition 2.9, $t/_{\Theta\langle{E\cup F},\emptyset\rangle}=t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\cap W\times W=t/_{\mathop{\leftrightarrow}^{*}_{E}}\cap W\times W=t/_{\Theta\langle{E,F\rangle}}$.
If $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}$, then $t/_{\Theta\langle{E\cup F},\emptyset\rangle}=t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\cap W\times W=t/_{\mathop{\leftrightarrow}^{*}_{F}}\cap W\times W=t/_{\Theta\langle{E,F\rangle}}$.
($\Leftarrow$) Let $t\in W$ and $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. By Proposition 4.36,
$s\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle E\cup F,\emptyset\rangle}$.
As $R\langle E,F\rangle$ and $R\langle F,E\rangle$ are total, by Lemma 4.41,
$s{\downarrow}_{R\langle E,F\rangle}=p/_{\Theta\langle E,F\rangle}$ for some $p\in W$,
and
$s{\downarrow}_{R\langle F,E\rangle}=q/_{\Theta\langle F,E\rangle}$ for some $q\in W$.
Consequently, by Proposition 4.34,
$s\mathop{\leftrightarrow}^{*}_{E\cup F}t$, $s\mathop{\leftrightarrow}^{*}_{E}p$ and $s\mathop{\leftrightarrow}^{*}_{F}q$.
Therefore $t\mathop{\leftrightarrow}^{*}_{E\cup F}s\mathop{\leftrightarrow}^{*}_{E}p$ and $t\mathop{\leftrightarrow}^{*}_{E\cup F}s\mathop{\leftrightarrow}^{*}_{E}q$. Therefore,
$p\in t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and $q\in t/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
By our assumption,
$t/_{\Theta\langle{E\cup F},\emptyset\rangle}=t/_{\Theta\langle{E,F\rangle}}$ or $t/_{\Theta\langle{E\cup F},\emptyset\rangle}=t/_{\Theta\langle{F},E\rangle})$.
First, suppose that $t/_{\Theta\langle{E\cup F},\emptyset\rangle}=t/_{\Theta\langle{E,F\rangle}}$. Therefore, we have $p\in t/_{\Theta\langle{E,F\rangle}}$. Thus, by $s\mathop{\leftrightarrow}^{*}_{E}p$, we have $s\mathop{\leftrightarrow}^{*}_{E}p\mathop{\leftrightarrow}^{*}_{E}t$. Accordingly, $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. By the definition of $s$,
$t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\subseteq t/_{\mathop{\leftrightarrow}^{*}_{E}}$.
Since $E\subseteq E\cup F$, $t/_{\mathop{\leftrightarrow}^{*}_{E}}\subseteq t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. Thus, $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$.
Second, suppose that $t/_{\Theta\langle{E\cup F},\emptyset\rangle}=t/_{\Theta\langle{F,E\rangle}}$. Then by a similar argument as above we can show that $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}$. ∎
Lemma 13.120.
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRSs $R\langle E,F\rangle$ and $R\langle F,E\rangle$ are total. Then we can decide in $O(n\,\mathrm{log}\,n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Proof.
By Lemma 9.81, $R\langle E\cup F,\emptyset\rangle$ is total as well. Consequently, by Lemma 9.85,
$\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if for each $p\in W$, $\left(p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{E}}\mbox{ or }p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{F}}\right)$.
Consequently, by Lemma 13.119,
$\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if for each $p\in W$, $\left(p/_{\Theta\langle{E\cup F},\emptyset\rangle}=p_{\Theta\langle{E,F\rangle}}\mbox{ or }p/_{\Theta\langle{E\cup F},\emptyset\rangle}=p/_{\Theta\langle{F},E\rangle}\right)$.
Justified by the above observations and by Statement 6.60, we compute as follows. Recall that $NUM(E\cup F,E)$ stored in a red-black tree $RBT3$, and $NUM(E\cup F,F)$ stored in a red-black tree $RBT4$. Alternately in a coordinated way, we traverse the red-black trees $RBT3$ and $RBT4$ in in-order traversal, For each $a\in C\langle E\cup F,\emptyset\rangle$, we read the pair $(a,k)$, where $k=|\{b\in C\langle E,F\rangle\mid b\subseteq a\,\}|$, and the pair $(a,\ell)$, where =|$\ell=|\{b\in C\langle F,E\rangle\mid b\subseteq a\,\}|$. In this way, we decide in $O(n)$ time whether each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class, see Definition 4.26. If the answer is ’yes’, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, otherwise, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. ∎
14 Main Case 3
In this main case, $E$ and $F$ are GTESs over a signature $\Sigma$ such that $\Sigma$ has a symbol of arity at least $2$, and $R\langle E,F\rangle$ or $R\langle F,E\rangle$ are total.
Theorem 14.121.
Let $\Sigma$ be a signature having a symbol of arity at least $2$, and let $E$ and $F$ be GTESs over $\Sigma$ such that GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ are total. We can decide in $O(n^{3})$ time whether
$\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Proof.
By Lemma 9.81, $R\langle E\cup F,\emptyset\rangle$ is total as well. Then by Lemma 9.85 and Proposition 9.84, and Proposition 9.82 we have the result. ∎
We present a faster and simpler decision algorithm. Without loss of generality, we may assume that $R\langle E,F\rangle$ is total from now on throughout this section. Example 3.21 and its sequels, Examples 5.54, 7.73, 10.100, 14.125, and 14.132 plus Example 3.22, and its sequels, Examples 5.55, 7.74, 10.101, 14.126, and 14.133 present this case. We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.47.
14.1 Partial Depth First Search on the Auxiliary dpwpa
We now define a partial depth first search on the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, which is a restricted and slightly modified verison of the depth first search on graphs [8]. We start the partial depth first search from the elements of $C(E\cup F,\emptyset)\setminus C(E,F)$.
Definition 14.122.
Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and let $E$ and $F$ be GTESs over $\Sigma$.
Algorithm Partial Depth First Search on Auxiliary dpwpa $AUX[E;F]$ (PDFS for short)
Input: the auxiliary dpwpa $AUX[E;F]$. Output: Boolean.
{ Output is true if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$, false otherwise. }
var $a:C\langle E\cup F,\emptyset\rangle$;
$\mathrm{PDFS}(G\langle E,F\rangle):\;\mathrm{Boolean}$
for each vertex $a\in C\langle E\cup F,\emptyset\rangle$ do
begin
if not $a.equal_{E}$ and not $a.equal_{F}$
[…]
{  if $a\not\in C\langle E,F\rangle$ and $a\not\in C\langle F,E\rangle$ then GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. }
$a.visited:=\mathrm{false}$
end
for each vertex $a\in C\langle E\cup F,\emptyset\rangle$ do
if not $a.equal_{E}$ and not $a.visited$
{ if $a\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$ and $a$ has not been visited yet }
${\bf then}$ ${\bf for}$ ${\bf each}$ vertex $b\in Adj(a)$ do
${\bf if}$ not $b.visited$
${\bf then}$ if not $\mathrm{PDFS-VISIT}(AUX[E;F],b)$ then
${\bf return}\;\mathrm{false}$
${\bf output}(\mathrm{true})$
${\bf function}\;\mathrm{PDFS-VISIT}(AUX[E;F],a):\;\mathrm{Boolean}$
$a.visited:=\mathrm{true}$;
if $a.keeps_{F}=\mathrm{true}$
${\bf then}$ begin
for each vertex $b\in Adj(a)$ do
${\bf if}$ not $b.visited$
${\bf then}$ ${\bf if}$ ${\bf not}$ $\mathrm{PDFS-VISIT}(AUX[E;F],b)$
${\bf then}$ ${\bf return}$ $\mathrm{false}$
end
${\bf else}\;{\bf return}\;\mathrm{false}$
${\bf output}(\mathrm{true})$
When Algorithm PDFS starts the search from a vertex $a\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$, it does not set the attribute $a.visited$ to true. Consequently if for this vertex $a\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$, there exists a cycle leading from the vertex $a$ to itself, then Algorithm PDFS visits vertex $a$ twice.
We run Algorithm PDFS on the auxiliary dpwpa $AUX[E;F]$. We shall show that PDFS on $AUX[E;F]$ outputs true if and only if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
Lemma 14.123.
[8] Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and let $E$ and $F$ be GTESs over $\Sigma$ such that $R\langle E,F\rangle$ is total. Let $a\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$.
(i) During the run of Algorithm PDFS on $AUX[E;F]$, Function $\mathrm{PDFS-VISIT}$ on $(AUX[E;F],a)$ visits at most once those vertices $b\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from $a$ to $b$.
(ii) During the run of Algorithm PDFS on $AUX[E;F]$, Function $\mathrm{PDFS-VISIT}$ visits at most once those vertices $b\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex $a\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to $b$.
Proof.
We obtain Statement (i) by direct inspection of Definition 14.122. Statement (i) implies Statement (ii). ∎
Proposition 14.124.
[8] Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and let $E$ and $F$ be GTESs over $\Sigma$ such that $R\langle E,F\rangle$ is total. Algorithm PDFS runs on $AUX[E;F]$ in $O(n)$ time.
Proof.
The running time of PDFS is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$ [8]. Recall that $n=size(E)+size(F)$. By Definition 4.26 and Statement 4.45, $|C\langle E\cup F,\emptyset\rangle|\leq|W|\leq n$. By Statement 6.63 $|A[E;F]|\leq n$. Hence $|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|\leq 2n$. Then we have the proposition. ∎
Example 14.125.
We now continue Examples 3.21, 5.54, 7.73, and 10.100. First Algorithm PDFS decides whether for each vertex $a\in C\langle E\cup F,\emptyset\rangle$, it holds that $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$, that is, $\left(a.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$. Recall that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$. As a result
(#$\left(\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}\;\,\mbox{and}\;\,\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}\right)=\mathrm{true}$.
Consequently, Algorithm PDFS outputs false.
Example 14.126.
We now continue Examples 3.22, 5.55, 7.74, and 10.101. Recall that there exist one path of positive length and two cycles in $AUX[E;F]$.
The first cycle: Its length is $1$, and it starts from the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
The path of positive length: Its length is $1$, and it starts from the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
The second cycle: Its length is $1$, and it starts from the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
First Algorithm PDFS decides whether for each $a\in C\langle E\cup F,\emptyset\rangle$, it holds that $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$, that is, $\left(a.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$. Observe that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$. Thus
(#$\left(\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}\;\mbox{or}\;\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}\right)=\mathrm{true}$.
Observe that \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$ and \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$. As a result
(\$$\left(\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}\;\mbox{or}\;\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}\right)=\mathrm{true}$.
Algorithm PDFS examines the vertices of $AUX[E;F]$, it looks for vertices from which it can start the search. To this end, Algorithm PDFS decides whether for each $a\in C\langle E\cup F,\emptyset\rangle$, it holds that $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$, that is, $\left(a.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$. First, Algorithm PDFS examines the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. It observes that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$. Then Algorithm PDFS starts the search from #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. It goes along the cycle (#$(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})$. Algorithm PDFS observes that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$. Second, Algorithm PDFS examines the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Since \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, Algorithm PDFS does not start the search starting from \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Consequently, Algorithm PDFS outputs true.
14.2 Decision Algorithm and its verification
We present our decision algorithm and show its correctnes.
Lemma 14.127.
Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$. Let $E$ and $F$ be GTESs over $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total. Then the following conditions are pairwise equivalent:
  • (i)
    $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$,
  • (ii)
    for each $t\in W$, $\left(t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}\mbox{ or }t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}\right)$.
  • (iii)
    For each $t\in W$, $\left(t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}\mbox{ or }t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle F,E\rangle}\right)$.
  • (iv)
    GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
  • (v)
    for each $a\in C\langle E\cup F,\emptyset\rangle$,
    • $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$ and
    • if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
  • (vi)
    for each $a\in C\langle E\cup F,\emptyset\rangle$,
    • $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$ and
    • if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
Proof.
$\mathrm{(i)}\Leftrightarrow\mathrm{(ii)}$: By Lemma 9.81 and Lemma 9.85.
$\mathrm{(ii)}\Leftrightarrow\mathrm{(iii)}$: By Lemma 9.86.
$\mathrm{(iii)}\Leftrightarrow\mathrm{(iv)}$: By Lemma 10.99.
$\mathrm{(iv)}\Leftrightarrow\mathrm{(v)}$: By Lemma 10.94.
$\mathrm{(v)}\Leftrightarrow\mathrm{(vi)}$: By Statement 4 in Lemma 8.76. ∎
Lemma 14.128.
Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and let $E$ and $F$ be GTESs over $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total.
1. If GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$, then Conditions (a) and (b) hold:
  • (a)
    For each vertex $a\in C\langle E\cup F,\emptyset\rangle$, $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$, and
  • (b)
    Algorithm PDFS on $AUX[E;F]$ visits exactly once each vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex $b\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to $a$. For each visited vertex $a$, $a.keeps_{F}=\mathrm{true}$.
  • (c)
    Algorithm PDFS on $AUX[E;F]$ outputs true.
2. If GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$, then Algorithm PDFS on $AUX[E;F]$ outputs false and Conditions (d) or (e) hold:
  • (d)
    There exists a vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$.
  • (e)
    Algorithm PDFS on $AUX[E;F]$ visits at most once each vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex $b\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to $a$. Moreover, Algorithm PDFS visits a vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that $a.keeps_{F}=\mathrm{false}$.
Proof.
By direct inspection of Definition 14.122 and by the equivalence $\mathrm{(iv)}\Leftrightarrow\mathrm{(vi)}$ in Lemma 14.127. ∎
Lemma 14.129.
Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and let $E$ and $F$ be GTESs over $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total. $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if Algorithm PDFS on $AUX[E;F]$ outputs true.
Proof.
By the equivalence $\mathrm{(i)}\Leftrightarrow\mathrm{(iv)}$ in Lemma 14.127 and by Lemma 14.128. ∎
Lemma 14.130.
Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and let $E$ and $F$ be GTESs over $\Sigma$. Assume that $R\langle E,F\rangle$ is total. Then we can decide in $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Proof.
We run Algorithm PDFS on $AUX[E;F]$ in $O(n)$ time by Lemma 14.124. By Lemma 14.129, Algorithm PDFS on the auxiliary dpwpa $AUX[E;F]$ outputs the answer to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. ∎
Lemma 14.131.
Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and let $E$ and $F$ be GTESs over $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total. Then we can decide in $O(n\,\mathrm{log}\,n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Proof.
Recall that we have run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.47, we got the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$. Then we run Algorithm PDFS on $AUX[E;F]$ in $O(n)$ time. Algorithm PDFS on input $AUX[E;F]$ outputs the answer to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Thence the complexity of our entire computation is $O(n\,\mathrm{log}\,n)$. ∎
Example 14.132.
We now continue Examples 3.21, 7.73, 10.100, and 14.125. By the last lines of Example 10.100, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Accordingly Condition (iii) in Lemma 14.127 does not hold. For this reason, by Lemma 14.127, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. By the last line of Example, 14.125 Algorithm PDFS outputs false to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. As a consequence, our decision algorithm also outputs false to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Example 14.133.
We now continue Examples 3.22, 5.55, 7.74, 10.101, and 14.126. By the last line of Example 10.101, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Therefore Condition (iii) in Lemma 14.127 holds. Consequently, by Lemma 14.127, $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. By the last line of Example, 14.126 Algorithm PDFS outputs true to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. As a consequence, our decision algorithm also outputs true to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
15 Main Case 4
In this main case, $E$ and $F$ are GTESs over a signature $\Sigma$ such that $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total. We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.47. Example 3.23 and its sequel Example 5.56, furthermore, Example 3.24 and its sequels Examples 5.57 and 15.134 demonstrate this case.
Example 15.134.
We continue Examples 3.24 and 5.57. We recall that
$f(\#,\mathsterling)\mathop{\leftrightarrow}^{*}_{E\cup F}f(\$,\flat)$,  $(f(\#,\mathsterling),f(\$,\flat))\not\in\mathop{\leftrightarrow}^{*}_{E}$,   and   $(f(\#,\mathsterling),f(\$,\flat))\not\in\mathop{\leftrightarrow}^{*}_{F}$.
The idea of the ()($\Rightarrow$)( ⇒ ) direction of the proof of the next theorem comes from the proof of the above observation.
Theorem 15.135.
Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$. Let $E$ and $F$ be GTESs over $\Sigma$ such that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total. Then
$\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$ or $F\subseteq\mathop{\leftrightarrow}^{*}_{E}$.
Proof.
()($\Rightarrow$)( ⇒ ) We proceed by contrapositive. Assume that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$. That is, there exists $l_{1}\doteq r_{1}\in E$ such that $(l_{1},r_{1})\not\in\mathop{\leftrightarrow}^{*}_{F}$ and there exists $l_{2}\doteq r_{2}\in F$ such that $(l_{2},r_{2})\not\in\mathop{\leftrightarrow}^{*}_{E}$. Then $l_{1}\neq r_{1}$ and $l_{2}\neq r_{2}$. Let $\sigma\in\Sigma_{k}$ with $k\geq 2$, and by Lemma 4.42, let $v\in T_{\Sigma}\setminus\bigcup W/_{\mathop{\leftrightarrow}^{*}_{E}}$, and $z\in T_{\Sigma}\setminus\bigcup W/_{\mathop{\leftrightarrow}^{*}_{F}}$. Then
$\sigma(\sigma(v,l_{1},\ldots,l_{1}),\sigma(z,l_{2},\ldots,l_{2}),v,\ldots,v)\mathop{\leftrightarrow}^{*}_{E\cup F}\sigma(\sigma(v,r_{1},\ldots,r_{1}),\sigma(z,r_{2},\ldots,r_{2}),v,\ldots,v)$.
We now show that
$(\sigma(\sigma(v,l_{1},\ldots,l_{1}),\sigma(z,l_{2},\ldots,l_{2}),v,\ldots,v),\sigma(\sigma(v,r_{1},\ldots,r_{1}),\sigma(z,r_{2},\ldots,r_{2}),v,\ldots,v))\not\in\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$,
and hence we have
$\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
To this end, first we show that
$(\sigma(\sigma(v,l_{1},\ldots,l_{1}),\sigma(z,l_{2},\ldots,l_{2}),v,\ldots,v),\sigma(\sigma(v,r_{1},\ldots,r_{1}),\sigma(z,r_{2},\ldots,r_{2}),v,\ldots,v))\not\in\mathop{\leftrightarrow}^{*}_{E}$.
As $l_{1}\neq r_{1}$ and $l_{2}\neq r_{2}$,
$\sigma(\sigma(v,l_{1},\ldots,l_{1}),\sigma(z,l_{2},\ldots,l_{2}),v,\ldots,v)\neq\sigma(\sigma(v,r_{1},\ldots,r_{1}),\sigma(z,r_{2},\ldots,r_{2}),v,\ldots,v)$.
We now show that
$(\sigma(\sigma(v,l_{1},\ldots,l_{1}),\sigma(z,l_{2},\ldots,l_{2}),v,\ldots,v),\sigma(\sigma(v,r_{1},\ldots,r_{1}),\sigma(z,r_{2},\ldots,r_{2}),v,\ldots,v))\not\in\mathop{\leftrightarrow}_{E}^{+}$.
To this purpose we now show the following statement.
Claim 15.136.
Let
$\sigma(\sigma(v,l_{1},\ldots,l_{1}),\sigma(z,l_{2},\ldots,l_{2}),v,\ldots,v)=s_{0}\mathop{\leftrightarrow}_{[\alpha_{1},p_{1}\doteq q_{1}]}s_{1}\mathop{\leftrightarrow}_{[\alpha_{2},p_{2}\doteq q_{2}]}\cdots\mathop{\leftrightarrow}_{[\alpha_{n},p_{n}\doteq q_{n}]}s_{n}=t,\;n\geq 1.$ (3)
[…]
Proof.
[…]
Base case: $n=1$. By the definition of $v$, $v\not\in\bigcup W/_{\mathop{\leftrightarrow}^{*}_{E}}$, hence
$v$ is not a subterm of $\sigma(\sigma(v,l_{1},\ldots,l_{1}),\sigma(z,l_{2},\ldots,l_{2}),v,\ldots,v)|_{\alpha_{1}}$,
thus $\alpha_{1}$ is not a prefix of $11$.
[…]
$t\neq\sigma(\sigma(v,r_{1},\ldots,r_{1}),\sigma(z,r_{2},\ldots,r_{2}),v,\ldots,v)$.
Then
$(\sigma(\sigma(v,l_{1},\ldots,l_{1}),\sigma(z,l_{2},\ldots,l_{2}),v,\ldots,v),\sigma(\sigma(v,r_{1},\ldots,r_{1}),\sigma(z,r_{2},\ldots,r_{2}),v,\ldots,v))\not\in\mathop{\leftrightarrow}^{*}_{E}$.
We can show similarly that
$(\sigma(\sigma(v,l_{1},\ldots,l_{1}),\sigma(z,l_{2},\ldots,l_{2}),v,\ldots,v)),\sigma(\sigma(v,r_{1},\ldots,r_{1}),\sigma(z,r_{2},\ldots,r_{2}),v,\ldots,v))\not\in\mathop{\leftrightarrow}^{*}_{F}$.
Consequently, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
()($\Leftarrow$)( ⇐ ) If $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{F}$, and hence $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. If $F\subseteq\mathop{\leftrightarrow}^{*}_{E}$, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}$, and hence $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. ∎
Consequence 15.137.
Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$. Let $E$ and $F$ be GTESs over $\Sigma$ such that $R\langle E,F\rangle$ and $R\langle F,E\rangle$ are not total. Then
$\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $E\subseteq\Theta\langle F,E\rangle$ or $F\subseteq\Theta\langle E,F\rangle$.
Proof.
By Definition 2.9, we have $\Theta\langle F,E\rangle=\mathop{\leftrightarrow}^{*}_{F}\cap W\times W$ and $\Theta\langle E,F\rangle=\mathop{\leftrightarrow}^{*}_{E}\cap W\times W$. Hence Theorem 15.135 implies our consequence. ∎
Lemma 15.138.
Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and let $E$ and $F$ be GTESs over $\Sigma$ such that $R\langle E,F\rangle$ and $R\langle F,E\rangle$ are not total. We can decide in $O(n\,\mathrm{log}\,n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Proof.
By Consequence 15.137, we proceed as follows. Recall that we have run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.47, we got the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$ in $O(n\,\mathrm{log}\,n)$ time.
First we decide whether $E\subseteq\Theta\langle F,E\rangle$. By Consequence 15.137 and Proposition 4.25,
$E\subseteq\Theta\langle F,E\rangle$ if and only if
for each $l\doteq r$ in $E$, $(l,r)\in\Theta\langle F,E\rangle$ if and only if
for each $(u,v)\in\tau\langle E,F\rangle$, $u/_{\rho\langle F,E\rangle}=v/_{\rho\langle F,E\rangle}$ if and only if
$\mathrm{FIND}_{\rho\langle F,E\rangle}(u)=\mathrm{FIND}_{\rho\langle F,E\rangle}(v)$.
We decide whether for each $(u,v)\in\tau\langle E,F\rangle$, $\mathrm{FIND}_{\rho\langle F,E\rangle}(u)=\mathrm{FIND}_{\rho\langle F,E\rangle}(v)$. If so, $E\subseteq\Theta\langle F,E\rangle$, otherwise $E\not\subseteq\Theta\langle F,E\rangle$.
Second, similarly to above, we decide whether $F\subseteq\Theta\langle E,F\rangle$.
We call above the functions $\mathrm{FIND}_{\rho\langle F,E\rangle}$ and $\mathrm{FIND}_{\rho\langle E,F\rangle}$ a total of at most $2n$ times, where $n=size(E)+size(F)$. These callings, by the main result of [16], require $O(n)$ time. As a result, the time complexity of our entire computation is $O(n\,\mathrm{log}\,n)$. ∎
16 Main Result
We now sum up our decidability and complexity results.
Theorem 16.139.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. It is decidable in $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n=size(E)+size(F)$.
Proof.
We run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.47, we get the following output in $O(n\,\mathrm{log}\,n)$ time:
  • the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$;
  • the set of nullary symbols $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$;
  • $INC(E\cup F,E)$ stored in a red-black tree saved in a variable $RBT1$;
  • $INC(E\cup F,F)$ stored in a red-black tree saved in a variable $RBT2$;
  • $NUM(E\cup F,E)$ stored in a red-black tree saved in a variable $RBT3$;
  • $NUM(E\cup F,F)$ stored in a red-black tree saved in a variable $RBT4$;
  • the GTRS $R\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT5$; each node in $RBT5$ contains a search key which is a rule in $R\langle E\cup F,\emptyset\rangle$, and satellite data $counter_{E}$ and $counter_{F}$;
  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT6$; for every ordered pair $(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
  • the GTRS $R\langle E,F\rangle$ stored in a red-black tree saved in a variable $RBT7$;
  • the GTRS $R\langle F,E\rangle$ stored in a red-black tree saved in a variable $RBT8$;
  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
By Lemma 4.49, we decide in $O(n)$ time whether $R\langle E,F\rangle$ is total, and whether $R\langle F,E\rangle$ is total.
  • If both of them are total, then we proceed as in main Case 2, see Lemma 13.120.
  • Otherwise,
    • if $\Sigma$ is a unary alphabet, then we proceed as in Main Case 1, see Lemma 12.113.
    • If $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total, then we proceed as in Main Case 3, see Lemma 14.131.
    • If $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total, then we proceed as in Main Case 4, see Lemma 15.138. ∎
By Theorem 16.139 and Proposition 9.82, we have the following consequence.
Consequence 16.140.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then we can decide in $O(n^{2})$ time whether there exists a GTES $H$ over $\Sigma$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n=size(E)+size(F)$. If the answer is ’yes’, then $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$ and $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Otherwise $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is not a congruence on the ground term algebra ${\cal T}(\Sigma)$.
Proof.
By Theorem 16.139, we decide in $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. By Proposition 9.82, if the answer is ’yes’, then there exists a GTES $H$ over $\Sigma$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, and $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$; otherwise $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is not a congruence on the ground term algebra ${\cal T}(\Sigma)$. ∎
17 Conclusion
We showed that for any GTESs $E$ and $F$ over a signature $\Sigma$, the following four statements are equivalent:
  • 1.
    $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
  • 2.
    $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
  • 3.
    There exists a GTES $H$ over $\Sigma$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
  • 4.
    $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$.
We showed that we can decide in $O(n^{2})$ time for given GTESs $E$ and $F$ over a signature $\Sigma$ whether Statement 1 above holds. Here $n=size(E)+size(F)$. We can construct the GTES $E\cup F$ in $O(n^{2})$ time.
We used the fast ground completion algorithm introducing additional constants of Fülöp and Vágvölgyi [2], which, given a GTESs $E$ and $F$ over a signature $\Sigma$ produces in $O(n\,\mathrm{log}\;n)$ time a finite set $C\langle E,F\rangle$ of new constants and a reduced GTRS $R\langle E,F\rangle$ over $\Sigma\cup C\langle E,F\rangle$ such that $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$. Along the proof, we distinguished four main cases. Main Case 2 overlaps with Main Case 1 and Main Case 3, and is a subcase of the union of Main Case 1 and Main Case 3.
Given GTES $E$ and GTES $F$, for every $p\in W$, we can construct the deterministic bottom-up tree automata ${\cal A}$, ${\cal B}$, and ${\cal C}$ such that ${\cal A}$ recognizes $p/_{\mathop{\leftrightarrow}^{*}_{E}}$, ${\cal B}$ recognizes $p/_{\mathop{\leftrightarrow}^{*}_{F}}$, and ${\cal C}$ recognizes $p/_{\mathop{\leftrightarrow}^{*}_{E}\cup F}$, see Proposition 2.10. Thence, in Main Cases 1, 2, and 3, we can reduce our decidability problem to deciding the equivalence of two deterministic bottom-up tree automata. Champavère et al [5] presented an algorithm for deciding inclusion between two tree languages recognized by given bottom-up tree automata. Applying their algorithm to our decision problem in Main Cases 1, 2, and 3 yields an alternative decision algorithm, which is more sophisticated than ours, and in general takes more time than ours, see Lemmas 12.103, 13.116, and 14.121.
We raise the following more general problem:
Problem 17.141.
Let $D$, $E$, and $F$ be GTESs over a signature $\Sigma$. Can we decide whether $\mathop{\leftrightarrow}^{*}_{D\cup E\cup F}=\mathop{\leftrightarrow}^{*}_{D}\cup\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$? If we can, then what is the time complexity of our decision algorithm?
One might think that we can generalize our constructions and decidability results in a straightforward way. However, the following example shows that this belief is not justified.
Example 17.142.
Consider the unary ranked alphabet $\Sigma=\Sigma_{0}=\{\,\#,\;\$,\;\mathsterling\,\}$. Let
$D=\{\,\#\doteq\$\,\}$,  $E=\{\,\#\doteq\mathsterling\,\}$,  $F=\{\,\$\doteq\mathsterling\,\}$,  $H=D\cup E\cup F$.
Then
  • $ST\langle H\rangle=\Sigma_{0}$,
  • $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{D}\cup\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$,
  • #$\#/_{\mathop{\leftrightarrow}^{*}_{D}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,   #$\#/_{\mathop{\leftrightarrow}^{*}_{E}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,   #$\#/_{\mathop{\leftrightarrow}^{*}_{F}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,
  • #$\#/_{\mathop{\leftrightarrow}^{*}_{D}},\;\#/_{\mathop{\leftrightarrow}^{*}_{E}},\;\#/_{\mathop{\leftrightarrow}^{*}_{F}}$ are pairwise incomparable for the inclusion relation, and
  • for each $t\in ST\langle H\rangle$,   $t/_{\mathop{\leftrightarrow}^{*}_{D}}\subset t/_{\mathop{\leftrightarrow}^{*}_{H}}$,   $t/_{\mathop{\leftrightarrow}^{*}_{E}}\subset t/_{\mathop{\leftrightarrow}^{*}_{H}}$,   and   $t/_{\mathop{\leftrightarrow}^{*}_{F}}\subset t/_{\mathop{\leftrightarrow}^{*}_{H}}$.
Then the generalization of Condition (ii) in Lemma 12.112 does not hold.

### v2

@@
 finitely generated congruence on the ground term algebra, reduced ground term rewrite system , bottom-up tree automaton , square time decision algorithm
-journal: some journal
+journal: Journal of Automata, Languages and Combinatorics
 […]
-Department of Foundations of Computer Science University of Szeged, addressline=Arpad ter 2, city=Szeged, postcode=H-6720, country=Hungary
-{highlights}
-We can decide in square time whether the union of two finitely generated congruences on the ground term algebra is also a congruence.
-If the answer is ’yes’, then the union of the two generator sets of ground equations generates the union of the congruences.
+Department of Foundations of Computer Science University of Szeged, addressline=Árpád tér 2, city=Szeged, postcode=H-6720, country=Hungary
 1 Introduction
-Vágvölgyi [1] showed that it is decidable for any given ground term equation systems (GTESs for short) $E$ and $F$ whether there is a GTES $H$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cap\mathop{\leftrightarrow}^{*}_{F}$.
-If the answer is yes, then we can construct such a GTES $H$.
+Kozen [1], Nelson and Oppen [2], and Tarjan [3] studied the word problem for ground term equation systems (GTESs) and applied the graph-based method of congruence closure.
+In order to decide for a given GTES $E$ and ground terms $s$ and $t$, whether $s\mathop{\leftrightarrow}^{*}_{E}t$ holds, they constructed the subterm graph of $E$ and $s$ and $t$ and computed the congruential consequences of $E$ on the vertices of the graph.
+Kozen [1] showed that the word problem for every GTES is decidable in polynomial time.
+Nelson and Oppen [2] and Tarjan [3] showed that the word problem for every GTES is decidable in $O(n\,\mathrm{log}\,n)$ time applying the congruence closure procedure on the subterm graph of the GTES $E$ and ground terms $s$ and $t$.
+Continuing this research line, Snyder [4] gave an efficient algorithm which, given a GTES $E$ over a signature $\Sigma$, produces an equivalent, reduced ground term rewriting system $R$ over $\Sigma$ in $O(n\,\mathrm{log}\,n)$ time.
+Then Fülöp and Vágvölgyi [5] gave a simple and efficient algorithm which, given a GTES $E$ over a signature $\Sigma$, produces a reduced ground term rewriting system (GTRS for short) $R$ over a signature containing $\Sigma$ such that two terms are congruent modulo $E$ if and only if their $R$-normal forms are equal.
+The algorithm calls a congruence closure procedure [3] on the subterm graph of the GTES $E$ and runs in $O(n\,\mathrm{log}\,n)$ time, where $n$ is the size of $E$.
+We say that a tree language is congruential if it is the union of finitely many congruence classes of a finitely generated congruence.
+Brainerd [6], Kozen [1] showed that congruential tree languages are the same as recognizable tree languages.
+Applying the above algorithm of Fülöp and Vágvölgyi [5], for a given GTES $E$ and ground terms $p_{1},\ldots,p_{k}$ of size $n$, a deterministic bottom-up tree automaton can be constructed in $O(n\,\mathrm{log}\,n)$ time which recognizes the congruential tree language $p_{1}/_{\mathop{\leftrightarrow}^{*}_{E}}\cup\cdots\cup p_{k}/_{\mathop{\leftrightarrow}^{*}_{E}}$ [7].
+Vágvölgyi [8] showed that it is decidable for any given GTESs $E$ and $F$ whether there is a GTES $H$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cap\mathop{\leftrightarrow}^{*}_{F}$.
+If the answer is “yes”, then we can construct such a GTES $H$.
 We show the dual to the above result, first we show that for any GTESs $E$ and $F$, $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra if and only if there exists a GTES $H$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
@@
 Thus, throughout the paper we consider signatures with constants.
-We now sketch the proof of our main result.
+We now recall the fast ground completion algorithm of Fülöp and Vágvölgyi [5].
 Let $E$ and $F$ be GTESs over $\Sigma$, and let $W$ denote the set of subterms appearing in the GTESs $E$ and $F$.
-Fülöp and Vágvölgyi [2] presented a fast ground completion algorithm which, given GTESs $E$ and $F$ over a signature $\Sigma$, produces in $O(n\,\mathrm{log}\;n)$ time a finite set $C\langle E,F\rangle$ of new constants and a reduced ground term rewrite system (GTRS for short) $R\langle E,F\rangle$ over $\Sigma$ extended with the new constants such that $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$.
-Here $n$ stands for the number of occurrences of symbols in $E$ plus number of occurrences of symbols in $F$.
+Fülöp and Vágvölgyi [5] presented a fast ground completion algorithm which, given GTESs $E$ and $F$ over a signature $\Sigma$, produces in $O(n\,\mathrm{log}\,n)$ time a finite set $C\langle E,F\rangle$ of new constants and a reduced GTRS $R\langle E,F\rangle$ over $\Sigma$ extended with the new constants such that $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$.
+Here $n$ stands for the number of occurrences of symbols in $E$ plus the number of occurrences of symbols in $F$.
 Furthermore, the set $C\langle E,F\rangle$ of new constants is the set of the equivalence classes of the restriction $\Theta\langle E,F\rangle$ of $\mathop{\leftrightarrow}^{*}_{E}$ to the set $W$.
-For this purpose we proceed as follows:
-[…]
-[…]
-We compute the following sets and relations and dpwpa in $O(n\,\mathrm{log}\,n)$ time:
-  • the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$;
-  • the set of nullary symbols $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$;
-  • the GTRS $R\langle E\cup F,\emptyset\rangle$ stored in a red-black tree;
-  • the GTRS $R\langle E,F\rangle$ stored in a red-black tree;
-  • the GTRS $R\langle F,E\rangle$ stored in a red-black tree;
-  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
-Then we decide in $O(n)$ time whether $R\langle E,F\rangle$ and $R\langle F,E\rangle$ are total. Accordingly, we distinguish the following four main cases.
+For this purpose we proceed as follows: In the first step in $O(n)$ time, we create a subterm directed, labeled, and acyclic multigraph (dag) $G$ for $E$ and $F$ [4].
+We consider the vertices in the subterm dag $G$ as terms in $W$.
+The relation $\tau\langle E,F\rangle$ on the vertices of $G$ corresponds to $E$.
+We construct the relation $\tau\langle E,F\rangle$.
+Then we run on $G$ a congruence closure algorithm for the relation $\tau\langle E,F\rangle$ in $O(n\,\mathrm{log}\,n)$ time [9].
+In this way, we obtain on $G$ the congruence closure $\rho\langle E,F\rangle$ of the relation $\tau\langle E,F\rangle$.
+Considering the vertices in the subterm dag $G$ as terms, $\rho\langle E,F\rangle$ yields the restriction $\Theta\langle E,F\rangle$ of $\mathop{\leftrightarrow}^{*}_{E}$ to $W$.
+Then using $\rho\langle E,F\rangle$, $\Theta\langle E,F\rangle$, and $C\langle E,F\rangle$, we produce the GTRS $R\langle E,F\rangle$ in $O(n)$ time.
+GTRS $R\langle E,F\rangle$ is the set of all rewrite rules $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$, where $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, and ${b}_{1},\ldots,{b}_{m},{b}\in C\langle E,F\rangle$, and $\{\,\sigma(t_{1},\ldots,t_{m})\mid t_{1}\in{b}_{1},\ldots,t_{m}\in{b}_{m}\,\}\subseteq{b}$.
+Fülöp and Vágvölgyi [5] showed that GTRS $R\langle E,F\rangle$ is reduced, and hence is convergent.
+We construct GTRSs $R\langle F,E\rangle$ and $R\langle E\cup F,\emptyset\rangle$ in a similar way.
+The intuitive idea of the paper is as follows.
+Let $E$ and $F$ be GTESs over $\Sigma$.
+For every subterm $t$ appearing in $E\cup F$, GTRS $R\langle E,F\rangle$ reduces all elements of the congruence class $t/_{\mathop{\leftrightarrow}^{*}_{\langle E,F\rangle}}$ to the constant (equivalence class) $t/_{\Theta\langle E,F\rangle}$, and GTRS $R\langle E\cup F,\emptyset\rangle$ reduces every element of the congruence class $t/_{\mathop{\leftrightarrow}^{*}_{\langle E\cup F,\emptyset\rangle}}$ to the constant (equivalence class) $t/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+For every subterm $t$ appearing in $E\cup F$, the congruence class $t/_{{\mathop{\leftrightarrow}^{*}_{\langle E\cup F,\emptyset\rangle}}}$ is equal to the congruence class $t/_{{\mathop{\leftrightarrow}^{*}_{\langle E,F\rangle}}}$ if and only if $t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}$ and for every $p\in T_{\Sigma}$ and reduction sequence ${\cal RS}1$ from $p$ to $t/_{\Theta\langle E\cup F,\emptyset\rangle}$ by $R\langle{E\cup F},\emptyset\rangle$, there exists a reduction sequence ${\cal RS}2$ from $p$ to $t/_{\Theta\langle E,F\rangle}$ by $R\langle E,F\rangle$, where each rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle{E\cup F},\emptyset\rangle$ applied at any node of $p$ along the reduction sequence ${\cal RS}1$ is matched with some rewrite rule $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}$ in $R\langle E,F\rangle$ applied by GTRS $R\langle E,F\rangle$ at the same node along the reduction sequence ${\cal RS}2$.
+By matching we mean that ${c}_{i}\subseteq{b}_{i}$ for all $i\in\{\,1,\ldots,m\,\}$, and ${c}\subseteq{b}$.
+For every ${b}\in C\langle E\cup F,\emptyset\rangle$, we say that GTRS $R\langle E,F\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant b{b} if the following holds.
+If GTRS $R\langle E\cup F,\emptyset\rangle$ applies a rewrite rule of the form $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$ along a reductions sequence from $p$ to $t/_{\Theta\langle E,F\rangle}$, then $R\langle E,F\rangle$ also applies a rewrite rule $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}$ for some ${c}_{1},\ldots,{c}_{m},c\in C\langle E,F\rangle$ at the same node, that is, $R\langle E,F\rangle$ also move forward along its reduction sequence.
+For every ${b}\in C\langle E\cup F,\emptyset\rangle$, we define that GTRS $R\langle F,E\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant b{b} in a similar way.
+For each constant ${d}\in C\langle E\cup F,\emptyset\rangle$, we keep track of all constants ${c}\in C\langle E\cup F,\emptyset\rangle$ which appear as a right-hand side of a rewrite rule applied along the reduction sequence from some ground term $t\in T_{\Sigma}$ to the constant d{d} by $R\langle E\cup F,\emptyset\rangle$.
+To this end, we construct a directed pseudograph without parallel arcs (dpwpa), $AUX[E;F]$ with vertices $C\langle E\cup F,\emptyset\rangle$.
+At each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, we indicate whether it is an element of $C\langle E,F\rangle$ whether it is an element of $C\langle F,E\rangle$, whether GTRS $R\langle E,F\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant b{b}, and whether GTRS $R\langle F,E\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant b{b}.
+Every directed arc (b,c)({b},{c}) is associated with some rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$ in $R\langle E\cup F,\emptyset\rangle$, where ${c}={b}_{i}$ for some $i\in\{\,1,\ldots,m\,\}$.
+That is, b{b} is the right-hand side of the rewrite rule, and c{c} appears on the left-hand side of the rewrite rule.
+We say that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${d}\in C\langle E\cup F,\emptyset\rangle$ if ${d}\in C\langle E,F\rangle$ and for all constants c{c} appearing in a reduction sequence from a ground term $p\in T_{\Sigma}$ to d{d} by GTRS $R\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing $c$.
+That is, for all constants c{c} with a path of positive length or a cycle from c{c} to d{d} in $AUX[E;F]$, $R\langle E,F\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing $c$.
+We define that GTRS $R\langle F,E\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${d}\in C\langle E\cup F,\emptyset\rangle$ in a similar way.
+During the run of the decision algorithm, we carry out a partial depth first search on the auxiliary dpwpa $AUX[E;F]$ to decide whether $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${d}\in C\langle E\cup F,\emptyset\rangle$, and to decide whether $R\langle F,E\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${d}\in C\langle E\cup F,\emptyset\rangle$.
+We say that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if for each constant ${d}\in C\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on d{d}.
+We show that for every subterm $t$ appearing in $E\cup F$, $\left(t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{{\mathop{\leftrightarrow}^{*}_{E}}}\mbox{ or }t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{{\mathop{\leftrightarrow}^{*}_{F}}}\right)$ if and only if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
+We consider the following four main cases.
 Main Case 1: $\Sigma$ is a unary signature.
-In this main case $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if for each $a\in C\langle E\cup F,\emptyset\rangle$, Condition (i) or Condition (ii) hold:
-  • (i)
-    • (a)
-      $a\in C\langle E,F\rangle$ and
-    • (b)
-      for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]\rangle=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
-  • (ii)
-    • (a)
-      $a\in C\langle F,E\rangle$ and
-    • (b)
-      for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
-We decide the above formulas by carrying out a nested partial depth first search on the auxiliary dpwpa $AUX[E;F]$.
-In the outer loop we carry out a partial depth first search on $AUX[E;F]$ and decide Conditions (i).(a) and (ii).(a).
-In the inner loop we visit those vertices which are positive-step reachable from the current vertex of the outer loop, and examine at each visited vertex $a$ the value of the Boolean attribute $keeps_{E}$ (respectively $keeps_{F}$), which tells us whether $R\langle E,F\rangle$ (resp.
-$R\langle F,E\rangle$) keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $a$.
-In this way we decide Condition (ii).(a) (resp.
-Condition (ii).(b)).
-In such manner, we decide in additional $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
+Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
 Main Case 2: Both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total.
 Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class.
-We decide in additional $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
-Main Case 3: $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total.
-Without loss of generality we may assume that $R\langle E,F\rangle$ is total.
-Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if for each $a\in C\langle E\cup F,\emptyset\rangle$,
-  • (1)
-    $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$ and
-  • (2)
-    if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
-In $O(n)$ time, we decide set memberships $a\in C\langle E,F\rangle$ and $a\in C\langle F,E\rangle$ for all $a\in C\langle E\cup F,\emptyset\rangle$.
-Therefore we decide in $O(n)$ time whether Condition (1) holds for all $a\in C\langle E\cup F,\emptyset\rangle$.
-We decide in $O(n)$ time whether Condition (2) holds for all $a\in C\langle E\cup F,\emptyset\rangle$ by carrying out a partial depth first search on the auxiliary dpwpa $AUX[E;F]$ by narrowing down the set of vertices from which we start a session of the search.
-We start the search sessions from those $C\langle E\cup F,\emptyset\rangle$ vertices $a$ which are not in $C\langle E,F\rangle$ and have not been visited previously.
-If we start a search session from a vertex $a$ and there exists a cycle leading from the vertex $a$ to itself, then along this cycle Algorithm PDFS visits vertex $a$ for a second time.
-During the current session of the search we find a vertex $b$ with false value for the attribute $keeps_{F}$ if and only if Condition (2) does not hold for the vertex $a\in C\langle E\cup F,\emptyset\rangle$ from which we started the current session of the search.
-We decide whether Condition (1) holds for all $a\in C\langle E\cup F,\emptyset\rangle$ and whether Conditions (2) holds for all $a\in C\langle E\cup F,\emptyset\rangle$ in $O(n)$ time, in this way, we decide in additional $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
+Main Case 3: $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$ and $R\langle E,F\rangle$ or $R\langle F,E\rangle$ is total.
+Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
 Main Case 4: $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total.
-Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$ or $F\subseteq\mathop{\leftrightarrow}^{*}_{E}$ if and only if $E\subseteq\Theta\langle F,E\rangle$ or $F\subseteq\Theta\langle E,F\rangle$.
-In this case, by direct inspection of the representations $\tau\langle E,F\rangle$ of $E$, $\tau\langle F,E\rangle$ of $F$, $\rho\langle E,F\rangle$ of ${\Theta\langle E,F\rangle}$ and $\rho\langle F,E\rangle$ of ${\Theta\langle F,E\rangle}$ in the subterm dag for $E$ and $F$, we decide in additional $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
-Summing up the previous thoughts, our decision algorithm goes as follows: We compute the following sets and relations and dpwpa in $O(n\,\mathrm{log}\,n)$ time:
-  • the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$;
-  • the set of nullary symbols $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$;
-  • the GTRS $R\langle E\cup F,\emptyset\rangle$ stored in a red-black tree;
-  • the GTRS $R\langle E,F\rangle$ stored in a red-black tree;
-  • the GTRS $R\langle F,E\rangle$ stored in a red-black tree;
-  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
-We decide whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total.
-  • If both of them are total, then we proceed as in main Case 2.
-  • Otherwise,
-    • if $\Sigma$ is a unary alphabet, then we procced as in Main Case 1.
-    • If $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total, the we proceed as in Main Case 3.
-    • If $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total, then we proceed as in Main Case 4.
+Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$ or $F\subseteq\mathop{\leftrightarrow}^{*}_{E}$.
 In Main Case 1, the overall complexity of our computation is $O(n^{2})$.
@@
 Algorithm concepts
-[…]
+In algorithmic analysis we use the RAM (random access machine) model [11, Section 2.2], [12, Section 2.1].
+For this model of computation, each simple operation (+,,,=,if,call)(+,\,*,\,–,\,=,\,if,\,call) takes exactly one time step.
+An array is a collection of items of same data type stored at contiguous memory locations.
+For the concept of a red-black tree, see [13, Section 13].
 Proposition 2.1.
-[8, Section 13] The search operation takes $O(\mathrm{log}\;n)$ time on a red-black tree, where $n$ is the size of the tree.
+[13, Section 13] The search operation takes $O(\mathrm{log}\,n)$ time on a red-black tree, where $n$ is the size of the tree.
 Proposition 2.2.
-[8, Section 13] The time-complexity of inserting $n\in{\mathbb{N}}$ elements into an empty red-black tree, not allowing duplicate values in the resulting red-black tree, is $O(n\,\mathrm{log}\;n)$.
+[13, Section 13] The time-complexity of inserting $n\in{\mathbb{N}}$ elements into an empty red-black tree, not allowing duplicate values in the resulting red-black tree, is $O(n\,\mathrm{log}\,n)$.
 Proof.
@@
 We pad the binary sequence $\alpha$ with leading zeros to length $\ell$ by producing the binary sequence $0^{\ell-k}\alpha$.
-We now recall the definition of the lexicographic order on $BIN$.
-For $\alpha,\beta\in BIN$, we say that $\alpha$ is lexicographically less than $\beta$, denoted by $\alpha<_{lex}\beta$ if either $\alpha$ is a proper prefix of $\beta$, or there exist $\gamma_{1},\gamma_{2},\gamma_{3}\in BIN$ such that $\alpha=\gamma_{1}0\gamma_{2}$ and $\beta=\gamma_{1}1\gamma_{3}$.
-[…]
-[…]
-  • Noetherian if there exists no infinite sequence of elements $a_{1},a_{2},a_{3},\ldots$ in $A$ such that $a_{1}{\rightarrow}a_{2}{\rightarrow}a_{3}{\rightarrow}\ldots$,
-  • confluent if for any elements $a_{1},a_{2},a_{3}$ in $A$, whenever $a_{1}{\rightarrow}^{*}a_{2}$ and $a_{1}{\rightarrow}^{*}a_{3}$, there exists an element $a_{4}$ in $A$ such that $a_{2}{\rightarrow}^{*}a_{4}$ and $a_{3}{\rightarrow}^{*}a_{4}$,
-  • convergent if it is Noetherian and confluent.
-[…]
-Proposition 2.3.
-[…]
-If $a\in A$ has a unique normalform, then we denote the latter by $a{\downarrow}$.
+We now recall the definition of the lexicographic order on $BIN$ [14, Definition 6.4.1.].
+Given two distinct binary strings, to determine whether one precedes the other in the lexicographic order we compare the numbers in the words.
+There are two possibilities:
+  1. 1.
+    The words have different lengths, and each number in the shorter word is identical to the corresponding number in the longer word.
+  2. 2.
+    The words have the same or different lengths, and at some position, the numbers in the words differ.
+If Condition 1 holds, then the shorter word precedes the longer.
+If Condition 2 holds, we locate the leftmost position $\ell$ at which the numbers differ.
+The order of the words is determined by the order of the numbers at position $\ell$.
+Definition 2.3 ([14, Definition 6.4.1.]).
+Let $\alpha=i_{1}i_{2}\ldots i_{m}$ and $\beta=j_{1}j_{2}\ldots j_{n}$ be binary strings. We say that $\alpha$ is lexicographically less than $\beta$ and write $\alpha<_{lex}\beta$ if
+  1. 1.
+    $m<n$ and $i_{k}=j_{k}$ for every $k=1,2,\ldots,m$ or
+  2. 2.
+    $i_{k}\neq j_{k}$ for some $k$ and for the smallest such $k$, we have $i_{k}<j_{k}$.
+Sets and Relations [15]
+For any sets $A$ and $B$, $A\subseteq B$ means that $A$ is a subset of $B$.
+Proper inclusion is denoted by $A\subset B$.
+The union of two sets $A$ and $B$ is denoted by $A\cup B$.
+The cardinality of a set $A$ is denoted $|A|$.
+A relation on a set $A$ is a subset $\rightarrow$} of $A\times A$ [15].
+We write ${b}{\rightarrow}{c}$ for $({b},{c})\in\,{\rightarrow}$.
+We denote by $\rightarrow$$^{*}$ the reflexive transitive closure of $\rightarrow$}, by $\leftrightarrow^{+}$ the transitive symmetric closure of $\rightarrow$}, and by $\leftrightarrow^{*}$ the reflexive transitive symmetric closure of $\rightarrow$}.
+Note that $\leftrightarrow^{*}$ is an equivalence relation on $A$.
+Let $\rightarrow$} be a relation on a set $A$.
+An element ${b}\in A$ is in normal form (irreducible for $\rightarrow$}) if there exists no ${c}\in A$ such that ${b}{\rightarrow}{c}$.
+We say that ${c}\in A$ is a normal form of ${b}\in A$ if ${b}{\rightarrow}^{*}{c}$ and c{c} is in normal form.
+A relation $\rightarrow$} on a set $A$ is
+  • 1.
+    normalizing if every element has a normal form,
+  • 2.
+    terminating if there exists no infinite sequence of elements ${b}_{1},{b}_{2},{b}_{3},\ldots$ in $A$ such that ${b}_{1}{\rightarrow}{b}_{2}{\rightarrow}{b}_{3}{\rightarrow}\ldots$,
+  • 3.
+    confluent if for any elements ${b}_{1},{b}_{2},{b}_{3}$ in $A$, whenever ${b}_{1}{\rightarrow}^{*}{b}_{2}$ and ${b}_{1}{\rightarrow}^{*}{b}_{3}$, there exists an element ${b}_{4}$ in $A$ such that ${b}_{2}{\rightarrow}^{*}{b}_{4}$ and ${b}_{3}{\rightarrow}^{*}{b}_{4}$,
+  • 4.
+    convergent if it is terminating and confluent.
+Any terminating relation is normalizing. Hence we have the following proposition.
+Proposition 2.4 ([15, Lemma 2.1.8]).
+For any convergent relation $\rightarrow$} on a set $A$, every ${b}\in A$ has a unique normal form, called the normal form of b{b}.
+If ${b}\in A$ has a unique normalform, then we denote the latter by ${b}{\downarrow}$.
 Let $\rho$ be an equivalence relation on $A$.
-Then for every $a\in A$, we denote by $a/_{\rho}$ the $\rho$ equivalence class containing $a$, i.e.
-$a/_{\rho}=\{\,b\mid a\rho b\,\}$.
-The classes associated with an equivalence relation $\rho$ are called equivalence classes.
+Then for every ${b}\in A$, we denote by ${b}/_{\rho}$ the $\rho$ equivalence class containing b{b}, i.e.
+${b}/_{\rho}=\{\,{c}\mid{b}\rho{c}\,\}$.
 For each $B\subseteq A$, let $B/_{\rho}=\{\,b/_{\rho}\mid b\in B\,\}$.
 The empty product $\Pi_{i=1}^{0}n_{i}=1$ over $\mathbb{N}$} is the result of multiplying no factors.
-It is by convention equal to $1$.
 The ceiling function maps a real number $x$ to the least integer greater than or equal to $x$, denoted $\lceil x\rceil$.
-Ground terms
-A signature $\Sigma$ is a finite set of symbols in which every element $\sigma$ has an arity, $\mathrm{arity}(\sigma)\in{\mathbb{N}}$.
-For each integer $m\in{\mathbb{N}}$, $\Sigma_{m}$ denotes the elements of $\Sigma$ which have arity $m$.
-We lose no generality assuming that $\Sigma_{0}\not=\emptyset$, see the Introduction.
+Ground terms [15]
+A signature $\Sigma$ is a finite set of symbols in which every element $\sigma$ is associated with a non-negative integer, the arity of $\sigma$ denoted by $\mathrm{arity}(\sigma)$.
+For each integer $k\in{\mathbb{N}}$, $\Sigma_{k}$ denotes the elements of $\Sigma$ which have arity $k$.
+We lose no generality assuming that $\Sigma_{0}\not$$\emptyset$, see the Introduction.
 If $\Sigma=\Sigma_{0}\cup\Sigma_{1}$, then $\Sigma$ is a unary signature.
@@
   • (ii)
-    $\sigma(t_{1},\ldots,t_{m})\in U$ whenever $\sigma\in\Sigma_{m}$ with $m\in{\mathbb{N}}$ and $t_{1},\ldots,t_{m}\in U$.
+    $\sigma(t_{1},\ldots,t_{m})\in U$ whenever $\sigma\in\Sigma_{m}$ with $m\geq 1$ and $t_{1},\ldots,t_{m}\in U$.
 Ground terms are also called trees. A set of ground terms is also called a tree language. For a ground term $t\in T_{\Sigma}$, the set $sub(t)$ of subterms of $t$ is defined by recursion as follows:
@@
 For a ground term $t\in T_{\Sigma}$, the set of positions $\mathrm{POS}(t)\subseteq{\mathbb{N}}^{*}$ of $t$ is defined by tree induction.
+  • 1.
     If $t\in\Sigma_{0},$ then $\mathrm{POS}(t)=\{\,\varepsilon\,\}$.
-  • If $t=\sigma(t_{1},\ldots,t_{m})$ with $\sigma\in\Sigma_{m},\ m>0,$ then $\mathrm{POS}(t)=\{\,\varepsilon\,\}\cup\{\,i\alpha\mid 1\leq i\leq m,\alpha\in\mathrm{POS}(t_{i})\,\}$.
-The size of a ground term $t\in T_{\Sigma}$ is $size(t)=|POS(t)|$.
-For a finite set $Z\subseteq T_{\Sigma}$ of ground terms, $size(Z)$ is the number of occurrences of symbols of $\Sigma$ in $Z$, i.e., $size(Z)=\sum_{t\in Z}size(t)$.
+  • 2.
+    If $t=\sigma(t_{1},\ldots,t_{m})$ with $\sigma\in\Sigma_{m}$, $m\geq 1$, then $\mathrm{POS}(t)=\{\,\varepsilon\,\}\cup\{\,i\alpha\mid 1\leq i\leq m,\alpha\in\mathrm{POS}(t_{i})\,\}$.
+The size of a ground term $t\in T_{\Sigma}$ is $\mathrm{size}(t)=|POS(t)|$.
+For a finite set $Z\subseteq T_{\Sigma}$ of ground terms, $\mathrm{size}(Z)$ is the number of occurrences of symbols of $\Sigma$ in $Z$, i.e., $\mathrm{size}(Z)=\sum_{t\in Z}\mathrm{size}(t)$.
 For any ground term $t\in T_{\Sigma}$ and position $\alpha\in\mathrm{POS}(t)$, we introduce the subterm $t|_{\alpha}\in T_{\Sigma}$ of $t$ at $\alpha$ as follows:
+  • 1.
     for $t\in\Sigma_{0}$, $t|_{\varepsilon}=t$;
+  • 2.
     for $t=\sigma(t_{1},\ldots,t_{m})$ with $m\geq 1$ and $\sigma\in\Sigma_{m}$ if $\alpha=\varepsilon$ then $t|_{\alpha}=t$, otherwise if $\alpha=i\beta$ with $1\leq i\leq m$, then $t|_{\alpha}=t_{i}|_{\beta}$.
@@
 The set of all $n$-contexts over $\Sigma$ is denoted by $CON_{\Sigma,n}$.
-For example, let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\$\,\}$, and $\Sigma_{2}=\{\,f,g\,\}$.
+For example, let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\mathdollar\,\}$, and $\Sigma_{2}=\{\,f,g\,\}$.
 Then $g(f(\diamondsuit,\#),g(\diamondsuit,\diamondsuit))\in CON_{\Sigma,3}$.
@@
 For example,
-$g(f(\diamondsuit,\#),f(\diamondsuit,\diamondsuit))[g(\$,\$),\#,\$]=g(f(g(\$,\$),\#),f(\#,\$))$.
+$)g(f(\diamondsuit,\#),f(\diamondsuit,\diamondsuit))[g(\mathdollar,\mathdollar),\#,\mathdollar]=g(f(g(\mathdollar,\mathdollar),\#),f(\#,\mathdollar$)).
 Let $\delta$ be a $1$-context over $\Sigma$, and let $\alpha\in POS(\delta)$ be such that $t|_{\alpha}=\diamondsuit$.
 Then we call $\alpha$ the address of $\diamondsuit$ in $\delta$, and $addr(\delta)$ stands for $\alpha$.
-Let $\delta$ be a $1$-context over $\Sigma$ such that $\delta\neq\diamondsuit$, and let $c\in\Sigma_{0}$ be a constant.
-Then we say that the ground term $\delta[c]$ is a proper extension of the constant $c\in\Sigma_{0}$.
+Let $\delta$ be a $1$-context over $\Sigma$ such that $\delta\neq\diamondsuit$, and let ${d}\in\Sigma_{0}$ be a constant.
+Then we say that the ground term $\delta[{d}]$ is a proper extension of the constant ${d}\in\Sigma_{0}$.
 Ground term rewrite systems and equation systems [4]
 Let $\Sigma$ be a signature, and $R$ be a finite subset of $T_{\Sigma}\times T_{\Sigma}$.
-The size $size(R)$ of $R$ is the number of occurrences of symbols of $\Sigma$ in $R$, ie., $size(R)=\sum_{(l,r)\in R}size(l)+size(r)$.
+The size of $R$ is denoted by $\mathrm{size}(R)$, and is the number of occurrences of symbols of $\Sigma$ in $R$, ie., $\mathrm{size}(R)=\sum_{(l,r)\in R}\mathrm{size}(l)+\mathrm{size}(r)$.
 Let $\Sigma\langle R\rangle$ denote the signature of those symbols which occur in $R$.
@@
 Moreover, we denote a GTES as $E$ rather than $R$ [4].
-Statement 2.4.
-Let $E$ and $F$ be GTESs over a signature $\Sigma$. We can construct the GTES $E\cup F$ in $O(n\,\mathrm{log}\,n)$ time.
+Let $E$ and $F$ be GTESs over a signature $\Sigma$.
+Then we define the GTES $E\cup F$ over $\Sigma$ as the union of the sets $E$ and $F$.
+Statement 2.5.
+Let $E$ and $F$ be GTESs over a signature $\Sigma$. We can construct the GTES $E\cup F$ over $\Sigma$ in $O(n\,\mathrm{log}\,n)$ time.
 Proof.
-In $O(n\,\mathrm{log}\,n)$ time we mergesort the GTES $E$ and the GTES $F$ [10, Section 2.2].
-Then in $O(n)$ time, scan through each of the sorted relations to produce the result.
+In $O(n\,\mathrm{log}\,n)$ time we mergesort the GTES $E$ and the GTES $F$ [16, Section 2.2].
+Then in $O(n)$ time, we concurrently scan through both of the sorted relations to produce the result.
 When the concurrent scan of both GTESs reveals the same equation in both relations, only one of the equations is retained.
@@
 The Boolean value $FLAG\langle E,F\rangle$ indicates whether $\Sigma=\Sigma\langle E,F\rangle$.
-Statement 2.6.
-Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=size(E)+size(F)$.
+Statement 2.7.
+Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=\mathrm{size}(E)+\mathrm{size}(F)$.
 We can count the elements of $\Sigma\langle E,F\rangle$ and compute the Boolean value $FLAG\langle E,F\rangle$ in $O(n)$ time.
@@
 $ST\langle E\rangle=\{\,t\in T_{\Sigma}\mid t\in sub(u)\mbox{ or }t\in sub(v)\mbox{ for some }u\doteq v\mbox{ in }E\,\}$
-is the set of subterms occurring $E$.
-Statement 2.7.
-1. Let $Z\subseteq T_{\Sigma}$ be a finite set of ground terms over a signature $\Sigma$. Then $|sub(Z)|\leq size(Z)$.
-2. Let $E$ be a GTES over a signature $\Sigma$. Then $|ST\langle E\rangle|\leq size(E)$.
-Proof.
-By definition. ∎
+is the set of subterms occurring $E$. The below claim is immediate.
+Statement 2.8.
+1. Let $Z\subseteq T_{\Sigma}$ be a finite set of ground terms over a signature $\Sigma$. Then $|sub(Z)|\leq\mathrm{size}(Z)$.
+2. Let $E$ be a GTES over a signature $\Sigma$. Then $|ST\langle E\rangle|\leq\mathrm{size}(E)$.
 Let $\Sigma$ be a signature.
-A $\Sigma$-algebra ${\cal A}$ consists of a carrier set $A$, and a mapping that associates with each symbol $\sigma\in\Sigma_{m}$ a mapping $\sigma^{\cal A}:A^{m}\rightarrow A$ for every $m\in{\mathbb{N}}$ [9].
-The $\Sigma$-ground-term algebra ${\cal T}(\Sigma)$ has $T_{\Sigma}$ as carrier set, and interprets the symbols $\sigma\in\Sigma_{m}$ for every $m\in{\mathbb{N}}$ as follows:
-$\sigma^{{\cal T}(\Sigma)}:T_{\Sigma}^{m}\rightarrow T_{\Sigma}:(t_{1},\ldots,t_{m})\mapsto\sigma(t_{1},\ldots,t_{m})$.
+A $\Sigma$-algebra $\cal A$} consists of a carrier set $A$, and a mapping that associates with each symbol $\sigma\in\Sigma_{k}$ a mapping $\sigma^{\cal A}:A^{k}\rightarrow A$ for every $k\in{\mathbb{N}}$ [15, Definition 3.2.1].
+The $\Sigma$-ground-term algebra ${\cal T}(\Sigma)$ has $T_{\Sigma}$ as carrier set, and interprets the symbols $\sigma\in\Sigma_{k}$ for every $k\in{\mathbb{N}}$ as follows:
+$\sigma^{{\cal T}(\Sigma)}:T_{\Sigma}^{k}\rightarrow T_{\Sigma}:(t_{1},\ldots,t_{k})\mapsto\sigma(t_{1},\ldots,t_{k})$.
 Let $E$ be a GTES over $\Sigma$.
@@
 It is well known that $\mathop{\leftrightarrow}^{*}_{E}$ is the least congruence on ${\cal T}(\Sigma)$ containing $E$.
-Let $R$ be a GTRS $R$ over $\Sigma$.
+Let $R$ be a GTRS over $\Sigma$.
 We say that GTRS $R$ is equivalent to the GTES $E$ if $\mathop{\leftrightarrow}^{*}_{R}=\mathop{\leftrightarrow}^{*}_{E}$.
 We say that GTRS $R$ reaches a ground term $s\in T_{\Sigma}$ from a ground term $t\in T_{\Sigma}$ if $t\mathop{\rightarrow}_{R}^{*}s$ holds.
-We say that GTRS $R$ is Noetherian, (confluent, etc.) if $\mathop{\rightarrow}_{R}$ is Noetherian (confluent, etc.).
-A term $t\in T_{\Sigma}$ is irreducible with respect to $R$ if it is irreducible with respect to $\mathop{\rightarrow}_{R}$.
+We say that GTRS $R$ is terminating, (confluent, etc.) if $\mathop{\rightarrow}_{R}$ is terminating (confluent, etc.).
+A term $t\in T_{\Sigma}$ is irreducible for $R$ if it is irreducible for $\mathop{\rightarrow}_{R}$.
 If a term $t\in T_{\Sigma}$ has a unique $\mathop{\rightarrow}_{R}$ normal form, then the latter is denoted by $t{\downarrow}_{R}$.
-A GTRS $R$ is reduced if for every rewrite rule $l\dot{\rightarrow}r$ in $R$, $l$ is irreducible with respect to $R-\{\,l\dot{\rightarrow}r\,\}$ and $r$ is irreducible for $R$.
+A GTRS $R$ is reduced if for every rewrite rule $l\dot{\rightarrow}r$ in $R$, $l$ is irreducible for $R-\{\,l\dot{\rightarrow}r\,\}$ and $r$ is irreducible for $R$.
 We recall the following important result.
-Proposition 2.8.
-[3] Every reduced GTRS $R$ is convergent.
+Proposition 2.9 ([4, Theorem 2.14]).
+Every reduced GTRS $R$ is convergent.
 Definition 2.10.
@@
 Bottom-up tree automata [17], [18]
-Let $\Sigma$ be a signature, a bottom-up tree automaton (bta for short) over $\Sigma$ is a quadruple ${\cal A}=(\Sigma,A,R,A_{f})$, where $A$ is a finite set of states of arity 00, $\Sigma\cap A=\emptyset$, $A_{f}(\subseteq A)$ is the set of final states, $R$ is a finite set of rewrite rules of the following form:
-$\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a$ with $m\in{\mathbb{N}}$, $\sigma\in\Sigma_{m}$, $a_{1},\ldots,a_{m},a\in A$.
+Let $\Sigma$ be a signature, a bottom-up tree automaton (bta for short) over $\Sigma$ is a quadruple ${\cal A}=(\Sigma,A,R,A_{f})$, where $A$ is a finite set of states of arity 0, $\Sigma\cap A=\emptyset$, $A_{f}(\subseteq A)$ is the set of final states, $R$ is a finite set of rewrite rules of the following form:
+$\sigma({b}_{1},\ldots,{b}_{k})\dot{\rightarrow}{b}$ with $k\in{\mathbb{N}}$, $\sigma\in\Sigma_{k}$, $b_{1},\ldots,b_{k},{b}\in A$.
 We consider $R$ as a GTRS over $\Sigma\cup A$.
-The size $size({\cal A})$ of $\cal A$ is $|A|+size(R)$.
-The tree language recognized by $\cal A$ is $L({\cal A})=\{\,t\in T_{\Sigma}\mid(\exists a\in A_{f})\;t\mathop{\rightarrow}_{R}^{*}a\}$.
-We say that ${\cal A}=(\Sigma,A,R,A_{f})$ is a deterministic bta if for any $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, $a_{1},\ldots,a_{m}\in A$, there is at most one rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R$.
+The size of $\cal A$ is denoted by $\mathrm{size}({\cal A})$, and is $|A|+\mathrm{size}(R)$.
+The tree language recognized by $\cal A$ is $L({\cal A})=\{\,t\in T_{\Sigma}\mid(\exists{b}\in A_{f})\;t\mathop{\rightarrow}_{R}^{*}{b}\}$.
+We say that ${\cal A}=(\Sigma,A,R,A_{f})$ is a deterministic bta if for any $\sigma\in\Sigma_{k}$, $k\in{\mathbb{N}}$, ${b}_{1},\ldots,{b}_{k}\in A$, there is at most one rule with left-hand side $\sigma({b}_{1},\ldots,{b}_{k})$ in $R$.
 Proposition 2.11.
-[13] Let $E$ and $F$ be GTESs over a signature $\Sigma$, and let $W=ST(E\cup F)$.
+[7] Let $E$ and $F$ be GTESs over a signature $\Sigma$, and let $W=ST\langle E\cup F\rangle$.
 For every $p\in W$, we can construct in $O(n\,\mathrm{log}\,n)$ time the deterministic bottom-up tree automaton ${\cal A}=(\Sigma,C\langle E,F\rangle,R\langle E,F\rangle,\{\,p/_{\Theta\langle E,F\rangle}\,\})$, and $\cal A$} recognizes the tree language $p/_{\mathop{\leftrightarrow}^{*}_{E}}$.
 Proposition 2.12.
-[5] Given two btas ${\cal A}=(\Sigma,A,R_{\cal A},A_{f})$ and ${\cal B}=(\Sigma,B,R_{\cal B},B_{f})$, we can decide whether $L({\cal A})\subseteq L({\cal B})$ in $O(size({\cal A})\cdot size({\cal B}))$ time.
+[10] Given two btas ${\cal A}=(\Sigma,A,R_{\cal A},A_{f})$ and ${\cal B}=(\Sigma,B,R_{\cal B},B_{f})$, we can decide whether $L({\cal A})\subseteq L({\cal B})$ in $O(\mathrm{size}({\cal A})\cdot\mathrm{size}({\cal B}))$ time.
 Directed pseudographs without parallel arcs [19]
+A directed pseudograph is a type of directed graph that allows for both parallel arcs (i.e.
+pairs of directed arcs with the same tail and the same head) and loops (i.e, directed arcs whose head and tail coincide) [19, top of page 4].
 A directed pseudograph without parallel arcs (dpwpa for short) $G$ is an ordered pair $(V,A)$, where $V$ is a finite set, and $A\subseteq V\times V$ is a relation on $V$.
@@
 We say that $w$ is from vertex $x_{1}$ to vertex $x_{k}$.
-The length of $w$ is the number $k-1$ of its arcs.
+The length of $w$ is the number k1k-1 of its arcs [19, page 12].
+A single vertex $x_{1}$ is considered as a walk from $x_{1}$ to $x_{1}$ of length 0.
 A trail is a walk $w$ in which all arcs are distinct.
@@
 We say that a vertex $x$ is reachable from a vertex $y$ if there exists a walk $w$ from vertex $y$ to vertex $x$.
-In particular, a vertex is reachable from itself.
+By definition, every vertex is reachable from itself.
 We say that a vertex $x$ is positive-step reachable from a vertex $y$ if there exists a walk $w$ of positive length from vertex $y$ to vertex $x$.
@@
 Congruence closure on subterm dags [4]
-Let $\Sigma$ be a signature.
-A subterm dag over $\Sigma$ is a directed, labeled and acyclic multigraph $G=(V,A,\lambda)$, where $V$ is the set of vertices, $A\subseteq V\times V$ is the multiset of directed arcs and $\lambda:V\rightarrow\Sigma$ is the labeling function such that if $outdegree(v)=m$ for some vertex $v\in V$, then $\lambda(v)\in\Sigma_{m}$.
-The multiset of directed arcs leaving a vertex is ordered.
-We denote by $v/i$ the $i$th successor of a vertex $v$.
-The vertices of $G$ can be interpreted as ground terms over $\Sigma$ by the function ^$\widehat{\;}:V\rightarrow W$ which is defined by recursion: for a vertex $v\in V$, we have $\widehat{v}=\lambda(v)$ if $outdegree(v)=0$ and we have $\widehat{v}=\sigma(\widehat{v/1},\ldots,\widehat{v/m})$ if $outdegree(v)=m$ for some $m>0$ and $\lambda(v)=\sigma$.
-Let $E$ be a GTES over $\Sigma$.
-We say that $G$ is a subterm dag for a GTES $E$ if   $\widehat{\;}$  is a bijection between $V$ and $ST(E)$.
-Let $E$ and $F$ be GTESs over $\Sigma$.
-We say that $G$ is a subterm dag for the GTES $E$ and GTES $F$ if   $\widehat{\;}$  is a bijection between $V$ and $ST(E)\cup ST(F)$.
+[…]
 Definition 2.13.
@@
 We now cite Proposition 3.4 in [4].
-Proposition 2.13.
-[3] Let $E$ be a GTES over a signature $\Sigma$, and $n=size(E)$. We can construct in $O(n)$ time a subterm dag for the GTES $E$ and the relation $\tau\langle E\rangle$.
+Proposition 2.14 ([4, Proposition 3.4]).
+Let $E$ be a GTES over a signature $\Sigma$, and $n=\mathrm{size}(E)$. We can construct in $O(n)$ time a subterm dag for the GTES $E$ and the relation $\tau\langle E\rangle$.
 By Proposition 2.14, we have the following result.
-Statement 2.14.
-Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=size(E)+size(F)$. We can construct in $O(n)$ time a subterm dag for the GTESs $E$ and $F$.
+Statement 2.15.
+Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=\mathrm{size}(E)+\mathrm{size}(F)$. We can construct in $O(n)$ time a subterm dag for the GTESs $E$ and $F$.
 Definition 2.16.
@@
 [9] Let $E$ be a GTES over $\Sigma$.
-There exists an algorithm which, given a subterm dag $G=(V,A,\lambda)$ for $E$, and a relation $\tau$ on $V$, produces the congruence closure $\rho$ of $\tau$ on $V$ in $O(n\,\mathrm{log}\,n)$ time, where $n=size(E)$.
+There exists an algorithm which, given a subterm dag $G=(V,A,\lambda)$ for $E$, and a relation $\tau$ on $V$, produces the congruence closure $\rho$ of $\tau$ on $V$ in $O(n\,\mathrm{log}\,n)$ time, where $n=\mathrm{size}(E)$.
 The algorithm gives each $\rho$-class a unique name in the set { $\{\,1,\ldots,|V/\rho|\,\}$, and also outputs an array of the cardinalities of all $\rho$-classes.
@@
 A possible implementation of such a function was described in [3].
-3 Running Examples
-We now start our eight running examples. Throughout this section, $E$ and $F$ are GTESs over a signature $\Sigma$, and $W$ stands for $ST\langle E\cup F\rangle$.
+3 Running Examples for the Main Cases
+Along the proof of our main result, we shall distinguish four main cases.
+We define these four main cases, and then we present eight running examples to illustrate our concepts and results.
+In each main case, a running example presents GTESs $E$ and $F$ such that $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ and a running example presents $E$ and $F$ such that $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
+From now on throughout the paper, $E$ and $F$ are GTESs over a signature $\Sigma$, and $W$ stands for $ST\langle E\cup F\rangle$.
+  • Main Case 1:
+    $\Sigma$ is a unary signature.
+Example 3.18 and its sequels, Examples 5.51, 7.72, 10.105.
+and 12.119, and 12.124, also Example 3.19 and its sequels, 5.52, 7.73, 10.106, 12.120, and 12.125 illustrate this case.
+  • Main Case 2:
+    Both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total.
+Example 3.20 and its sequels Examples 5.53, 13.127, further Example 3.21 and its sequels, Examples 5.54, 13.128, exemplify this case.
+  • Main Case 3:
+    $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total.
+Example 3.22 and its sequels, Examples 5.55, 7.74, 10.110, 14.135, and 14.142, moreover Example 3.23, and its sequels, Examples 5.56, 7.75, 10.111, 14.136, and 14.143 present this case.
+  • Main Case 4:
+    $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total.
+Example 3.24 and its sequel Example 5.57, furthermore, Example 3.25 and its sequels Example 5.58 and Example 15.144 demonstrate this case.
+Main Case 2 overlaps with Main Case 1 and Main Case 3, and is a subcase of the union of Main Case 1 and Main Case 3.
+For each running example, we compute the sets $ST\langle E\rangle$, $ST\langle F\rangle$, $W$ and the relations $\mathop{\leftrightarrow}^{*}_{E}$, $\mathop{\leftrightarrow}^{*}_{F}$, $\mathop{\leftrightarrow}^{*}_{E\cup F}$, $\Theta\langle E,F\rangle$, $\Theta\langle F,E\rangle$, $\Theta\langle E\cup F,\emptyset\rangle$, and state whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ or $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
+Recall that $E$ and $F$ are GTESs over a signature $\Sigma$, and $W$ stands for $ST\langle E\cup F\rangle$.
 Example 3.18.
-Let $\Sigma=\Sigma_{0}\cup\Sigma_{1}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{1}=\{\,f,\;g\,\}$, $E=\{\,f(\#)\doteq g(\#)\,\}$, and $F=\{\,f(\$)\doteq g(\$)\,\}$. Observe the following equations:
+[…]
 $\mathop{\leftrightarrow}^{*}_{E}=\{\,(\delta(s),\delta(t))\mid\delta\in CON_{\Sigma,1},s,t\in\{\,f(\#),\;g(\#)\,\}\,\}$,
-$\mathop{\leftrightarrow}^{*}_{F}=\{\,(\delta(s),\delta(t))\mid\delta\in CON_{\Sigma,1},s,t\in\{\,f(\$),\;g(\$)\,\}\,\}$,
+[…]
 $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$,
-$ST\langle E\rangle=\{\,\#,\;f(\#),\;g(\#)\,\}$,   $ST\langle F\rangle=\{\,\$,\;f(\$),\;g(\$)\,\}$, and $W=\{\,\#,\;\$,\;f(\#),\;f(\$),\;g(\#),\;g(\$)\,\}$.
+[…]
 […]
@@
 Example 3.19.
-Let $\Sigma=\Sigma_{0}\cup\Sigma_{1}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{1}=\{\,f\,\}$, $E=\{\,f(f(\#))\doteq\#\,\}$, and $F=\{\,f(f(f(\#)))\doteq\#\,\}$.
+Let $\Sigma=\Sigma_{0}\cup\Sigma_{1}$, $\Sigma_{0}=\{\,\#,\;\mathdollar\,\}$, $\Sigma_{1}=\{\,f\,\}$, $E=\{\,f(f(\#))\doteq\#\,\}$, and $F=\{\,f(f(f(\#)))\doteq\#\,\}$.
 For every $n\in{\mathbb{N}}$, we define $f^{n}(\#)$ as follows:
@@
 Observe the following equations:
-$\mathop{\leftrightarrow}^{*}_{E}=\{\,f^{m}(\#),f^{n}(\#))\mid m,n\in{\mathbb{N}},\;m-n\mbox{ is divisible by }2\,\}$,
-$\mathop{\leftrightarrow}^{*}_{F}=\{\,f^{m}(\#),f^{n}(\#))\mid m,n\in{\mathbb{N}},\;m-n\mbox{ is divisible by }3\,\}$,
+$\mathop{\leftrightarrow}^{*}_{E}=\{\,(f^{m}(\#),f^{n}(\#))\mid m,n\in{\mathbb{N}},\;m-n\mbox{ is divisible by }2\,\}$,
+$\mathop{\leftrightarrow}^{*}_{F}=\{\,(f^{m}(\#),f^{n}(\#))\mid m,n\in{\mathbb{N}},\;m-n\mbox{ is divisible by }3\,\}$,
 $\mathop{\leftrightarrow}^{*}_{E\cup F}=T_{\Sigma}\times T_{\Sigma}$,
@@
 $ST\langle E\rangle=\{\,\#,\;f(\#),\;f^{2}(\#)\,\}$,  $ST\langle F\rangle=\{\,\#,\;f(\#),\;f^{2}(\#),\;f^{3}(\#)\,\}$, and $W=ST\langle F\rangle$.
-The equivalence classes of $\Theta\langle E,F\rangle$ are {#$\{\,\#,\;f^{2}(\#)\}$, $\{\,f(\#),\;f^{3}(\#)\,\}$.
-The equivalence classes of $\Theta\langle F,E\rangle$ are {#$\{\,\#,\;f^{3}(\#)\}$, $\{\,f(\#)\,\}$,   $\{\,f^{2}(\#)\,\}$.
-The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is {#$\{\,\#,\;f(\#),\;f^{2}(\#),\;f^{3}(\#)\,\}$.
+The equivalence classes of $\Theta\langle E,F\rangle$ are $\{\,\#,\;f^{2}(\#)\}$, $\{\,f(\#),\;f^{3}(\#)\,\}$.
+The equivalence classes of $\Theta\langle F,E\rangle$ are $\{\,\#,\;f^{3}(\#)\}$, $\{\,f(\#)\,\}$,   $\{\,f^{2}(\#)\,\}$.
+The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is $\{\,\#,\;f(\#),\;f^{2}(\#),\;f^{3}(\#)\,\}$.
 We note that
-#$\#/_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and #$\#/_{\Theta\langle F,E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
-Consequently,
-for each $t\in W$, $\left(t/_{\Theta\langle{E\cup F},\emptyset\rangle}\neq t/_{\Theta\langle E,F\rangle}\mbox{ and }t/_{\Theta\langle{E\cup F},\emptyset\rangle}\neq t/_{\Theta\langle{F},E\rangle}\right)$.
+for each $t\in W$, $\left(t/_{\Theta\langle E,F\rangle}\subset t/_{\Theta\langle{E\cup F},\emptyset\rangle}\mbox{ and }t/_{\Theta\langle{F},E\rangle}\subset t/_{\Theta\langle{E\cup F},\emptyset\rangle}\right)$.
 Example 3.20.
-Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\,\$,\;\mathsterling\,\}$, $\Sigma_{2}=\{\,f\,\}$.
-Intuitively, we now define GTESs $E$ and $F$ such that Statements 1 – 4 hold.
+[…]
 1. $\mathop{\leftrightarrow}^{*}_{E}$ has four congruence classes:
+  • 1.
+    The singleton $\{\,\#\,\}$. Because there is no equation in $E$ with left-hand side or right-hand side #\#. We consider #\# as the representative of this class.
+  • 2.
+    The set of all ground terms which are different from the constant #\#, and contain only the symbols $f$ and #\#.
+Applying the first three equations, $E$ rewrites each ground term $t$ of height $2$ in the above set to f(#,#)f(\#,\#).
+Hence for each ground term $t$ in this set, we have $t\mathop{\leftrightarrow}^{*}_{E}f(\#,\#)$.
+We consider f(#,#)f(\#,\#) as the representative of this class.
+  • 3.
 […]
-  • the set of all ground terms which are different from the constant ##\##, and contain only the symbols $f$ and ##\##,
-  • the set of all ground terms containing only the symbols $f$ and \$currency-dollar\\$\$, and
+  • 4.
 […]
-2. We obtain GTES $F$ from GTES $E$ by swapping the constants ##\## and \$currency-dollar\\$\$.
+[…]
 3. $\mathop{\leftrightarrow}^{*}_{F}$ has four congruence classes:
+  • 1.
 […]
-  • the set of all ground terms which are different from the constant \$currency-dollar\\$\$, and contain only the symbols $f$ and \$currency-dollar\\$\$,
-  • the set of all ground terms containing only the symbols $f$ and ##\##, and
+  • 2.
+[…]
+  • 3.
+    the set of all ground terms containing only the symbols $f$ and #\#, and
+  • 4.
 […]
 4. $\mathop{\leftrightarrow}^{*}_{E\cup F}$ has three congruence classes:
-  • the set of all ground terms which contain only the symbols $f$ and \$currency-dollar\\$\$,
-  • the set of all ground terms containing only the symbols $f$ and ##\##, and
+  • 1.
+[…]
+  • 2.
+    the set of all ground terms containing only the symbols $f$ and #\#, and
+  • 3.
 […]
 Let
-$E=\{\,f(\#,f(\#,\#))\doteq f(\#,\#),\;f(f(\#,\#),\#)\doteq f(\#,\#),\;f(f(\#,\#),f(\#,\#))\doteq f(\#,\#),\;\newline f(\$,\$)\doteq\$,\;\newline f(\#,\$)\doteq\mathsterling,\;f(\#,\mathsterling)\doteq\mathsterling,\;\newline f(\$,\#)\doteq\mathsterling,\;f(\$,\mathsterling)\doteq\mathsterling,\;\newline f(\mathsterling,\#)\doteq\mathsterling,\;f(\mathsterling,\$)\doteq\mathsterling,\;f(\mathsterling,\mathsterling)\doteq\mathsterling\newline f(\$,f(\#,\#))\doteq\mathsterling,\;f(\mathsterling,f(\#,\#))\doteq\mathsterling,\;\newline f(f(\#,\#),\$)\doteq\mathsterling,\;f(f(\#,\#),\mathsterling)\doteq\mathsterling\,\}$
-$F=\{\,f(\$,f(\$,\$))\doteq f(\$,\$),\;f(f(\$,\$),\$)\doteq f(\$,\$),\;f(f(\$,\$),f(\$,\$))\doteq f(\$,\$),\;\newline f(\#,\#)\doteq\#,\;\newline f(\$,\#)\doteq\mathsterling,\;f(\$,\mathsterling)\doteq\mathsterling,\;\newline f(\#,\$)\doteq\mathsterling,\;f(\#,\mathsterling)\doteq\mathsterling,\;\newline f(\mathsterling,\#)\doteq\mathsterling,\;f(\mathsterling,\$)\doteq\mathsterling,\;f(\mathsterling,\mathsterling)\doteq\mathsterling\newline f(\#,f(\$,\$))\doteq\mathsterling,\;f(\mathsterling,f(\$,\$))\doteq\mathsterling,\;\newline f(f(\$,\$),\#)\doteq\mathsterling,\;f(f(\$,\$),\mathsterling)\doteq\mathsterling\,\}$.
-GTES $E$ is symmetric to GTES $F$. Note the following equations:
-  • $ST\langle E\rangle=\{\,\#,\;\$,\;\mathsterling,\;\newline f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;\newline f(f(\#,\#),f(\#,\#)),\;f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;\newline f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling)\,\}$,
-  • $ST\langle F\rangle=\{\,\$,\;\#,\;\mathsterling,\;\newline f(\$,\$),\;f(\$,\#),\;f(\$,\mathsterling),\;\newline f(\#,\$),\;f(\#,\#),\;f(\#,\mathsterling),\;\newline f(\mathsterling,\$),\;f(\mathsterling,\#),\;f(\mathsterling,\mathsterling),\;\newline f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$)),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;\newline f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}$,
-  • $W=\{\,\#,\;\$,\;\mathsterling,\;\newline f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;\newline f(f(\#,\#),f(\#,\#)),\;f(\$,f(\$,\$)),\;\newline f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$)),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}$.
-The equivalence classes of $\Theta\langle E,F\rangle$ are
-{#},{\$$\{\,\#\,\},\;\newline \{\,\$,\;f(\$,\$)\,\},\;\newline \{\,f(\#,\#),\;f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#))\,\},\newline \;\{\,\mathsterling,\;f(\#,\$),\;f(\#,\mathsterling),\;f(\$,\#),\;f(\$,\mathsterling),\;f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}$.
-The equivalence classes of $\Theta\langle F,E\rangle$ are
-{#$\{\,\#,\;f(\#,\#)\,\},\;\newline \{\,\$\,\},\newline \{\,f(\$,\$),\;f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$))\},\;\newline \{\,\mathsterling,\;f(\#,\$),\;f(\#,\mathsterling),\;f(\$,\#),\;f(\$,\mathsterling),\;f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}$.
-The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is $W$.
-$\mathop{\leftrightarrow}^{*}_{E}=\{\,(\#,\#)\,\}\cup(T_{\{\,\#,f\,\}}\setminus\{\,\#\,\})\times(T_{\{\,\#,f\,\}}\setminus\{\,\#\,\})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\$,f\,\}})\right)\times\left(T_{\Sigma}\setminus(T_{\{\,\#,f\,\}})\cup T_{\{\,\$,f\,\}})\right)\cup\newline (T_{\{\,\$,f\,\}}\times T_{\{\,\$,f\,\}})$.
+$E=\{\,f(\#,f(\#,\#))\doteq f(\#,\#),\;f(f(\#,\#),\#)\doteq f(\#,\#),\;f(f(\#,\#),f(\#,\#))\doteq f(\#,\#),\;\newline f(\mathdollar,\mathdollar)\doteq\mathdollar,\;\newline f(\#,\mathdollar)\doteq\mathsterling,\;f(\#,\mathsterling)\doteq\mathsterling,\;\newline f(\mathdollar,\#)\doteq\mathsterling,\;f(\mathdollar,\mathsterling)\doteq\mathsterling,\;\newline f(\mathsterling,\#)\doteq\mathsterling,\;f(\mathsterling,\mathdollar)\doteq\mathsterling,\;f(\mathsterling,\mathsterling)\doteq\mathsterling,\newline f(\mathdollar,f(\#,\#))\doteq\mathsterling,\;f(\mathsterling,f(\#,\#))\doteq\mathsterling,\;\newline f(f(\#,\#),\mathdollar)\doteq\mathsterling,\;f(f(\#,\#),\mathsterling)\doteq\mathsterling\,\}$,
+[…]
+Note the following equations:
+  • 1.
+    $ST\langle E\rangle=\{\,\#,\;\mathdollar,\;\mathsterling,\;\newline f(\#,\#),\;f(\#,\mathdollar),\;f(\#,\mathsterling),\;\newline f(\mathdollar,\#),\;f(\mathdollar,\mathdollar),\;f(\mathdollar,\mathsterling),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\mathdollar),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;\newline f(f(\#,\#),f(\#,\#)),\;f(\mathdollar,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;\newline f(f(\#,\#),\mathdollar),\;f(f(\#,\#),\mathsterling)\,\}$,
+  • 2.
+[…]
+  • 3.
+[…]
+$\mathop{\leftrightarrow}^{*}_{E}=\{\,(\#,\#)\,\}\cup(T_{\{\,\#,f\,\}}\setminus\{\,\#\,\})\times(T_{\{\,\#,f\,\}}\setminus\{\,\#\,\})\cup(T_{\{\,\mathdollar,f\,\}}\times T_{\{\,\mathdollar,f\,\}})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\mathdollar,f\,\}})\right)\times\left(T_{\Sigma}\setminus(T_{\{\,\#,f\,\}})\cup T_{\{\,\mathdollar,f\,\}})\right)$.
+$\mathop{\leftrightarrow}^{*}_{E}$ has the congruence classes $\{\,\#\,\}$,   $T_{\{\,\#,f\,\}}\setminus\{\,\#\,\}$,   $T_{\{\,\mathdollar,f\,\}}$,   and $T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\mathdollar,f\,\}})$.
+Then
+  • 1.
+    $T_{\{\,\#,f\,\}}$ is the set of all ground terms containing only the symbols $f$ and #\#.
+  • 2.
+[…]
+  • 3.
+[…]
+  • 4.
 […]
 […]
-$\mathop{\leftrightarrow}^{*}_{F}=(T_{\{\,\#,f\,\}})\times T_{\{\,\#,f\,\}})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\$,f\,\}}\cup T_{\{\,\#,f\,\}}\right)\times\left(T_{\Sigma}\setminus(T_{\{\,\$,f\,\}}\cup T_{\{\,\#,f\,\}}))\right)\cup\newline \{\,(\$,\$)\,\}\cup\left(T_{\{\,\$,f\,\}}\setminus\{\,\$\,\}\right)\times\left(T_{\{\,\$,f\,\}}\setminus\{\,\$\,\}\right)$.
 […]
-$\mathop{\leftrightarrow}^{*}_{E\cup F}=(T_{\{\,\#,f\,\}})\times T_{\{\,\#,f\,\}})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\$,f\,\}}\cup T_{\{\,\#,f\,\}})\right)\times\left((T_{\Sigma}\setminus(T_{\{\,\$,f\,\}}\cup T_{\{\,\#,f\,\}})\right)\cup\newline (T_{\{\,\$,f\,\}}\times T_{\{\,\$,f\,\}})$.
-$\mathop{\leftrightarrow}^{*}_{E\cup F}$ has the congruence classes $T_{\{\,\#,f\,\}}$,   $T_{\{\,\$,f\,\}}$,   and $T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\$,f\,\}})$.
+$\mathop{\leftrightarrow}^{*}_{E\cup F}=(T_{\{\,\#,f\,\}})\times T_{\{\,\#,f\,\}})\cup(T_{\{\,\mathdollar,f\,\}}\times T_{\{\,\mathdollar,f\,\}})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\mathdollar,f\,\}})\right)\times\left((T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\mathdollar,f\,\}})\right)$.
+$\mathop{\leftrightarrow}^{*}_{E\cup F}$ has the congruence classes $T_{\{\,\#,f\,\}}$,   $T_{\{\,\mathdollar,f\,\}}$,   and $T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\mathdollar,f\,\}})$.
 Observe that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$ and $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}=\mathop{\leftrightarrow}^{*}_{E\cup F}$.
+The equivalence classes of $\Theta\langle E,F\rangle$ are
+[…]
+The equivalence classes of $\Theta\langle F,E\rangle$ are
+[…]
+The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is $W$.
 Example 3.21.
 […]
-Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\$,\;\mathsterling,\;\flat\,\}$, $\Sigma_{2}=\{\,f\,\}$,
-$E=\{\,\#\doteq\$,\;\#\doteq\mathsterling,\;\#\doteq f(\#,\#),\;\#\doteq f(\#,\flat),\;\#\doteq f(\flat,\#)\,\}$,
+[…]
+$E=\{\,\#\doteq\mathdollar,\;\#\doteq\mathsterling,\;\#\doteq f(\#,\#),\;\#\doteq f(\#,\flat),\;\#\doteq f(\flat,\#)\,\}$,
 and
-$F=\{\,\mathsterling\doteq\flat,\newline \mathsterling\doteq f(\#,\#),\;\mathsterling\doteq f(\#,\$),\;\mathsterling\doteq f(\#,\mathsterling),\;\newline \mathsterling\doteq f(\$,\#),\;\mathsterling\doteq f(\$,\$),\;\mathsterling\doteq f(\$,\mathsterling),\;\newline \mathsterling\doteq f(\mathsterling,\#),\;\mathsterling\doteq f(\mathsterling,\$),\;\mathsterling\doteq f(\mathsterling,\mathsterling)\,\}$.
+$F=\{\,\mathsterling\doteq\flat,\newline \mathsterling\doteq f(\#,\#),\;\mathsterling\doteq f(\#,\mathdollar),\;\mathsterling\doteq f(\#,\mathsterling),\;\newline \mathsterling\doteq f(\mathdollar,\#),\;\mathsterling\doteq f(\mathdollar,\mathdollar),\;\mathsterling\doteq f(\mathdollar,\mathsterling),\;\newline \mathsterling\doteq f(\mathsterling,\#),\;\mathsterling\doteq f(\mathsterling,\mathdollar),\;\mathsterling\doteq f(\mathsterling,\mathsterling)\,\}$.
 We note that for each ground term $t\in T_{\Sigma}$ with $height(t)\geq 1$,
@@
 $\mathop{\leftrightarrow}^{*}_{E}=\{\,(\flat,\flat)\,\}\cup((T_{\Sigma}\setminus\{\,\flat\,\})\times(T_{\Sigma}\setminus\{\,\flat\,\}))$,
-$\mathop{\leftrightarrow}^{*}_{F}=\{\,(\#,\#),\;(\$,\$)\,\}\cup((T_{\Sigma}\setminus\{\,\#,\;\$\,\})\times(T_{\Sigma}\setminus\{\,\#,\;\$\,\}))$, and
+$)\mathop{\leftrightarrow}^{*}_{F}=\{\,(\#,\#),\;(\mathdollar,\mathdollar)\,\}\cup((T_{\Sigma}\setminus\{\,\#,\;\mathdollar\,\})\times(T_{\Sigma}\setminus\{\,\#,\;\mathdollar\,\}$)), and
 $\mathop{\leftrightarrow}^{*}_{E\cup F}=T_{\Sigma}\times T_{\Sigma}$.
-Note that (#$(\#,\flat)\not\in\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Consequently,
+Note that $(\#,\flat)\not\in\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Consequently,
 $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$.
 We notice the following equations:
-$W=\{\,\#,\;\$,\;\mathsterling,\;\flat,\;\newline f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;f(\$,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;f(\mathsterling,\flat),\;\newline f(\flat,\#),\;f(\flat,\$),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\}$.
+$W=\{\,\#,\;\mathdollar,\;\mathsterling,\;\flat,\;\newline f(\#,\#),\;f(\#,\mathdollar),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\mathdollar,\#),\;f(\mathdollar,\mathdollar),\;f(\mathdollar,\mathsterling),\;f(\mathdollar,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\mathdollar),\;f(\mathsterling,\mathsterling),\;f(\mathsterling,\flat),\;\newline f(\flat,\#),\;f(\flat,\mathdollar),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\}$.
 The equivalence classes of $\Theta\langle E,F\rangle$ are
-{#,\$,£$\{\,\#,\;\$,\;\mathsterling,\;\newline \,f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;f(\$,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\flat),f(\mathsterling,\mathsterling),\;\newline f(\flat,\#),\;f(\flat,\$),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\},\;\newline \{\,\flat\,\}$.
+$\{\,\#,\;\mathdollar,\;\mathsterling,\;\newline \,f(\#,\#),\;f(\#,\mathdollar),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\mathdollar,\#),\;f(\mathdollar,\mathdollar),\;f(\mathdollar,\mathsterling),\;f(\mathdollar,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\mathdollar),\;f(\mathsterling,\flat),f(\mathsterling,\mathsterling),\;\newline f(\flat,\#),\;f(\flat,\mathdollar),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\},\;\newline \{\,\flat\,\}$.
 The equivalence classes of $\Theta\langle F,E\rangle$ are
-{£$\{\,\mathsterling,\;\flat,\;\newline \,f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;f(\$,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\flat),f(\mathsterling,\mathsterling),\;\newline f(\flat,\#),\;f(\flat,\$),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\},\newline \{\,\#\,\},\{\,\$\,\}$.
+$\{\,\mathsterling,\;\flat,\;\newline \,f(\#,\#),\;f(\#,\mathdollar),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\mathdollar,\#),\;f(\mathdollar,\mathdollar),\;f(\mathdollar,\mathsterling),\;f(\mathdollar,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\mathdollar),\;f(\mathsterling,\flat),f(\mathsterling,\mathsterling),\;\newline f(\flat,\#),\;f(\flat,\mathdollar),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\},\newline \{\,\#\,\},\{\,\mathdollar\,\}$.
 The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is $W$. We note that
-#$\#/_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and #$\#/_{\Theta\langle{F},E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
+$_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and $_{\Theta\langle{F},E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
 Example 3.22.
-Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{2}=\{\,f\,\}$,
-$E=\{\,f(\#,\#)\doteq\$,\;f(\#,\$)\doteq\$,\;f(\$,\#)\doteq\$,\;f(\$,\$)\doteq\$\,\}$ and $F=\{\,f(\#,\#)\doteq\#\,\}$.
+Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\mathdollar\,\}$, $\Sigma_{2}=\{\,f\,\}$,
+[…]
 Note the following equations:
-$ST\langle E\rangle=\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$,  $ST\langle F\rangle=\{\,\#,\;f(\#,\#)\,\}$,  $W=ST\langle E\rangle$.
 […]
-The equivalence classes of $\Theta\langle F,E\rangle$ are {#$\{\,\#,\;f(\#,\#)\,\},\;\{\,\$\,\},\;\{\,f(\#,\$)\,\},\;\{\,f(\$,\#)\,\},\;\{\,f(\$,\$)\,\}$.
-The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is {#,\$$\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$.
+[…]
+[…]
+[…]
 Observe that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$.
-Since (#,\$$(\#,\$)\not\in\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, and (#,\$$(\#,\$)\in\mathop{\leftrightarrow}^{*}_{E\cup F}$, we have $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$.
+Since $(\#,\mathdollar)\not\in\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, and $(\#,\mathdollar)\in\mathop{\leftrightarrow}^{*}_{E\cup F}$, we have $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$.
 Example 3.23.
-Let $\Sigma=\Sigma_{0}\cup\Sigma_{1}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{1}=\{\,g\,\}$, $\Sigma_{2}=\{\,f\,\}$,
-$E=\{\,f(\#,\#)\doteq g(\#),\;f(\#,g(\#))\doteq g(\#),\;f(g(\#),\#)\doteq g(\#),\;f(g(\#),g(\#))\doteq g(\#),\;g(g(\#))\doteq g(\#),\;\newline f(\#,\$)\doteq\$,\;f(\$,\#)\doteq\$,\;f(\$,\$)\doteq\$,\;g(\$)\doteq\$\,\}$.
+[…]
+[…]
 $F=\{\,f(\#,\#)\doteq\#,\;g(\#)\doteq\#\,\}$.
-Intuitively, $F$ can simulate the first five equations of $E$, that is, the congruence generated by the first five equations of $E$ is a proper subset of $\mathop{\leftrightarrow}^{*}_{F}$.
-We need the constant \$currency-dollar\\$\$ and the last four equations of $E$ to ensure that $\mathop{\leftrightarrow}^{*}_{E}\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$.
-Note the following equations:
-  • $ST\langle E\rangle=\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\#),\;g(\$),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(g(\#))\,\}$,
+[…]
+  • 1.
+    $ST\langle E\rangle=\{\,\#,\;\mathdollar,\;f(\#,\#),\;f(\#,\mathdollar),\;f(\mathdollar,\#),\;f(\mathdollar,\mathdollar),\;g(\#),\;g(\mathdollar),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(g(\#))\,\}$,
+  • 2.
     $ST\langle F\rangle=\{\,\#,\;f(\#,\#),\;g(\#)\,\}$,
+  • 3.
     $W=ST\langle E\rangle$.
 The equivalence classes of $\Theta\langle E,F\rangle$ are
-{#},$\{\,\#\,\},\{\,f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\},\;\newline \{\,\$,\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\$),\;g(g(\$))\,\}$,
+[…]
 The equivalence classes of $\Theta\langle F,E\rangle$ are
-{#$\{\,\#,\;f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\}$,
-{\$},$\{\,\$\,\},\;\{\,f(\#,\$)\,\},\;\{\,f(\$,\#)\,\},\;\{\,f(\$,\$)\,\},\;\{\,g(\$)\,\},\;\{\,g(g(\$))\,\}$.
+$\{\,\#,\;f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\}$,
+[…]
 The equivalence classes of $\Theta\langle E\cup F,\emptyset\rangle$ are
-{#$\{\,\#,\;f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\},\;\newline \{\,\$,\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\$),\;g(g(\$))\,\}$.
+[…]
 Observe that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$.
-\$$\$/_{\mathop{\leftrightarrow}^{*}_{E}}=\$/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=\{\,t\in T_{\Sigma}\mid\mbox{ there exists }\alpha\in POS(t)\mbox{ such that }t|_{\alpha}=\$\,\}$.
-#$\#/_{\mathop{\leftrightarrow}^{*}_{F}}=\#/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=\{\,t\in T_{\Sigma}\mid\mbox{ there does not exist }\alpha\in POS(t)\mbox{ such that }t|_{\alpha}=\$\,\}$.
+$\mathdollar/_{\mathop{\leftrightarrow}^{*}_{E}}=\mathdollar/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=\{\,t\in T_{\Sigma}\mid\mbox{ there exists }\alpha\in POS(t)\mbox{ such that }t|_{\alpha}=\mathdollar\,\}$.
+[…]
 Accordingly $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
 Example 3.24.
-Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{2}=\{\,f\,\}$,
-$E=\{\,f(\#,\#)\doteq f(\#,\$),\;f(\#,\#)\doteq f(\$,\#),\;f(\#,\#)\doteq f(\$,\$)\,\}$ and $F=\{\,\#\doteq\$\,\}$.
+Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\mathdollar\,\}$, $\Sigma_{2}=\{\,f\,\}$,
+[…]
 Observe the following equations and inclusion:
-$ST\langle E\rangle=\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$,  $ST\langle F\rangle=\{\,\#,\$\,\}$,   $W=ST\langle E\rangle$,  $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$,   $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{F}$.
+[…]
 Since $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$, each equivalence class of $\Theta\langle E,F\rangle$ is a subset of some equivalence class of $\Theta\langle F,E\rangle$.
-The equivalence classes of $\Theta\langle E,F\rangle$ are {#},{\$},$\{\,\#\,\},\;\{\,\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$.
-The equivalence classes of $\Theta\langle F,E\rangle$ are {#,\$},$\{\,\#,\;\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$.
-The equivalence classes of $\Theta\langle E\cup F,\emptyset\rangle$ are {#,\$},$\{\,\#,\;\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$.
+[…]
+[…]
+[…]
 Example 3.25.
-Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\flat,\;\$,\;\mathsterling\,\}$, $\Sigma_{2}=\{\,f\,\}$,
-$E=\{\,\#\doteq\$\,\}$, and $F=\{\,\mathsterling\doteq\flat\,\}$.
-We remark that $f(\#,\mathsterling)\mathop{\leftrightarrow}^{*}_{E\cup F}f(\$,\flat)$,  $(f(\#,\mathsterling),f(\$,\flat))\not\in\mathop{\leftrightarrow}^{*}_{E}$,   and $(f(\#,\mathsterling),f(\$,\flat))\not\in\mathop{\leftrightarrow}^{*}_{F}$.
-Thence
+[…]
+$E=\{\,\#\doteq\mathdollar\,\}$, and $F=\{\,\mathsterling\doteq\flat\,\}$.
+[…]
 $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$.
 Observe the following equations:
-$ST\langle E\rangle=\{\,\#,\;\$\,\}$,   $ST\langle F\rangle=\{\,\mathsterling,\;\flat\,\}$,   $W=\{\,\#,\;\$,\;\mathsterling,\;\flat\,\}$.
+[…]
 […]
@@
 4 Fast Ground Completion Introducing Additional Constants
-From now on throughout this section, let $\Sigma$ be a signature, $E$ and $F$ be GTESs over $\Sigma$, $W=ST(E\cup F)$, and $n=size(E)+size(F)$.
-We recall the fast ground completion algorithm of Fülöp and Vágvölgyi [2].
+From now on throughout the paper, let $\Sigma$ be a signature, $E$ and $F$ be GTESs over $\Sigma$, $W=ST\langle E\cup F\rangle$, and $n=\mathrm{size}(E)+\mathrm{size}(F)$.
+We recall the fast ground completion algorithm of Fülöp and Vágvölgyi [5, Algorithm 3.6.].
 Its input is $\Sigma$, $E$, and $F$.
-Their fast ground completion algorithm [2] produces in $O(n\,\mathrm{log}\;n)$ time a finite set of new constants and a reduced GTRS $R\langle E,F\rangle$ over $\Sigma$ extended with the new constants such that $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$.
+Their fast ground completion algorithm [5, Algorithm 3.6.] produces in $O(n\,\mathrm{log}\,n)$ time a finite set of new constants and a reduced GTRS $R\langle E,F\rangle$ over $\Sigma$ extended with the new constants such that $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$.
 Their algorithm calls a congruence closure algorithm on the subterm dag for $E\cup F$ [9], and runs in $O(n\,\mathrm{log}\,n)$ time by Proposition 2.17.
-We now cite the Claim on [17, p. 167].
-Proposition 4.25.
-[17] $\Theta\langle E,F\rangle=\{\,(\widehat{x},\widehat{y})\mid(x,y)\in\rho\langle E,F\rangle\,\}$.
+We now cite the Claim on [1, p. 167].
+Proposition 4.26 ([1, Claim on page 167]).
+$\Theta\langle E,F\rangle=\{\,(\widehat{x},\widehat{y})\mid(x,y)\in\rho\langle E,F\rangle\,\}$.
 Definition 4.27.
@@
 The elements of $C\langle E,F\rangle$ are viewed as constants, and we define the signature $\Sigma\cup C\langle E,F\rangle$.
-Statement 4.27.
-$|C\langle E,F\rangle|\leq|W|\leq size(W)\leq n$.
-Proof.
-By Definition 4.26. ∎
-Definition 4.28.
-[2] Let
+By Definition 4.27, we have the following statement.
+Statement 4.28.
+$|C\langle E,F\rangle|\leq|W|\leq\mathrm{size}(W)\leq n$.
+Definition 4.29 ([5, Section 3]).
+Let
 $R\langle E,F\rangle=\newline \{\,\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\sigma(t_{1},\ldots,t_{m})/_{\Theta\langle E,F\rangle}\mid\sigma\in\Sigma_{m},m\in{\mathbb{N}},t_{1},\ldots,t_{m}\in T_{\Sigma},\sigma(t_{1},\ldots,t_{m})\in W\,\}$
-be a GTRS over the ranked alphabet $\Sigma\cup C\langle E,F\rangle$.
+be a GTRS over the signature $\Sigma\cup C\langle E,F\rangle$.
 Note that for any $\sigma(t_{1},\ldots,t_{m})\in W$ with $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, and $t_{1},\ldots,t_{m}\in T_{\Sigma}$ there exists at most one rewrite rule with left-hand side $\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})$ in $R\langle E,F\rangle$.
 Definition 4.30.
-We say that GTRS $R\langle E,F\rangle$ is total if for any $\sigma\in\Sigma_{m}$ with $m\in{\mathbb{N}}$, and $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$, there exists a rewrite rule in GTRS $R\langle E,F\rangle$ with left-hand side $\sigma(a_{1},\ldots,a_{m})$.
+We say that GTRS $R\langle E,F\rangle$ is total if for any $\sigma\in\Sigma_{m}$ with $m\in{\mathbb{N}}$, and ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$, there exists a rewrite rule in GTRS $R\langle E,F\rangle$ with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$.
 Fülöp and Vágvölgyi [5] showed the following properties of $R\langle E,F\rangle$.
-Proposition 4.30.
-[2] $R\langle E,F\rangle$ is reduced.
+Proposition 4.31 ([5, Proposition 3.1]).
+$R\langle E,F\rangle$ is reduced.
 By Proposition 4.31 and Proposition 2.9, we have the following result.
@@
 [5] $R\langle E,F\rangle$ is convergent.
-Proposition 4.32.
-[2] For every $t\in W$, we have $t\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
+Proposition 4.33 ([5, Proposition 3.2]).
+For every $t\in W$, we have $t\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
 Proposition 4.34.
@@
 Accordingly $t/_{\Theta\langle E,F\rangle}$ is a $\mathop{\rightarrow}_{R\langle E,F\rangle}$ normal form of $t$.
-By Proposition 2.3, Proposition 4.31 $t\in W$ has a unique normal form, $t{\downarrow}_{R\langle E,F\rangle}$.
+By Proposition 2.4 and Proposition 4.32, $t\in W$ has a unique normal form, $t{\downarrow}_{R\langle E,F\rangle}$.
 Then we have $t{\downarrow}_{R\langle E,F\rangle}=t/_{\Theta\langle E,F\rangle}$.
 ∎
+Proposition 4.35 ([5, Lemma 3.3]).
+For all $s\in T_{\Sigma}$ and $t\in W$,
+if $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$, then $s\mathop{\leftrightarrow}^{*}_{E}t$.
+Proposition 4.36 ([5, Theorem 3.4]).
+[5] $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$.
+Proposition 4.37 ([5, Theorem 3.5]).
+For all $s,t\in T_{\Sigma}$,
+$s\mathop{\leftrightarrow}^{*}_{E}t$ if and only if $s{\downarrow}_{R\langle E,F\rangle}=t{\downarrow}_{R\langle E,F\rangle}$.
+Proposition 4.34 and Proposition 4.37 imply the following proposition.
 Proposition 4.38.
-[2] For all $s\in T_{\Sigma}$ and $t\in W$,
-if $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$, then $s\mathop{\leftrightarrow}^{*}_{E}t$.
-Proposition 4.35.
-[2] For all $s,t\in T_{\Sigma}$,
-$s\mathop{\leftrightarrow}^{*}_{E}t$ if and only if $s{\downarrow}_{R\langle E,F\rangle}=t{\downarrow}_{R\langle E,F\rangle}$.
-Proposition 4.36.
-[2] For all $s\in T_{\Sigma}$ and $t\in W$,
+For all $s\in T_{\Sigma}$ and $t\in W$,
 if $s\mathop{\leftrightarrow}^{*}_{E}t$, then $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
-Proposition 4.37.
-[2] $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$.
 By Proposition 4.35 and Proposition 4.38, we have the following result.
@@
 $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$ if and only if $s\mathop{\leftrightarrow}^{*}_{E}t$.
-Statement 4.39.
+Statement 4.40.
 If $R\langle E,F\rangle=\emptyset$, then $R\langle E,F\rangle$ is not total.
@@
 By Proposition 4.37, $\sigma(t_{1},\ldots,t_{m}){\downarrow}_{R\langle E,F\rangle}=s{\downarrow}_{R\langle E,F\rangle}$.
-By Proposition 4.33 $s{\downarrow}_{R\langle E,F\rangle}=s/_{\Theta\langle E,F\rangle}$.
+By Proposition 4.34, $s{\downarrow}_{R\langle E,F\rangle}=s/_{\Theta\langle E,F\rangle}$.
 Thus $\sigma(t_{1},\ldots,t_{m}){\downarrow}_{R\langle E,F\rangle}=s/_{\Theta\langle E,F\rangle}$.
 Therefore,
-$\sigma(t_{1},\ldots,t_{m})\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\sigma(t_{1}{\downarrow}_{R\langle E,F\rangle},\ldots,t_{m}{\downarrow}_{R\langle E,F\rangle})=$ (by Proposistion 4.32)
+$\sigma(t_{1},\ldots,t_{m})\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\sigma(t_{1}{\downarrow}_{R\langle E,F\rangle},\ldots,t_{m}{\downarrow}_{R\langle E,F\rangle})=$ (by Proposition 4.33)
 $\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})\mathop{\rightarrow}_{R\langle E,F\rangle}s/_{\Theta\langle E,F\rangle}$    (by direct inspection of the form of the rewrite rules of $R\langle E,F\rangle$).
@@
 Then $s=\sigma$ for some $\sigma\in\Sigma_{0}$.
-By our assumption that $R\langle E,F\rangle$ is a total GTES and by Definitions 4.28 and 4.29, there is a rewrite rule $\sigma\dot{\rightarrow}\sigma/_{\Theta\langle E,F\rangle}$ in $R\langle E,F\rangle$.
+By our assumption that $R\langle E,F\rangle$ is a total GTES and by Definitions 4.29 and 4.30, there is a rewrite rule $\sigma\dot{\rightarrow}\sigma/_{\Theta\langle E,F\rangle}$ in $R\langle E,F\rangle$.
 Then $\sigma\in W$ and $s{\downarrow}_{R\langle E,F\rangle}=\sigma/_{\Theta\langle E,F\rangle}$.
-[…]
+Induction Step: Let $k\in{\mathbb{N}}$, and let us assume that we have shown the lemma for all integers less than or equal to $k$.
+Let $height(s)=k+1$.
+Then $s=\sigma(s_{1},\ldots,s_{m})$ for some $\sigma\in\Sigma_{m}$, $m\geq 1$, and $s_{1},\ldots,s_{m}\in T_{\Sigma}$.
+By the induction hypothesis, for each $i\in\{\,1,\ldots,m\,\}$, there exist $t_{i}\in W$ such that $s_{i}{\downarrow}_{R\langle E,F\rangle}=t_{i}/_{\Theta\langle E,F\rangle}$.
+By our assumption that $R\langle E,F\rangle$ is a total GTRS, there is a rewrite rule with left-hand side $\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})$ in $R\langle E,F\rangle$.
+By Definition 4.29, this rule is of the form
 $\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\sigma(p_{1},\ldots,p_{m})/_{\Theta\langle E,F\rangle}$
-[…]
+in $R\langle E,F\rangle$ for some $p_{i}\in t_{i}/_{\Theta\langle E,F\rangle}$ for every $i\in\{\,1,\ldots,m\,\}$. Moreover $\sigma(p_{1},\ldots,p_{m})\in W$. Then
 $s=\sigma(s_{1},\ldots,s_{m})\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}s(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})\mathop{\rightarrow}_{R\langle E,F\rangle}\sigma(p_{1},\ldots,p_{m})/_{\Theta\langle E,F\rangle}$.
-By direct inspection of Definition 4.28, we get that $p_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$.
+By direct inspection of Definition 4.29, we get that $\sigma(p_{1},\ldots,p_{m})/_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$.
 Thus $s{\downarrow}_{R\langle E,F\rangle}=\sigma(p_{1},\ldots,p_{m})/_{\Theta\langle E,F\rangle}$.
 ∎
+The following proposition is a consequence of Lemma 4.41 and Lemma 4.42 and Proposition 4.37.
 Proposition 4.43.
 $\bigcup W/_{\mathop{\leftrightarrow}^{*}_{E}}=T_{\Sigma}$ if and only if $R\langle E,F\rangle$ is total.
+Statement 4.44.
+GTRS $R\langle E,F\rangle$ reaches a constant ${b}\in C\langle E,F\rangle$ from a ground term $t$ over the signature $\Sigma\cup C\langle E,F\rangle$ if and only if $t{\downarrow}_{R\langle E,F\rangle}={b}$.
 Proof.
-It is a consequence of Lemma 4.40 and Lemma 4.41 and Proposition 4.35. ∎
-Statement 4.43.
-GTRS $R\langle E,F\rangle$ reach a constant $a\in C\langle E,F\rangle$ from a ground term $t$ over the ranked alphabet $\Sigma\cup C\langle E,F\rangle$ if and only if $t{\downarrow}_{R\langle E,F\rangle}=a$.
-Proof.
-()($\Rightarrow$)( ⇒ ) The constant $a\in C\langle E,F\rangle$ is irreducibe for $R\langle E,F\rangle$ by direct inspection of Definition 4.28.
+$(\Rightarrow)$ The constant ${b}\in C\langle E,F\rangle$ is irreducibe for $R\langle E,F\rangle$ by direct inspection of Definition 4.29.
 By Proposition 4.32, $R\langle E,F\rangle$} is convergent.
-Accordingly $a\in C\langle E,F\rangle$ is the unique $R\langle E,F\rangle$-normal form of $t$.
-()($\Leftarrow$)( ⇐ ) This direction is trivial. ∎
+Accordingly ${b}\in C\langle E,F\rangle$ is the unique $R\langle E,F\rangle$-normal form of $t$.
+$(\Leftarrow)$ This direction is trivial. ∎
 Definition 4.45.
 Let
-$BSTEP\langle E,F\rangle=\{\,(a,b)\in C\langle E,F\rangle\times C\langle E,F\rangle\mid\mbox{ there exists a rewrite rule }\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E,F\rangle\mbox{ and }1\leq i\leq m\mbox{ such that }b=a_{i}\,\}$.
+[…]
 We call an element of $BSTEP\langle E,F\rangle$ a transition backstep.
-Statement 4.45.
-Let $n=size(E)+size(F)$. Then
-1. $|R\langle E,F\rangle|\leq|W|\leq size(W)\leq size(E\cup F)\leq size(E)+size(F)=n$.
-2. $size(R\langle E,F\rangle)\leq size(W)+|W|\leq 2\cdot n$  and
-3. $|BSTEP\langle E,F\rangle|\leq size(R\langle E,F\rangle)\leq 2\cdot n$.
+Statement 4.46.
+Let $n=\mathrm{size}(E)+\mathrm{size}(F)$. Then
+1. $|R\langle E,F\rangle|\leq|W|\leq\mathrm{size}(W)\leq size(E\cup F)\leq\mathrm{size}(E)+\mathrm{size}(F)=n$.
+2. $\mathrm{size}(R\langle E,F\rangle)\leq\mathrm{size}(W)+|W|\leq 2n$  and
+3. $|BSTEP\langle E,F\rangle|\leq\mathrm{size}(R\langle E,F\rangle)\leq 2n$.
 Proof.
-By direct inspection of Definition 4.28 and by Statement 2.7 we have Statement 1.
-[…]
-By direct inspection of Definition 4.44 and by Statement 2 we have Statement 3.
+By direct inspection of Definition 4.29 and by Statement 2.8 we have Statement 1.
+We now show Statement 2.
+By Definition 4.29, for each rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E,F\rangle$, there exists a ground term $\sigma(t_{1},\ldots,t_{m})\in W$ such that ${b}_{i}=t_{i}/_{\Theta\langle E,F\rangle}$ for every $i\in\{\,1,\ldots,m\,\}$.
+Then we assign to the occurrence of ${b}_{i}$ the root symbol of the ground term $t_{i}$ for every $i\in\{\,1,\ldots,m\,\}$.
+Consequently $\mathrm{size}(R\langle E,F\rangle)\leq\mathrm{size}(W)+|W|$.
+By Statement 1, $\mathrm{size}(W)+|W|\leq 2n$.
+By direct inspection of Definition 4.45 and by Statement 2 we have Statement 3. ∎
 We now adopt the fast ground completion algorithm of Fülöp and Vágvölgyi [5] that produces the GTRS $R\langle E,F\rangle$.
-Definition 4.46.
-Algorithm Fast Ground Completion Introducing Additional Constants (FGC for short) [2]
+Definition 4.47 ([5, Algorithm 3.6]).
+Algorithm Fast Ground Completion Introducing Additional Constants (FGC for short)
 Input: A signature $\Sigma$, GTES $E$ and GTES $F$ over $\Sigma$.
 Output:
+  • 1.
     the relation $\rho\langle E,F\rangle$,
+  • 2.
     the set of nullary symbols $C\langle E,F\rangle$,
+  • 3.
     the GTRS $R\langle E,F\rangle$ stored without duplicates in a red-black tree $RBT1$, and
+  • 4.
     the relation $BSTEP\langle E,F\rangle$ stored without duplicates in a red-black tree $RBT2$.
@@
 The binary sequence obtained in this way is called the binary representation of the symbol $\sigma\in\Sigma\langle E,F\rangle$, and is denoted as $\phi(\sigma)$.
-Similarly, for each constant $a\in C\langle E,F\rangle$, we pad the binary form of its number with leading zeros to length $\ell$.
-The binary sequence obtained in this way is called the binary representation of the constant $a\in C\langle E,F\rangle$, and is denoted as $\phi(a)$.
-The binary representation of a rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E,F\rangle$ is the string $\phi(\sigma)\phi(a_{1}),\ldots,\phi(a_{m})\phi(a)$.
+Similarly, for each constant ${b}\in C\langle E,F\rangle$, we pad the binary form of its number with leading zeros to length $\ell$.
+The binary sequence obtained in this way is called the binary representation of the constant ${b}\in C\langle E,F\rangle$, and is denoted as $\phi({b})$.
+The binary representation of a rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E,F\rangle$ is the string $\phi(\sigma)\phi({b}_{1}),\ldots,\phi({b}_{m})\phi({b})$.
 We store the binary representations of the rewrite rules in $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT1$ in lexicographic order.
-The binary representation of an ordered pair $(a,b)\in BSTEP\langle E,F\rangle$ is the ordered pair $(\phi(a),\phi(b))$.
+The binary representation of an ordered pair $({b},{c})\in BSTEP\langle E,F\rangle$ is the ordered pair $(\phi({b}),\phi({c}))$.
 We store the binary representations of the elements of $BSTEP\langle E,F\rangle$ without duplicates in the red-black tree $RBT2$ in lexicographic order.
-Here for every ordered pair $(a,b)\in BSTEP\langle E,F\rangle$, the search key is the first component $\phi(a)$, the second component, $\phi(b)$ is satellite data.
-var $i$, $m$: ${\mathbb{N}}$; $\sigma$: $\Sigma$; $a,a_{1},\ldots,a_{mar_{\Sigma}}:C\langle E,F\rangle$; $RBT1$, $RBT2$: red-black tree;
+Here for every ordered pair $({b},{c})\in BSTEP\langle E,F\rangle$, the search key is the first component $\phi({b})$, the second component, $\phi({c})$ is satellite data.
+var $i$, $m$: $\mathbb{N}$}; $\sigma$: $\Sigma$; ${b},{b}_{1},\ldots,{b}_{mar_{\Sigma}}:C\langle E,F\rangle$; $RBT1$, $RBT2$: red-black tree;
 1. Create a subterm dag $G=(V,A,\lambda)$ for $E$ and $F$.
@@
 $m:=outdegree(x);$
-$a:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x)$;
+${b}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x)$;
 for each $1\leq i\leq m$ do
-$a_{i}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x/i)$;
-$R\langle E,F\rangle:=R\langle E,F\rangle\cup\{\,\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\,\}$;
-$BSTEP\langle E,F\rangle:=BSTEP\langle E,F\rangle\cup\{\,(a_{1},a),\ldots,(a_{m},a)\,\}$;
+${b}_{i}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x/i)$;
+$R\langle E,F\rangle:=R\langle E,F\rangle\cup\{\,\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\,\}$;
+$BSTEP\langle E,F\rangle:=BSTEP\langle E,F\rangle\cup\{\,({b}_{1},{b}),\ldots,({b}_{m},{b})\,\}$;
 end
-By Statement 4.45 and Proposition 2.2, the time complexity of storing $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT1$ is $O(n\,\mathrm{log}\,n)$, where $n=size(E)+size(F)$.
-By Statement 4.45 and Proposition 2.2, the time complexity of storing $BSTEP\langle E,F\rangle$ without duplicates in the red-black tree $RBT2$ is $O(n\,\mathrm{log}\,n)$, where $n=size(E)+size(F)$.
-Proposition 4.47.
-[2] Algorithm FGC takes as input a signature $\Sigma$, and GTESs $E$ and $F$ over $\Sigma$. Let $n=size(E)+size(F)$.
+By Statement 4.46 and Proposition 2.2, the time complexity of storing $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT1$ is $O(n\,\mathrm{log}\,n)$, and the time complexity of storing $BSTEP\langle E,F\rangle$ without duplicates in the red-black tree $RBT2$ is $O(n\,\mathrm{log}\,n)$, where $n=\mathrm{size}(E)+\mathrm{size}(F)$.
+Proposition 4.48 ([5, Algorithm 3.6]).
+Algorithm FGC takes as input a signature $\Sigma$, and GTESs $E$ and $F$ over $\Sigma$. Let $n=\mathrm{size}(E)+\mathrm{size}(F)$.
 1. Algorithm FGC produces in $O(n\,\mathrm{log}\,n)$ time the relation $\rho\langle E,F\rangle$ and the set of nullary symbols $C\langle E,F\rangle$.
 2.
-In $O(n\,\mathrm{log}\,n)$ time, Algorithm FGC computes the GTRS $R\langle E,F\rangle$ over $\Sigma\cup C\langle E,F\rangle$, and stores $R\langle E,F\rangle$ without duplicates in the lexical order $<_{lex}$ on binary sequences in the red-black tree $RBT1$.
+In $O(n\,\mathrm{log}\,n)$ time, Algorithm FGC computes the GTRS $R\langle E,F\rangle$ over $\Sigma\cup C\langle E,F\rangle$, and stores $R\langle E,F\rangle$ without duplicates in the lexicographic order $_{lex}$ on binary sequences in the red-black tree $RBT1$.
 3.
@@
 Hence we can construct the relation $BSTEP\langle E,F\rangle$ in $O(n\,\mathrm{log}\,n)$ time as well.
-In addition to the result of Fülöp and Vágvölgyi [2], we now compute the cost of storing the GTRS $R\langle E,F\rangle$.
-and the cost of storing the relation $BSTEP\langle E,F\rangle$.
+In addition to the result of Fülöp and Vágvölgyi [5], we now compute the cost of storing the GTRS $R\langle E,F\rangle$, and the cost of storing the relation $BSTEP\langle E,F\rangle$.
 First we show Statement 1.
@@
 Algorithm FGC produces the GTRS $R\langle E,F\rangle$ apart from that, for every equivalence class $\vartheta\in\Theta\langle E,F\rangle$, Algorithm FGC writes the unique name for $\vartheta$ [5].
-By Statement 4.45, $size(R\langle E,F\rangle)\leq 2\cdot|W|\leq 2\cdot n$.
+By Statement 4.46, $\mathrm{size}(R\langle E,F\rangle)\leq 2\cdot|W|\leq 2n$.
 Then by Proposition 2.2, the time-complexity of inserting the elements of $R\langle E\cup F,\emptyset\rangle$ into an empty red-black tree, not allowing duplicate values in the resulting red-black tree $RBT1$, is $O(n\,\mathrm{log}\,n)$,
 We now show Statement 3.
-In the relation $BSTEP\langle E,F\rangle\subseteq C\langle E,F\rangle\times C\langle E,F\rangle$, we collect all pairs $(a_{i},a)$ such that there is a rewrite rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a$ in $R\langle E,F\rangle$.
-By Statement 6.63, $|BSTEP\langle E,F\rangle|\leq n$.
+In the relation $BSTEP\langle E,F\rangle\subseteq C\langle E,F\rangle\times C\langle E,F\rangle$, we collect all pairs $({b}_{i},{b})$ such that there is a rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$ in $R\langle E,F\rangle$.
+By Statement 4.46, $|BSTEP\langle E,F\rangle|\leq 2n$.
 Then by Proposition 2.2, the time-complexity of inserting the elements of $BSTEP\langle E,F\rangle$ into an empty red-black tree, not allowing duplicate values in the resulting red-black tree, is $O(n\,\mathrm{log}\,n)$.
+∎
 We give the GTRS $R\langle E,F\rangle$ as input to an algorithm by giving the red-black tree $RBT1$ output by Algorithm FGC storing the GTRS $R\langle E,F\rangle$.
@@
 Definition 4.49.
-Let $E$ and $F$ be GTESs over a ranked alphabet $\Sigma$.
+Let $E$ and $F$ be GTESs over a signature $\Sigma$.
 Algorithm Deciding whether GTRS $R\langle E,F\rangle$ is Total (DTOT for short).
@@
 If $R\langle E,F\rangle=\emptyset$, then return $\mathrm{false}$.
-Otherwise, traverse the red-black tree $RBT$ in in-order traversal, and read all binary sequences stored in the red-black tree $RBT$ in the lexical order $<_{lex}$ on binary sequences.
+Otherwise, traverse the red-black tree $RBT$ in in-order traversal, and read all binary sequences stored in the red-black tree $RBT$ in the lexicographic order $_{lex}$ on binary sequences.
 For every $m\in{\mathbb{N}}$, $\sigma\in\Sigma\langle R\langle E,F\rangle_{m}$, read the rules with $\sigma$ appering on the left-hand side one after another.
@@
 begin
-count all rules with $\sigma$ appearing on the left-hand side, and store the result in the variable $count$;
+count all rules with $\sigma$ appearing on the left-hand side, and store the result in the variable countcount;
 $m:=\mathrm{arity}(\sigma)$;
@@
 Lemma 4.50.
-Algorithm DTOT decides in $O(n)$ time whether $R\langle E,F\rangle$ is total, where $n=size(E)+size(F)$.
+Algorithm DTOT decides in $O(n)$ time whether $R\langle E,F\rangle$ is total, where $n=\mathrm{size}(E)+\mathrm{size}(F)$.
 Proof.
-First, assume that $R\langle E,F\rangle=\emptyset$.
-Recall that there exists a constant $\sigma\in\Sigma_{0}$.
-Then there does not exist a rewrite rule with left-hand side $\sigma$ in $R\langle E,F\rangle$.
-Consequently, $R\langle E,F\rangle$ is not total.
-Second, assume that $R\langle E,F\rangle\neq\emptyset$.
+If $R\langle E,F\rangle=\emptyset$, then by Statement 4.40, we are done.
+Assume that $R\langle E,F\rangle\neq\emptyset$.
 If $FLAG\langle E,F\rangle=\mathrm{false}$, that is, $|\Sigma\langle E,F\rangle|<|\Sigma|$, then $R\langle E,F\rangle$ is not total.
 Assume that $FLAG\langle E,F\rangle=\mathrm{true}$, that is, $|\Sigma\langle E,F\rangle|=|\Sigma|$.
-Then for each symbol $\sigma\in\Sigma$, we count the rewrite rules in $R\langle E,F\rangle$ with $\sigma$ appearing in the left-hand side and store the result in the variable $count$.
-If there exists a symbol $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, such that for $\sigma$, $count<|\langle C\langle E,F\rangle\rangle|^{m}$, then there exist constants $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$ such that there does not exist a rewrite rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R\langle E,F\rangle$.
+Then for each symbol $\sigma\in\Sigma$, we count the rewrite rules in $R\langle E,F\rangle$ with $\sigma$ appearing in the left-hand side and store the result in the variable countcount.
+If there exists a symbol $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, such that for $\sigma$, $count<|\langle C\langle E,F\rangle\rangle|^{m}$, then there exist constants ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$ such that there does not exist a rewrite rule with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$ in $R\langle E,F\rangle$.
 Therefore, $R\langle E,F\rangle$ is not total.
-Conversely if $R\langle E,F\rangle$ is not total, then there exists a symbol $\sigma\in\Sigma_{m}$ and constants $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$ such that there does not exist a rewrite rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R\langle E,F\rangle$.
+Conversely if $R\langle E,F\rangle$ is not total, then there exists a symbol $\sigma\in\Sigma_{m}$ and constants ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$ such that there does not exist a rewrite rule with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$ in $R\langle E,F\rangle$.
 Then for $\sigma$, $count<|\langle C\langle E,F\rangle\rangle|^{m}$.
@@
 Assume that $|\Sigma\langle E,F\rangle|=|\Sigma|$ and $R\langle E,F\rangle\neq\emptyset$.
-Then Algorithm DTOT traverse the red-black tree $RBT$ in in-order traversal, and it reads all binary sequences stored in the red-black tree $RBT$ in the lexical order $<_{lex}$ on binary sequences.
-For every $\sigma\in\Sigma\langle R\langle E,F\rangle$, it reads the rules with $\sigma$ appering on the left-hand side one after another, and counts the rewrite rules in $R\langle E,F\rangle$ with $\sigma$ appearing in the left-hand side for all symbols $\sigma\in\Sigma$.
+Then Algorithm DTOT traverse the red-black tree $RBT$ in in-order traversal, and it reads all binary sequences stored in the red-black tree $RBT$ in the lexicographic order $_{lex}$ on binary sequences.
+For every $\sigma\in\Sigma\langle R\langle E,F\rangle\rangle$, it reads the rules with $\sigma$ appering on the left-hand side one after another, and counts the rewrite rules in $R\langle E,F\rangle$ with $\sigma$ appearing in the left-hand side for all symbols $\sigma\in\Sigma$.
 This takes $O(n)$ time.
@@
 Otherwise, for every symbol $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, we carry out $m$ multiplications when computing $|C\langle E,F\rangle|^{m}$.
-Because each symbol $\sigma\in\Sigma\langle R\langle E,F\rangle_{m}$, $m\in{\mathbb{N}}$, has $m$ constants in its arguments in the left-hand side of some rewrite rule in $R\langle E,F\rangle$.
-The number of multiplications $\leq$ $\sum_{m\in{\mathbb{N}}}\sum_{\sigma\in\Sigma\langle R\langle E,F\rangle_{m}}m\leq$ $size(R\langle E,F\rangle)\leq n$.
-by Statement 4.45.
-Consequently, the number of multiplications is less than or equal to $n$.
+Because each symbol $\sigma\in\Sigma\langle R\langle E,F\rangle\rangle_{m}$, $m\in{\mathbb{N}}$, has $m$ constants in its arguments in the left-hand side of some rewrite rule in $R\langle E,F\rangle$.
+The number of multiplications is less than or equal to $\sum_{m\in{\mathbb{N}}}\sum_{\sigma\in\Sigma\langle R\langle E,F\rangle\rangle_{m}}m\leq$ $\mathrm{size}(R\langle E,F\rangle)\leq 2n$ by Statement 4.46.
+Consequently, the number of multiplications is less than or equal to $2n$.
 The number of comparisons of nonnegative integers is $|\Sigma|\leq n$.
@@
 We now illustrate the concepts of Section 4 by continuing our running examples.
-As in Section 3, throughout this section $E$ and $F$ are GTESs over a signature $\Sigma$, and $W$ stands for $ST\langle E\cup F\rangle$.
+We compute the sets of constants $C\langle E,F\rangle$, $C\langle F,E\rangle$, $C\langle E\cup F,\emptyset\rangle$, and the GTRSs $R\langle E,F\rangle$, $R\langle F,E\rangle$, $R\langle E\cup F,\emptyset\rangle$.
+Recall that $E$ and $F$ are GTESs over a signature $\Sigma$, and $W$ stands for $ST\langle E\cup F\rangle$.
 Example 5.51.
@@
 […]
-$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;g(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow},\;g(\$)/_{\Theta\langle E,F\rangle}\;\}$,
+$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;f(\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\mathdollar)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;g(\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow},\;g(\mathdollar)/_{\Theta\langle E,F\rangle}\;\}$,
 […]
-$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;g(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E,F\rangle}\;\}$,
+$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle F,E\rangle},\;f(\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar)/_{\Theta\langle F,E\rangle},\;\newline g(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle F,E\rangle},\;g(\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar)/_{\Theta\langle F,E\rangle}\;\}$,
 […]
-$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline g(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},g(\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
+$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline g(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},g(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
 Example 5.52.
@@
 $C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;f(\#),\;f^{2}(\#),\;f^{3}(\#)\,\}\,\}=\{\,W\,\}$,
-$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
-We note that $R\langle E,F\rangle$, $R\langle F,E\rangle$, and $R\langle E\cup F,E\rangle$ are total.
+$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
+We note that $R\langle E,F\rangle$, $R\langle F,E\rangle$, and $R\langle E\cup F,\emptyset\rangle$ are total.
 Example 5.53.
 We continue Example 3.20.
-$C\langle E,F\rangle=\{\,\{\,\#\,\},\;\newline \{\,\$,\;f(\$,\$),\;f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$))\,\},\;\newline \{\,f(\#,\#),\;f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#))\,\},\newline \;\{\,\mathsterling,\;f(\#,\$),\;f(\#,\mathsterling),\;f(\$,\#),\;f(\$,\mathsterling),\;f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}\,\}$,
-$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;f(\#/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\#/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\$/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\$/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\;f(\$/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\mathsterling/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\#/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;\newline f(f(\#,\#)/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;\,f(f(\#,\#)/_{\Theta\langle E,F\rangle},f(\#,\#))/_{\Theta\langle E,F\rangle}\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle}\;\;\newline f(\$/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;f(\mathsterling/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(f(\#,\#)/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;f(f(\#,\#)/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle}\,\}$.
-$C\langle F,E\rangle=\{\,\{\,\$\,\},\;\newline \{\,\#,\;f(\#,\#),\,f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#))\,\},\;\newline \{\,f(\$,\$),\;f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$))\,\},\newline \;\{\,\mathsterling,\;f(\$,\#),\;f(\$,\mathsterling),\;f(\#,\$),\;f(\#,\mathsterling),\;f(\mathsterling,\$),\;f(\mathsterling,\#),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling)\,\}\,\}$,
-$R\langle F,E\rangle=\{\,\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\;\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\$/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;f(\$/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\$/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\;f(\#/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\mathsterling/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\$/_{\Theta\langle F,E\rangle},f(\$,\$)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;\newline f(f(\$,\$)/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;\,f(f(\$,\$)/_{\Theta\langle F,E\rangle},f(\$,\$))/_{\Theta\langle F,E\rangle}\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle}\;\;\newline f(\#/_{\Theta\langle F,E\rangle},f(\$,\$)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;f(\mathsterling/_{\Theta\langle F,E\rangle},f(\$,\$)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(f(\$,\$)/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;f(f(\$,\$)/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$.
-$C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;\$,\;\mathsterling,\;\newline f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#)),\;\;\newline f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$)),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}\,\}$,
-$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
+[…]
+$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;f(\#/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\#/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\mathdollar/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\mathdollar/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\;f(\mathdollar/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\mathsterling/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\#/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;\newline f(f(\#,\#)/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;\,f(f(\#,\#)/_{\Theta\langle E,F\rangle},f(\#,\#))/_{\Theta\langle E,F\rangle}\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle}\;\;\newline f(\mathdollar/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\mathsterling/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(f(\#,\#)/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(f(\#,\#)/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle}\,\}$.
+$C\langle F,E\rangle=\{\,\{\,\mathdollar\,\},\;\newline \{\,\#,\;f(\#,\#),\,f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#))\,\},\;\newline \{\,f(\mathdollar,\mathdollar),\;f(\mathdollar,f(\mathdollar,\mathdollar)),\;f(f(\mathdollar,\mathdollar),\mathdollar),\;f(f(\mathdollar,\mathdollar),f(\mathdollar,\mathdollar))\,\},\newline \;\{\,\mathsterling,\;f(\mathdollar,\#),\;f(\mathdollar,\mathsterling),\;f(\#,\mathdollar),\;f(\#,\mathsterling),\;f(\mathsterling,\mathdollar),\;f(\mathsterling,\#),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\mathdollar,\mathdollar)),\;f(\mathsterling,f(\mathdollar,\mathdollar)),\;f(f(\mathdollar,\mathdollar),\#),\;f(f(\mathdollar,\mathdollar),\mathsterling),\;\newline f(\mathdollar,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\mathdollar),\;f(f(\#,\#),\mathsterling)\,\}\,\}$,
+$R\langle F,E\rangle=\{\,\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle F,E\rangle},\;\;\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathdollar/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\;\;f(\mathdollar/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathdollar/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\#/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\;f(\#/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\mathsterling/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\mathdollar/_{\Theta\langle F,E\rangle},f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\;\;\newline f(f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\;\;\,f(f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},f(\mathdollar,\mathdollar))/_{\Theta\langle F,E\rangle}\dot{\rightarrow}f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle}\;\;\newline f(\#/_{\Theta\langle F,E\rangle},f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\mathsterling/_{\Theta\langle F,E\rangle},f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$.
+[…]
+$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
 We observe that $R\langle E,F\rangle$, $R\langle F,E\rangle$, and $R\langle E\cup F,\emptyset\rangle$ are total.
@@
 $C\langle E,F\rangle=\{\,\#/_{\Theta\langle E,F\rangle},\;\flat/_{\Theta\langle E,F\rangle}\,\}$,
-$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\flat\dot{\rightarrow}\flat/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\flat/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;f(\flat/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#\,\}$.
-$C\langle F,E\rangle=\{\,\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle}\,\}$,
-$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\flat\dot{\rightarrow}\mathsterling_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\newline f(\$/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\$/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\$/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathsterling/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathsterling/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$.
+$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\flat\dot{\rightarrow}\flat/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\flat/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;f(\flat/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#\,\}$.
+$C\langle F,E\rangle=\{\,\#/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle}\,\}$,
+$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle F,E\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\flat\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\newline f(\mathdollar/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathdollar/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathdollar/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathsterling/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathsterling/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$.
 $C\langle E\cup F,\emptyset\rangle=\{\,\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
-$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\flat\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
+$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\flat\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
 We note that $R\langle E,F\rangle$, $R\langle F,E\rangle$, and $R\langle E\cup F,\emptyset\rangle$ are total.
@@
 We now continue Example 3.22.
-$C\langle E,F\rangle=\{\,\#/_{\Theta\langle E,F\rangle},\;\$/_{\Theta\langle E,F\rangle}\,\}$.
-$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\$/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle}\;\}$.
+$C\langle E,F\rangle=\{\,\#/_{\Theta\langle E,F\rangle},\;\mathdollar/_{\Theta\langle E,F\rangle}\,\}$.
+$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\newline f(\mathdollar/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;f(\mathdollar/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle}\;\}$.
 Observe that $R\langle E,F\rangle$ is total.
-$C\langle F,E\rangle=\{\,\{\,\#,\;f(\#,\#)\,\},\;\{\,\$\,\},\;\{\,f(\#,\$)\,\},\;\{\,f(\$,\#)\,\},\;\{\,f(\$,\$)\,\}\,\}$.
-$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},f(\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\$)/_{\Theta\langle F,E\rangle},\;\newline f(\$/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\#)/_{\Theta\langle F,E\rangle},\;f(\$/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle}\,\}$.
+[…]
+$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},f(\#/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\mathdollar)/_{\Theta\langle F,E\rangle},\;\newline f(\mathdollar/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\#)/_{\Theta\langle F,E\rangle},\;f(\mathdollar/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle}\,\}$.
 Note that $R\langle F,E\rangle$ is not total.
-$C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}\,\}$.
-$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,E\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
+[…]
+$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
 We see that $R\langle E\cup F,\emptyset\rangle$ is total.
@@
 We now continue Example 3.23. Recall that the set of equivalence classes of $\Theta\langle E,F\rangle$ is:
-$C\langle E,F\rangle=\{\,\{\,\#\,\},\;\{\,f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\},\;\newline \{\,\$,\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\$),\;g(g(\$))\,\}\,\}$.
-That is, $C\langle E,F\rangle=\{\,\#/_{\Theta\langle E,F\rangle},\;\$/_{\Theta\langle E,F\rangle},\,f(\#,\#)/_{\Theta\langle E,F\rangle}\,\}$.
-$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline f(g(\#)/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;f(g(\#)/_{\Theta\langle E,F\rangle},g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;g(g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\$/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;g(\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\$)/_{\Theta\langle E,F\rangle},\;g(g(\$)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle}$.
+[…]
+That is, $C\langle E,F\rangle=\{\,\#/_{\Theta\langle E,F\rangle},\;\mathdollar/_{\Theta\langle E,F\rangle},\,f(\#,\#)/_{\Theta\langle E,F\rangle}\,\}$.
+$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline f(g(\#)/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;f(g(\#)/_{\Theta\langle E,F\rangle},g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline g(\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;g(g(\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\#)/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;f(\mathdollar/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\newline f(\mathdollar/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;g(\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}g(\mathdollar)/_{\Theta\langle E,F\rangle},\;g(g(\mathdollar)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle}$.
 Note that $R\langle E,F\rangle$ is total.
 Recall that the set of equivalence classes of $\Theta\langle F,E\rangle$ is:
-$C\langle F,E\rangle=\{\,\#,\;f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\}$,
-{\$},$\{\,\$\,\},\;\{\,g(\$)\,\},\;\{\,f(\#,\$)\,\},\;\{\,f(\$,\#)\,\},\;\{\,f(\$,\$)\,\},\;\{\,g(g(\$))\,\}$.
+$C\langle F,E\rangle=\{\,\{\,\#,\;f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\}$,
+[…]
 That is,
-$C\langle F,E\rangle=\{\,\#/_{\Theta\langle F,E\rangle},\;\$/_{\Theta\langle F,E\rangle},\;g(\$)/_{\Theta\langle E,E\rangle},\;f(\#,\$)/_{\Theta\langle E,F\rangle},\;f(\$,\#)/_{\Theta\langle E,F\rangle},\;\newline f(\$,\$)/_{\Theta\langle E,F\rangle},\;g(g(\$))/_{\Theta\langle E,F\rangle}\;\,\}$.
-$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},g(\#)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\newline f(g(\#)/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(g(\#)/_{\Theta\langle F,E\rangle},g(\#)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\newline g(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;g(\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}g(\$)/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\$)/_{\Theta\langle F,E\rangle},\,f(\$/_{\Theta\langle F,E\rangle},\#_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\#)/_{\Theta\langle F,E\rangle}\newline f(\$/_{\Theta\langle F,E\rangle},\$_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;g(g(\$)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}g(g(\$))/_{\Theta\langle F,E\rangle}\,\}$.
+$C\langle F,E\rangle=\{\,\{\,\#/_{\Theta\langle F,E\rangle},\;\mathdollar/_{\Theta\langle F,E\rangle},\;g(\mathdollar)/_{\Theta\langle F,E\rangle},\;f(\#,\mathdollar)/_{\Theta\langle F,E\rangle},\;f(\mathdollar,\#)/_{\Theta\langle F,E\rangle},\;\newline f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\;g(g(\mathdollar))/_{\Theta\langle F,E\rangle}\,\}\,\}$.
+$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},g(\#)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\newline f(g(\#)/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(g(\#)/_{\Theta\langle F,E\rangle},g(\#)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\newline g(\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;g(\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}g(\mathdollar)/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\mathdollar)/_{\Theta\langle F,E\rangle},\,f(\mathdollar/_{\Theta\langle F,E\rangle},\#_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\#)/_{\Theta\langle F,E\rangle}\newline f(\mathdollar/_{\Theta\langle F,E\rangle},\mathdollar_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\;g(g(\mathdollar)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}g(g(\mathdollar))/_{\Theta\langle F,E\rangle}\,\}$.
 Note that $R\langle F,E\rangle$ is not total.
 Recall that the set of equivalence classes of $\Theta\langle E\cup F,\emptyset\rangle$ is:
-$C\langle E\cup F,\emptyset\rangle=\ \{\,\#,\;f(\#,\#),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(\#),\;g(g(\#))\,\},\;\newline \{\,\$,\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\$),\;g(g(\$))\,\}$.
-That is, $C\langle E\cup F,\emptyset\rangle=\{\,\#/_{\Theta\langle E\cup F,\emptyset\rangle},\$/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
-$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\$_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\$/_{\Theta\langle E\cup F,\emptyset\rangle},\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline g(\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
+[…]
+That is, $C\langle E\cup F,\emptyset\rangle=\{\,\#/_{\Theta\langle E\cup F,\emptyset\rangle},\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
+$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\mathdollar_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline g(\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;g(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
 Observe that $R\langle E\cup F,\emptyset\rangle$ is total.
@@
 We continue Example 3.24.
-$C\langle E,F\rangle=\{\,\{\,\#\,\},\;\{\,\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}\,\}$.
-$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\newline f(\$/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle}\,\},$
-$C\langle F,E\rangle=\{\,\{\,\#,\;\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}\,\}$,
-$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle F,E\rangle}\,\}$,
+[…]
+$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\newline f(\mathdollar/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;f(\mathdollar/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle}\,\},$
+[…]
+$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle F,E\rangle}\,\}$,
 $C\langle E\cup F,\emptyset\rangle=C\langle F,E\rangle$,   $R\langle E\cup F,\emptyset\rangle=R\langle F,E\rangle$.
@@
 We now continue Example 3.25.
-$C\langle E,F\rangle=\{\,\{\,\#,\;\$\,\},\;\{\,\mathsterling\,\},\;\{\,\flat\,\}\,\}$,
-$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\flat\dot{\rightarrow}\flat/_{\Theta\langle E,F\rangle}\,\}$,
-$C\langle F,E\rangle=\{\,\{\,\#\,\},\;\{\,\$\,\},\;\{\,\mathsterling,\;\flat\,\}\,\}$,
-$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\flat\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$,
-$C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;\$\,\},\;\,\{\,\mathsterling,\;\flat\,\}\,\}$, and
-$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E\cup F,\emptyset\rangle},\;\flat\dot{\rightarrow}\mathsterling/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
+[…]
+$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\flat\dot{\rightarrow}\flat/_{\Theta\langle E,F\rangle}\,\}$,
+[…]
+$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle F,E\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\flat\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$,
+[…]
+$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E\cup F,\emptyset\rangle},\;\flat\dot{\rightarrow}\mathsterling/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
 Observe that GTRS $R\langle E,F\rangle$, GTRS $R\langle F,E\rangle$, and GTRS $R\langle E\cup F,\emptyset\rangle$ are not total.
 6 The Auxiliary dpwpa for the GTESs $E$ and $F$
-From now on throughout the paper, $E$ and $F$ are GTESs over a signature $\Sigma$, $n=size(E)+size(F)$, and $W$ stands for $ST\langle E\cup F\rangle$.
-We define the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F]\rangle]$ for the GTESs $E$ and $F$.
-To simplify our notations, each vertex of the dpwpa $AUX[E;F]\rangle$ is a constant $a\in C\langle E\cup F,\emptyset\rangle$, i.e, an equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$.
-Furthermore, we study $R\langle{E\cup F},\emptyset\rangle$ reachability of a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$.
+Recall that $E$ and $F$ are GTESs over a signature $\Sigma$, $n=\mathrm{size}(E)+\mathrm{size}(F)$, and $W$ stands for $ST\langle E\cup F\rangle$.
+We define the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$.
+To simplify our notations, each vertex of the dpwpa $AUX[E;F]$ is a constant ${b}\in C\langle E\cup F,\emptyset\rangle$, i.e, an equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$.
+Furthermore, we study the $R\langle{E\cup F},\emptyset\rangle$ reachability of a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$.
 Definition 6.59.
 Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then let
-$INC(E\cup F,E)=\{\,(a,b)\mid a\in C\langle E\cup F,\emptyset\rangle,b\in C\langle E,F\rangle,\mbox{ and }b\subseteq a\,\}$,
-$NUM(E\cup F,E)=\{\,(a,k)\mid a\in C\langle E\cup F,\emptyset\rangle,k=|\{b\in C\langle E,F\rangle\mid b\subseteq a\,\}|\,\}$,
-$INC(E\cup F,F)=\{\,(a,b)\mid a\in C\langle E\cup F,\emptyset\rangle,b\in C\langle F,E\rangle\mbox{ and }b\subseteq a\,\}$, and
-$NUM(E\cup F,F)=\{\,(a,k)\mid a\in C\langle E\cup F,\emptyset\rangle,k=|\{b\in C\langle F,E\rangle\mid b\subseteq a\,\}|\,\}$.
+$INC(E\cup F,E)=\{\,({b},{c})\mid{b}\in C\langle E\cup F,\emptyset\rangle,{c}\in C\langle E,F\rangle,\mbox{ and }{c}\subseteq{b}\,\}$,
+$NUM(E\cup F,E)=\{\,({b},k)\mid{b}\in C\langle E\cup F,\emptyset\rangle,k=|\{{c}\in C\langle E,F\rangle\mid{c}\subseteq{b}\,\}|\,\}$,
+$INC(E\cup F,F)=\{\,({b},{c})\mid{b}\in C\langle E\cup F,\emptyset\rangle,{c}\in C\langle F,E\rangle\mbox{ and }{c}\subseteq{b}\,\}$, and
+$NUM(E\cup F,F)=\{\,({b},k)\mid{b}\in C\langle E\cup F,\emptyset\rangle,k=|\{{c}\in C\langle F,E\rangle\mid{c}\subseteq{b}\,\}|\,\}$.
 Intuitively, $INC(E\cup F,E)$ shows the inclusions between the elements of $C\langle E\cup F,\emptyset\rangle$ and $C\langle E,F\rangle$.
-$NUM(E\cup F,E)$ consists of all pairs $(a,k)$, where the second component $k$ is the number of all pairs in $INC(E\cup F,E)$, with first component $a$.
-Statement 6.59.
+$NUM(E\cup F,E)$ consists of all pairs (b,k)({b},k), where the second component $k$ is the number of all pairs in $INC(E\cup F,E)$, with first component b{b}.
+Similar statements hold true for $INC(E\cup F,F)$ and $NUM(E\cup F,F)$ as well.
+Statement 6.60.
 Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then $|INC(E\cup F,E)|\leq n$ and $|INC(E\cup F,F)|\leq n$.
 Proof.
-By Statement 4.27, $|C\langle E,F\rangle|\leq n$.
-For each $b\in C\langle E,F\rangle$, there exists at most one $a\in C\langle E\cup F,\emptyset\rangle$ such that $(a,b)\in INC(E\cup F,E)$.
+By Statement 4.28, $|C\langle E,F\rangle|\leq n$.
+For each ${c}\in C\langle E,F\rangle$, there exists at most one ${b}\in C\langle E\cup F,\emptyset\rangle$ such that $({b},{c})\in INC(E\cup F,E)$.
 Therefore, $|INC(E\cup F,E)|\leq|C\langle E,F\rangle|\leq n$.
 We can show similarly that $|INC(E\cup F,F)|\leq|C\langle F,E\rangle|\leq n$. ∎
-Statement 6.60.
+We get the following statement by direct inspection of Definition 6.59.
+Statement 6.61.
 1.
-For every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,b)\in INC(E\cup F,E)$ if and only if $a\in C\langle E,F\rangle$.
-Similarly, for every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,b)\in INC(E\cup F,F)$ if and only if $a\in C\langle F,E\rangle$.
+For every ${b}\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $({b},{c})\in INC(E\cup F,E)$ if and only if ${b}\in C\langle E,F\rangle$.
+Similarly, for every ${b}\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $({b},{c})\in INC(E\cup F,F)$ if and only if ${b}\in C\langle F,E\rangle$.
 2.
-For every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,k)\in NUM(E\cup F,E)$.
-Here $k=1$ if and only if $a\in C\langle E,F\rangle$.
-Similarly, for every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,\ell)\in NUM(E\cup F,F)$.
-Here $\ell=1$ if and only if $a\in C\langle F,E\rangle$.
-Proof.
-By direct inspection of Definition 6.58
+For every ${b}\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $({b},k)\in NUM(E\cup F,E)$.
+Here $k=1$ if and only if ${b}\in C\langle E,F\rangle$.
+Similarly, for every ${b}\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $({b},\ell)\in NUM(E\cup F,F)$.
+Here $\ell=1$ if and only if ${b}\in C\langle F,E\rangle$.
 Definition 6.62.
-Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $a\in C\langle E\cup F,\emptyset\rangle$.
-We say that GTRS $R\langle E,F\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant $a\in C\langle E\cup F,\emptyset\rangle$ if
-for every rewrite rule of the form $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle{E\cup F},\emptyset\rangle$ with $m\in{\mathbb{N}}$,
-[…]
-there exists a rewrite rule in $R\langle E,F\rangle$ of the form $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b$ for some $b\in C\langle E,F\rangle$.
-Note that for the constant $b\in C\langle E,F\rangle$ appearing in the definition above, we have $b\subseteq a$.
-Definition 6.61 also defines the notion that GTRS $R\langle F,E\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant $a\in C\langle E\cup F,\emptyset\rangle$.
+Let $E$ and $F$ be GTESs over a signature $\Sigma$, and ${b}\in C\langle E\cup F,\emptyset\rangle$.
+We say that GTRS $R\langle E,F\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant ${b}\in C\langle E\cup F,\emptyset\rangle$ if
+for every rewrite rule of the form $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle{E\cup F},\emptyset\rangle$ with
+$m\in{\mathbb{N}}$, $\sigma\in\Sigma^{(m)}$, and ${b}_{1},\ldots,{b}_{m}\in C\langle E\cup F,\emptyset\rangle$,
+for every ${c}_{1},\ldots,{c}_{m}\in C\langle E,F\rangle$ with ${c}_{i}\subseteq{b}_{i}$ for every $i\in\{\,1,\ldots,m\,\}$,
+there exists a rewrite rule in $R\langle E,F\rangle$ of the form $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}$ for some ${c}\in C\langle E,F\rangle$.
+Note that for the constant ${c}\in C\langle E,F\rangle$ appearing in the definition above, we have ${c}\subseteq{b}$.
+Definition 6.62 also defines the notion that GTRS $R\langle F,E\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant ${b}\in C\langle E\cup F,\emptyset\rangle$.
 Definition 6.63.
 Let $E$ and $F$ be GTESs over a signature $\Sigma$.
-The auxiliary dpw for the GTESs $E$ and $F$ is $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, where $A[E;F]=BSTEP\langle E\cup F,\emptyset\rangle$.
+The auxiliary dpwpa for the GTESs $E$ and $F$ is $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, where $A[E;F]=BSTEP\langle E\cup F,\emptyset\rangle$.
 We assign the attributes
@@
 to the vertices of $C\langle E\cup F,\emptyset\rangle$. Furthermore, we assign the satellite data $counter_{E},counter_{F}:{\mathbb{N}}$ to the rewrite rules of $R\langle E\cup F,\emptyset\rangle$.
-For each vertex $a\in$, our decision algorithm computes the values of the above attributes such that the following conditions hold:
+For each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, our decision algorithm computes the values of the above attributes such that the following conditions hold:
   • 1.
-    $a.equal_{E}=\mathrm{true}$ if and only if $a\in C\langle E,F\rangle$.
+    ${b}.equal_{E}=\mathrm{true}$ if and only if ${b}\in C\langle E,F\rangle$.
   • 2.
-    $a.equal_{F}=\mathrm{true}$ if and only if $a\in C\langle F,E\rangle$.
+    ${b}.equal_{F}=\mathrm{true}$ if and only if ${b}\in C\langle F,E\rangle$.
   • 3.
-    $a.keeps_{E}=\mathrm{true}$ if and only if $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $a$.
+    ${b}.keeps_{E}=\mathrm{true}$ if and only if $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing b{b}.
   • 4.
-    $a.keeps_{F}=\mathrm{true}$ if and only if $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $a$.
-[…]
-Statement 6.63.
-Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then $|A[E;F]|\leq n$.
+    ${b}.keeps_{F}=\mathrm{true}$ if and only if $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing b{b}.
+Recall that for each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, list Adj(b)Adj({b}) consists of all vertices ${c}\in C\langle E\cup F,\emptyset\rangle$ such that there exists a rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$ in $R\langle E\cup F,\emptyset\rangle$ with $\sigma\in\Sigma_{m}$, $m\geq 1$, and ${c}={b}_{i}$ for some $i\in\{\,1,\ldots,m\,\}$.
+Statement 6.64.
+Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then $|A[E;F]|\leq 2n$.
 Proof.
-By Definition 6.62, Statement 4.45, we have $|A[E;F]|=|BSTEP\langle E\cup F,\emptyset\rangle|\leq n$. ∎
-Intuitively if $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$, then there exists a $1$-context $\zeta$ over $\Sigma$ such that such that
+By Definition 6.63, Statement 4.46, we have $|A[E;F]|=|BSTEP\langle E\cup F,\emptyset\rangle|\leq 2n$. ∎
+Intuitively, the basic idea of the proof of the following lemma is similar to that of the pumping lemma for regular languages [20], context-free languages [20], and regular tree languages [18], [21].
+The main difference from the proof of the pumping lemma is that we iterate contracting the ground term rather than pumping it.
+We stop when we get such a small ground term that we cannot contract it any more.
+If $R\langle E\cup F,\emptyset\rangle$ reaches a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$, then there exists a $1$-context $\zeta$ over $\Sigma$ such that such that
+  • 1.
     $\zeta\neq\diamondsuit$,
-  • $R\langle E\cup F,\emptyset\rangle$ reach $a$ from $\zeta[b]$, and
-  • $R\langle E\cup F,\emptyset\rangle$ reduces the subterms of $\zeta[b]$ at different prefixes of $\mathrm{addr}(\zeta)$ – except possibly for the pair of $\varepsilon$ and $\mathrm{addr}(\zeta)$ – to different constants in $C\langle E\cup F,\emptyset\rangle$.
+  • 2.
+    $R\langle E\cup F,\emptyset\rangle$ reaches b{b} from $\zeta[{c}]$, and
+  • 3.
+    $R\langle E\cup F,\emptyset\rangle$ reduces the subterms of $\zeta[{c}]$ at different prefixes of $\mathrm{addr}(\zeta)$ – except possibly for the pair of $\varepsilon$ and $\mathrm{addr}(\zeta)$ – to different constants in $C\langle E\cup F,\emptyset\rangle$.
 Thus the length of $\mathrm{addr}(\zeta)$ is less than equal to $|C\langle E\cup F,\emptyset\rangle|$.
@@
 Let $E$ and $F$ be GTESs over a signature $\Sigma$.
-Let $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$.
+Let $R\langle E\cup F,\emptyset\rangle$ reach a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$.
 Then there exists a $1$-context $\zeta$ over $\Sigma$, where
-  • (a)
+  • (b)
     $\zeta\neq\diamondsuit$,
-  • (b)
+  • (c)
     $k\geq 1$ denotes $\mathrm{length}(\mathrm{addr}(\zeta))$,
   • (c)
-[…]
+    for every $i\in\{\,0,\ldots,k\,\}$, $\alpha_{i}$ denotes the prefix of $\mathrm{addr}(\zeta)$ of length $i$
 such that
   • (d)
-    $\zeta[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=a$,
+    $\zeta[{c}]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}={b}$,
   • (e)
-    $\zeta[b]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}\neq\zeta[b]|_{\alpha_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$ for any $0\leq i<j\leq k$ such that $1\leq i$ or $j\leq k-1$, and
+    $\zeta[{c}]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}\neq\zeta[{c}]|_{\alpha_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$ for any $0\leq i<j\leq k$ such that $1\leq i$ or $j\leq k-1$, and
   • (f)
@@
 Proof.
-Intuitively, the basic idea of the proof is similar to that of the pumping lemma for regular languages [18], context-free languages [18], and regular tree languages [12], [19].
-The main difference from the proof of the pumping lemma is that we iterate contracting the ground term rather than pumping it.
-We stop when we get such a small ground term that we cannot contract it any more.
 It is sufficient to prove that there exists a $1$-context $\zeta$ over $\Sigma$ such that Conditions (a)–(e) hold.
 Because for any $1$-context $\zeta$ over $\Sigma$ satisfying Conditions (a)–(e), Condition (f) holds as well.
-By Statement 4.43, there exists a $1$-context $\delta$ over $\Sigma$ such that $\delta\neq\diamondsuit$ and $\delta[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=a$, i.e.
+By Statement 4.44, there exists a $1$-context $\delta$ over $\Sigma$ such that $\delta\neq\diamondsuit$ and $\delta[{c}]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}={b}$, i.e.
 Conditions (a)–(d) hold for $\zeta=\delta$.
@@
 Case 1: Conditions (e) holds for $\zeta=\delta$. Then we are done.
-[…]
-$\left(0\leq i<j\leq k-1\mbox{ or }1\leq i<j\leq k\right)$ and $\zeta[b]|_{\beta_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=\zeta[b]|_{\beta_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$.
+Case 2: Condition (e) does not hold for $\zeta=\delta$.
+For every $i\in\{\,0,\ldots,n\,\}$, let $\beta_{i}$ denote the prefix of $\mathrm{addr}(\delta)$ of length $i$.
+There exist $i,j\in\{\,0,\ldots,k\,\}$ such that
+$\left(0\leq i<j\leq k-1\mbox{ or }1\leq i<j\leq k\right)$ and $\zeta[{c}]|_{\beta_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=\zeta[{c}]|_{\beta_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$.
 We define $\delta^{\prime}\in CON_{\Sigma,1}$ from $\delta$ by replacing the subterm $\delta|_{\beta_{i}}$ at $\beta_{i}$ by the subterm $\delta|_{\beta_{j}}$ at $\beta_{j}$. Then
-$\delta^{\prime}\neq\diamondsuit$, $\mathrm{length}(\mathrm{addr}(\delta^{\prime}))<\mathrm{length}(\mathrm{addr}(\delta))$, and $\delta^{\prime}[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=a$ i.e.
+$\delta^{\prime}\neq\diamondsuit$, $\mathrm{length}(\mathrm{addr}(\delta^{\prime}))<\mathrm{length}(\mathrm{addr}(\delta))$, and $\delta^{\prime}[{c}]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}={b}$ i.e.
 Conditions (a)–(d) hold for $\zeta=\delta^{\prime}$.
@@
 ∎
-Intuitively, we now consider two constants, $a$ and $b$ in $C\langle E\cup F,\emptyset\rangle$ and a sequence of rewrite rules in $R\langle{E\cup F},\emptyset\rangle$ of which consecutive applications leads us from $b$ to $a$.
-Let $a,b\in C\langle E\cup F,\emptyset\rangle$, $k\geq 1$, and for each $j=1,\ldots,k$, let
-  • $\sigma_{j}(a_{j1},\ldots,a_{jm_{j}})\dot{\rightarrow}a_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$, be a rewrite rule and
-[…]
+Intuitively, we now consider two constants, b{b} and c{c} in $C\langle E\cup F,\emptyset\rangle$ and a sequence of rewrite rules in $R\langle{E\cup F},\emptyset\rangle$ of which consecutive applications leads us from c{c} to b{b}.
+Let ${b},{c}\in C\langle E\cup F,\emptyset\rangle$, $k\geq 1$, and for each $j=1,\ldots,k$, let
+  • 1.
+    $\sigma_{j}({b}_{j1},\ldots,{b}_{jm_{j}})\dot{\rightarrow}{b}_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$, be a rewrite rule and
+  • 2.
+    $i_{j}\in\{\,1,\ldots,m_{j}\,\}$ be an integer
 such that
-  • $b=a_{1i_{1}}$,
-  • for each $j=1,\ldots,k-1$, $a_{j}=a_{(j+1)i_{j+1}}$, and
-  • $a=a_{k}$.
+  • 1.
+    ${c}={b}_{1i_{1}}$,
+  • 2.
+    for each $j=1,\ldots,k-1$, ${b}_{j}={b}_{(j+1)i_{j+1}}$, and
+  • 3.
+    ${b}={b}_{k}$.
 The above sequence of rewrite rules is connected in the sense that we can apply them one after another, the right-hand side of each rule appears on the left-hand side of the next rule.
-Furthermore, the consecutive applications of these rewrite rules leads us from the constant $b$ to the constant $a$ since $b$ appears on the left-hand side of the first rule, and $a$ is the right-hand side of the last rule.
+Furthermore, the consecutive applications of these rewrite rules leads us from the constant c{c} to the constant b{b} since c{c} appears on the left-hand side of the first rule, and b{b} is the right-hand side of the last rule.
 We can apply the above rewrite rules one after another in the following way:
-$b$ appears in the left-hand side of the first rewrite rule,
-we apply the first rewrite rule to its left-hand side, $\sigma_{1}(a_{11},\ldots,a_{1m_{1}})$, we get the right-hand side $a_{1}$, which appears on the left-hand side of the second rewrite rule,
-we apply the second rewrite rule to its left-hand side, $\sigma_{2}(a_{21},\ldots,a_{2m_{2}})$, we get the right-hand side $a_{2}$, which appears on the left-hand side of the third rewrite rule,
-we apply the third rewrite rule to its left-hand side, $\sigma_{3}(a_{31},\ldots,a_{3m_{3}})$, we get the right-hand side $a_{3}$, which appears on the left-hand side of the fourth rewrite rule, and son on.
-Finally, the right-hand side $a_{k-1}$ of the last but one rewrite rule appears in the left-hand side of the last rewrite rule,
-we apply the last rewrite rule to its left-hand side, $\sigma_{k}(a_{k1},\ldots,a_{km_{k}})$, we get its right-hand side $a_{k}=a$.
-We now state formally and show the following intuitive statement: if $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$, then there exists a connected sequence of rewrite rules in $R\langle E\cup F,\emptyset\rangle$ that leads us from the constant $b$ to the constant $a$.
-Moreover, $b$ and the intermediate constants between $b$ and $a$, appearing as the right-hand sides of the applied rewrite rules except for the last one, are pairwise different.
+c{c} appears in the left-hand side of the first rewrite rule,
+we apply the first rewrite rule to its left-hand side, $\sigma_{1}({b}_{11},\ldots,{b}_{1m_{1}})$, we get the right-hand side ${b}_{1}$, which appears on the left-hand side of the second rewrite rule,
+we apply the second rewrite rule to its left-hand side, $\sigma_{2}({b}_{21},\ldots,{b}_{2m_{2}})$, we get the right-hand side ${b}_{2}$, which appears on the left-hand side of the third rewrite rule, and son on.
+Finally, the right-hand side ${b}_{k-1}$ of the last but one rewrite rule appears in the left-hand side of the last rewrite rule,
+we apply the last rewrite rule to its left-hand side, $\sigma_{k}({b}_{k1},\ldots,{b}_{km_{k}})$, we get its right-hand side ${b}_{k}={b}$.
+We now state formally and show the following intuitive statement: if $R\langle E\cup F,\emptyset\rangle$ reaches a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$, then there exists a connected sequence of rewrite rules in $R\langle E\cup F,\emptyset\rangle$ that leads us from the constant c{c} to the constant b{b}.
+Moreover, c{c} and the intermediate constants between c{c} and b{b}, appearing as the right-hand sides of the applied rewrite rules except for the last one, are pairwise different.
 Lemma 6.66.
-[…]
-  • $b=a_{1i_{1}}$,
-  • for each $j=1,\ldots,k-1$, $a_{j}=a_{(j+1)i_{j+1}}$,
-  • $b\not\in\{\,a_{1},\ldots,a_{k-1}\,\}$ and for any $11\leq i<j\leq k-$ 1, $a_{i}\neq a_{j}$, and
-  • $a=a_{k}$.
+Let $E$ and $F$ be GTESs over a signature $\Sigma$.
+Let $R\langle E\cup F,\emptyset\rangle$ reach a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$.
+Then there exists an integer $1\leq k\leq|C\langle E\cup F,\emptyset\rangle|$ and for each $j=1,\ldots,k$, there exists a rewrite rule $\sigma_{j}({b}_{j1},\ldots,{b}_{jm_{j}})\dot{\rightarrow}{b}_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$, and an integer $i_{j}\in\{\,1,\ldots,m_{j}\,\}$ such that
+  • 1.
+    ${c}={b}_{1i_{1}}$,
+  • 2.
+    for each $j=1,\ldots,k-1$, ${b}_{j}={b}_{(j+1)i_{j+1}}$,
+  • 3.
+    ${c}\not\in\{\,{b}_{1},\ldots,{b}_{k-1}\,\}$ and for any $1\leq i<j\leq k-1$, ${b}_{i}\neq{b}_{j}$, and
+  • 4.
+    ${b}={b}_{k}$.
 Proof.
 By Lemma 6.65, there exists a $1$-context $\zeta$ over $\Sigma$, where
-  • (a)
+  • (b)
     $\zeta\neq\diamondsuit$,
-  • (b)
+  • (c)
     $k\geq 1$ denotes $\mathrm{length}(\mathrm{addr}(\zeta))$,
   • (c)
-[…]
+    for every $i\in\{\,0,\ldots,k\,\}$, $\alpha_{i}$ denotes the prefix of $\mathrm{addr}(\zeta)$ of length $i$
 such that
   • (d)
-    $\zeta[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=a$,
+    $\zeta[{c}]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}={b}$,
   • (e)
-    $\zeta[b]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}\neq\zeta[b]|_{\alpha_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$ for any $0\leq i<j\leq k$ such that $1\leq i$ or $j\leq k-1$, and
+    $\zeta[{c}]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}\neq\zeta[{c}]|_{\alpha_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$ for any $0\leq i<j\leq k$ such that $1\leq i$ or $j\leq k-1$, and
   • (f)
     $k\leq|C\langle E\cup F,\emptyset\rangle|$.
-As $\mathrm{addr}(\xi)=\alpha_{k}$, we have $\xi[b]|_{\alpha_{k}}=b$.
-For each $i\in\{1,\ldots,k-1\,\}$, let $a_{i}=\zeta[b]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$.
-Since $b=\zeta[b]|_{\alpha_{k}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$, by Condition (e) we have
-$b\not\in\{\,a_{1},\ldots,a_{k-1}\,\}$ and for any $11\leq i<j\leq k-$ 1, $a_{i}\neq a_{j}$.
+As $\mathrm{addr}(\xi)=\alpha_{k}$, we have $\xi[{c}]|_{\alpha_{k}}={c}$.
+For each $i\in\{1,\ldots,k-1\,\}$, let ${b}_{i}=\zeta[{c}]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$.
+Since ${c}=\zeta[{c}]|_{\alpha_{k}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$, by Condition (e) we have
+${c}\not\in\{\,{b}_{1},\ldots,{b}_{k-1}\,\}$ and for any $1\leq i<j\leq k-1$, ${b}_{i}\neq{b}_{j}$.
 Then consider a sequence of $R\langle{E\cup F},\emptyset\rangle$ reductions, where we indicate the rewrite rules applied at the positions $\alpha_{k-1}$, $\alpha_{k-2}$, $\ldots$, $\alpha_{0}$:
-$\zeta[b]=t_{0}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{1}\mathop{\rightarrow}_{[\alpha_{k-1},\sigma_{1}(a_{11},\ldots,a_{1m_{1}})\dot{\rightarrow}a_{1}]}t_{1}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{2}\mathop{\rightarrow}_{[\alpha_{k-2},\sigma_{2}(a_{21},\ldots,a_{2m_{2}})\dot{\rightarrow}a_{2}]}t_{2}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}\newline \cdots\mathop{\rightarrow}_{[\alpha_{1},\sigma_{k-1}(a_{(k-1)1},\ldots,a_{(k-1)m_{1}})\dot{\rightarrow}a_{k-1}]}t_{k-1}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{k}\mathop{\rightarrow}_{[\alpha_{0},\sigma_{k}(a_{k1},\ldots,a_{km_{k}})\dot{\rightarrow}a_{k}]}a_{k}=a$.
+$\zeta[{c}]=t_{0}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{1}\mathop{\rightarrow}_{[\alpha_{k-1},\sigma_{1}({b}_{11},\ldots,{b}_{1m_{1}})\dot{\rightarrow}{b}_{1}]}t_{1}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{2}\mathop{\rightarrow}_{[\alpha_{k-2},\sigma_{2}({b}_{21},\ldots,{b}_{2m_{2}})\dot{\rightarrow}{b}_{2}]}t_{2}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}\newline \cdots\mathop{\rightarrow}_{[\alpha_{1},\sigma_{k-1}({b}_{(k-1)1},\ldots,{b}_{(k-1)m_{1}})\dot{\rightarrow}{b}_{k-1}]}t_{k-1}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{k}\mathop{\rightarrow}_{[\alpha_{0},\sigma_{k}({b}_{k1},\ldots,{b}_{km_{k}})\dot{\rightarrow}{b}_{k}]}{b}_{k}={b}$.
 Here
+  • 1.
     $s_{i}\in T_{\Sigma\cup C\langle E\cup F,\emptyset\rangle}$ for every $i\in\{1,\ldots,k\,\}$, and $t_{i}\in T_{\Sigma\cup C\langle E\cup F,\emptyset\rangle}$ for every $i\in\{0,\ldots,k-1\,\}$,
-  • for each $i\in\{1,\ldots,k\,\}$, $a_{i}=\zeta[b]|_{\alpha_{k-i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$,
-[…]
-  • for each $j\in\{1,\ldots,k\,\}$, $\sigma_{j}(a_{j1},\ldots,a_{jm_{j}})\dot{\rightarrow}a_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$,
-[…]
-  • $a=a_{k}$. ∎
+  • 2.
+    for each $i\in\{1,\ldots,k\,\}$, ${b}_{i}=\zeta[{c}]|_{\alpha_{k-i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$,
+  • 3.
+    ${c}={b}_{1i_{1}}$ for some $i_{1}\in\{\,1,\ldots,m_{1}\,\}$,
+  • 4.
+    for each $j\in\{1,\ldots,k\,\}$, $\sigma_{j}({b}_{j1},\ldots,{b}_{jm_{j}})\dot{\rightarrow}{b}_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$,
+  • 5.
+    for each $j=1,\ldots,k-1$, there exists $i_{j+1}\in\{\,1,\ldots,m_{j+1}\,\}$ such that ${b}_{j}={b}_{(j+1)i_{j+1}}$, and
+  • 6.
+    ${b}={b}_{k}$. ∎
+We have the following result by Lemma 6.66, Statement 4.44, and by the definition of $A[E;F]$ in Definition 6.63.
 Lemma 6.67.
 Let $E$ and $F$ be GTESs over a signature $\Sigma$.
-If $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$, then there exists a path of positive length or a cycle from the vertex $a$ to the vertex $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
-Proof.
-We have the lemma by Lemma 6.65, Statement 4.43, and by the definition of $A[E;F]$ in Definition 6.62. ∎
+If $R\langle E\cup F,\emptyset\rangle$ reaches a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$, then there exists a path of positive length or a cycle from the vertex b{b} to the vertex c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
 We now show the converse of Lemma 6.67.
-Intuitively, if there exists a connected sequence of rewrite rules in $R\langle E\cup F,\emptyset\rangle$ that leads us from a constant $b\in C\langle E\cup F,\emptyset\rangle$ to a constant $a\in C\langle E\cup F,\emptyset\rangle$, then $R\langle E\cup F,\emptyset\rangle$ reach the constant $a$ from a proper extension of the constant $b$.
-Lemma 6.67.
-[…]
-$\sigma(s_{1},\ldots,s_{i-1},a_{i},s_{i+1},\ldots,s_{m})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\sigma(a_{1},\ldots,a_{m})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a$.
-Proof.
-By Definition 4.26, Definition 4.28, and Proposition 4.32 we have the lemma. ∎
+Intuitively, if there exists a connected sequence of rewrite rules in $R\langle E\cup F,\emptyset\rangle$ that leads us from a constant ${c}\in C\langle E\cup F,\emptyset\rangle$ to a constant ${b}\in C\langle E\cup F,\emptyset\rangle$, then $R\langle E\cup F,\emptyset\rangle$ reaches the constant b{b} from a proper extension of the constant c{c}.
+By Definition 4.27, Definition 4.29, and Proposition 4.33 we have the following lemma.
 Lemma 6.68.
 Let $E$ and $F$ be GTESs over a signature $\Sigma$.
-Let $k\geq 1$, and $(a_{k},a_{k-1}),\,(a_{k-1},a_{k-2}),\,\ldots,\,(a_{1},a_{0})\in A[E;F]$.
-Then there exists a $1$-context $\delta$ over $\Sigma$ such that $\delta\neq\diamondsuit$ and $\delta[a_{0}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}a_{k}$.
+Let $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle{E\cup F},\emptyset\rangle$ be a rewrite rule, where $m\geq 1$, ${b}_{1},\ldots,{b}_{m},{b}\in C\langle E\cup F,\emptyset\rangle$.
+Let $i\in\{\,1,\ldots,m\,\}$.
+Then there exist ground terms $s_{1},\ldots,s_{i-1},s_{i+1},\ldots,s_{m}\in T_{\Sigma}$ such that
+$\sigma(s_{1},\ldots,s_{i-1},{b}_{i},s_{i+1},\ldots,s_{m})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\sigma({b}_{1},\ldots,{b}_{m})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{b}$.
+Lemma 6.69.
+Let $E$ and $F$ be GTESs over a signature $\Sigma$.
+Let $k\geq 1$, and $({b}_{k},{b}_{k-1}),\,({b}_{k-1},{b}_{k-2}),\,\ldots,\,({b}_{1},{b}_{0})\in A[E;F]$.
+Then there exists a $1$-context $\delta$ over $\Sigma$ such that $\delta\neq\diamondsuit$ and $\delta[{b}_{0}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}{b}_{k}$.
 Proof.
-By Definition 6.62 and Definition 4.44, for each $j=1,\ldots,k$, there exists a rewrite rule $\sigma_{j}(a_{j1},\ldots,a_{jm_{j}})\dot{\rightarrow}a_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$, such that
-[…]
+By Definition 6.63 and Definition 4.45, for each $j=1,\ldots,k$, there exists a rewrite rule $\sigma_{j}({b}_{j1},\ldots,{b}_{jm_{j}})\dot{\rightarrow}{b}_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$, such that
+${b}_{0}\in\{\,{b}_{11},\ldots,{b}_{1m_{1}}\,\}$ and for each $j=1,\ldots,k-1$, there exists $i_{j+1}\in\{\,1,\ldots,m_{j+1}\,\}$ such that ${b}_{j}={b}_{(j+1)i_{j+1}}$.
 We proceed by induction on $k$.
@@
 By the induction hypothesis, there exists a $1$-context $\delta$ over $\Sigma$ such that
-$\delta\neq\diamondsuit$ and $\delta[a_{0}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}a_{k}$.
+$\delta\neq\diamondsuit$ and $\delta[{b}_{0}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}{b}_{k}$.
 By Lemma 6.68, there exist ground terms $s_{1},\ldots,s_{i-1},s_{i+1},\ldots,s_{m_{k+1}}\in T_{\Sigma}$ such that
-$\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},a_{k},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})=\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},a_{(k+1)i_{k+1}},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\newline \sigma_{k+1}(a_{(k+1)1},\ldots,a_{(k+1)i_{k+1}-1},a_{(k+1)i_{k+1}},a_{(k+1)i_{k+1}+1},\ldots,a_{(k+1)m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a_{k+1}$.
+$\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},{b}_{k},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})=\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},{b}_{(k+1)i_{k+1}},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\newline \sigma_{k+1}({b}_{(k+1)1},\ldots,{b}_{(k+1)i_{k+1}-1},{b}_{(k+1)i_{k+1}},{b}_{(k+1)i_{k+1}+1},\ldots,{b}_{(k+1)m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{b}_{k+1}$.
 Thence we have
-$\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},\delta[a_{0}],s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},a_{k},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})=\newline \sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},a_{(k+1)i_{k+1}},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\newline \sigma_{k+1}(a_{(k+1)1},\ldots,a_{(k+1)i_{k+1}-1},a_{(k+1)i_{k+1}},a_{(k+1)i_{k+1}+1},\ldots,a_{(k+1)m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a_{k+1}$.
+$\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},\delta[{b}_{0}],s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},{b}_{k},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})=\newline \sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},{b}_{(k+1)i_{k+1}},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\newline \sigma_{k+1}({b}_{(k+1)1},\ldots,{b}_{(k+1)i_{k+1}-1},{b}_{(k+1)i_{k+1}},{b}_{(k+1)i_{k+1}+1},\ldots,{b}_{(k+1)m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{b}_{k+1}$.
 ∎
+We have the following statement by Lemma 6.69.
 Lemma 6.70.
 Let $E$ and $F$ be GTESs over a signature $\Sigma$.
-If there exists a path of positive length or a cycle from a vertex $a$ to a vertex $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E\cup F,\emptyset\rangle$ reach the constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of the constant $b\in C\langle E\cup F,\emptyset\rangle$.
-Proof.
-We have the lemma by Lemma 6.68. ∎
+If there exists a path of positive length or a cycle from a vertex b{b} to a vertex c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E\cup F,\emptyset\rangle$ reaches the constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of the constant ${c}\in C\langle E\cup F,\emptyset\rangle$.
+By Lemma 6.67 and Lemma 6.70, we have the following result.
 Lemma 6.71.
 Let $E$ and $F$ be GTESs over a signature $\Sigma$.
-$R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$ if and only if there is a path of positive length or a cycle from the vertex $a$ to the vertex $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
-Proof.
-By Lemma 6.66 and Lemma 6.69, we have the lemma. ∎
+$R\langle E\cup F,\emptyset\rangle$ reaches a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$ if and only if there is a path of positive length or a cycle from the vertex b{b} to the vertex c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
 7 Running Examples Further Continued
@@
 Example 7.72.
-We now continue Examples 3.17 and 5.50. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F]$ for the GTESs $E$ and $F$. Recall that
+We now continue Examples 3.18 and 5.51. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Recall that
 […]
-Observe that $A[E;F]=\{\,(f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$/_{\Theta\langle E\cup F,\emptyset\rangle})\,\}$.
+Observe that $A[E;F]=\{\,(f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\,\}$.
   • 1.
@@
   • 2.
-    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$,  \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, and
-    $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$.
-    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,  \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and
-    $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$.
+    $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$,  $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, and
+    $f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$.
+    $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,  $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and
+    $f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$.
   • 3.
-    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$,  \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}).keeps_{E}=\mathrm{true}$, and
-    $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}).keeps_{E}=\mathrm{true}$.
-    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$,  \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$, and
-    $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
+    $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$,  $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}).keeps_{E}=\mathrm{true}$, and
+    $f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}).keeps_{E}=\mathrm{true}$.
+    $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$,  $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$,   $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$, and
+    $f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
 Example 7.73.
-We now continue Examples 3.18 and 5.51. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Recall that
+We now continue Examples 3.19 and 5.52. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Recall that
 $C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;f(\#),\;f^{2}(\#),\;f^{3}(\#)\,\}\,\}=\{\,W\,\}$.
@@
   • 2.
-    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$   and   #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$.
+    $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$   and   $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$.
   • 3.
-    #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{false}$   and   #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
+    $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{false}$   and   $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
 Example 7.74.
-We now continue Example 3.21 and its sequel 5.54. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. By definition,
-$C\langle E\cup F,\emptyset\rangle,A[E;F]\rangle=C\langle E\cup F,\emptyset\rangle$.
+We now continue Example 3.22 and its sequel 5.55. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Recall that
+[…]
 Observe that
@@
 By direct inspection of $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$, we get that
-#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}\not\in C\langle E,F\rangle$ hence #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$,
-#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}\not\in C\langle F,E\rangle$ hence #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$.
-By direct inspection of $C\langle E,F\rangle$, $R\langle E,F\rangle$, $C\langle E\cup F,\emptyset\rangle$, and $R\langle E\cup F,\emptyset\rangle$, we get that $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$, hence #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$.
-The rewrite rule $f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ is in $R\langle E\cup F,\emptyset\rangle$ and $f(\#,\$)/_{\Theta\langle F,\emptyset\rangle}\subseteq\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-However, in $R\langle F,E\rangle$ there does not exist a rewrite rule with left-hand side $f(f(\#,\$)/_{\Theta\langle F,\emptyset\rangle},f(\#,\$)/_{\Theta\langle F,\emptyset\rangle})$.
-Hence $R\langle F,E\rangle$ does not keep up with $R\langle E\cup F,\emptyset\rangle$ writing #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
+$_{\Theta\langle E\cup F,\emptyset\rangle}\not\in C\langle E,F\rangle$ hence $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$,
+$_{\Theta\langle E\cup F,\emptyset\rangle}\not\in C\langle F,E\rangle$ hence $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$.
+By direct inspection of $C\langle E,F\rangle$, $R\langle E,F\rangle$, $C\langle E\cup F,\emptyset\rangle$, and $R\langle E\cup F,\emptyset\rangle$, we get that $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $_{\Theta\langle E\cup F,\emptyset\rangle}$, hence $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$.
+The rewrite rule $f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ is in $R\langle E\cup F,\emptyset\rangle$ and $f(\#,\mathdollar)/_{\Theta\langle F,E\rangle}\subseteq\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+However, in $R\langle F,E\rangle$ there does not exist a rewrite rule with left-hand side $f(f(\#,\mathdollar)/_{\Theta\langle F,E\rangle},f(\#,\mathdollar)/_{\Theta\langle F,E\rangle})$.
+Hence $R\langle F,E\rangle$ does not keep up with $R\langle E\cup F,\emptyset\rangle$ writing $_{\Theta\langle E\cup F,\emptyset\rangle}$, and $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
 Example 7.75.
-We now continue Examples 3.22 and 5.55. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Observe that
-$A[E;F]=\{\,(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(\$/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(\$/_{\Theta\langle E\cup F,\emptyset\rangle},\$/_{\Theta\langle E\cup F,\emptyset\rangle})\,\}$.
+We now continue Examples 3.23 and 5.56. We construct the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$. Observe that
+$A[E;F]=\{\,(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle}),\;(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle},\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\,\}$.
 By direct inspection of $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$, we get that
-#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,
-\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$,
-We now show that GTRS $R\langle F,\emptyset\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
-  • #$\#/_{\Theta\langle F,E\rangle}=\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-Furthermore, #$\#/_{\Theta\langle F,E\rangle}$ is the only element of $C\langle F,E\rangle$ which is a subset of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-  • For the rewrite rule #˙#$\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\in R\langle E\cup F,\emptyset\rangle$, there is a rewrite rule with left-hand side ##\## in $R\langle F,E\rangle$.
+$_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$,
+$\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$,
+We now show that GTRS $R\langle F,E\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing $_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
+  • 1.
+    $_{\Theta\langle F,E\rangle}=\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+Furthermore, $_{\Theta\langle F,E\rangle}$ is the only element of $C\langle F,E\rangle$ which is a subset of $_{\Theta\langle E\cup F,\emptyset\rangle}$.
+  • 2.
+    For the rewrite rule $\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\in R\langle E\cup F,\emptyset\rangle$, there is a rewrite rule with left-hand side #\# in $R\langle F,E\rangle$.
+  • 3.
     For the rewrite rule $f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\in R\langle E\cup F,\emptyset\rangle$, there is a rewrite rule with left-hand side $f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})$ in $R\langle F,E\rangle$.
-Consequently, GTRS $R\langle F,E\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing ##\##, and
-#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
-We now show that GTRS $R\langle F,E\rangle$ does not keep up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
-  • $g(\$/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\$/_{\Theta\langle E\cup F,\emptyset\rangle}\in R\langle E\cup F,\emptyset\rangle$,
-  • $g(g(\$))/_{\Theta\langle F,E\rangle})\subseteq\$/_{\Theta\langle E\cup F,\emptyset\rangle})$, and
-  • in $R\langle E\cup F,\emptyset\rangle$, there does not exist a rewrite rule with left-hand side $g(g(g(\$))/_{\Theta\langle F,E\rangle})$.
-Thus GTRS $R\langle F,E\rangle$ does not keep up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$, and
-\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
+Consequently, GTRS $R\langle F,E\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing #\#, and
+$_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
+We now show that GTRS $R\langle F,E\rangle$ does not keep up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that
+  • 1.
+    $g(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}\in R\langle E\cup F,\emptyset\rangle$,
+  • 2.
+    $g(g(\mathdollar))/_{\Theta\langle F,E\rangle})\subseteq\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle})$, and
+  • 3.
+    in $R\langle E\cup F,\emptyset\rangle$, there does not exist a rewrite rule with left-hand side $g(g(g(\mathdollar))/_{\Theta\langle F,E\rangle})$.
+Thus GTRS $R\langle F,E\rangle$ does not keep up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$, and
+$\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
 8 Constructing the Auxiliary dpwpa for the GTESs $E$ and $F$
-We construct the sets and relations and dpwpa, which we introduced in Section 6 and our decision procedure needs.
-[…]
+We adopt and slightly modify the fast ground completion algorithm of Fülöp and Vágvölgyi in Definition 4.29.
+This way, we produce the GTRSs $R\langle E,F\rangle$, GTRS $R\langle F,E\rangle$, and $R\langle E\cup F,\emptyset\rangle$, and construct the sets and relations and dpwpa, which we introduced in Section 6 and our decision procedure needs.
+We define a mapping $counter_{E}:R\langle E\cup F,\emptyset\rangle\rightarrow{\mathbb{N}}$, to each rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$, $m\in{\mathbb{N}}$, in $R\langle E\cup F,\emptyset\rangle$, we assign the number of rules $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}$ in $R\langle E,F\rangle$, where ${c}_{i}\subseteq{b}_{i}$ for every $i\in\{\,1,\ldots,m\,\}$.
+In pseudocode we treat the nonnegative integer $counter_{E}(\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b})$ as satellite data of the rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$, and denote it as $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}.counter_{E}$.
+In this way, we store the mapping $counter_{E}$.
+We define the mapping $counter_{F}:R\langle E\cup F,\emptyset\rangle\rightarrow{\mathbb{N}}$ and treat $counter_{F}(\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b})$ in a similar way.
 Definition 8.76.
@@
 Output:
+  1. 1.
     the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$;
+  2. 2.
     the set of nullary symbols $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$;
-  • $INC(E\cup F,E)$ stored without duplicates in a red-black tree saved in a variable $RBT1$, for every ordered pair$(a,b)\in INC\langle E\cup F,E\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
-  • $INC(E\cup F,F)$ stored in a red-black tree saved without duplicates in a variable $RBT2$, for every ordered pair$(a,b)\in INC(E\cup F,F)$, the search key is the first component $a$, the second component, $b$ is satellite data;
-  • $NUM(E\cup F,E)$ stored in a red-black tree saved without duplicates in a variable $RBT3$, for every ordered pair$(a,k)\in NUM(E\cup F,E)$, the search key is the first component $a$, the second component, $k$ is satellite data;
-  • $NUM(E\cup F,F)$ stored in a red-black tree saved without duplicates in a variable $RBT4$ for every ordered pair$(a,k)\in NUM(E\cup F,F)$, the search key is the first component $a$, the second component, $k$ is satellite data;
+  3. 3.
+[…]
+  4. 4.
+[…]
+  5. 5.
+[…]
+  6. 6.
+[…]
+  7. 7.
     the GTRS $R\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved without duplicates in a variable $RBT5$; each node in $RBT5$ contains a serach key which is a rule in $R\langle E\cup F,\emptyset\rangle$, and satellite data $counter_{E}$ and $counter_{F}$;
-  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved without duplicates in a variable $RBT6$; for every ordered pair$(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
+  8. 8.
+[…]
+  9. 9.
     the GTRS $R\langle E,F\rangle$ stored in a red-black tree saved without duplicates in a variable $RBT7$;
+  10. 10.
     the GTRS $R\langle F,E\rangle$ stored in a red-black tree saved without duplicates in a variable $RBT8$;
-  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$;
+  11. 11.
+    the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes ${b}.equal_{E}$, ${b}.equal_{F}$, ${b}.keeps_{E}$, and ${b}.keeps_{F}$;
 var $i$, $m$: integer; $\sigma$: symbol in $\Sigma$;
-$a,a_{1},\ldots,a_{mar_{\Sigma}}:C\langle E\cup F,\emptyset\rangle$;  $b,b_{1},\ldots,b_{mar_{\Sigma}}:C\langle E,F\rangle$;   $c,c_{1},\ldots,c_{mar_{\Sigma}}:C\langle F,E\rangle$;
+${b},{b}_{1},\ldots,{b}_{mar_{\Sigma}}:C\langle E\cup F,\emptyset\rangle$;  ${c},{c}_{1},\ldots,{c}_{mar_{\Sigma}}:C\langle E,F\rangle$;   ${d},{d}_{1},\ldots,{d}_{mar_{\Sigma}}:C\langle F,E\rangle$;
 $RBT1$, $RBT2$, $RBT3$, $RBT4$, $RBT6$, $RBT5$, $RBT7$, $RBT8$: red-black trees.
-1. Create a subterm dag $G=(V,A,\lambda)$ for $E$ and $F$.
+(Here comes the main part of the algorithm CAD.
+We run the fast ground completion algorithm in Definition 4.29 for each of GTESs $E$, $F$, and $E\cup F$, in this way we compute data items 1, 2, 7, 9, and 10 above.
+During the run of the algorithm CAD, it computes also the other data items listed above.
+Our algorithm consists of six steps:
+  1. 1.
+    We create a subterm dag $G=(V,A,\lambda)$ for $E$ and $F$.
+  2. 2.
+    We compute the relations $INC(E\cup F,E)$, $INC(E\cup F,F)$, $NUM(E\cup F,E)$, $NUM(E\cup F,F)$, $BSTEP\langle E\cup F,\emptyset\rangle$, the GTRSs $R\langle E\cup F,\emptyset\rangle$, $R\langle E,F\rangle$, $R\langle F,E\rangle$, the values of the attributes $equal_{E}$, $equal_{F}$ for every vertex in $C\langle E,F\rangle$, and the satellite data $counter_{E}$, and $counter_{F}$ for every rule in $R\langle E\cup F,\emptyset\rangle$.
+  3. 3.
+    We compute the relations $NUM(E\cup F,E)$ and $NUM(E\cup F,F)$.
+  4. 4.
+    We create the adjacency-list representation of the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$.
+  5. 5.
+    We compute the values of the attribute $keeps_{E}$.
+  6. 6.
+    We compute the values of the attribute $keeps_{F}$.)
+1.{ Create a subterm dag $G=(V,A,\lambda)$ for $E$ and $F$. }
 By Proposition 2.17, compute the congruence closure $\rho\langle E,F\rangle$ of the relation $\tau\langle E,F\rangle=\{\,(u,v)\mid(\hat{u},\hat{v})\in E\,\}.$
 $C\langle E,F\rangle:=\{\,x/_{\rho\langle E,F\rangle}\mid x\in V\,\}$.
-By Proposition 2.16, compute the congruence closure $\rho\langle F,E\rangle$ of the relation $\tau\langle F,E\rangle=\{\,(u,v)\mid(\hat{u},\hat{v})\in W\,\}.$
+By Proposition 2.17, compute the congruence closure $\rho\langle F,E\rangle$ of the relation $\tau\langle F,E\rangle=\{\,(u,v)\mid(\hat{u},\hat{v})\in F\,\}.$
 $C\langle F,E\rangle:=\{\,x/_{\rho\langle F,E\rangle}\mid x\in V\,\}$.
@@
 end
-for each $c\in C\langle E\cup F,\emptyset\rangle$ do
+for each ${d}\in C\langle E\cup F,\emptyset\rangle$ do
 begin
-$c.equal_{E}=\mathrm{\mathrm{false}};$;
-$c.equal_{F}=\mathrm{\mathrm{false}};$
+${d}.equal_{E}=\mathrm{\mathrm{false}};$
+${d}.equal_{F}=\mathrm{\mathrm{false}};$
 end
@@
 $m:=outdegree(x);$
-$a:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x)$;
-$b:=\mathrm{FIND}_{\rho\langle F,E\rangle}(x)$;
-$c:=\mathrm{FIND}_{\rho\langle E\cup F,\emptyset\rangle}(x)$;
-$INC(E\cup F,E):=INC(E\cup F,E)\cup\{\,(c,a)\,\}$;
-{ store $INC(E\cup F,E)$ without duplicates in the red-black tree $RBT1$.}
-$INC(E\cup F,F):=INC(E\cup F,F(\cup\{\,(c,b)\,\}$;
-{ store $INC(E\cup F,F)$ without duplicates in the red-black tree $RBT2$.}
-if $|a|=|c|$
-${\bf then}$ $c.equal_{E}:=\mathrm{\mathrm{true}};$
-if $|b|=|c|$
-${\bf then}$ $c.equal_{F}:=\mathrm{\mathrm{true}};$
-{ By Proposition 2.16 we know the cardinalities of all $W/_{\Theta\langle E,F\rangle}$-classes, $W/_{\Theta\langle F,E\rangle}$-classes, and $W/_{\Theta\langle E\cup F,\emptyset\rangle}$-classes. }
+${b}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x)$;
+${c}:=\mathrm{FIND}_{\rho\langle F,E\rangle}(x)$;
+${d}:=\mathrm{FIND}_{\rho\langle E\cup F,\emptyset\rangle}(x)$;
+$INC(E\cup F,E):=INC(E\cup F,E)\cup\{\,({d},{b})\,\}$;
+(store $INC(E\cup F,E)$ without duplicates in the red-black tree $RBT1$.)
+$INC(E\cup F,F):=INC(E\cup F,F)\cup\{\,({d},{c})\,\}$;
+(store $INC(E\cup F,F)$ without duplicates in the red-black tree $RBT2$.)
+if |b|=|d||{b}|=|{d}|
+$\bf then$} ${d}.equal_{E}:=\mathrm{\mathrm{true}};$
+if |c|=|d||{c}|=|{d}|
+$\bf then$} ${d}.equal_{F}:=\mathrm{\mathrm{true}};$
+(By Proposition 2.17 we know the cardinalities of all $W/_{\Theta\langle E,F\rangle}$-classes, $W/_{\Theta\langle F,E\rangle}$-classes, and $W/_{\Theta\langle E\cup F,\emptyset\rangle}$-classes.)
 for each $1\leq i\leq m$ do
 begin
-$a_{i}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x/i)$;
-$b_{i}:=\mathrm{FIND}_{\rho\langle F,E\rangle}(x/i)$;
-$c_{i}:=\mathrm{FIND}_{\rho\langle E\cup F,\emptyset\rangle}(x/i)$;
+${b}_{i}:=\mathrm{FIND}_{\rho\langle E,F\rangle}(x/i)$;
+${c}_{i}:=\mathrm{FIND}_{\rho\langle F,E\rangle}(x/i)$;
+${d}_{i}:=\mathrm{FIND}_{\rho\langle E\cup F,\emptyset\rangle}(x/i)$;
 end
-if $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c\not\in R\langle E\cup F,\emptyset\rangle$
+if $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}\not\in R\langle E\cup F,\emptyset\rangle$
 then begin
-$R\langle E\cup F,\emptyset\rangle:=R\langle E\cup F,\emptyset\rangle\cup\{\,\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c\,\}$;
-{ Store $R\langle E\cup F,\emptyset\rangle$ without duplicates in the red-black tree $RBT5$. }
-$\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}:=0$;
-$\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{F}:=0$;
-$BSTEP\langle E\cup F,\emptyset\rangle:=BSTEP\langle E\cup F,\emptyset\rangle\cup\{\,(c_{1},c),\ldots,(c_{m},c)\,\}$;
-{ Store $BSTEP\langle E\cup F,\emptyset\rangle$ without duplicates in the red-black tree $RBT6$}
+$R\langle E\cup F,\emptyset\rangle:=R\langle E\cup F,\emptyset\rangle\cup\{\,\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}\,\}$;
+(Store $R\langle E\cup F,\emptyset\rangle$ without duplicates in the red-black tree $RBT5$.)
+$\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{E}:=0$;
+$\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{F}:=0$;
+$BSTEP\langle E\cup F,\emptyset\rangle:=BSTEP\langle E\cup F,\emptyset\rangle\cup\{\,({d}_{1},{d}),\ldots,({d}_{m},{d})\,\}$;
+(Store $BSTEP\langle E\cup F,\emptyset\rangle$ without duplicates in the red-black tree $RBT6$)
 end
-if $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\not\in R\langle E,F\rangle$
+if $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\not\in R\langle E,F\rangle$
 then begin
-$R\langle E,F\rangle:=R\langle E,F\rangle\cup\{\,\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\,\}$;
-{ Store $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT7$. }
-$\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}:=\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}+1$;
+$R\langle E,F\rangle:=R\langle E,F\rangle\cup\{\,\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\,\}$;
+(Store $R\langle E,F\rangle$ without duplicates in the red-black tree $RBT7$.)
+$\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{E}:=\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{E}+1$;
 end
-if $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b\not\in R\langle F,E\rangle$
+if $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}\not\in R\langle F,E\rangle$
 then begin
-$R\langle F,E\rangle:=R\langle F,E\rangle\cup\{\,\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b\,\}$;
-{ Store $R\langle F,E\rangle$ without duplicates in the red-black tree $RBT8$. }
-$\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{F}:=\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{F}+1$;
+$R\langle F,E\rangle:=R\langle F,E\rangle\cup\{\,\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}\,\}$;
+(Store $R\langle F,E\rangle$ without duplicates in the red-black tree $RBT8$.)
+$\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{F}:=\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{F}+1$;
 end
 end
-3. { Computing the relations $NUM(E\cup F,E)$, $NUM(E\cup F,F)$. }
+3. { Computing the relations $NUM(E\cup F,E)$ and $NUM(E\cup F,F)$. }
 Read the nodes of the red-black tree $RBT1$ in search key order, during this process create the red-black tree $RBT3$ containing without duplicates the relation $NUM(E\cup F,E)$;
@@
 We build the adjacency-list representation of the dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ by reading in search key order the nodes of the black-red tree $RBT6$ storing the relation $BSTEP\langle E\cup F,\emptyset\rangle$.
-Recall that for every ordered pair $(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data.
+Recall that for every ordered pair $({b},{c})\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component b{b}, the second component, c{c} is satellite data.
 5. { Computing the values of the attribute $keeps_{E}$. }
-for each $a\in C\langle E\cup F,\emptyset\rangle$ do
-$a.keeps_{E}:=\mathrm{true}$;
-for each rewrite rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle$ with $m\in{\mathbb{N}}$ do
-if $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a.counter_{E}\neq\Pi_{i=1}^{m}|a_{i}/_{\Theta\langle E,F\rangle}|$
-then $a.keeps_{E}:=\mathrm{false}$
+for each ${b}\in C\langle E\cup F,\emptyset\rangle$ do
+${b}.keeps_{E}:=\mathrm{true}$;
+for each rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E\cup F,\emptyset\rangle$ with $m\in{\mathbb{N}}$ do
+if $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}.counter_{E}\neq\Pi_{i=1}^{m}|{b}_{i}/_{\Theta\langle E,F\rangle}|$
+then ${b}.keeps_{E}:=\mathrm{false}$
 6. { Computing the values of the attribute $keeps_{F}$. }
-for each $a\in C\langle E\cup F,\emptyset\rangle$ do
-$a.keeps_{F}:=\mathrm{true}$;
-for each rewrite rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle$ with $m\in{\mathbb{N}}$ do
-if $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a.counter_{F}\neq\Pi_{i=1}^{m}|a_{i}/_{\Theta\langle F,E\rangle}|$
-then $a.keeps_{F}:=\mathrm{false}$
+for each ${b}\in C\langle E\cup F,\emptyset\rangle$ do
+${b}.keeps_{F}:=\mathrm{true}$;
+for each rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E\cup F,\emptyset\rangle$ with $m\in{\mathbb{N}}$ do
+if $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}.counter_{F}\neq\Pi_{i=1}^{m}|{b}_{i}/_{\Theta\langle F,E\rangle}|$
+then ${b}.keeps_{F}:=\mathrm{false}$
 Lemma 8.77.
@@
 1. Algorithm CAD on $\Sigma$, $E$, and $F$ produces
+  • 1.
     the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$;
+  • 2.
     the set of nullary symbols $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$;
+  • 3.
     $INC(E\cup F,E)$ stored without duplicate values in a red-black tree saved in a variable $RBT1$;
+  • 4.
     $INC(E\cup F,F)$ stored without duplicate values in a red-black tree saved in a variable $RBT2$;
+  • 5.
     $NUM(E\cup F,E)$ stored without duplicate values in a red-black tree saved in a variable $RBT3$;
+  • 6.
     $NUM(E\cup F,F)$ stored without duplicate values in a red-black tree saved in a variable $RBT4$;
+  • 7.
     the GTRS $R\langle E\cup F,\emptyset\rangle$ stored without duplicate values in a red-black tree saved in a variable $RBT5$, each node in $RBT5$ contains a search key which is a rule in $R\langle E\cup F,\emptyset\rangle$, and satellite data $counter_{E}$ and $counter_{F}$;
-  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored without duplicate values in a red-black tree saved in a variable $RBT6$; for every ordered pair$(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
+  • 8.
+[…]
+  • 9.
     the GTRS $R\langle E,F\rangle$ stored without duplicate values in a red-black tree saved in a variable $RBT7$;
+  • 10.
     the GTRS $R\langle F,E\rangle$ stored without duplicate values in a red-black tree saved in a variable $RBT8$;
+  • 11.
     the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$.
 2.
-Algorithm CAD on $\Sigma$, $E$, and $F$ decides for every constant $a\in C\langle E\cup F,\emptyset\rangle$, whether $a\in C\langle E,F\rangle$ and whether $a\in C\langle F,E\rangle$, and sets the value of the attribute $a.equal_{E}$ to true if and only if $a\in C\langle E,F\rangle$, and sets the value of the attribute $a.equal_{F}$ to true if and only if $a\in C\langle F,E\rangle$.
-[…]
-[…]
+Algorithm CAD on $\Sigma$, $E$, and $F$ decides for every constant ${b}\in C\langle E\cup F,\emptyset\rangle$, whether ${b}\in C\langle E,F\rangle$ and whether ${b}\in C\langle F,E\rangle$, and sets the value of the attribute ${b}.equal_{E}$ to true if and only if ${b}\in C\langle E,F\rangle$, and sets the value of the attribute ${b}.equal_{F}$ to true if and only if ${b}\in C\langle F,E\rangle$.
+3.
+Algorithm CAD on $\Sigma$, $E$, and $F$ decides for any ${d}\in R\langle{E\cup F},\emptyset\rangle$ and any rewrite rule $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$ with right-hand side d{d}, whether for all ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$ such that ${b}_{i}\subseteq{d}_{i}$ for every $i\in\{\,1,\ldots,m\,\}$, there exists a rewrite rule in $R\langle E,F\rangle$ with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$.
+If so, then it sets the value of the attribute ${d}.keeps_{E}$ to $\mathrm{true}$, otherwise it sets the value of the attribute ${d}.keeps_{E}$ to $\mathrm{false}$.
+4.
+Algorithm CAD on $\Sigma$, $E$, and $F$ decides for any ${d}\in R\langle{E\cup F},\emptyset\rangle$ and any rewrite rule $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$ with right-hand side d{d}, whether for all ${c}_{1},\ldots,{c}_{m}\in C\langle F,E\rangle$ such that ${c}_{i}\subseteq{d}_{i}$ for every $i\in\{\,1,\ldots,m\,\}$, there exists a rewrite rule in $R\langle F,E\rangle$ with left-hand side $\sigma({c}_{1},\ldots,{c}_{m})$.
+If so, then it sets the value of the attribute ${d}.keeps_{F}$ to $\mathrm{true}$, otherwise it sets the value of the attribute ${d}.keeps_{F}$ to $\mathrm{false}$.
 Proof.
@@
 By Proposition 2.17, we know the cardinality of all $W/_{\Theta\langle E,F\rangle}$-classes, $W/_{\Theta\langle F,E\rangle}$-classes, and $W/_{\Theta\langle E\cup F,\emptyset\rangle}$-classes.
-For all $a\in R\langle E,F\rangle$ and $c\in R\langle E\cup F,\emptyset\rangle$ such that $a\cap c\neq\emptyset$, we have $a=c$ if and only if $|a|=|c|$.
+For all ${b}\in R\langle E,F\rangle$ and ${d}\in R\langle E\cup F,\emptyset\rangle$ such that ${b}\cap{d}\neq\emptyset$, we have b=d{b}={d} if and only if |b|=|d||{b}|=|{d}|.
 Hence we have Statement 2.
 We now show Statement 3.
-For every rewrite rule $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c\in R\langle E\cup F,\emptyset\rangle$, we decide whether for all $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$ with $a_{1}\subseteq c_{1},\ldots,a_{m}\subseteq c_{m}$, there exists a rewrite rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R\langle E,F\rangle$.
-When constructing the GTRS $R\langle E,F\rangle$, whenever we add to $R\langle E,F\rangle$ a new rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$, where $a_{1}\subseteq a_{1},\ldots,a_{m}\subseteq c_{m}$, we add $1$ to the attribute $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}$.
-In this way, we count the rules in $R\langle E,F\rangle$ with left-hand side $\sigma(a_{1},\ldots,a_{m})$, where $a_{1}\subseteq c_{1},\ldots,a_{m}\subseteq c_{m}$, and store the result in the attribute $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}$.
-Then $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c.counter_{E}=\Pi_{i=1}^{m}|c_{i}/_{\Theta\langle E,F\rangle}|$ if and only if for all $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$ with $a_{1}\subseteq c_{1},\ldots,a_{m}\subseteq c_{m}$, there exists a rewrite rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R\langle E,F\rangle$.
-When Algorithm CAD terminates, for every vertex $c\in C\langle E\cup F,\emptyset\rangle$,
-$c.keeps_{E}=\mathrm{true}$ if and only if
-for every rewrite rule of the form $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c\in R\langle{E\cup F},\emptyset\rangle$,
-[…]
-there exists a rewrite rule in $R\langle E,F\rangle$ of the form $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a$ for some $a\in C\langle E,F\rangle$.
+For every rewrite rule $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}\in R\langle E\cup F,\emptyset\rangle$, we decide whether for all ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$ with ${b}_{1}\subseteq{d}_{1},\ldots,{b}_{m}\subseteq{d}_{m}$, there exists a rewrite rule with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$ in $R\langle E,F\rangle$.
+When constructing the GTRS $R\langle E,F\rangle$, whenever we add to $R\langle E,F\rangle$ a new rule with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$, where ${b}_{1}\subseteq{b}_{1},\ldots,{b}_{m}\subseteq{d}_{m}$, we add $1$ to the attribute $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{E}$.
+In this way, we count the rules in $R\langle E,F\rangle$ with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$, where ${b}_{1}\subseteq{d}_{1},\ldots,{b}_{m}\subseteq{d}_{m}$, and store the result in the attribute $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{E}$.
+Then $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}.counter_{E}=\Pi_{i=1}^{m}|{d}_{i}/_{\Theta\langle E,F\rangle}|$ if and only if for all ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$ with ${b}_{1}\subseteq{d}_{1},\ldots,{b}_{m}\subseteq{d}_{m}$, there exists a rewrite rule with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$ in $R\langle E,F\rangle$.
+When Algorithm CAD terminates, for every vertex ${d}\in C\langle E\cup F,\emptyset\rangle$,
+${d}.keeps_{E}=\mathrm{true}$ if and only if
+for every rewrite rule of the form $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$,
+for every ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$ with ${b}_{i}\subseteq{d}_{i}$ for all $i\in\{\,1,\ldots,m\,\}$,
+there exists a rewrite rule in $R\langle E,F\rangle$ of the form $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$ for some ${b}\in C\langle E,F\rangle$.
 The proof of Satement 4 is similar to that of Statement 3. ∎
+Statement 8.78.
+1. We can decide in $O(n)$ time whether for all ${b}\in C\langle E\cup F,\emptyset\rangle$, ${b}\in C\langle E,F\rangle$.
+2. We can decide in $O(n)$ time whether for all ${b}\in C\langle E\cup F,\emptyset\rangle$, ${b}\in C\langle F,E\rangle$.
+Proof.
+First we show Statement 1.
+We read each nodes of the red-black tree $RBT3$ in lexicographic order.
+For every ${b}\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $({b},k)\in NUM(E\cup F,E)$.
+If $k=1$, then by Statement 6.61, ${b}\in C\langle E,F\rangle$.
+Otherwise, ${b}\not\in C\langle E,F\rangle$.
+We can show Statement 2 similarly to Statement 1. ∎
 Lemma 8.79.
-Let $E$ and $F$ be GTESs over a signature $\Sigma$. Algorithm CAD runs in $O(n\,\mathrm{log}\,n)$ time on $\Sigma$, $E$, and $F$, where $n=size(E)+size(F)$.
+Let $E$ and $F$ be GTESs over a signature $\Sigma$. Algorithm CAD runs in $O(n\,\mathrm{log}\,n)$ time on $\Sigma$, $E$, and $F$, where $n=\mathrm{size}(E)+\mathrm{size}(F)$.
 Proof.
-The proof consists of showing the following nine statements:
-1. The time complexity of constructing the red-black trees $RBT1$ and $RBT2$ is $O(n)$.
-2.
-The time complexity of constructing the relation $\tau\langle E\cup F,\emptyset\rangle$ and the GTES $E\cup F$ stored without duplicate values in a red-black tree saved in the variable $RBT5$, is $O(n^{2})$ time.
-3. The time complexity of computing the value attributes $a.equal_{E}$ and $a.equal_{F}$ for all constants $a\in C\langle E\cup F,\emptyset\rangle$, is $O(n)$.
-4. The time complexity of constructing the following data is $O(n\,\mathrm{log}\,n)$.
-  • the GTRS $R\langle E\cup F,\emptyset\rangle$ stored without duplicate values in a red-black tree saved in the variable $RBT5$, each node in $RBT5$ contains a search key which is a rule in $R\langle E\cup F,\emptyset\rangle$;
-  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in the variable $RBT6$;
-  • the GTRS $R\langle E,F\rangle$ stored without duplicate values in a red-black tree saved in the variable $RBT7$;
-  • the GTRS $R\langle F,E\rangle$ stored without duplicate values in a red-black tree saved in the variable $RBT8$;
-5. The time complexity of constructing
-  • $NUM(E\cup F,E)$ stored without duplicate values in a red-black tree saved in a variable $RBT3$, and
-  • $NUM(E\cup F,F)$ stored without duplicate values in a red-black tree saved in a variable $RBT4$
-is $O(n)$.
-6. The time complexitity of computing the values of the satellite data $counter_{E}$, and $counter_{F}$ for all rules in $R\langle E\cup F,\emptyset\rangle$ is $O(n)$.
-7. The time complexity of constructing the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ is $O(n)$.
-8. The time complexitity of computing the values of the attributes $equal_{E}$ and $equal_{F}$ of all constants $a\in C\langle E\cup F,\emptyset\rangle$ is $O(n)$.
-9. The time complexitity of computing the values of the attributes $keeps_{E}$ and $keeps_{F}$ of all constants $a\in C\langle E\cup F,\emptyset\rangle$ is $O(n\,\mathrm{log}\,n)$ time.
-First we now show Statement 1.
-By Statement 6.59, $INC(E\cup F,E)leqn$.
+We show Statements 8.808.88.
+Statement 8.80.
+The time complexity of constructing the red-black trees $RBT1$ and $RBT2$ is $O(n)$.
+Proof.
+By Statement 6.60, $INC(E\cup F,E)\leq n$.
 Consequently, when we construct the red-black tree $RBT1$, we carry out at most $n$ insertion.
@@
 Similarly, we construct the red-black tree $RBT2$ in $O(n\,\mathrm{log}\,n)$ time.
-Second, by Statement 2.4, we have Statement 2.
-Third, we carry out $2n$ comparisons of nonnegative numbers in $O(n)$ time.
-Fourth, By Statement 2 and Statement 3 of 4.47, we have Statement 4.
-Fifth, we now show Statement 5.
+∎
+By Statement 2.5, we have the following result.
+Statement 8.81.
+The time complexity of constructing the relation $\tau\langle E\cup F,\emptyset\rangle$ and the GTES $E\cup F$ stored without duplicate values in a red-black tree saved in the variable $RBT5$, is $O(n^{2})$ time.
+We carry out $2n$ comparisons of nonnegative numbers in $O(n)$ time. Hence we have the following fact.
+Statement 8.82.
+The time complexity of computing the value attributes ${b}.equal_{E}$ and ${b}.equal_{F}$ for all constants ${b}\in C\langle E\cup F,\emptyset\rangle$, is $O(n)$.
+By Statement 2 and Statement 3 of 4.48, we have the following fact.
+Statement 8.83.
+The time complexity of constructing the following data is $O(n\,\mathrm{log}\,n)$.
+  • 1.
+    the GTRS $R\langle E\cup F,\emptyset\rangle$ stored without duplicate values in a red-black tree saved in the variable $RBT5$, each node in $RBT5$ contains a search key which is a rule in $R\langle E\cup F,\emptyset\rangle$;
+  • 2.
+    $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in the variable $RBT6$;
+  • 3.
+    the GTRS $R\langle E,F\rangle$ stored without duplicate values in a red-black tree saved in the variable $RBT7$;
+  • 4.
+    the GTRS $R\langle F,E\rangle$ stored without duplicate values in a red-black tree saved in the variable $RBT8$;
+Statement 8.84.
+The time complexity of constructing
+  • 1.
+    $NUM(E\cup F,E)$ stored without duplicate values in a red-black tree saved in a variable $RBT3$, and
+  • 2.
+    $NUM(E\cup F,F)$ stored without duplicate values in a red-black tree saved in a variable $RBT4$
+is $O(n)$.
+Proof.
 When we construct the red-black tree $RBT3$, we read the whole red-black tree $RBT1$ in search key order in $O(n)$ time.
 Each modification takes constant time.
-By Statement 6.59, $|INC(E\cup F,E)|\leq n$.
+By Statement 6.60, $|INC(E\cup F,E)|\leq n$.
 Consequently, we construct the red-black tree $RBT3$ in $O(n)$ time.
 Similarly, we construct the red-black tree $RBT4$ in $O(n)$ time.
-Sixth, we now show Statement 6.
+∎
+Statement 8.85.
+The time complexitity of computing the values of the satellite data $counter_{E}$, and $counter_{F}$ for all rules in $R\langle E\cup F,\emptyset\rangle$ is $O(n)$.
+Proof.
 When Algorithm CAD computes the values of the satellite data $counter_{E}$ and $counter_{F}$ for all rules of the GTRS $R\langle E\cup F,\emptyset\rangle$, it carries out at most $2n$ additions, since $|C\langle E,F\rangle|\leq|W|\leq n$ and $|C\langle F,E\rangle|\leq|W|\leq n$.
-Seventh we now show Statement 7.
+∎
+Statement 8.86.
+The time complexity of constructing the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ is $O(n)$.
+Proof.
 We build the adjacency-list representation of the dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ by reading in search key order the nodes of the black-red tree $RBT6$ storing the relation $BSTEP\langle E\cup F,\emptyset\rangle$.
-Recall that for every ordered pair $(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data.
+Recall that for every ordered pair $({b},{c})\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component b{b}, the second component, c{c} is satellite data.
 Then the time complexity of building the adjacency-list representation of the dpwpa $AUX[E;F]$ is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$.
-By Statement 4.27 and Statement 4.45 $|C\langle E\cup F,\emptyset\rangle|\leq|W|\leq n$.
-By Statement 6.63, $|A[E;F]|=|BSTEP\langle E\cup F,\emptyset\rangle|\leq size(E)+size(F)=n$.
+By Statement 4.28, $|C\langle E\cup F,\emptyset\rangle|\leq|W|\leq n$.
+By Statement 6.64, $|A[E;F]|=|BSTEP\langle E\cup F,\emptyset\rangle|\leq\mathrm{size}(E)+size(F)=2n$.
 Therefore the time complexity of building the adjacency-list representation of the dpwpa $AUX[E;F]$ is $O(n)$.
-Eighth, Statement 8.78 implies Statement 8.
-Nineth, we now show Statement 9.
-We now estimate the number of operations when computing the values of the attributes $a.keeps_{E}$ for all $a\in C\langle E\cup F,\emptyset\rangle$.
-For each rewrite rule $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle$, Algorithm CAD carries out $m$ searches in the red-black tree $RBT3$ and less than or equal to $m$ multiplications.
-By Statement 4.45, $size(R\langle E\cup F,\emptyset\rangle)\leq 2\cdot n$.
-In this way, when we compute the values of the attribute $keeps_{E}$ for all rewrite rules $\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle$, the number of searches in the red-black tree $RBT3$ is less than equal to $n$.
+∎
+Statement 8.78 implies the following fact.
+Statement 8.87.
+The time complexitity of computing the values of the attributes $equal_{E}$ and $equal_{F}$ of all constants ${b}\in C\langle E\cup F,\emptyset\rangle$ is $O(n)$.
+Statement 8.88.
+The time complexitity of computing the values of the attributes $keeps_{E}$ and $keeps_{F}$ of all constants ${b}\in C\langle E\cup F,\emptyset\rangle$ is $O(n\,\mathrm{log}\,n)$ time.
+Proof.
+We now estimate the number of operations when computing the values of the attributes ${b}.keeps_{E}$ for all ${b}\in C\langle E\cup F,\emptyset\rangle$.
+For each rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E\cup F,\emptyset\rangle$, Algorithm CAD carries out $m$ searches in the red-black tree $RBT3$ and less than or equal to $m$ multiplications.
+By Statement 4.46, $\mathrm{size}(R\langle E\cup F,\emptyset\rangle)\leq 2n$.
+In this way, when we compute the values of the attribute $keeps_{E}$ for all rewrite rules $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E\cup F,\emptyset\rangle$, the number of searches in the red-black tree $RBT3$ is less than equal to $n$.
 Consequently, by Proposition 2.1, the searches in the red-black tree $RBT3$ take $O(n\,\mathrm{log}\,n)$ time.
-Furthermore,
-the number of multiplications $\leq\sum_{\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E\cup F,\emptyset\rangle}m\leq size(R\langle E\cup F,\emptyset\rangle)\leq 2n$
-by Statement 4.45. Thus, the number of multiplications is less than or equal to $n$.
-For each $a\in C\langle E\cup F,\emptyset\rangle$, the number of comparisons of nonnegative integers during computing the value of the attribute $a.keeps_{E}$ is equal to the number of rules in $R\langle E\cup F,\emptyset\rangle$ with right-hand side $a$.
-Consequently, the number of comparisons of nonnegative integers during computing the values of the attributes $a.keeps_{E}$ for all $a\in C\langle E\cup F,\emptyset\rangle$ is equal to $|R\langle E\cup F,\emptyset\rangle|$.
-By Statement 4.45
-$|R\langle E\cup F,\emptyset\rangle|\leq n$.
-Consequently the time complexity of computing for all $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attribute $a.keeps_{E}$ is $O(n\,\mathrm{log}\,n)$.
-Similarly, the time complexity of computing for all $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attribute $a.keeps_{F}$ is $O(n\,\mathrm{log}\,n)$.
-The Lemma follows from Statements 1, 2, 3, 4, 5, 6, 7, 8, and 9. ∎
-Statement 8.78.
-1. We can decide in $O(n)$ time whether for all $a\in C\langle E\cup F,\emptyset\rangle$, $a\in C\langle E,F\rangle$.
-2. We can decide in $O(n)$ time whether for all $a\in C\langle E\cup F,\emptyset\rangle$, $a\in C\langle F,E\rangle$.
-Proof.
-First we show Statement 1.
-We read each nodes of the red-black tree $RBT3$ in lexicographic order.
-For every $a\in C\langle E\cup F,\emptyset\rangle$, there exists a unique ordered pair $(a,k)\in NUM(E\cup F,E)$.
-If $k=1$, then by Statement 6.60, $a\in C\langle E,F\rangle$.
-Otherwise, $a\not\in C\langle E,F\rangle$.
-We can show Statement 2 similarly to Statement 1. ∎
+Furthermore, the number of multiplications is less than or equal to $\sum_{\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E\cup F,\emptyset\rangle}m\leq\mathrm{size}(R\langle E\cup F,\emptyset\rangle)\leq 2n$ by Statement 4.46.
+Thus, the number of multiplications is less than or equal to $2n$.
+For each ${b}\in C\langle E\cup F,\emptyset\rangle$, the number of comparisons of nonnegative integers during computing the value of the attribute ${b}.keeps_{E}$ is equal to the number of rules in $R\langle E\cup F,\emptyset\rangle$ with right-hand side b{b}.
+Consequently, the number of comparisons of nonnegative integers during computing the values of the attributes ${b}.keeps_{E}$ for all ${b}\in C\langle E\cup F,\emptyset\rangle$ is equal to $|R\langle E\cup F,\emptyset\rangle|$.
+By Statement 4.46 $|R\langle E\cup F,\emptyset\rangle|\leq n$.
+Consequently the time complexity of computing for all ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attribute ${b}.keeps_{E}$ is $O(n\,\mathrm{log}\,n)$.
+Similarly, the time complexity of computing for all ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attribute ${b}.keeps_{F}$ is $O(n\,\mathrm{log}\,n)$.
+∎
+The Lemma follows from Statements 8.808.88. ∎
 9 Preparatory Results
@@
 Proposition 9.89.
-Let $A$ be a set, $a\in A$, and let $\rho,\tau,\omega$ be equivalence relations on $A$ such that $a/_{\rho}\subset a/_{\omega}$ and $a/_{\tau}\subset a/_{\omega}$.
-Then $a/_{\rho}\cup a/_{\tau}\subset a/_{\omega}$.
+Let $A$ be a set, ${b}\in A$, and let $\rho,\tau,\omega$ be equivalence relations on $A$ such that ${b}/_{\rho}\subset{b}/_{\omega}$ and ${b}/_{\tau}\subset{b}/_{\omega}$.
+Then ${b}/_{\rho}\cup{b}/_{\tau}\subset{b}/_{\omega}$.
 Proof.
-There exist $b,c\in A$, such that $b\in a/_{\omega}\setminus a/_{\rho}$ and $c\in a/_{\omega}\setminus a/_{\tau}$. We now distinguish three cases.
-Case 1: $(b,c)\in\rho$. Then by the definition of $b$, $(a,c)\not\in\rho$. By the definition of $c$, $(a,c)\not\in\tau$. Thence $(a,c)\not\in\rho\cup\tau$.
-Case 2: $(b,c)\in\tau$. Then by the definition of $c$, $(a,b)\not\in\tau$. By the definition of $b$, $(a,b)\not\in\rho$. Consequently $(a,b)\not\in\rho\cup\tau$.
-Case 3: $(b,c)\not\in\rho\cup\tau$. Then by the definition of $b$ and $c$, $(b,c)\in\omega$, and hence $(b,c)\in\omega\setminus(\rho\cup\tau)$.
-In all three cases we have $a/_{\rho}\cup a/_{\tau}\subset a/_{\omega}$. ∎
+There exist ${c},{d}\in A$, such that ${c}\in{b}/_{\omega}\setminus{b}/_{\rho}$ and ${d}\in{b}/_{\omega}\setminus{b}/_{\tau}$. We now distinguish three cases.
+Case 1: $({c},{d})\in\rho$. Then by the definition of c{c}, $({b},{d})\not\in\rho$. By the definition of d{d}, $({b},{d})\not\in\tau$. Thence $({b},{d})\not\in\rho\cup\tau$.
+Case 2: $({c},{d})\in\tau$. Then by the definition of d{d}, $({b},{c})\not\in\tau$. By the definition of c{c}, $({b},{c})\not\in\rho$. Consequently $({b},{c})\not\in\rho\cup\tau$.
+Case 3: $({c},{d})\not\in\rho\cup\tau$. Then by the definition of c{c} and d{d}, $({c},{d})\in\omega$, and hence $({c},{d})\in\omega\setminus(\rho\cup\tau)$.
+In all three cases we have ${b}/_{\rho}\cup{b}/_{\tau}\subset{b}/_{\omega}$. ∎
 Proposition 9.90.
-Let $A$ be a set, $a\in A$, and let $\rho,\tau,\omega$ be equivalence relations on $A$ such that $a/_{\rho}\cup a/_{\tau}=a/_{\omega}$. Then $a/_{\rho}=a/_{\omega}$ or $a/_{\tau}=a/_{\omega}$.
+Let $A$ be a set, ${b}\in A$, and let $\rho,\tau,\omega$ be equivalence relations on $A$ such that ${b}/_{\rho}\cup{b}/_{\tau}={b}/_{\omega}$.
+Then ${b}/_{\rho}={b}/_{\omega}$ or ${b}/_{\tau}={b}/_{\omega}$.
 Proof.
-By contradiction, assume that $a/_{\rho}\neq a/_{\omega}$ and $a/_{\tau}\neq a/_{\omega}$.
-Then $a/_{\rho}\subset a/_{\omega}$ and $a/_{\tau}\subset a/_{\omega}$.
-Consequently, by Proposition 9.79, $a/_{\rho}\cup a/_{\tau}\subset a/_{\omega}$.
+By contradiction, assume that ${b}/_{\rho}\neq{b}/_{\omega}$ and ${b}/_{\tau}\neq{b}/_{\omega}$.
+Then ${b}/_{\rho}\subset{b}/_{\omega}$ and ${b}/_{\tau}\subset{b}/_{\omega}$.
+Consequently, by Proposition 9.89, ${b}/_{\rho}\cup{b}/_{\tau}\subset{b}/_{\omega}$.
 ∎
@@
 Proof.
-[…]
+Let $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, and ${b}_{1},\ldots,{b}_{m}\in C\langle E\cup F,\emptyset\rangle$.
+Then there exist $t_{1},\ldots,t_{m}\in W$ such that for every $i\in\{\,1,\ldots,m\,\}$, $t_{i}/_{\Theta\langle E\cup F,\emptyset\rangle}={b}_{i}$.
+As GTRS $R\langle E,F\rangle$ is total, there exists a rewrite rule $\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})\dot{\rightarrow}t/_{\Theta\langle E,F\rangle}$ in $R\langle E,F\rangle$.
+Accordingly, by the definition of $R\langle E,F\rangle$, $\sigma(t_{1},\ldots,t_{m})\in W$ and $\sigma(t_{1},\ldots,t_{m})\mathop{\leftrightarrow}^{*}_{E}t$.
+Then $\sigma(t_{1},\ldots,t_{m})\mathop{\leftrightarrow}^{*}_{E\cup F}t$.
+For every $i\in\{\,1,\ldots,m\,\}$, $t_{i}/_{\Theta\langle E\cup F,\emptyset\rangle}\in C\langle E\cup F,\emptyset\rangle$.
+Thus, by the definition of $R\langle E\cup F,\emptyset\rangle$, the rewrite rule $\sigma(t_{1}/_{\Theta\langle E\cup F,\emptyset\rangle},\ldots,t_{m}/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}t/_{\Theta\langle E\cup F,\emptyset\rangle}$ is in $R\langle E\cup F,\emptyset\rangle$.
+∎
 Proposition 9.92.
@@
 $(2\Rightarrow 3)$ For the GTES $E\cup F$, we have $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
-$(3\Rightarrow 4)$ $\mathop{\leftrightarrow}^{*}_{H}$ is a congrunce on the ground term algebra ${\cal T}(\Sigma)$, hence $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$.
+$(3\Rightarrow 4)$ $\mathop{\leftrightarrow}^{*}_{H}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$, hence $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$.
 $(4\Rightarrow 1)$ $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$.
@@
 $s=q_{0}\mathop{\leftrightarrow}_{[\alpha_{1},(l_{1},r_{1})]}q_{1}\mathop{\leftrightarrow}_{[\alpha_{2},(l_{2},r_{2})]}q_{2}\mathop{\leftrightarrow}_{[\alpha_{3},(l_{3},r_{3})]}\cdots\mathop{\leftrightarrow}_{[\alpha_{n},(l_{n},r_{n})]}q_{n}=t$ for some $n\in{\mathbb{N}}$,
-[…]
+where for every $i\in\{\,0,\ldots,n\,\}$, $q_{i}\in T_{\Sigma}$ and for every $i\in\{\,1,\ldots,n\,\}$, $l_{i}\doteq r_{i}\in E\cup F$. We now show that
 $s\mathop{\leftrightarrow}^{*}_{E}t$ or $s\mathop{\leftrightarrow}^{*}_{F}t$.
-[…]
+If $n=0$, then $s=t$ and hence $s\mathop{\leftrightarrow}^{*}_{E}t$.
+Let us assume that $n\geq 1$.
+Let $\alpha_{i}$ be a shortest word in the set $\{\,\alpha_{1},\alpha_{2},\ldots,\alpha_{n}\,\}$, where $i\in\{\,1,\ldots,n\,\}$.
+Then there exist $\delta\in CON_{\Sigma,1}$, $v,z\in T_{\Sigma}$ such that
+  • 1.
     $addr(\delta)=\alpha_{i}$,
+  • 2.
     $s=\delta[v]$, $t=\delta[z]$, and
+  • 3.
     $q_{i-1}=\delta[l_{i}]$, $q_{i}=\delta[r_{i}]$.
@@
 $s=\delta[v]\mathop{\leftrightarrow}^{*}_{E}\delta[z]=t$ or $s=\delta[v]\mathop{\leftrightarrow}^{*}_{F}\delta[z]=t$. ∎
+By Proposition 2.11 and Proposition 2.12 we have the following result.
 Proposition 9.94.
-Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=size(E)+size(F)$. Let $W=ST\langle E\cup F\rangle$. We can decide in $O(n^{3})$ time whether for each $p\in W$,
+Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=\mathrm{size}(E)+\mathrm{size}(F)$. Let $W=ST\langle E\cup F\rangle$. We can decide in $O(n^{3})$ time whether for each $p\in W$,
 $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{E}}$ or $p/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=p/_{\mathop{\leftrightarrow}^{*}_{F}}$.
-Proof.
-By Proposition 2.10 and Proposition 2.11 we have the result. ∎
 Lemma 9.95.
@@
 $s{\downarrow}_{R\langle E,F\rangle}=t/_{\Theta\langle E,F\rangle}$ for some $t\in W$.
-By Propositon 4.34, $s\mathop{\leftrightarrow}^{*}_{E}t$. Then
+By Propositon 4.35, $s\mathop{\leftrightarrow}^{*}_{E}t$. Then
 $t\mathop{\leftrightarrow}^{*}_{E}s\mathop{\leftrightarrow}^{*}_{E\cup F}p$.
@@
 10 GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$
-We now introduce the notion of GTRS $R\langle E,F\rangle$ simulating the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant $a\in C\langle E\cup F,\emptyset\rangle$, then the concept of GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
+We now introduce the notion of GTRS $R\langle E,F\rangle$ simulating the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${b}\in C\langle E\cup F,\emptyset\rangle$.
+We study the case when GTRS $R\langle E,F\rangle$ is total.
+Furthermore, we introduce the concept of GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
 Then we describe the connection between the latter concept and the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
 Definition 10.98.
-Let $E$ and $F$ be GTESs over a signature $\Sigma$, and let $a\in C\langle E\cup F,\emptyset\rangle$.
-We say that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$ if Conditions 1 and 2 hold:
+Let $E$ and $F$ be GTESs over a signature $\Sigma$, and let ${b}\in C\langle E\cup F,\emptyset\rangle$.
+We say that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${b}\in C\langle E\cup F,\emptyset\rangle$ if Conditions 1 and 2 hold:
   • 1.
-    $a\in C\langle E,F\rangle$.
+    ${b}\in C\langle E,F\rangle$.
   • 2.
-    For every $b\in C\langle E\cup F,\emptyset\rangle$, if $R\langle{E\cup F},\emptyset\rangle$ reach $a$ from a proper extension of the constant $b$, then
-    for every rewrite rule of the form $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b\in R\langle{E\cup F},\emptyset\rangle$ with $m\in{\mathbb{N}}$,
-[…]
-    there exists $c\in C\langle E,F\rangle$ such that the rewrite rule $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c$ is in $R\langle E,F\rangle$.
-Statement 10.89.
-$c\subseteq b$ for $c,b$ appearing in Condition 2 in Definition 10.88.
-Intuitively, in Condition 2 in Definition 10.88, the rewrite rule $\sigma(c_{1},\ldots,c_{m})\dot{\rightarrow}c$ in $R\langle E,F\rangle$ fits to the rewrite rule $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b\in R\langle{E\cup F},\emptyset\rangle$.
-Statement 10.90.
-Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total, and let $a\in C\langle E\cup F,\emptyset\rangle$.
-GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$ if and only if $a\in C\langle E,F\rangle$.
-Proof.
-By direct inspection of Definition 4.29 and Definition 10.88. ∎
+    For every ${c}\in C\langle E\cup F,\emptyset\rangle$, if $R\langle{E\cup F},\emptyset\rangle$ reaches b{b} from a proper extension of the constant c{c}, then
+    for every rewrite rule of the form $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}\in R\langle{E\cup F},\emptyset\rangle$ with $m\in{\mathbb{N}}$,
+    for all ${d}_{1},\ldots,{d}_{m}\in C\langle E,F\rangle$ with ${d}_{i}\subseteq{c}_{i}$ for all $i\in\{\,1,\ldots,m\,\}$,
+    there exists ${d}\in C\langle E,F\rangle$ such that the rewrite rule $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}$ is in $R\langle E,F\rangle$.
+Statement 10.99.
+${d}\subseteq{c}$ for d,c{d},{c} appearing in Condition 2 in Definition 10.98.
+Intuitively, in Condition 2 in Definition 10.98, the rewrite rule $\sigma({d}_{1},\ldots,{d}_{m})\dot{\rightarrow}{d}$ in $R\langle E,F\rangle$ matches the rewrite rule $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}\in R\langle{E\cup F},\emptyset\rangle$.
+By direct inspection of Definition 4.30 and Definition 10.98, we get the following statement.
+Statement 10.100.
+Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total, and let ${b}\in C\langle E\cup F,\emptyset\rangle$.
+GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${b}\in C\langle E\cup F,\emptyset\rangle$ if and only if ${b}\in C\langle E,F\rangle$.
 Definition 10.101.
 Let $E$ and $F$ be GTESs over a signature $\Sigma$.
-We say that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if for each constant $a\in C\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on $a$.
-Statement 10.92.
+We say that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if for each constant ${b}\in C\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on b{b}.
+We have the following statement by direct inspection of Definition 10.101 and Statement 10.100.
+Statement 10.102.
 Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ is total.
-GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant $a\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
-1. $a\in C\langle E,F\rangle$.
-2. $a\in C\langle F,E\rangle$ and GTRS $R\langle F,E\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on $a$.
-Proof.
-We have the satement by direct inspection of Definition 10.91 and Statement 10.90. ∎
+GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant ${b}\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
+1. ${b}\in C\langle E,F\rangle$.
+2. ${b}\in C\langle F,E\rangle$ and GTRS $R\langle F,E\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on b{b}.
+By Definitions 10.98 and 10.101, and Lemma 6.71, and Definition 6.62, we have the following lemma.
 Lemma 10.103.
 Let $E$ and $F$ be GTESs over a signature $\Sigma$.
-GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant $a\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
+GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant ${b}\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
 1.
-$a\in C\langle E,F\rangle$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
+${b}\in C\langle E,F\rangle$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing c{c}.
 2.
-$a\in C\langle F,E\rangle$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
-Proof.
-By Definitions 10.88 and 10.91, and Lemma 6.70, and Definition 6.61, we have the lemma. ∎
+${b}\in C\langle F,E\rangle$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing c{c}.
+By Statement 10.102, Definition 10.98, Lemma 6.71, and Definition 6.62 we have the following lemma.
 Lemma 10.104.
 Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that $R\langle E,F\rangle$ is total.
-GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant $a\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
-1. $a\in C\langle E,F\rangle$.
+GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if and only if for each constant ${b}\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
+1. ${b}\in C\langle E,F\rangle$.
 2.
-$a\in C\langle F,E\rangle$ and for each constant $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
-Proof.
-By Statement 10.92, Definition 10.88, Lemma 6.70, and Definition 6.61 we have the lemma. ∎
+${b}\in C\langle F,E\rangle$ and for each constant ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing c{c}.
 Example 10.105.
 We now continue Example 3.18, Example 5.51, and Example 7.72. There exist two paths of positive length in $AUX[E;F]$. There do not exist cycles in $AUX[E;F]$.
-The length of the first path is $1$, and it starts from the vertex $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+The length of the first path is $1$, and it starts from the vertex $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$.
+By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $_{\Theta\langle E\cup F,\emptyset\rangle}$.
 Recall that
-$f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$   and   #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$.
-The length of the second path is $1$, and it starts from the vertex $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach $f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+$f(\#)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$   and   $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{true}$.
+The length of the second path is $1$, and it starts from the vertex $f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$.
 Recall that
-$f(\$)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$   and   \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
+$f(\mathdollar)/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$   and   $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
 […]
@@
 There exists one cycle in $AUX[E;F]$.
-It is of length $1$, and it starts from the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+It is of length $1$, and it starts from the vertex $_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$.
+By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $_{\Theta\langle E\cup F,\emptyset\rangle}$.
 Recall that
-#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and
-#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
-Observe that neither Condition 1 in nor Condition 2 in Lemma 10.93 holds for the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+$_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and
+$_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{E}=\mathrm{false}$, $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
+Observe that neither Condition 1 in nor Condition 2 in Lemma 10.103 holds for the vertex $_{\Theta\langle E\cup F,\emptyset\rangle}$.
 Thus by Lemma 10.103, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
-Intuitively, assume that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a vertex $c\in C\langle E\cup F,\emptyset\rangle$ of $AUX[E;F]$.
-Consider a term $\delta[a_{1},\ldots,a_{k}]$ for some $\delta\in CON_{\Sigma,k}$, $a_{1},\ldots,a_{k}\in C\langle{E\cup F},\emptyset\rangle$ and a sequence of reductions
-(i)   $\delta[a_{1},\ldots,a_{k}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}c$.
-[…]
-(ii)   $\delta[b_{1},\ldots,b_{k}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}d$
-for some $d\in C\langle E,F\rangle$.
+Intuitively, assume that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a vertex ${d}\in C\langle E\cup F,\emptyset\rangle$ of $AUX[E;F]$.
+Consider a term $\delta[{b}_{1},\ldots,{b}_{k}]$ for some $\delta\in CON_{\Sigma,k}$, ${b}_{1},\ldots,{b}_{k}\in C\langle{E\cup F},\emptyset\rangle$ and a sequence of reductions
+(i)   $\delta[{b}_{1},\ldots,{b}_{k}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}{d}$.
+Furthermore, consider any constants ${c}_{1},\ldots,{c}_{k}\in C\langle E,F\rangle$ such that ${c}_{j}\subseteq{b}_{j}$ for every $j\in\{\,1,\ldots,k\,\}$.
+We substitute ${c}_{j}$ for ${b}_{j}$ in $\delta[{b}_{1},\ldots,{b}_{k}]$, in this way, we get the term $\delta[{c}_{1},\ldots,{c}_{k}]$.
+Since GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$, we construct a fitting sequence of reductions
+(ii)   $\delta[{c}_{1},\ldots,{c}_{k}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}{e}$
+for some ${e}\in C\langle E,F\rangle$.
 We obtain (ii) following step by step the reductions in (i).
 For each reduction of (i), we apply a fitting rule at the same position in (ii).
-Then for each reduction step, for the two terms $s$ and $t$ obtained in (i) and (ii), respectively, each constant $b\in C\langle E,F\rangle$ appearing in $t$ corresponds to the constant $a\in C\langle E\cup F,\emptyset\rangle$ appearing at the same position in $s$.
-Moreover $b$, as congruence class, is subset of the corresponding constant $a$.
-Finally, constant $d$ corresponds to constant $c$, and $d\subseteq c$.
+Then for each reduction step, for the two terms $s$ and $t$ obtained in (i) and (ii), respectively, each constant ${c}\in C\langle E,F\rangle$ appearing in $t$ corresponds to the constant ${b}\in C\langle E\cup F,\emptyset\rangle$ appearing at the same position in $s$.
+Moreover c{c}, as congruence class, is subset of the corresponding constant b{b}.
+Finally, constant e{e} corresponds to constant d{d}, and ${e}\subseteq{d}$.
 Formally, we state and prove the following result.
 Lemma 10.107.
-Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant $c\in C\langle E\cup F,\emptyset\rangle$.
+Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${d}\in C\langle E\cup F,\emptyset\rangle$.
 Let Conditions (a) and (b) hold.
   • (a)
-    $\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-2}[a_{(n-2)1},\ldots,a_{(n-2)k_{n-2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[a_{11},\ldots,a_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}c$,
-[…]
+    $\delta_{n}[{b}_{n1},\ldots,{b}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{b}_{(n-1)1},\ldots,{b}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-2}[{b}_{(n-2)1},\ldots,{b}_{(n-2)k_{n-2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{b}_{11},\ldots,{b}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{d}$,
+    where $n\geq 1$, for every $i\in\{\,1,\ldots,n\,\}$, $\delta_{i}\in CON_{\Sigma,{k_{i}}}$, $k_{i}\in{\mathbb{N}}$, and ${b}_{i1},\ldots,{b}_{i{k_{i}}}\in C\langle{E\cup F},\emptyset\rangle$.
   • (b)
-[…]
-[…]
-$\delta_{n}[b_{n_{1}},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-2}[b_{(n-2)1},\ldots,b_{(n-2)k_{n-2}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\newline \delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}d$.
+    For every $j\in\{\,1,\ldots,k_{n}\,\}$, ${c}_{nj}\in C\langle E,F\rangle$ and ${c}_{nj}\subseteq{b}_{nj}$.
+Then for any $i\in\{\,1,\ldots,n-1\,\}$, $j\in\{\,1,\ldots,k_{i}\,\}$, there exist ${c}_{ij}\in C\langle E,F\rangle$ with ${c}_{ij}\subseteq{b}_{ij}$, and there exists ${e}\in C\langle E,F\rangle$ with ${e}\subseteq{d}$ such that the following sequence of reductions holds:
+$\delta_{n}[{c}_{n_{1}},\ldots,{c}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[{c}_{(n-1)1},\ldots,{c}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-2}[{c}_{(n-2)1},\ldots,{c}_{(n-2)k_{n-2}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\newline \delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}{e}$.
 Proof.
@@
 Base Case: $n=1$. Assume that Conditions (a) and (b) hold. Then $\delta_{1}=\sigma(\diamondsuit,\ldots,\diamondsuit)$ for some $\sigma\in\Sigma_{k_{1}}$, $k_{1}\in{\mathbb{N}}$, where
-$\delta_{1}[a_{11},\ldots,a_{1k_{1}}]=\sigma(a_{11},\ldots,a_{1k_{1}})$,   $\sigma(a_{11},\ldots,a_{1k_{1}})\dot{\rightarrow}c\in R\langle{E\cup F},\emptyset\rangle$,   and   $\sigma(a_{11},\ldots,a_{1k_{1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}c$.
-[…]
-$\sigma(b_{11},\ldots,b_{1k_{1}})\dot{\rightarrow}b\in R\langle E,F\rangle$, where $b\subseteq c$.
-Since $\delta_{1}[b_{11},\ldots,b_{1k_{1}}]=\sigma(b_{11},\ldots,b_{1k_{1}})$, we have
-$\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}b$, where $b\in C\langle E,F\rangle$ and $b\subseteq c$.
+$\delta_{1}[{b}_{11},\ldots,{b}_{1k_{1}}]=\sigma({b}_{11},\ldots,{b}_{1k_{1}})$,   $\sigma({b}_{11},\ldots,{b}_{1k_{1}})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$,   and   $\sigma({b}_{11},\ldots,{b}_{1k_{1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{d}$.
+By Condition (b), for every $j\in\{\,1,\ldots,k_{1}\,\}$, ${c}_{1j}\in C\langle E,F\rangle$ and ${c}_{1j}\subseteq{b}_{1j}$.
+We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${d}\in C\langle E\cup F,\emptyset\rangle$.
+Consequently by the rewrite rule $\sigma({b}_{11},\ldots,{b}_{1k_{1}})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$ and Definition 10.98, and Statement 10.99, there exists a rewrite rule
+$\sigma({c}_{11},\ldots,{c}_{1k_{1}})\dot{\rightarrow}{c}\in R\langle E,F\rangle$, where ${c}\subseteq{d}$.
+Since $\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]=\sigma({c}_{11},\ldots,{c}_{1k_{1}})$, we have
+$\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}{c}$, where ${c}\in C\langle E,F\rangle$ and ${c}\subseteq{d}$.
 Induction Step: Let $n\geq 1$, and assume that the statement of the claim is true for $n$.
@@
   • ($\mathrm{a}^{\prime}$)
-    $\delta_{n+1}[a_{(n+1)1},\ldots,a_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[a_{11},\ldots,a_{1k_{0}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}c$,
-[…]
+    $\delta_{n+1}[{b}_{(n+1)1},\ldots,{b}_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n}[{b}_{n1},\ldots,{b}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{b}_{(n-1)1},\ldots,{b}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{b}_{21},\ldots,{b}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{b}_{11},\ldots,{b}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{d}$,
+    where for every $i\in\{\,0,\ldots,n+1\,\}$, $\delta_{i}\in CON_{\Sigma,{k_{i}}}$, $k_{i}\in{\mathbb{N}}$, and ${b}_{i1},\ldots,{b}_{i{k_{i}}}\in C\langle{E\cup F},\emptyset\rangle$.
   • ($\mathrm{b}^{\prime}$)
-[…]
+    For every $j\in\{\,1,\ldots,k_{n+1}\,\}$, ${c}_{(n+1)j}\in C\langle E,F\rangle$ and ${c}_{(n+1)j}\subseteq{b}_{(n+1)j}$.
 Then Conditions (c)–(f) hold:
   • (c)
-[…]
+    We obtain $\delta_{n+1}$ from $\delta_{n}$ by replacing the $j$th occurrence from left to right of the hole $\diamondsuit$ by the context $\sigma(\diamondsuit,\ldots,\diamondsuit)$, where $j\in\{\,1,\ldots,k_{n}\,\}$, $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, and $k_{n}=k_{n+1}-m+1$.
   • (d)
-    $\delta_{n+1}[a_{(n+1)1},\ldots,a_{(n+1)k_{n+1}}]=\delta_{n}[a_{(n+1)1},\ldots,a_{(n+1)(j-1)},\sigma(a_{(n+1)j},\ldots,a_{(n+1)(j+m-1)}),a_{(n+1)(j+m)},\ldots,a_{(n+1)k_{n+1}}]$.
+    $\delta_{n+1}[{b}_{(n+1)1},\ldots,{b}_{(n+1)k_{n+1}}]=\delta_{n}[{b}_{(n+1)1},\ldots,{b}_{(n+1)(j-1)},\sigma({b}_{(n+1)j},\ldots,{b}_{(n+1)(j+m-1)}),{b}_{(n+1)(j+m)},\ldots,{b}_{(n+1)k_{n+1}}]$.
   • (e)
-    For every $\ell\in\{1,\ldots,j-1\,\}$, $a_{n\ell}=a_{(n+1)\ell}$ and for every $\ell\in\{j+1,\ldots,k_{n}\,\}$, $a_{n\ell}=a_{(n+1)(\ell+m-1)}$.
+    For every $\ell\in\{1,\ldots,j-1\,\}$, ${b}_{n\ell}={b}_{(n+1)\ell}$ and for every $\ell\in\{j+1,\ldots,k_{n}\,\}$, ${b}_{n\ell}={b}_{(n+1)(\ell+m-1)}$.
   • (f)
-    In the last reduction step of (a’) we apply the rule $\sigma(a_{(n+1)j},\ldots,a_{(n+1)(j+m-1)})\dot{\rightarrow}a_{nj}\in R\langle{E\cup F},\emptyset\rangle$.
-By Condition ($\mathrm{a}^{\prime}$), $R\langle E\cup F,\emptyset\rangle$ reach $c$ from a proper extension of $a_{nj}$.
-By Condition ($\mathrm{b}^{\prime}$), for every $\ell\in\{\,j,\ldots,j+m-1\,\}$, $b_{(n+1)\ell}\in C\langle E,F\rangle$ and $b_{(n+1)\ell}\subseteq a_{n\ell}$.
-We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $c\in C\langle E\cup F,\emptyset\rangle$.
+    In the last reduction step of (a’) we apply the rule $\sigma({b}_{(n+1)j},\ldots,{b}_{(n+1)(j+m-1)})\dot{\rightarrow}{b}_{nj}\in R\langle{E\cup F},\emptyset\rangle$.
+By Condition ($\mathrm{a}^{\prime}$), $R\langle E\cup F,\emptyset\rangle$ reaches d{d} from a proper extension of ${b}_{nj}$.
+By Condition ($\mathrm{b}^{\prime}$), for every $\ell\in\{\,j,\ldots,j+m-1\,\}$, ${c}_{(n+1)\ell}\in C\langle E,F\rangle$ and ${c}_{(n+1)\ell}\subseteq{b}_{n\ell}$.
+We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${d}\in C\langle E\cup F,\emptyset\rangle$.
 Therefore by the rewrite rule appearing in Condition (f) there exists a rewrite rule
-$\sigma(b_{(n+1)j},\ldots,b_{(n+1)(j+m-1)})\dot{\rightarrow}b$ in $R\langle E,F\rangle$ with $b\subseteq a_{nj}$.
-For every $\ell\in\{1,\ldots,j-1\,\}$, let $b_{n\ell}=b_{(n+1)\ell}$,
-let $b_{nj}=b$,
-for every $\ell\in\{j+1,\ldots,k_{n}\,\}$, let $b_{n\ell}=b_{(n+1)(\ell+j-1)}$.
+$\sigma({c}_{(n+1)j},\ldots,{c}_{(n+1)(j+m-1)})\dot{\rightarrow}{c}$ in $R\langle E,F\rangle$ with ${c}\subseteq{b}_{nj}$.
+For every $\ell\in\{1,\ldots,j-1\,\}$, let ${c}_{n\ell}={c}_{(n+1)\ell}$,
+let ${c}_{nj}={c}$,
+for every $\ell\in\{j+1,\ldots,k_{n}\,\}$, let ${c}_{n\ell}={c}_{(n+1)(\ell+j-1)}$.
 Thence
-for every { $\ell\in\{\,1,\ldots,k_{n}\,\}$, $b_{n\ell}\subseteq a_{n\ell}$ and $\delta_{n+1}[b_{(n+1)1},\ldots,b_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]$.
-By the sequence of reductions $(a^{\prime})$ we have
-$\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{1}[a_{11},\ldots,a_{1k_{0}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}c$.
+for every { $\ell\in\{\,1,\ldots,k_{n}\,\}$, ${c}_{n\ell}\subseteq{b}_{n\ell}$ and $\delta_{n+1}[{c}_{(n+1)1},\ldots,{c}_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n}[{c}_{n1},\ldots,{c}_{nk_{n}}]$.
+By the sequence of reductions ($\mathrm{a}^{\prime}$) we have
+$\delta_{n}[{b}_{n1},\ldots,{b}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{b}_{(n-1)1},\ldots,{b}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{b}_{21},\ldots,{b}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{1}[{b}_{11},\ldots,{b}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{d}$.
 Consequently by the induction hypothesis,
-[…]
-there exists $d\in C\langle E,F\rangle$ with $d\subseteq c$
+for any $i\in\{\,1,\ldots,n-1\,\}$ and { $\ell\in\{\,1,\ldots,k_{i}\,\}$, there exist ${c}_{i\ell}\in C\langle E,F\rangle$, with ${c}_{i\ell}\subseteq{b}_{i\ell}$, and
+there exists ${e}\in C\langle E,F\rangle$ with ${e}\subseteq{d}$
 such that the following sequence of reductions holds:
-$\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}d$.
+$\delta_{n}[{c}_{n1},\ldots,{c}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[{c}_{(n-1)1},\ldots,{c}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}{e}$.
 Then we have
-$\delta_{n+1}[b_{(n+1)1},\ldots,b_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\newline \cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}d$.∎
+$\delta_{n+1}[{c}_{(n+1)1},\ldots,{c}_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n}[{c}_{n1},\ldots,{c}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[{c}_{(n-1)1},\ldots,{c}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\newline \cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}{e}$.∎
+Lemma 10.107 implies the following lemma.
 Lemma 10.108.
-Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant $c\in C\langle E\cup F,\emptyset\rangle$.
-Let $s\in T_{\Sigma}$ and $s\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}c$.
-Then there exists $d\in C\langle E,F\rangle$ with $d\subseteq c$ such that $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}d$.
-Proof.
-Lemma 10.97 implies the lemma. ∎
+Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${d}\in C\langle E\cup F,\emptyset\rangle$.
+Let $s\in T_{\Sigma}$ and $s\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}{d}$.
+Then there exists ${e}\in C\langle E,F\rangle$ with ${e}\subseteq{d}$ such that $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}{e}$.
 Lemma 10.109.
@@
 Proof.
-()($\Rightarrow$)( ⇒ ) Let $t\in W$. Then by our assumption,
-$t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}\mbox{ or }t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle F,E\rangle}$.
+$(\Rightarrow)$ Let $t\in W$. Then by our assumption,
 […]
+Assume that $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{{\mathop{\leftrightarrow}^{*}_{E}}}$.
+The case $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{{\mathop{\leftrightarrow}^{*}_{F}}}$ is similar.
+We now show that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on $t/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+To this end, let ${d}\in C\langle E\cup F,\emptyset\rangle$ be such that $R\langle{E\cup F},\emptyset\rangle$ reaches $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ from a proper extension of the constant d{d}.
+Then there exists $\delta\in CON_{\Sigma,1}$ such that $\delta[{d}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
+Let $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$ be a rewrite rule with ${b}_{1},\ldots,{b}_{m}\in C\langle{E\cup F},\emptyset\rangle$, and let ${c}_{1},\ldots,{c}_{m}\in C\langle E,F\rangle$ with ${c}_{i}\subseteq{b}_{i}$ for all $i\in\{\,1,\ldots,m\,\}$.
+We now show that there exists a rewrite rule in $R\langle E,F\rangle$ with left-hand side $\sigma({c}_{1},\ldots,{c}_{m})$.
+Let $t_{1}\in{c}_{1},\ldots,t_{m}\in{c}_{m}$.
+By Definition 4.27, we have
 $t_{1},\ldots,t_{m}\in W$,
-$t_{1}/_{\Theta\langle E,F\rangle}=b_{1},\ldots,t_{m}/_{\Theta\langle E,F\rangle}=b_{m}$,
-$t_{1}/_{\Theta\langle{E\cup F},\emptyset\rangle}=a_{1},\ldots,t_{m}/_{\Theta\langle{E\cup F},\emptyset\rangle}=a_{m}$.
-Hence $\sigma(t_{1}/_{\Theta\langle{E\cup F},\emptyset\rangle},\ldots,t_{m}/_{\Theta\langle{E\cup F},\emptyset\rangle})\dot{\rightarrow}c\in R\langle{E\cup F},\emptyset\rangle$.
+$t_{1}/_{\Theta\langle E,F\rangle}={c}_{1},\ldots,t_{m}/_{\Theta\langle E,F\rangle}={c}_{m}$,
+$t_{1}/_{\Theta\langle{E\cup F},\emptyset\rangle}={b}_{1},\ldots,t_{m}/_{\Theta\langle{E\cup F},\emptyset\rangle}={b}_{m}$.
+Hence $\sigma(t_{1}/_{\Theta\langle{E\cup F},\emptyset\rangle},\ldots,t_{m}/_{\Theta\langle{E\cup F},\emptyset\rangle})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$.
 Then by Proposition 4.33,
-$\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\delta[\sigma(a_{1},\ldots,a_{m})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[c]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
+$\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\delta[\sigma({b}_{1},\ldots,{b}_{m})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[{d}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
 By Proposition 4.35, $\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\leftrightarrow}^{*}_{E\cup F}t$.
@@
 By Proposition 4.38,
-$\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t_{\Theta\langle E,F\rangle}$.
-By Proposition 4.31, $R\langle E,F\rangle$ is convergent. Observe that $t_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$. Thence we have
+$\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
+By Proposition 4.32, $R\langle E,F\rangle$ is convergent. Observe that $t/_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$. Thence we have
 $\delta[\sigma(t_{1},\ldots,t_{m})]{\downarrow}_{R\langle E,F\rangle}=t/_{\Theta\langle E,F\rangle}$.
-[…]
+By the set memberships $t_{1},\ldots,t_{m}\in W$ and by Proposition 4.33, we have $t_{j}\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t_{j}/_{\Theta\langle{E},F\rangle}$ for every $j\in\{\,1,\ldots,m\,\}$.
+Then
 $\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\delta[\sigma(t_{1}/_{\Theta\langle{E},F\rangle},\ldots,t_{m}/_{\Theta\langle{E},F\rangle})]$.
@@
 Therefore, there exists a rewrite rule with left-hand side $\sigma(t_{1}/_{\Theta\langle{E},F\rangle},\ldots,t_{m}/_{\Theta\langle{E},F\rangle})$ in $R\langle{E},F\rangle$.
-()($\Leftarrow$)( ⇐ ) Let $t\in W$.
+$(\Leftarrow)$ Let $t\in W$.
 We assumed that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
@@
 Consequently by Proposition 4.38, $s\mathop{\rightarrow}_{R\langle E\cup F,E\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
-By Lemma 10.98, there exists $d\in C\langle E,F\rangle$ such that $d\subseteq t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}d$.
-Then $d=p/_{\Theta\langle E,F\rangle}$ for some $p\in W$.
+By Lemma 10.108, there exists ${e}\in C\langle E,F\rangle$ such that ${e}\subseteq t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}{e}$.
+Then ${e}=p/_{\Theta\langle E,F\rangle}$ for some $p\in W$.
 By Proposition 4.35, $s\mathop{\leftrightarrow}^{*}_{E\cup F}t$ and $s\mathop{\leftrightarrow}^{*}_{E}p$.
 Then $s\mathop{\leftrightarrow}^{*}_{E\cup F}p$.
-Thus $d=p/_{\Theta\langle E,F\rangle}\subseteq t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}$.
-Consequently $d=t/_{\Theta\langle E,F\rangle}$.
+Thus ${e}=p/_{\Theta\langle E,F\rangle}\subseteq t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}$.
+Consequently ${e}=t/_{\Theta\langle E,F\rangle}$.
 Hence $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$.
@@
 Example 10.110.
-We continue Example 3.21 and its sequels, Examples 5.54, 7.73.
+We continue Example 3.22 and its sequels, Examples 5.55, 7.74.
 There does not exist a path of positive length in $AUX[E;F]$.
 There exists one cycle in $AUX[E;F]$.
-The cycle is of length $1$, it starts from the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-Recall that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$ and #$\#/_{\Theta\langle F,E\rangle}.keeps_{F}=\mathrm{false}$.
+The cycle is of length $1$, it starts from the vertex $_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$.
+By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $_{\Theta\langle E\cup F,\emptyset\rangle}$.
+Recall that $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$ and $_{\Theta\langle F,E\rangle}.keeps_{F}=\mathrm{false}$.
 Then the formula
-#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$ or #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$
-does not hold. Thence the following statement does not hold: for each $a\in C\langle E\cup F,\emptyset\rangle$,
-  • $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$ and
-  • if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
+$_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$ or $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$
+does not hold. Thence the following statement does not hold: for each ${b}\in C\langle E\cup F,\emptyset\rangle$,
+  • 1.
+    ${b}.equal_{E}=\mathrm{true}$ or ${b}.equal_{F}=\mathrm{true}$ and
+  • 2.
+    if ${b}\not\in C\langle E,F\rangle$, then for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
 By Lemma 10.104, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
 Example 10.111.
-We continue Example 3.22, and its sequels, Examples 5.55, and 7.74. There exist one path of positive length and two cycles in $AUX[E;F]$.
-The first cycle: Its length is $1$, and it starts from the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+We continue Example 3.23, and its sequels, Examples 5.56, and 7.75. There exist one path of positive length and two cycles in $AUX[E;F]$.
+The first cycle: Its length is $1$, and it starts from the vertex $_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$.
+By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $_{\Theta\langle E\cup F,\emptyset\rangle}$.
 Recall that
-#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
-The path of positive length: Its length is $1$, and it starts from the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+$_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$, and $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
+The path of positive length: Its length is $1$, and it starts from the vertex $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$.
+By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $_{\Theta\langle E\cup F,\emptyset\rangle}$.
 Recall that
-\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
-The second cycle: Its length is $1$, and it starts from the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+$\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
+The second cycle: Its length is $1$, and it starts from the vertex $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$.
 Recall that
-\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
-Observe that the following statement holds: for each $a\in C\langle E\cup F,\emptyset\rangle$,
-  • $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$ and
-  • if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
+$\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
+Observe that the following statement holds: for each ${b}\in C\langle E\cup F,\emptyset\rangle$,
+  • 1.
+    ${b}.equal_{E}=\mathrm{true}$ or ${b}.equal_{F}=\mathrm{true}$ and
+  • 2.
+    if ${b}\not\in C\langle E,F\rangle$, then for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
 By Lemma 10.104, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
 11 Main Cases of the Decision Algorithm
-We now define the four main cases of our decision algorithm. We show that we can decide which Main Cases hold in $O(n\,\mathrm{log}\,n)$ time.
+We now recall the four main cases of our decision algorithm. We show that we can decide in $O(n\,\mathrm{log}\,n)$ time which main cases hold.
   • Main Case 1:
     $\Sigma$ is a unary signature.
-Example 3.17 and its sequels, Examples 5.50, 7.71, 10.95.
-and 12.109, and 12.114, also Example 3.18 and its sequels, 5.51, 7.72, 10.96, 12.110, and 12.115 illustrate this case.
   • Main Case 2:
     Both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total.
-Example 3.19 and its sequels Examples 5.52, 13.117, further Example 3.20 and its sequels, Examples 5.53, 13.118, exemplify this case.
   • Main Case 3:
     $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total.
-Example 3.21 and its sequels, Examples 5.54, 7.73, 10.100, 14.125, and 14.132, moreover Example 3.22, and its sequels, Examples 5.55, 7.74, 10.101, 14.126, and 14.133 present this case.
   • Main Case 4:
     $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total.
-Example 3.23 and its sequel Example 5.56, furthermore, Example 3.24 and its sequels Example 5.57 and Example 15.134 demonstrate this case.
-Main Case 2 overlaps with Main Case 1 and Main Case 3, and is a subcase of the union of Main Case 1 and Main Case 3.
+Recall that Main Case 2 overlaps with Main Case 1 and Main Case 3, and is a subcase of the union of Main Case 1 and Main Case 3.
 Lemma 11.112.
@@
 We run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.48, we get the following output in $O(n\,\mathrm{log}\,n)$ time:
+  • 1.
     the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$;
+  • 2.
     the set of nullary symbols $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$;
+  • 3.
     $INC(E\cup F,E)$ stored in a red-black tree saved in a variable $RBT1$;
+  • 4.
     $INC(E\cup F,F)$ stored in a red-black tree saved in a variable $RBT2$;
+  • 5.
     $NUM(E\cup F,E)$ stored in a red-black tree saved in a variable $RBT3$;
+  • 6.
     $NUM(E\cup F,F)$ stored in a red-black tree saved in a variable $RBT4$;
+  • 7.
     the GTRS $R\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT5$; each node in $RBT5$ contains a search key which is a rule in $R\langle E\cup F,\emptyset\rangle$, and satellite data $counter_{E}$ and $counter_{F}$;
-  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT6$; for every ordered pair $(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
+  • 8.
+    $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT6$; for every ordered pair $({b},{c})\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component b{b}, the second component, c{c} is satellite data;
+  • 9.
     the GTRS $R\langle E,F\rangle$ stored in a red-black tree saved in a variable $RBT7$;
+  • 10.
     the GTRS $R\langle F,E\rangle$ stored in a red-black tree saved in a variable $RBT8$;
-  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
+  • 11.
+    the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes ${b}.equal_{E}$, ${b}.equal_{F}$, ${b}.keeps_{E}$, and ${b}.keeps_{F}$.
 By Lemma 4.50, we decide in $O(n)$ time whether $R\langle E,F\rangle$ is total, and whether $R\langle F,E\rangle$ is total.
@@
 In this main case, $\Sigma$ is a unary signature, and $E$ and $F$ are GTESs over $\Sigma$.
+We define the Algorithm Nested Partial Depth First Search (NPDFS for short) on the Auxiliary dpwpa $AUX[E;F]$ for the GTESs $E$ and $F$.
+NPDFS runs in $O(n^{2})$ time.
+Furthermore, $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle)$ if and only if NPDFS outputs true.
+Thus, we can decide in $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
+First we note that we can also apply the decision algorithm of Champavère et al [10].
+Lemma 9.93, Proposition 9.94, and Proposition 9.92 imply the following theorem.
 Theorem 12.113.
@@
 We can decide in $O(n^{3})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
-Proof.
-Lemma 9.83, Proposition 9.84, and Proposition 9.82 imply the Theorem. ∎
 We present a faster and simpler algorithm for solving our decidability problem.
 We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.48.
-Example 3.17 and its sequel, Example 5.50, furthermore Example 3.18 and its sequel, 5.51 illustrate this main case.
-We assume that we have run Algorithm CAD on $\Sigma$, $E$, and $F$, and study its output, see Proposition 4.47.
+Example 3.18 and its sequel, Example 5.51, furthermore Example 3.19 and its sequel, 5.52 illustrate this main case.
 Lemma 12.114.
 Let $E$ and $F$ be GTESs over a unary signature $\Sigma$.
-Let $a\in C\langle E\cup F,\emptyset\rangle$ be such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle F$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$.
+Let ${b}\in C\langle E\cup F,\emptyset\rangle$ be such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${b}\in C\langle E\cup F,\emptyset\rangle$.
 Assume that Conditions (a) and (b) hold:
   • (a)
-    $\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[a_{11},\ldots,a_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a$,
-[…]
+    $\delta_{n}[{b}_{n1},\ldots,{b}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{b}_{(n-1)1},\ldots,{b}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{b}_{21},\ldots,{b}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{b}_{21},\ldots,{b}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{b}_{11},\ldots,{b}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{b}$,
+    where $n\geq 1$ and for every $i\in\{\,1,\ldots,n\,\}$, $\left(k_{i}\in{\mathbb{N}},\;\delta_{i}\in CON_{\Sigma,k_{i}},\mbox{ and }{b}_{i1},\ldots,{b}_{ik_{i}}\in C\langle{E\cup F},\emptyset\rangle\right)$.
   • (b)
-[…]
-[…]
-$\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}b$,
+    ${c}_{n1},\ldots,{c}_{nk_{n}}\in C\langle E,F\rangle$ such that ${c}_{ni}\subseteq{b}_{ni}$ for every $i\in\{\,1,\ldots,k_{n}\,\}$.
+Then there exists ${c}\in C\langle E,F\rangle$ with ${c}\subseteq{b}$ and for every $j\in\{\,1,\ldots,n\,\}$, there exists ${c}_{j1},\ldots,{c}_{jk_{j}}\in C\langle E,F\rangle$ such that ${c}_{ji}\subseteq{b}_{ji}$ for every $i\in\{\,1,\ldots,k_{j}\,\}$, and the following sequence of reductions holds:
+$\delta_{n}[{c}_{n1},\ldots,{c}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{c}_{(n-1)1},\ldots,{c}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{c}_{21},\ldots,{c}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{c}_{21},\ldots,{c}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{c}$,
 Proof.
@@
 Base Case: $n=1$.
-Then $\delta_{1}=\sigma(\diamondsuit,\ldots,\diamondsuit)$ for some $\sigma\in\Sigma_{k_{1}}$, $\delta_{1}[a_{11},\ldots,{k_{1}}]=\sigma(a_{11},\ldots,{k_{1}})$, and $\sigma(a_{11},\ldots,a_{1k_{1}})\dot{\rightarrow}a\in R\langle{E\cup F},\emptyset\rangle$.
-We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$.
+Then $\delta_{1}=\sigma(\diamondsuit,\ldots,\diamondsuit)$ for some $\sigma\in\Sigma_{k_{1}}$, $\delta_{1}[{b}_{11},\ldots,{k_{1}}]=\sigma({b}_{11},\ldots,{k_{1}})$, and $\sigma({b}_{11},\ldots,{b}_{1k_{1}})\dot{\rightarrow}{b}\in R\langle{E\cup F},\emptyset\rangle$.
+We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${b}\in C\langle E\cup F,\emptyset\rangle$.
 Hence by Condition (b), and the rewrite rule above there exists a rewrite rule
-$\sigma(b_{11},\ldots,b_{1k_{1}})\dot{\rightarrow}b\in R\langle E,F\rangle$
-for some $b\in C\langle E,F\rangle$ with $b\subseteq a$. Then we have $\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}b$.
+$\sigma({c}_{11},\ldots,{c}_{1k_{1}})\dot{\rightarrow}{c}\in R\langle E,F\rangle$
+for some ${c}\in C\langle E,F\rangle$ with ${c}\subseteq{b}$. Then we have $\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}{c}$.
 Induction Step: Let $n\geq 1$, and assume that the statement of the lemma is true for $n$.
@@
   • ($\mathrm{a}^{\prime}$)
-    $\delta_{n+1}[a_{(n+1)1},\ldots,a_{(n+1)k_{(n+1)}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[a_{11},\ldots,a_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a$,
+    $\delta_{n+1}[{b}_{(n+1)1},\ldots,{b}_{(n+1)k_{(n+1)}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n}[{b}_{n1},\ldots,{b}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{b}_{(n-1)1},\ldots,{b}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{b}_{21},\ldots,{b}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{b}_{11},\ldots,{b}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{b}$,
+    where $n\geq 1$ and for every $i\in\{\,1,\ldots,n+1\,\}$, $\left(k_{i}\in{\mathbb{N}},\;\delta_{i}\in CON_{\Sigma,k_{i}},\mbox{ and }{b}_{i1},\ldots,{b}_{ik_{i}}\in C\langle{E\cup F},\emptyset\rangle\right)$.
+  • ($\mathrm{b}^{\prime}$)
+    ${c}_{(n+1)1},\ldots,{c}_{(n+1)k_{n+1}}\in C\langle E,F\rangle$ such that ${c}_{(n+1)i}\subseteq{b}_{(n+1)i}$ for every $i\in\{\,1,\ldots,k_{n+1}\,\}$.
 […]
-  • ($\mathrm{b}^{\prime}$)
-[…]
-Then we obtain $\delta_{n+1}$ from $\delta_{n}$ replacing the $\ell$th occurrence from left-to right of $\diamondsuit$ with the term $\sigma(\diamondsuit,\ldots,\diamondsuit)$ for some $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$.
-Consequently,
 $k_{n+1}-m+1=k_{n}$,
-$\delta_{n+1}[a_{(n+1)1},\ldots,a_{(n+1)k_{(n+1)}}]=\delta_{n}[a_{(n+1)1},\ldots,a_{(n+1)(\ell-1)},\sigma(a_{(n+1)\ell},\ldots,a_{(n+1)(\ell+m)}),a_{(n+1)(\ell+1)},\ldots a_{(n+1)k_{(n+1)}}]$,
-$\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]=\delta_{n}[a_{(n+1)1},\ldots,a_{(n+1)(\ell-1)},a_{n,\ell},a_{(n+1)(\ell+1)},\ldots a_{(n+1)k_{(n+1)}}]$,
-$a_{n1}=a_{(n+1)1},\ldots,a_{n(\ell-1)}=a_{(n+1)(\ell-1)},a_{n\ell}=a_{n\ell}a_{n(\ell+1)}=a_{(n+1)(\ell-m+1)},\ldots,a_{nk_{n}}=a_{(n+1)k_{(n+1)}}$
+$\delta_{n+1}[{b}_{(n+1)1},\ldots,{b}_{(n+1)k_{(n+1)}}]=\delta_{n}[{b}_{(n+1)1},\ldots,{b}_{(n+1)(\ell-1)},\sigma({b}_{(n+1)\ell},\ldots,{b}_{(n+1)(\ell+m)}),{b}_{(n+1)(\ell+1)},\ldots{b}_{(n+1)k_{(n+1)}}]$,
+$\delta_{n}[{b}_{n1},\ldots,{b}_{nk_{n}}]=\delta_{n}[{b}_{(n+1)1},\ldots,{b}_{(n+1)(\ell-1)},{b}_{n,\ell},{b}_{(n+1)(\ell+1)},\ldots{b}_{(n+1)k_{(n+1)}}]$,
+${b}_{n1}={b}_{(n+1)1},\ldots,{b}_{n(\ell-1)}={b}_{(n+1)(\ell-1)},{b}_{n(\ell+1)}={b}_{(n+1)(\ell-m+1)},\ldots,{b}_{nk_{n}}={b}_{(n+1)k_{(n+1)}}$
 and
-$\sigma(a_{(n+1)\ell},\ldots,a_{(n+1)(\ell+m)})\dot{\rightarrow}a_{n\ell}\in R\langle{E\cup F},\emptyset\rangle.$ (1)
-We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a\in C\langle E\cup F,\emptyset\rangle$.
-By Condition ($\mathrm{a}^{\prime}$), $R\langle E\cup F,\emptyset\rangle$ reach $a$ from a proper extension of $a_{n\ell}$.
+$\sigma({b}_{(n+1)\ell},\ldots,{b}_{(n+1)(\ell+m)})\dot{\rightarrow}{b}_{n\ell}\in R\langle{E\cup F},\emptyset\rangle.$ (1)
+We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${b}\in C\langle E\cup F,\emptyset\rangle$.
+By Condition ($\mathrm{a}^{\prime}$), $R\langle E\cup F,\emptyset\rangle$ reaches b{b} from a proper extension of ${b}_{n\ell}$.
 Therefore by the rewrite rule (1) and Condition ($\mathrm{b}^{\prime}$), there exists a rewrite rule
-$\sigma(b_{(n+1)\ell},\ldots,b_{(n+1)(\ell+m)})\dot{\rightarrow}b_{n\ell}\in R\langle{E\cup F},\emptyset\rangle.$ (2)
-for some $b_{n\ell}\in C\langle E\cup F,E\rangle$ with $b_{n\ell}\subseteq a_{n\ell}$. Let
-$b_{n1}=b_{(n+1)1},\ldots,b_{n(l-1)}=b_{(n+1)(l-1)},\ldots,b_{n(l+1)}=b_{(n+1)(l-m+1)},\ldots,b_{nk_{n}}=b_{(n+1)k_{(n+1)}}$
-By the induction hypothesis, there exists $b\in C\langle E,F\rangle$ with $b\subseteq a$ such that the following sequence of reductions holds:
-$\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}b$,
+$\sigma({c}_{(n+1)\ell},\ldots,{c}_{(n+1)(\ell+m)})\dot{\rightarrow}{c}_{n\ell}\in R\langle{E\cup F},\emptyset\rangle.$ (2)
+for some ${c}_{n\ell}\in C\langle E\cup F,E\rangle$ with ${c}_{n\ell}\subseteq{b}_{n\ell}$. Let
+${c}_{n1}={c}_{(n+1)1},\ldots,{c}_{n(l-1)}={c}_{(n+1)(l-1)},\ldots,{c}_{n(l+1)}={c}_{(n+1)(l-m+1)},\ldots,{c}_{nk_{n}}={c}_{(n+1)k_{(n+1)}}$
+By the induction hypothesis, there exists ${c}\in C\langle E,F\rangle$ with ${c}\subseteq{b}$ such that the following sequence of reductions holds:
+$\delta_{n}[{c}_{n1},\ldots,{c}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{c}_{(n-1)1},\ldots,{c}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{c}_{21},\ldots,{c}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{c}$,
 In this way
-$\delta_{n+1}[b_{(n+1)1},\ldots,b_{(n+1)k_{(}n+1)}]=\newline \delta_{n}[b_{(n+1)1},\ldots,b_{(n+1)(\ell-1)},\sigma(b_{(n+1)\ell},\ldots,b_{(n+1)(\ell+m)}),b_{(n+1)(\ell+1)},\ldots b_{(n+1)k_{(n+1)}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{n}[b_{(n+1)1},\ldots,b_{(n+1)(\ell-1)},b_{n,\ell},b_{(n+1)(\ell+1)},\ldots b_{(n+1)k_{(n+1)}}]=\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \cdots\newline \mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[b_{21},\ldots,b_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}b$, ∎
+$\delta_{n+1}[{c}_{(n+1)1},\ldots,{c}_{(n+1)k_{(}n+1)}]=\newline \delta_{n}[{c}_{(n+1)1},\ldots,{c}_{(n+1)(\ell-1)},\sigma({c}_{(n+1)\ell},\ldots,{c}_{(n+1)(\ell+m)}),{c}_{(n+1)(\ell+1)},\ldots{c}_{(n+1)k_{(n+1)}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{n}[{c}_{(n+1)1},\ldots,{c}_{(n+1)(\ell-1)},{c}_{n,\ell},{c}_{(n+1)(\ell+1)},\ldots{c}_{(n+1)k_{(n+1)}}]=\delta_{n}[{c}_{n1},\ldots,{c}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{n-1}[{c}_{(n-1)1},\ldots,{c}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{c}_{21},\ldots,{c}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{c}$.
+∎
 Lemma 12.115.
@@
 Proof.
-()($\Rightarrow$)( ⇒ ) Let $t\in W$.
+$(\Rightarrow)$ Let $t\in W$.
 Then by our assumption, $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$ or $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}$.
@@
 The case $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{F}}$ is similar.
-Let $b\in C\langle E\cup F,\emptyset\rangle$ be such that $R\langle{E\cup F},\emptyset\rangle$ reach $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ from a proper extension of the constant $b$.
-Then there exists $\delta\in C_{\Sigma,1}$ such that $\delta[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
-Consider any rewrite rule $\sigma(b_{1},\ldots,b_{m})\dot{\rightarrow}b$ in $R\langle E\cup F,\emptyset\rangle$.
-with right-hand side $b$.
+Let ${c}\in C\langle E\cup F,\emptyset\rangle$ be such that $R\langle{E\cup F},\emptyset\rangle$ reaches $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ from a proper extension of the constant c{c}.
+Then there exists $\delta\in CON_{\Sigma,1}$ such that $\delta[{c}]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
+Consider any rewrite rule $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}$ in $R\langle E\cup F,\emptyset\rangle$.
+with right-hand side c{c}.
 We now distinguish two cases depending on the value of $m$.
 Case 1: $m=1$.
-Then we consider the rewrite rule $\sigma(b_{1})\dot{\rightarrow}b\in R\langle{E\cup F},\emptyset\rangle$.
-Then let $c_{1}\in C\langle E,F\rangle$ with $c_{1}\subseteq b_{1}$, and let $t_{1}\in T_{\Sigma}$ such that $t_{1}\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}c_{1}$.
+Then we consider the rewrite rule $\sigma({c}_{1})\dot{\rightarrow}{c}\in R\langle{E\cup F},\emptyset\rangle$.
+Then let ${d}_{1}\in C\langle E,F\rangle$ with ${d}_{1}\subseteq{c}_{1}$, and let $t_{1}\in T_{\Sigma}$ such that $t_{1}\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}{d}_{1}$.
 Then we have
-$\delta[\sigma(t_{1})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\delta[\sigma(b_{1})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[b]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+$\delta[\sigma(t_{1})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\delta[\sigma({c}_{1})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[{c}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle E\cup F,\emptyset\rangle}$.
 By Proposition 4.35, $\delta[\sigma(t_{1})]\mathop{\leftrightarrow}^{*}_{E\cup F}t$.
@@
 Hence
-$t_{1}\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}c_{1}$,
-$\sigma(t_{1})\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\sigma(c_{1})\mathop{\rightarrow}_{R\langle E,F\rangle}c$ for some $c\in C\langle E,F\rangle$, where the rewrite rule $\sigma(c_{1})\dot{\rightarrow}c$ is in $R\langle E,F\rangle$, and
-$\delta[\sigma(t_{1})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\delta[\sigma(c_{1})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\delta[c]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
-Case 2: $m=0$. Then we consider the rewrite rule $\sigma\dot{\rightarrow}b\in R\langle{E\cup F},\emptyset\rangle$. Then we have
-$\delta[\sigma]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[b]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+$t_{1}\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}{d}_{1}$,
+$\sigma(t_{1})\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\sigma({d}_{1})\mathop{\rightarrow}_{R\langle E,F\rangle}{d}$ for some ${d}\in C\langle E,F\rangle$, where the rewrite rule $\sigma({d}_{1})\dot{\rightarrow}{d}$ is in $R\langle E,F\rangle$, and
+$\delta[\sigma(t_{1})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\delta[\sigma({d}_{1})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\delta[{d}]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
+Case 2: $m=0$. Then we consider the rewrite rule $\sigma\dot{\rightarrow}{c}\in R\langle{E\cup F},\emptyset\rangle$. Then we have
+$\delta[\sigma]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[{c}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle E\cup F,\emptyset\rangle}$.
 By Proposition 4.35, $\delta[\sigma]\mathop{\leftrightarrow}^{*}_{E\cup F}t$.
@@
 Therefore
-$\sigma\mathop{\rightarrow}_{R\langle E,F\rangle}c$ for some $c\in C\langle E,F\rangle$, where the rewrite rule $\sigma\dot{\rightarrow}c$ is in $R\langle E,F\rangle$, and
-$\delta[\sigma]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta[c]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
-()($\Leftarrow$)( ⇐ ) Let $t\in W$ and $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$.
+$\sigma\mathop{\rightarrow}_{R\langle E,F\rangle}{d}$ for some ${d}\in C\langle E,F\rangle$, where the rewrite rule $\sigma\dot{\rightarrow}{d}$ is in $R\langle E,F\rangle$, and
+$\delta[\sigma]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta[{d}]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
+$(\Leftarrow)$ Let $t\in W$ and $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$.
 Then $s\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
@@
 The case $R\langle F,E\rangle$ simulates $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ is similar.
-By Lemma 12.104, there exists $d\in C\langle E,F\rangle$ with $d\subseteq t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ such that $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}d$.
-By Proposition 4.38, $d=t/_{\Theta\langle E,F\rangle}$.
+By Lemma 12.114, there exists ${e}\in C\langle E,F\rangle$ with ${e}\subseteq t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ such that $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}{e}$.
+By Proposition 4.39, ${e}=t/_{\Theta\langle E,F\rangle}$.
 For this reason $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$.
@@
 Algorithm NPDFS calls the function NPDFS-VISIT, and in this way it implements the outer loop, which is a prefix of a depth first search on $AUX[E;F]$, and can terminate before visiting all vertices of $AUX[E;F]$ in case that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
+  • 1.
     Algorithm NPDFS calls the functions INNER-LOOP-E, INNER-LOOP-F, and NPDFS-VISIT.
+  • 2.
     Function NPDFS-VISIT calls the function INNER-LOOP-E, INNER-LOOP-F, and NPDFS-VISIT.
-In this way, algorithm NPDFS decides the following formula: For each $a\in C\langle E\cup F,\emptyset\rangle$,
-  • $a.equal_{E}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{E}=\mathrm{true}$, or
-  • $a.equal_{F}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
-Function INNER-LOOP-E is an inner loop inside the body of the outer loop, it also implements a prefix of a depth first search on those vertices $b$ which are positive-step reachable from the current node $a$ of the outer loop, it can terminate before visiting all vertices of $AUX[E;F]$ which are positive-step reachable from $a$ in case that it visits a vertex $b$ such that $b.keeps_{E}=\mathrm{false}$.
-It checks whether for each vertex $b$ which are positive-step reachable from the current node $a$ of the outer loop, $b.keeps_{E}=\mathrm{true}$.
+In this way, algorithm NPDFS decides the following formula: For each ${b}\in C\langle E\cup F,\emptyset\rangle$,
+  • 1.
+    ${b}.equal_{E}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{E}=\mathrm{true}$, or
+  • 2.
+    ${b}.equal_{F}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
+Function INNER-LOOP-E is an inner loop inside the body of the outer loop, it also implements a prefix of a depth first search on those vertices c{c} which are positive-step reachable from the current node b{b} of the outer loop, it can terminate before visiting all vertices of $AUX[E;F]$ which are positive-step reachable from b{b} in case that it visits a vertex c{c} such that ${c}.keeps_{E}=\mathrm{false}$.
+It checks whether for each vertex c{c} which are positive-step reachable from the current node b{b} of the outer loop, ${c}.keeps_{E}=\mathrm{true}$.
 INNER-LOOP-E decides the following formula:
-For each vertex $b\in C\langle E\cup F,\emptyset\rangle$ which is positive-step reachable from the current node $a$ of the outer loop, $b.keeps_{E}=\mathrm{true}$.
-Function INNER-LOOP-F is an inner loop inside the body of the outer loop, it also implements a prefix of a depth first search on those vertices $b$ which are positive-step reachable from the current node $a$ of the outer loop, it can terminate before visiting all vertices of $AUX[E;F]$ which are positive-step reachable from $a$ in case that in case that it visits a vertex $b$ such that $b.keeps_{F}=\mathrm{false}$.
-It checks whether for each vertex $b$ which are positive-step reachable from the current node $a$ of the outer loop, $b.keeps_{F}=\mathrm{true}$.
+For each vertex ${c}\in C\langle E\cup F,\emptyset\rangle$ which is positive-step reachable from the current node b{b} of the outer loop, ${c}.keeps_{E}=\mathrm{true}$.
+Function INNER-LOOP-F is an inner loop inside the body of the outer loop, it also implements a prefix of a depth first search on those vertices c{c} which are positive-step reachable from the current node b{b} of the outer loop, it can terminate before visiting all vertices of $AUX[E;F]$ which are positive-step reachable from b{b} in case that in case that it visits a vertex c{c} such that ${c}.keeps_{F}=\mathrm{false}$.
+It checks whether for each vertex c{c} which are positive-step reachable from the current node b{b} of the outer loop, ${c}.keeps_{F}=\mathrm{true}$.
 INNER-LOOP-F decides the following formula:
-For each vertex $b\in C\langle E\cup F,\emptyset\rangle$ which is positive-step reachable from the current node $a$ of the outer loop, $b.keeps_{F}=\mathrm{true}$.
+For each vertex ${c}\in C\langle E\cup F,\emptyset\rangle$ which is positive-step reachable from the current node b{b} of the outer loop, ${c}.keeps_{F}=\mathrm{true}$.
 Definition 12.116.
@@
 }
-var $a:C\langle E\cup F,\emptyset\rangle$;
-for each vertex $a\in C\langle E\cup F,\emptyset\rangle$ do
+var ${b}:C\langle E\cup F,\emptyset\rangle$;
+for each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ do
 begin
-$a.visited\_outer\_loop:=\mathrm{false}$
-$a.visited\_inner\_loop_{E}:=\mathrm{false}$
-$a.visited\_inner\_loop_{F}:=\mathrm{false}$
+${b}.visited\_outer\_loop:=\mathrm{false}$
+${b}.visited\_inner\_loop_{E}:=\mathrm{false}$
+${b}.visited\_inner\_loop_{F}:=\mathrm{false}$
 end
-for each vertex $a\in C\langle E\cup F,\emptyset\rangle$ do
-if not $a.visited\_outer\_loop$
+for each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ do
+if not b.visited_outer_loop{b}.visited\_outer\_loop
 then begin
-$a.visited\_outer\_loop:=\mathrm{true}$
-${\bf if}$ not ($a.equal_{E}$ or $a.equal_{F}$)
+${b}.visited\_outer\_loop:=\mathrm{true}$
+$\bf if$} not ( ${b}.equal_{E}$ or ${b}.equal_{F}$)
 then return $\mathrm{false}$
-${\bf for}$ ${\bf each}$ vertex $b\in Adj(a)$ do
-${\bf if}$ not (($a.equal_{E}$ and $\mathrm{INNER-LOOP-E}(AUX[E;F],b)$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],b)$
-or ($a.equal_{F}$ and $\mathrm{INNER-LOOP-F}(AUX[E;F],b)$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],b))$
+$\bf for$} $\bf each$} vertex ${c}\in Adj({b})$ do
+$\bf if$} not $($${b}.equal_{E}$ and $\mathrm{INNER-LOOP-E}(AUX[E;F],{c})$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],{c})$
+or (${b}.equal_{F}$ and $\mathrm{INNER-LOOP-F}(AUX[E;F],{c})$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],{c}))$
 then return $\mathrm{false}$
@@
 output $(\mathrm{true})$
-function $\mathrm{NPDFS-VISIT}(AUX[E;F],a):\;\mathrm{Boolean}$
-if not $a.visited\_outer\_loop$
+function $\mathrm{NPDFS-VISIT}(AUX[E;F],{b}):\;\mathrm{Boolean}$
+if not b.visited_outer_loop{b}.visited\_outer\_loop
 then begin
-$a.visited\_outer\_loop:=\mathrm{true}$
-${\bf if}$ not ($a.equal_{E}$ or $a.equal_{F}$)
+${b}.visited\_outer\_loop:=\mathrm{true}$
+$\bf if$} not (${b}.equal_{E}$ or ${b}.equal_{F}$)
 then return $\mathrm{false}$
-${\bf for}$ ${\bf each}$ vertex $b\in Adj(a)$ do
-${\bf if}$ not (($a.equal_{E}$ and $\mathrm{INNER-LOOP-E}(AUX[E;F],b)$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],b)$
-or ($a.equal_{F}$ and $\mathrm{INNER-LOOP-F}(AUX[E;F],b)$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],b))$
+$\bf for$} $\bf each$} vertex ${c}\in Adj({b})$ do
+$\bf if$} not $($${b}.equal_{E}$ and $\mathrm{INNER-LOOP-E}(AUX[E;F],{c})$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],{c})$
+or (${b}.equal_{F}$ and $\mathrm{INNER-LOOP-F}(AUX[E;F],{c})$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],{c}))$
 then return $\mathrm{false}$
@@
 output $(\mathrm{true})$
-function $\mathrm{INNER-LOOP-E}(AUX[E;F],a):\;\mathrm{Boolean}$
-$a.visited\_inner\_loop_{E}:=\mathrm{true}$;
-if $a.keeps_{E}$
-${\bf then}$ for each vertex $b\in Adj(a)$ do
+function $\mathrm{INNER-LOOP-E}(AUX[E;F],{b}):\;\mathrm{Boolean}$
+${b}.visited\_inner\_loop_{E}:=\mathrm{true}$;
+if ${b}.keeps_{E}$
+$\bf then$} for each vertex ${c}\in Adj({b})$ do
 […]
 ${\bf then}\;{\bf begin}$
-$a.visited\_inner\_loop_{E}:=\mathrm{false}$;
+${b}.visited\_inner\_loop_{E}:=\mathrm{false}$;
 ${\bf return}\;\mathrm{false}$
@@
 ${\bf else}\;{\bf begin}$
-$a.visited\_inner\_loop_{E}:=\mathrm{false}$;
+${b}.visited\_inner\_loop_{E}:=\mathrm{false}$;
 ${\bf return}\;\mathrm{false}$
 end
-$a.visited\_inner\_loop_{E}:=\mathrm{false}$;
+${b}.visited\_inner\_loop_{E}:=\mathrm{false}$;
 ${\bf output}(\mathrm{true})$
-function $\mathrm{INNER-LOOP-F}(AUX[E;F],a):\;\mathrm{Boolean}$
-$a.visited\_inner\_loop_{F}:=\mathrm{true}$;
-if $a.keeps_{F}$
-${\bf then}$ for each vertex $b\in Adj(a)$ do
+function $\mathrm{INNER-LOOP-F}(AUX[E;F],{b}):\;\mathrm{Boolean}$
+${b}.visited\_inner\_loop_{F}:=\mathrm{true}$;
+if ${b}.keeps_{F}$
+$\bf then$} for each vertex ${c}\in Adj({b})$ do
 […]
 ${\bf then}\;{\bf begin}$
-$a.visited\_inner\_loop_{F}:=\mathrm{false}$;
+${b}.visited\_inner\_loop_{F}:=\mathrm{false}$;
 ${\bf return}\;\mathrm{false}$
@@
 ${\bf else}\;{\bf begin}$
-$a.visited\_inner\_loop_{F}:=\mathrm{false}$;
+${b}.visited\_inner\_loop_{F}:=\mathrm{false}$;
 ${\bf return}\;\mathrm{false}$
 end
-$a.visited\_inner\_loop_{F}:=\mathrm{false}$;
+${b}.visited\_inner\_loop_{F}:=\mathrm{false}$;
 ${\bf output}(\mathrm{true})$
+By direct inspection of Definition 12.116, we get the following result.
 Lemma 12.117.
-Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. NPDFS outputs true on $AUX[E;F]$ if and only if for each $a\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
+Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. NPDFS outputs true on $AUX[E;F]$ if and only if for each ${b}\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
   • 1.
-    $a.equal_{E}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{E}=\mathrm{true}$.
+    ${b}.equal_{E}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{E}=\mathrm{true}$.
   • 2.
-    $a.equal_{F}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
-Proof.
-By direct inspection of Definition 12.106. ∎
+    ${b}.equal_{F}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
+We get the following lemma by direct inspection of Definition 12.116.
 Lemma 12.118.
@@
   • 1.
-    Algorithm NPDFS on $AUX[E;F]$ in the outer loop visits once each vertex $a\in C\langle E\cup F,\emptyset\rangle$.
+    Algorithm NPDFS on $AUX[E;F]$ in the outer loop visits once each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$.
   • 2.
-    For each vertex $a\in C\langle E\cup F,\emptyset\rangle$ visited in the outer loop,
-    • (a)
-      $a.equal_{E}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $b$ to $a$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{E}=\mathrm{true}$, or
+    For each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ visited in the outer loop,
     • (b)
-      $a.equal_{F}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $b$ to $a$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
+      ${b}.equal_{E}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from c{c} to b{b} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{E}=\mathrm{true}$, or
+    • (c)
+      ${b}.equal_{F}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from c{c} to b{b} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
   • 3.
@@
   • 4.
-    Algorithm NPDFS on $AUX[E;F]$ visits at most once each vertex $a\in C\langle E\cup F,\emptyset\rangle$ in the outer loop.
+    Algorithm NPDFS on $AUX[E;F]$ visits at most once each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ in the outer loop.
   • 5.
-    There exists a vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that Algorithm NPDFS on $AUX[E;F]$ visits $a\in C\langle E\cup F,\emptyset\rangle$ in the outer loop, and Conditions 2.(a) and 2.(b) are false for $a$.
+    There exists a vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ such that Algorithm NPDFS on $AUX[E;F]$ visits ${b}\in C\langle E\cup F,\emptyset\rangle$ in the outer loop, and Conditions 2.(a) and 2.(b) are false for b{b}.
   • 6.
     Algorithm NPDFS on $AUX[E;F]$ outputs false.
-Proof.
-By direct inspection of Definition 12.106. ∎
 Example 12.119.
@@
 Example 12.120.
-We now continue Examples 3.18, 5.51, 7.72, and 10.96.
+We now continue Examples 3.19, 5.52, 7.73, and 10.106.
 We run Algorithm NPDFS on $AUX[E;F]$.
@@
 Proof.
-Recall that $n=size(E)+size(F)$.
-By Statement 4.27, $|C\langle E\cup F,\emptyset\rangle|\leq n$.
-By Statement 6.63, $|A[E;F]|\leq n$.
-The running time of depth first search on $AUX[E;F]$ is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$ [8, Section 22.3].
+Recall that $n=\mathrm{size}(E)+\mathrm{size}(F)$.
+By Statement 4.28, $|C\langle E\cup F,\emptyset\rangle|\leq n$.
+By Statement 6.64, $|A[E;F]|\leq 2n$.
+The running time of depth first search on $AUX[E;F]$ is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$ [13, Section 22.3].
 Consequently, the running time of depth first search on $AUX[E;F]$ is $O(n)$.
@@
 Consequently, the outer loop on $AUX[E;F]$ takes $O(n)$ time.
-Function INNER-LOOP-E on $AUX[E;F]$ and any vertex $a$ implements a prefix of a depth first search on $AUX[E;F]$ starting from $a$.
+Function INNER-LOOP-E on $AUX[E;F]$ and any vertex b{b} implements a prefix of a depth first search on $AUX[E;F]$ starting from b{b}.
 Accordingly the running time of function INNER-LOOP-E is $O(n)$.
 Similarly, the running time of function INNER-LOOP-F is $O(n)$.
-By Statement 4.27, $|C\langle E\cup F,\emptyset\rangle|\leq n$, therefore INNER-LOOP-E and INNER-LOOP-F run at most $n$ times.
+By Statement 4.28, $|C\langle E\cup F,\emptyset\rangle|\leq n$, therefore INNER-LOOP-E and INNER-LOOP-F run at most $n$ times.
 Accordingly, Algorithm NPDFS runs on $AUX[E;F]$ in $O(n^{2})$ time.
@@
   • (iii)
-    GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle)$.
+    GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
   • (iv)
-    For each $a\in C\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $a$.
+    For each ${b}\in C\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant b{b}.
   • (v)
-    For each $a\in C\langle E\cup F,\emptyset\rangle$,
-    • $a.equal_{E}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{E}=\mathrm{true}$, or
-    • $a.equal_{F}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
+    For each ${b}\in C\langle E\cup F,\emptyset\rangle$,
+    • (a)
+      ${b}.equal_{E}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{E}=\mathrm{true}$, or
+    • (b)
+      ${b}.equal_{F}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
   • (vi)
@@
 Recall that we have run Algorithm CAD on $\Sigma$, $E$, and $F$.
-By Proposition 4.47, we have obtained in $O(n\,\mathrm{log}\,n)$ time the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
+By Proposition 4.48, we have obtained in $O(n\,\mathrm{log}\,n)$ time the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes ${b}.equal_{E}$, ${b}.equal_{F}$, ${b}.keeps_{E}$, and ${b}.keeps_{F}$.
 3.
@@
 Example 12.124.
-We now continue Examples 3.17, 5.50, 7.71, 10.95.
-and 12.109.
+We now continue Examples 3.18, 5.51, 7.72, 10.105.
+and 12.119.
 Recall that we run Algorithm NPDFS on the auxiliary dpwpa $AUX[E;F]$, and Algorithm NPDFS outputs true to the question whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
@@
 Example 12.125.
-We now continue Examples 3.18, 5.51, 7.72, 10.96, and 12.110.
+We now continue Examples 3.19, 5.52, 7.73, 10.106, and 12.120.
 Recall that we run Algorithm NPDFS on the auxiliary dpwpa $AUX[E;F]$, and Algorithm NPDFS outputs false to the question whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
@@
 In this main case, $E$ and $F$ are GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total.
+We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.48.
+We showed that $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class.
+By direct inspection of $INC(E\cup F,E)$ stored in a red-black tree and $INC(E\cup F,F)$ stored in a red-black tree we decide in $O(n)$ time whether each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class.
+Thence we decide in $O(n\,\mathrm{log}\,n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
+Example 3.20 and its sequels, Examples 5.53 and 13.127, in addition Example 3.21 and its sequels, Examples 5.54 and 13.128 exemplify this case.
+First we note that we can also apply the decision algorithm of Champavère et al [10].
+Lemma 9.97 and Proposition 9.94, and Proposition 9.92 imply the following theorem.
 Theorem 13.126.
@@
 $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
-Proof.
-Lemma 9.87 and Proposition 9.84, and Proposition 9.82 imply the Theorem. ∎
 We present a faster and simpler algorithm for solving our decidability problem.
-We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.47.
-We showed that $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class.
-By direct inspection of $INC(E\cup F,E)$ stored in a red-black tree, and $INC(E\cup F,F)$ stored in a red-black tree we decide in $O(n)$ time whether each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class.
-Thence we decide in $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
-Example 3.19 and its sequels, Examples 5.52, 13.117, in addition Example 3.20 and its sequels, Examples 5.53, 13.118.
-exemplify this case.
 Example 13.127.
-We now continue Examples 3.19 and 5.52. Recall that
+We now continue Examples 3.20 and 5.53. Recall that
+  • 1.
     both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total,
-  • #$\#/_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
-  • \$$\$/_{\Theta\langle E,F\rangle}=\$/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
-  • £$\mathsterling/_{\Theta\langle E,F\rangle}=\mathsterling/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
-  • #$\#/_{\Theta\langle F,E\rangle}=\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
-  • \$$\$/_{\Theta\langle F,E\rangle}\subset\$/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
-  • £$\mathsterling/_{\Theta\langle F,E\rangle}=\mathsterling/_{\Theta\langle{E\cup F},\emptyset\rangle}$, and
+  • 2.
+    $_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
+  • 3.
+    $\mathdollar/_{\Theta\langle E,F\rangle}=\mathdollar/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
+  • 4.
+    $\mathsterling/_{\Theta\langle E,F\rangle}=\mathsterling/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
+  • 5.
+    $_{\Theta\langle F,E\rangle}=\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
+  • 6.
+    $\mathdollar/_{\Theta\langle F,E\rangle}\subset\mathdollar/_{\Theta\langle{E\cup F},\emptyset\rangle}$,
+  • 7.
+    $\mathsterling/_{\Theta\langle F,E\rangle}=\mathsterling/_{\Theta\langle{E\cup F},\emptyset\rangle}$, and
+  • 8.
     $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
 Example 13.128.
-We now continue Examples 3.20 and 5.53. Recall that
+We now continue Examples 3.21 and 5.54. Recall that
+  • 1.
     both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total,
+  • 2.
     the only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is $W$,
-  • #$\#/_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and #$\#/_{\Theta\langle{F},E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$, and
+  • 3.
+    $_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and $_{\Theta\langle{F},E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$, and
+  • 4.
     $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$.
@@
 $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if for each $p\in W$, $\left(p/_{\Theta\langle{E\cup F},\emptyset\rangle}=p_{\Theta\langle{E,F\rangle}}\mbox{ or }p/_{\Theta\langle{E\cup F},\emptyset\rangle}=p/_{\Theta\langle{F},E\rangle}\right)$.
-Justified by the above observations and by Statement 6.60, we compute as follows.
+Justified by the above observations and by Statement 6.61, we compute as follows.
 Recall that $NUM(E\cup F,E)$ stored in a red-black tree $RBT3$, and $NUM(E\cup F,F)$ stored in a red-black tree $RBT4$.
-Alternately in a coordinated way, we traverse the red-black trees $RBT3$ and $RBT4$ in in-order traversal, For each $a\in C\langle E\cup F,\emptyset\rangle$, we read the pair $(a,k)$, where $k=|\{b\in C\langle E,F\rangle\mid b\subseteq a\,\}|$, and the pair $(a,\ell)$, where =|$\ell=|\{b\in C\langle F,E\rangle\mid b\subseteq a\,\}|$.
+Alternately in a coordinated way, we traverse the red-black trees $RBT3$ and $RBT4$ in in-order traversal, For each ${b}\in C\langle E\cup F,\emptyset\rangle$, we read the pair (b,k)({b},k), where $k=|\{{c}\in C\langle E,F\rangle\mid{c}\subseteq{b}\,\}|$, and the pair $({b},\ell)$, where $\ell=|\{{c}\in C\langle F,E\rangle\mid{c}\subseteq{b}\,\}|$.
 In this way, we decide in $O(n)$ time whether each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class, see Definition 4.27.
-If the answer is ’yes’, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, otherwise, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
+If the answer is “yes”, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, otherwise, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
 ∎
@@
 In this main case, $E$ and $F$ are GTESs over a signature $\Sigma$ such that $\Sigma$ has a symbol of arity at least $2$, and $R\langle E,F\rangle$ or $R\langle F,E\rangle$ are total.
+Without loss of generality, we may assume that $R\langle E,F\rangle$ is total from now on throughout this section.
+Example 3.22 and its sequels, Examples 5.55, 7.74, 10.110, 14.135, and 14.142 plus Example 3.23, and its sequels, Examples 5.56, 7.75, 10.111, 14.136, and 14.143 present this case.
+We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.48.
+First we note that we can also apply the decision algorithm of Champavère et al [10].
 Theorem 14.131.
@@
 We present a faster and simpler decision algorithm.
-Without loss of generality, we may assume that $R\langle E,F\rangle$ is total from now on throughout this section.
-Example 3.21 and its sequels, Examples 5.54, 7.73, 10.100, 14.125, and 14.132 plus Example 3.22, and its sequels, Examples 5.55, 7.74, 10.101, 14.126, and 14.133 present this case.
-We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.47.
 14.1 Partial Depth First Search on the Auxiliary dpwpa
@@
 { Output is true if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$, false otherwise. }
-var $a:C\langle E\cup F,\emptyset\rangle$;
+var ${b}:C\langle E\cup F,\emptyset\rangle$;
 $\mathrm{PDFS}(G\langle E,F\rangle):\;\mathrm{Boolean}$
-for each vertex $a\in C\langle E\cup F,\emptyset\rangle$ do
+for each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ do
 begin
-if not $a.equal_{E}$ and not $a.equal_{F}$
+if not ${b}.equal_{E}$ and not ${b}.equal_{F}$
 […]
-{  if $a\not\in C\langle E,F\rangle$ and $a\not\in C\langle F,E\rangle$ then GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
+{  if ${b}\not\in C\langle E,F\rangle$ and ${b}\not\in C\langle F,E\rangle$ then GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
 }
-$a.visited:=\mathrm{false}$
+${b}.visited:=\mathrm{false}$
 end
-for each vertex $a\in C\langle E\cup F,\emptyset\rangle$ do
-if not $a.equal_{E}$ and not $a.visited$
-{ if $a\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$ and $a$ has not been visited yet }
-${\bf then}$ ${\bf for}$ ${\bf each}$ vertex $b\in Adj(a)$ do
-${\bf if}$ not $b.visited$
-${\bf then}$ if not $\mathrm{PDFS-VISIT}(AUX[E;F],b)$ then
+for each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ do
+if not ${b}.equal_{E}$ and not b.visited{b}.visited
+{ if ${b}\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$ and b{b} has not been visited yet }
+$\bf then$} $\bf for$} $\bf each$} vertex ${c}\in Adj({b})$ do
+$\bf if$} not c.visited{c}.visited
+$\bf then$} if not $\mathrm{PDFS-VISIT}(AUX[E;F],{c})$ then
 ${\bf return}\;\mathrm{false}$
 ${\bf output}(\mathrm{true})$
-${\bf function}\;\mathrm{PDFS-VISIT}(AUX[E;F],a):\;\mathrm{Boolean}$
-$a.visited:=\mathrm{true}$;
-if $a.keeps_{F}=\mathrm{true}$
+${\bf function}\;\mathrm{PDFS-VISIT}(AUX[E;F],{b}):\;\mathrm{Boolean}$
+${b}.visited:=\mathrm{true}$;
+if ${b}.keeps_{F}=\mathrm{true}$
 $\bf then$} begin
-for each vertex $b\in Adj(a)$ do
-${\bf if}$ not $b.visited$
-${\bf then}$ ${\bf if}$ ${\bf not}$ $\mathrm{PDFS-VISIT}(AUX[E;F],b)$
+for each vertex ${c}\in Adj({b})$ do
+$\bf if$} not c.visited{c}.visited
+$\bf then$} $\bf if$} $\bf not$} $\mathrm{PDFS-VISIT}(AUX[E;F],{c})$
 $\bf then$} $\bf return$} $\mathrm{false}$
@@
 ${\bf output}(\mathrm{true})$
-When Algorithm PDFS starts the search from a vertex $a\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$, it does not set the attribute $a.visited$ to true.
-Consequently if for this vertex $a\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$, there exists a cycle leading from the vertex $a$ to itself, then Algorithm PDFS visits vertex $a$ twice.
+When Algorithm PDFS starts the search from a vertex ${b}\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$, it does not set the attribute b.visited{b}.visited to true.
+Consequently if for this vertex ${b}\in C\langle E\cup F,\emptyset\rangle\setminus C(E,F)$, there exists a cycle leading from the vertex b{b} to itself, then Algorithm PDFS visits vertex b{b} twice.
 We run Algorithm PDFS on the auxiliary dpwpa $AUX[E;F]$.
@@
 [13] Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and let $E$ and $F$ be GTESs over $\Sigma$ such that $R\langle E,F\rangle$ is total.
-Let $a\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$.
-(i) During the run of Algorithm PDFS on $AUX[E;F]$, Function $\mathrm{PDFS-VISIT}$ on $(AUX[E;F],a)$ visits at most once those vertices $b\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from $a$ to $b$.
-(ii) During the run of Algorithm PDFS on $AUX[E;F]$, Function $\mathrm{PDFS-VISIT}$ visits at most once those vertices $b\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex $a\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to $b$.
+Let ${b}\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$.
+(i) During the run of Algorithm PDFS on $AUX[E;F]$, Function $\mathrm{PDFS-VISIT}$ on (AUX[E;F],b)(AUX[E;F],{b}) visits at most once those vertices ${c}\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from b{b} to c{c}.
+(ii) During the run of Algorithm PDFS on $AUX[E;F]$, Function $\mathrm{PDFS-VISIT}$ visits at most once those vertices ${c}\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex ${b}\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to c{c}.
 Proof.
@@
 The running time of PDFS is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$ [13].
-Recall that $n=size(E)+size(F)$.
-By Definition 4.26 and Statement 4.45, $|C\langle E\cup F,\emptyset\rangle|\leq|W|\leq n$.
-By Statement 6.63 $|A[E;F]|\leq n$.
+Recall that $n=\mathrm{size}(E)+\mathrm{size}(F)$.
+By Definition 4.27 and Statement 4.46, $|C\langle E\cup F,\emptyset\rangle|\leq|W|\leq n$.
+By Statement 6.64 $|A[E;F]|\leq 2n$.
 Hence $|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|\leq 3n$.
@@
 Example 14.135.
-We now continue Examples 3.21, 5.54, 7.73, and 10.100.
-First Algorithm PDFS decides whether for each vertex $a\in C\langle E\cup F,\emptyset\rangle$, it holds that $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$, that is, $\left(a.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$.
-Recall that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$.
+We now continue Examples 3.22, 5.55, 7.74, and 10.110.
+First Algorithm PDFS decides whether for each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, it holds that ${b}\in C\langle E,F\rangle$ or ${b}\in C\langle F,E\rangle$, that is, $\left({b}.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$.
+Recall that $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$.
 As a result
-(#$\left(\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}\;\,\mbox{and}\;\,\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}\right)=\mathrm{true}$.
+$\left(\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}\;\,\mbox{and}\;\,\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}\right)=\mathrm{true}$.
 Consequently, Algorithm PDFS outputs false.
 Example 14.136.
-We now continue Examples 3.22, 5.55, 7.74, and 10.101. Recall that there exist one path of positive length and two cycles in $AUX[E;F]$.
-The first cycle: Its length is $1$, and it starts from the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+We now continue Examples 3.23, 5.56, 7.75, and 10.111. Recall that there exist one path of positive length and two cycles in $AUX[E;F]$.
+The first cycle: Its length is $1$, and it starts from the vertex $_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$.
+By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant $_{\Theta\langle E\cup F,\emptyset\rangle}$.
 Recall that
-#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
-The path of positive length: Its length is $1$, and it starts from the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+$_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$, $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$, and $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
+The path of positive length: Its length is $1$, and it starts from the vertex $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$.
+By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant $_{\Theta\langle E\cup F,\emptyset\rangle}$.
 Recall that
-\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
-The second cycle: Its length is $1$, and it starts from the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+$\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
+The second cycle: Its length is $1$, and it starts from the vertex $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ and leads to $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of the constant $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$.
 Recall that
-\$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
-First Algorithm PDFS decides whether for each $a\in C\langle E\cup F,\emptyset\rangle$, it holds that $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$, that is, $\left(a.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$.
-Observe that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$.
+$\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$, and $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{false}$.
+First Algorithm PDFS decides whether for each ${b}\in C\langle E\cup F,\emptyset\rangle$, it holds that ${b}\in C\langle E,F\rangle$ or ${b}\in C\langle F,E\rangle$, that is, $\left({b}.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$.
+Observe that $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$.
 Thus
-(#$\left(\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}\;\mbox{or}\;\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}\right)=\mathrm{true}$.
-Observe that \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$ and \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$. As a result
-(\$$\left(\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}\;\mbox{or}\;\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}\right)=\mathrm{true}$.
+$\left(\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}\;\mbox{or}\;\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}\right)=\mathrm{true}$.
+Observe that $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$ and $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$. As a result
+$\left(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}\;\mbox{or}\;\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}\right)=\mathrm{true}$.
 Algorithm PDFS examines the vertices of $AUX[E;F]$, it looks for vertices from which it can start the search.
-To this end, Algorithm PDFS decides whether for each $a\in C\langle E\cup F,\emptyset\rangle$, it holds that $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$, that is, $\left(a.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$.
-First, Algorithm PDFS examines the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-It observes that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$.
-Then Algorithm PDFS starts the search from #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-It goes along the cycle (#$(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})$.
-Algorithm PDFS observes that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
-Second, Algorithm PDFS examines the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$.
-Since \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, Algorithm PDFS does not start the search starting from \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+To this end, Algorithm PDFS decides whether for each ${b}\in C\langle E\cup F,\emptyset\rangle$, it holds that ${b}\in C\langle E,F\rangle$ or ${b}\in C\langle F,E\rangle$, that is, $\left({b}.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$.
+First, Algorithm PDFS examines the vertex $_{\Theta\langle E\cup F,\emptyset\rangle}$.
+It observes that $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$.
+Then Algorithm PDFS starts the search from $_{\Theta\langle E\cup F,\emptyset\rangle}$.
+It goes along the cycle $(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})$.
+Algorithm PDFS observes that $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$.
+Second, Algorithm PDFS examines the vertex $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$.
+Since $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, Algorithm PDFS does not start the search starting from $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$.
 Consequently, Algorithm PDFS outputs true.
@@
   • (v)
-    for each $a\in C\langle E\cup F,\emptyset\rangle$,
-    • $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$ and
-    • if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
+    for each ${b}\in C\langle E\cup F,\emptyset\rangle$,
+    • (a)
+      ${b}\in C\langle E,F\rangle$ or ${b}\in C\langle F,E\rangle$ and
+    • (b)
+      if ${b}\not\in C\langle E,F\rangle$, then for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing c{c}.
   • (vi)
-    for each $a\in C\langle E\cup F,\emptyset\rangle$,
-    • $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$ and
-    • if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
+    for each ${b}\in C\langle E\cup F,\emptyset\rangle$,
+    • (a)
+      ${b}.equal_{E}=\mathrm{true}$ or ${b}.equal_{F}=\mathrm{true}$ and
+    • (b)
+      if ${b}\not\in C\langle E,F\rangle$, then for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
 Proof.
@@
 $\mathrm{(v)}\Leftrightarrow\mathrm{(vi)}$: By Statement 4 in Lemma 8.77. ∎
+We get the following lemma by direct inspection of Definition 14.132 and by the equivalence $\mathrm{(iv)}\Leftrightarrow\mathrm{(vi)}$ in Lemma 14.137.
 Lemma 14.138.
@@
   • (a)
-    For each vertex $a\in C\langle E\cup F,\emptyset\rangle$, $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$, and
+    For each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, ${b}.equal_{E}=\mathrm{true}$ or ${b}.equal_{F}=\mathrm{true}$, and
   • (b)
-    Algorithm PDFS on $AUX[E;F]$ visits exactly once each vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex $b\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to $a$.
-For each visited vertex $a$, $a.keeps_{F}=\mathrm{true}$.
+    Algorithm PDFS on $AUX[E;F]$ visits exactly once each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex ${c}\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to b{b}.
+For each visited vertex b{b}, ${b}.keeps_{F}=\mathrm{true}$.
   • (c)
@@
   • (d)
-    There exists a vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$.
+    There exists a vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ such that ${b}.equal_{E}=\mathrm{true}$ or ${b}.equal_{F}=\mathrm{true}$.
   • (e)
-    Algorithm PDFS on $AUX[E;F]$ visits at most once each vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex $b\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to $a$.
-Moreover, Algorithm PDFS visits a vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that $a.keeps_{F}=\mathrm{false}$.
-Proof.
-By direct inspection of Definition 14.122 and by the equivalence $\mathrm{(iv)}\Leftrightarrow\mathrm{(vi)}$ in Lemma 14.127. ∎
+    Algorithm PDFS on $AUX[E;F]$ visits at most once each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex ${c}\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to b{b}.
+Moreover, Algorithm PDFS visits a vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ such that ${b}.keeps_{F}=\mathrm{false}$.
+We get the following result by the equivalence $\mathrm{(i)}\Leftrightarrow\mathrm{(iv)}$ in Lemma 14.137 and by Lemma 14.138.
 Lemma 14.139.
@@
 $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if Algorithm PDFS on $AUX[E;F]$ outputs true.
-Proof.
-By the equivalence $\mathrm{(i)}\Leftrightarrow\mathrm{(iv)}$ in Lemma 14.127 and by Lemma 14.128. ∎
 Lemma 14.140.
@@
 Recall that we have run Algorithm CAD on $\Sigma$, $E$, and $F$.
-By Proposition 4.47, we got the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
+By Proposition 4.48, we got the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes ${b}.equal_{E}$, ${b}.equal_{F}$, ${b}.keeps_{E}$, and ${b}.keeps_{F}$.
 Then we run Algorithm PDFS on $AUX[E;F]$ in $O(n)$ time.
@@
 Example 14.142.
-We now continue Examples 3.21, 7.73, 10.100, and 14.125.
+We now continue Examples 3.22, 7.74, 10.110, and 14.135.
 By the last lines of Example 10.110, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
@@
 For this reason, by Lemma 14.137, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
-By the last line of Example, 14.125 Algorithm PDFS outputs false to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
+By the last line of Example, 14.135 Algorithm PDFS outputs false to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
 As a consequence, our decision algorithm also outputs false to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
 Example 14.143.
-We now continue Examples 3.22, 5.55, 7.74, 10.101, and 14.126.
+We now continue Examples 3.23, 5.56, 7.75, 10.111, and 14.136.
 By the last line of Example 10.111, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
@@
 Consequently, by Lemma 14.137, $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
-By the last line of Example, 14.126 Algorithm PDFS outputs true to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
+By the last line of Example, 14.136 Algorithm PDFS outputs true to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
 As a consequence, our decision algorithm also outputs true to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
@@
 We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.48.
-Example 3.23 and its sequel Example 5.56, furthermore, Example 3.24 and its sequels Examples 5.57 and 15.134 demonstrate this case.
+We show that $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $E\subseteq\Theta\langle F,E\rangle$ or $F\subseteq\Theta\langle E,F\rangle$.
+By direct inspection of the relations $\tau\langle E,F\rangle$ and $\rho\langle F,E\rangle$, we decide whether $E\subseteq\Theta\langle F,E\rangle$.
+Similarly, by direct inspection of the relations $\tau\langle F,E\rangle$ and $\rho\langle E,F\rangle$, we decide whether $E\subseteq\Theta\langle F,E\rangle$.
+Example 3.24 and its sequel Example 5.57, furthermore, Example 3.25 and its sequels Examples 5.58 and 15.144 demonstrate this case.
 Example 15.144.
-We continue Examples 3.24 and 5.57. We recall that
-$f(\#,\mathsterling)\mathop{\leftrightarrow}^{*}_{E\cup F}f(\$,\flat)$,  $(f(\#,\mathsterling),f(\$,\flat))\not\in\mathop{\leftrightarrow}^{*}_{E}$,   and   $(f(\#,\mathsterling),f(\$,\flat))\not\in\mathop{\leftrightarrow}^{*}_{F}$.
-The idea of the ()($\Rightarrow$)( ⇒ ) direction of the proof of the next theorem comes from the proof of the above observation.
+We continue Examples 3.25 and 5.58. We recall that
+[…]
+The idea of the $(\Rightarrow)$ direction of the proof of the next theorem comes from the proof of the above observation.
 Theorem 15.145.
@@
 Proof.
-()($\Rightarrow$)( ⇒ ) We proceed by contrapositive.
+$(\Rightarrow)$ We proceed by contrapositive.
 Assume that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$.
@@
 $\sigma(\sigma(v,l_{1},\ldots,l_{1}),\sigma(z,l_{2},\ldots,l_{2}),v,\ldots,v)=s_{0}\mathop{\leftrightarrow}_{[\alpha_{1},p_{1}\doteq q_{1}]}s_{1}\mathop{\leftrightarrow}_{[\alpha_{2},p_{2}\doteq q_{2}]}\cdots\mathop{\leftrightarrow}_{[\alpha_{n},p_{n}\doteq q_{n}]}s_{n}=t,\;n\geq 1.$ (3)
-[…]
+be a sequence of $E$-rewrite steps. Then for each $i\in\{\,1,\ldots,n\,\}$, $\alpha_{i}$ is not a prefix of $11$.
 Proof.
-[…]
+By induction on $n$, we show that for each $i\in\{\,1,\ldots,n\,\}$, $\alpha_{i}$ is not a prefix of $11$.
 Base Case: $n=1$. By the definition of $v$, $v\not\in\bigcup W/_{\mathop{\leftrightarrow}^{*}_{E}}$, hence
@@
 thus $\alpha_{1}$ is not a prefix of $11$.
-[…]
-[…]
+Induction Step: Let $n\geq 2$ and let us assume that we have shown the statement for n1n-1.
+Then by the induction hypothesis, for each $i\in\{\,1,\ldots,n-1\,\}$, $\alpha_{i}$ is not a prefix of $11$.
+Accordingly we have $v=s_{0}|_{11}\mathop{\leftrightarrow}^{*}_{E}s_{n}|_{11}$.
+Consequently, by $v\not\in\bigcup W/_{\mathop{\leftrightarrow}^{*}_{E}}$, we have $s_{n}|_{11}\not\in\bigcup W/_{\mathop{\leftrightarrow}^{*}_{E}}$.
+Thus $s_{n}|_{11}$ is not a subterm of $p_{n}\in W$.
+As a consequence, $\alpha_{n}$ is not a prefix of $11$.
+Then the induction step is complete.
+∎
+Consider the sequence (3) of $E$ rewrite steps.
+By Claim 15.146, for each $i\in\{\,1,\ldots,n\,\}$, $\alpha_{i}$ is not a prefix of $11$.
+Consequently $t=\sigma(\sigma(t_{1},t_{2},\ldots t_{k}),p_{1},\ldots,p_{k-1})$ for some $p_{1},\ldots,p_{k-1}\in T_{\Sigma}$, and for each $j\in\{\,2,\ldots,k\,\}$, $l_{1}\mathop{\leftrightarrow}^{*}_{E}t_{j}$.
+Recall that $(l_{1},r_{1})\not\in\mathop{\leftrightarrow}^{*}_{E}$.
+Therefore for each $j\in\{\,2,\ldots,k\,\}$, $t_{j}\neq r_{1}$.
+Hence
 $t\neq\sigma(\sigma(v,r_{1},\ldots,r_{1}),\sigma(z,r_{2},\ldots,r_{2}),v,\ldots,v)$.
@@
 Consequently, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
-()($\Leftarrow$)( ⇐ ) If $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{F}$, and hence $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
+$(\Leftarrow)$ If $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{F}$, and hence $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
 If $F\subseteq\mathop{\leftrightarrow}^{*}_{E}$, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}$, and hence $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
 ∎
-Consequence 15.137.
+Consequence 15.147.
 Let $\Sigma$ be a signature such that $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$. Let $E$ and $F$ be GTESs over $\Sigma$ such that $R\langle E,F\rangle$ and $R\langle F,E\rangle$ are not total. Then
@@
 Proof.
-By Consequence 15.137, we proceed as follows.
+By Consequence 15.147, we proceed as follows.
 Recall that we have run Algorithm CAD on $\Sigma$, $E$, and $F$.
 By Proposition 4.48, we got the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$ in $O(n\,\mathrm{log}\,n)$ time.
-First we decide whether $E\subseteq\Theta\langle F,E\rangle$. By Consequence 15.137 and Proposition 4.25,
+First we decide whether $E\subseteq\Theta\langle F,E\rangle$. By Proposition 4.26,
 $E\subseteq\Theta\langle F,E\rangle$ if and only if
@@
 Second, similarly to above, we decide whether $F\subseteq\Theta\langle E,F\rangle$.
-We call above the functions $\mathrm{FIND}_{\rho\langle F,E\rangle}$ and $\mathrm{FIND}_{\rho\langle E,F\rangle}$ a total of at most $2n$ times, where $n=size(E)+size(F)$.
+By Consequence 15.147
+$\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $E\subseteq\Theta\langle F,E\rangle$ or $F\subseteq\Theta\langle E,F\rangle$.
+We call above the functions $\mathrm{FIND}_{\rho\langle F,E\rangle}$ and $\mathrm{FIND}_{\rho\langle E,F\rangle}$ a total of at most $2n$ times, where $n=\mathrm{size}(E)+\mathrm{size}(F)$.
 These callings, by the main result of [3], require $O(n)$ time.
@@
 Let $E$ and $F$ be GTESs over a signature $\Sigma$.
-It is decidable in $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n=size(E)+size(F)$.
+It is decidable in $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n=\mathrm{size}(E)+\mathrm{size}(F)$.
 Proof.
 We run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.48, we get the following output in $O(n\,\mathrm{log}\,n)$ time:
+  • 1.
     the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$;
+  • 2.
     the set of nullary symbols $C\langle E,F\rangle$, $C\langle F,E\rangle$, and $C\langle E\cup F,\emptyset\rangle$;
+  • 3.
     $INC(E\cup F,E)$ stored in a red-black tree saved in a variable $RBT1$;
+  • 4.
     $INC(E\cup F,F)$ stored in a red-black tree saved in a variable $RBT2$;
+  • 5.
     $NUM(E\cup F,E)$ stored in a red-black tree saved in a variable $RBT3$;
+  • 6.
     $NUM(E\cup F,F)$ stored in a red-black tree saved in a variable $RBT4$;
+  • 7.
     the GTRS $R\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT5$; each node in $RBT5$ contains a search key which is a rule in $R\langle E\cup F,\emptyset\rangle$, and satellite data $counter_{E}$ and $counter_{F}$;
-  • $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT6$; for every ordered pair $(a,b)\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component $a$, the second component, $b$ is satellite data;
+  • 8.
+    $BSTEP\langle E\cup F,\emptyset\rangle$ stored in a red-black tree saved in a variable $RBT6$; for every ordered pair $({b},{c})\in BSTEP\langle E\cup F,\emptyset\rangle$, the search key is the first component b{b}, the second component, c{c} is satellite data;
+  • 9.
     the GTRS $R\langle E,F\rangle$ stored in a red-black tree saved in a variable $RBT7$;
+  • 10.
     the GTRS $R\langle F,E\rangle$ stored in a red-black tree saved in a variable $RBT8$;
-  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
+  • 11.
+    the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes ${b}.equal_{E}$, ${b}.equal_{F}$, ${b}.keeps_{E}$, and ${b}.keeps_{F}$.
 By Lemma 4.50, we decide in $O(n)$ time whether $R\langle E,F\rangle$ is total, and whether $R\langle F,E\rangle$ is total.
+  • 1.
     If both of them are total, then we proceed as in main Case 2, see Lemma 13.130.
+  • 2.
     Otherwise,
+    • (a)
       if $\Sigma$ is a unary alphabet, then we proceed as in Main Case 1, see Lemma 12.123.
+    • (b)
       If $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total, then we proceed as in Main Case 3, see Lemma 14.141.
+    • (c)
       If $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total, then we proceed as in Main Case 4, see Lemma 15.148. ∎
 By Theorem 16.149 and Proposition 9.92, we have the following consequence.
-Consequence 16.140.
+Consequence 16.150.
 Let $E$ and $F$ be GTESs over a signature $\Sigma$.
-Then we can decide in $O(n^{2})$ time whether there exists a GTES $H$ over $\Sigma$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n=size(E)+size(F)$.
-If the answer is ’yes’, then $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$ and $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
+Then we can decide in $O(n^{2})$ time whether there exists a GTES $H$ over $\Sigma$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n=\mathrm{size}(E)+\mathrm{size}(F)$.
+If the answer is “yes”, then $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra ${\cal T}(\Sigma)$ and $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
 Otherwise $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is not a congruence on the ground term algebra ${\cal T}(\Sigma)$.
@@
 By Theorem 16.149, we decide in $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
-By Proposition 9.82, if the answer is ’yes’, then there exists a GTES $H$ over $\Sigma$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, and $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$; otherwise $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is not a congruence on the ground term algebra ${\cal T}(\Sigma)$.
+By Proposition 9.92, if the answer is “yes”, then there exists a GTES $H$ over $\Sigma$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, and $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$; otherwise $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is not a congruence on the ground term algebra ${\cal T}(\Sigma)$.
 ∎
@@
 We showed that we can decide in $O(n^{2})$ time for given GTESs $E$ and $F$ over a signature $\Sigma$ whether Statement 1 above holds.
-Here $n=size(E)+size(F)$.
+Here $n=\mathrm{size}(E)+\mathrm{size}(F)$.
 We can construct the GTES $E\cup F$ in $O(n^{2})$ time.
@@
 Main Case 2 overlaps with Main Case 1 and Main Case 3, and is a subcase of the union of Main Case 1 and Main Case 3.
-Given GTES $E$ and GTES $F$, for every $p\in W$, we can construct the deterministic bottom-up tree automata ${\cal A}$, ${\cal B}$, and ${\cal C}$ such that ${\cal A}$ recognizes $p/_{\mathop{\leftrightarrow}^{*}_{E}}$, ${\cal B}$ recognizes $p/_{\mathop{\leftrightarrow}^{*}_{F}}$, and ${\cal C}$ recognizes $p/_{\mathop{\leftrightarrow}^{*}_{E}\cup F}$, see Proposition 2.10.
-Thence, in Main Cases 1, 2, and 3, we can reduce our decidability problem to deciding the equivalence of two deterministic bottom-up tree automata.
-Champavère et al [5] presented an algorithm for deciding inclusion between two tree languages recognized by given bottom-up tree automata.
-Applying their algorithm to our decision problem in Main Cases 1, 2, and 3 yields an alternative decision algorithm, which is more sophisticated than ours, and in general takes more time than ours, see Lemmas 12.103, 13.116, and 14.121.
+[…]
 We raise the following more general problem:
-Problem 17.141.
+Problem 17.151.
 Let $D$, $E$, and $F$ be GTESs over a signature $\Sigma$.
@@
 Example 17.152.
-Consider the unary ranked alphabet $\Sigma=\Sigma_{0}=\{\,\#,\;\$,\;\mathsterling\,\}$. Let
-$D=\{\,\#\doteq\$\,\}$,  $E=\{\,\#\doteq\mathsterling\,\}$,  $F=\{\,\$\doteq\mathsterling\,\}$,  $H=D\cup E\cup F$.
+[…]
+$D=\{\,\#\doteq\mathdollar\,\}$,  $E=\{\,\#\doteq\mathsterling\,\}$,  $F=\{\,\mathdollar\doteq\mathsterling\,\}$,  $H=D\cup E\cup F$.
 Then
+  • 1.
     $ST\langle H\rangle=\Sigma_{0}$,
+  • 2.
     $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{D}\cup\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$,
-  • #$\#/_{\mathop{\leftrightarrow}^{*}_{D}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,   #$\#/_{\mathop{\leftrightarrow}^{*}_{E}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,   #$\#/_{\mathop{\leftrightarrow}^{*}_{F}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,
-  • #$\#/_{\mathop{\leftrightarrow}^{*}_{D}},\;\#/_{\mathop{\leftrightarrow}^{*}_{E}},\;\#/_{\mathop{\leftrightarrow}^{*}_{F}}$ are pairwise incomparable for the inclusion relation, and
+  • 3.
+    $_{\mathop{\leftrightarrow}^{*}_{D}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,   $_{\mathop{\leftrightarrow}^{*}_{E}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,   $_{\mathop{\leftrightarrow}^{*}_{F}}\subset\#/_{\mathop{\leftrightarrow}^{*}_{H}}$,
+  • 4.
+    $_{\mathop{\leftrightarrow}^{*}_{D}},\;\#/_{\mathop{\leftrightarrow}^{*}_{E}},\;\#/_{\mathop{\leftrightarrow}^{*}_{F}}$ are pairwise incomparable for the inclusion relation, and
+  • 5.
     for each $t\in ST\langle H\rangle$,   $t/_{\mathop{\leftrightarrow}^{*}_{D}}\subset t/_{\mathop{\leftrightarrow}^{*}_{H}}$,   $t/_{\mathop{\leftrightarrow}^{*}_{E}}\subset t/_{\mathop{\leftrightarrow}^{*}_{H}}$,   and   $t/_{\mathop{\leftrightarrow}^{*}_{F}}\subset t/_{\mathop{\leftrightarrow}^{*}_{H}}$.